A Complete Bibliography of Publications in the Journal of Cell Biology: 1975–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

−196 [GBM78]. 10 [LKKA79, SBGG78, WCC79]. 2 [Col79, DF77, JJ78]. 25 [OO79]. 5 [BAB79, GPPR76, KSS76]. ^{+} [DJM77, Kyt76a, Kyt76b, SHC^{+}79]. 2+ [BSU79, CGH79a, LFR78, RGS^{+}76, Täl78]. ^{45} [BSU79]. α [FS76]. β [NZD^{+}79].

-activated [RGS^{+}76]. -ATPase [BTS^{+}78, DJM77, SHC^{+}79]. -bromodeoxyuridine [BAB79, GPPR76, KSS76]. -bungarotoxin [FS76]. -dependent [MCP^{+}78]. -free [CGH79a]. -isoproterenol [MC79b]. -lipotropin [NZD^{+}79]. -nm [LKKA79, OO79, SBGG78, WCC79]. -phenylindole [Col79, JJ78]. -pump [MED77]. -yl [DF77].

1 [FGG78]. 100 [JF75, LG75b]. 100-Angstrom [JF75]. 125 [SFS77]. 125I [CC76, MF75, SC78, WKBL75a, BSHP79, GES79, HWAS79, HS79b, NSH78, PSC^{+}76]. 125I-diiodofluorescein [GES79]. 125I-insulin [BSHP79]. 125I-labeled [CC76, NSH78, PSC^{+}76]. 125I-ligands [HWAS79, HS79b]. 13 [SSSW77]. 13437 [LCK75]. 17beta [KAD75]. 18S [Leó76, PH75c].

2 [BE78, GS75]. 2-3H [HW75]. 21 [WG78]. 23187 [FS77, Sch75]. 28S [Leó76, PH75c].

38 [HMT79]. 3H [DH76, EKS^{+}79, FBF79, HW75, MR75a, MF75, NEK78b, NEK78a, SB76, SP76, GKS76, MZ79]. 3H-amino [MZ79]. 3T3 [AB76, ABL76, AB77, AB79, AL78, Bel77, EP79b, GDR77, GR78, GRW75, HT79a, NSP76, SVW79, TZ77].

4-glucan [SM75b]. 42-348 [MR76b]. 431 [CMC79, HMC79a, HMC79b]. 450 [FKNMT79, IP78a, MRDC77, MFKT78]. 4’6 [Col79, JJ78]. 4’6-diamidino- [Col79, JJ78]. 4S [Leó76].

5 [CDH^{+}79, DMMW76, RBC77, Sto75, WJTR77, ZJSF76]. 5-bisphosphate [SDTD^{+}79]. 5’-cyclic [DMMW76]. 5’-monophosphate [Sto75, WJTR77, ZJSF76]. 5’-nucleotidase [CDH^{+}79, RBC77]. 5S [Leó76].

60 [NCGC79].

75S [DAF77].

80S [KAB75]. 86Rubidium [Smi77].

929 [THC76]. 98S [MB77b].

A-23187 [FS77, Sch75]. A-431 [CMC79, HMC79a, HMC79b]. A-induced [ZS79b]. A-mediated [Let79]. A-mutant [SAPA77]. A-sepharose [LFR78]. A-stimulated [PELG76]. A.B.C. [RD78]. A23187 [ODS79, Sch76]. Ability [Zig77, KS79]. Absence [Cav75, Cav76, BBMB77, BSM79, CH76, JG76, VJGG79]. absorption [CB75c]. absorptive [RR77a]. Acanthamoeba [OT76, OT78, Pol76, RHHB76]. acanthias [EKS^{+}79]. acanthocephali [FMW75, FWB78, MW76]. accelerated [ES79]. accessory [CKCK78]. accretion [MKM76]. accumulate [Bon75a, Bon75b, BSM79]. accumulating [AV78]. accumulation [BGK^{+}79, CI78, DMS^{+}76, JM77, PJ76a, QLHTB79, VL78]. accumulations [JWJ75]. Acetabularia [AMBM76, FSSZ75, FKS76]. acetate [GHW75]. Acetobacter [Zaa79]. acetylated [BGK^{+}79]. Acetylcholine [BSM79, DF75, GHC78, Blo79, CP77d, FD78, FS76, FF79a, HS79a, JL79, KHV^{+}78, LC77, MM79, RV78, SRR79, VD76, WJB78]. Acetylcholinesterase [SRKM78, BG76a, CBM78, DK78a]. acetyltransferase [GNB^{+}77]. achieved [LF76]. acid [BHP75, BH78, DMB78, DM79, FBF79, GCY79, HO75, HB79, KP77c, KP77b, LLL75, RBK78, SWF78, SGV75, SDTD^{+}79, SS78d, VMR79]. acid-envelope [HO75]. acid-stained [BHP75]. acidic [SHWH^{+}77]. acids [CB75c, MZ79, SFFB75, SSC76b]. acinar [CW78b, CW78c, DPM76, HO77, IP78b, KG78, MCP^{+}78, OH78, TP76]. acrosomal [Til75, TKST78, TCT79, TK79]. acrosome [CB76, FOPY77, SS79a]. across [BD75a, BD75b, NSBS77a, NSBS77b]. ACTH [HMPB76, KSS76, SLR77]. ACTH-dependent [KSS76]. ACTH-stimulated [HMPB76]. Actin [CM78, FC77, GRW75, Kan76, KH75, Laz76, MSP78, MC78a, PCJG79, RWLC78, SS79d, TD75, Til75, TM76b, BRR78, BR79, BC79, BW78, BS79b, BS79c, BS77c, Bur78, CCTT77, CLM77, Con79, DSTW77, EL78, FTBS78, Fuk78, GP78, HS76a, HA77, HP79, HL79, IO79, Kan75, KC78, KW76, KR79a, KR79b, LTMS79, Lem79, MZ78, MT75, MCM77, OLK77, PH75b, Pol76, RGK78, SMG77, SO77, SH76, Til76a, Til76b, TKST78, Til78, TK79, WT79, Wei76b, Wei77a, WWLS79, WRP75]. actin-binding [HS76a, SH76]. actin-containing [MCM77]. Actin-like [KH75, SMG77]. actin-related [Wei76b, Wei77a]. Actinin [CP79b, FPP78, Laz76]. actinlike [MD75]. actinomycin [BDCD75, STF75]. actins [HSRP79]. Action [GMDT75, THC76, KPF75, TM79]. actions [ZJSF76]. activated [MBV77, RGS^{+}76]. activating [GJRR78]. Activation [BTS^{+}78, GW75, GJ78, Mil78, Sch75, SM77]. activator [Qui76, RP77, RR77b]. Active [PP79, CSP79, CGH79a, HD77, OB78, Pro75, RTC76]. activities [Ern75, MCC79, SM75b]. activity [AJ78, And77, BPS78, BS77e, CF76a, CF76b, CG78, FK78, FFSH77, GNB^{+}77, HLS75, JRFS75, KCH79, LJ75, Moo76a, PS79, RG78, RHDL77, SHH76, SRR76, SFFB75, Smi77, SC75a, SVW79, SS78f, TWC75, Von79, VMR79, WNQK78, WGC76, YGH78]. actomere [Til78]. Actomyosin [KNL76, MC79a]. adaptation [SPM79, ZS79a]. adaptations [SPM79]. adapted [KKKS76]. addition [BHR77, HBR77, Nan75]. adenine [RMP75, WF76]. adenine-thymine-rich [WF76]. adenocarcinoma [MHC77]. Adenosine [LG75b, ZJSF76, DMMW76, Gra75a, Gra75b, PW78, RBC77, RW75b, Sto75, WJTR77]. adenovirus [HMDH77]. adenylate [CF76a, CF76b, LJ75, RBC77]. adenylyl [PW78]. adhere [MSK^{+}79]. adherence [BB79a]. Adhesion [Jul78, MSS75, PT77, RTBE78, RGE78, BE78, CRBH78, GR78, GL79c, GM76, HLG76, Hel79, Let79, MB76, McG76, Moo76a, PDD78, Sne76a, Sne76b, UIOO76, WDB79]. adhesion-promoting [Moo76a]. adhesive [GM76, Mar77, MB76, McG76, Tak77, YYP77]. adipose [CPO77]. adjacent [NS77]. administered [OH78]. adrenal [CB75b, FFHL78, GLR77]. adrenals [KSS76]. adrenergic [BH79, Lan78]. adrenocortical [HMPB76, SLR77]. adsorption [FYS78]. Adult [WDMR77, BWDP77, DWM75]. affect [SC78]. affecting [GS75, HRL77, WGB77]. afferent [Hir77]. affinity [FKP76]. after [ALMS76, Bal79, BDCD75, BR77, EL75, ES76, FWS75, FS76, FS77, HA76, HW75, HB79, Hir77, Lar75, MG75b, NUY^{+}77, NBB77b, ODS79, Per77, SMM78, SO75, SB76, SD76, Smi77, WDM^{+}75, YB75]. against [OAB77]. age [SMJ78]. agent [DF77, Svo78]. agents [EP79a, OAB76, RW75a]. agglutinability [CT76]. agglutination [BGG^{+}75, FGG78, GL79c, vVRN76]. agglutinin [DMKR75, MJM76, SC78]. aggregates [CBM78, GD75, GDC76, GDR77, GR78, Til76b]. aggregation [BP77, DR77a, LF76, RD78, SS78d]. agilis [BMP76]. aging [LLMW77, SSR76]. Agonist [LS79a]. Agonist-induced [LS79a]. albino [KP75]. albumin [BCFS75, PSZK78b, RBM^{+}78, ST77b]. albumin-embedded [PSZK78b]. Alcian [FMR79]. alcohols [DTSM76]. aldolase [MCM77]. alga [AMBM76, BMP76, FSSZ75, SSF76, SSC76a, SSC76b, SCS76]. algae [KW76]. alkali [SM75b]. alkali-insoluble [SM75b]. allow [Bon75a]. allows [GES79]. Alpha [FPP78, BBDW76, CF79, CP79a, CP79b, FWS75, JL79, Laz76, WMP79, SB76]. Alpha-actinin [FPP78, CP79b, Laz76]. alpha-bungarotoxin [CF79, CP79a, FWS75, SB76]. alter [GPPR76]. alteration [PE79]. Alterations [PBH78, HWLC78, PS79, SFFB75, Til75, VJGG79, JLA^{+}76]. Altered [Han75a]. alternate [BK75]. alters [SM75b]. alveolar [BCFS75, EH77]. alveoli [Wil77b]. Ambystoma [LMPI76]. amebas [EWR77, EL78]. ameboid [HT79b, TRH76]. americana [Fus75a]. amine [WJB78]. amino [LLL75, SDTD^{+}79, MZ79, RBK78]. aminobutyric [FBF79]. Amniotic [CBH^{+}78]. amoeba [JJ75, RH75, MG75b]. amoeboid [CT77]. among [FJD76, GL77, GL79b, HL75, HWAS79, KS79, RTBE78, RGE78]. Amount [Cle75, JHW78, MKK79, Slu76]. amp [KP77a, AFB79, CCH78, DDS79, DS79, GDR77, GHMD77, HMPB76, MOR78, MC79b, WP75, vVRN76]. amp-induced [KP77a]. amphibian [AHB79, STF76, WJ76, ZM76]. amplification [Cav75, Cav76, DBB^{+}79]. amylase [FS77, GST79, PBJ77, WL76]. analog [PW78, WF75]. Analogous [BO78]. analogues [Gra75a, Gra75b]. analyses [Sne76a]. Analysis [MS76, TSPH78, AG75, BLTP^{+}76, Bou79, BCHN75, Bry76, CGK75, Edd75a, EL76, EP79b, EB78, FJD76, FPC76, GESG75, GTH76, JSGR78, Kri75, LLC79, MRG77, MBM^{+}79, MH75, MC78b, MR76b, MDR75, PSZK78b, RCM77, RHHB76, SMS78, SS79b, SMJ78, SHWPJ78, SB76, SSC76a, SSS77, SSS79, SBGG78, SBC76, TCPM77, WC75, vVRN76, EW77a, MCPG77]. Analytical [FRT^{+}76, RFB^{+}76, PD75]. analyzing [SS76a]. anaphase [Fus75b, IR78, MPHMT77]. Anatomy [CAS78, LG75a]. anchorage [KS79]. Andrenocorticotropin [NZD^{+}79]. anesthetic [EP79a]. anesthetic-induced [EP79a]. anesthetics [BFPU78b, NSP76]. Angstrom [JF75]. angular [AB79]. animal [BRBM77, Gre75, LE78a, LE78b, LZK75, SMJ78]. Anionic [KF79, AMS76, AH77, Far78, GTH75, HM75, SD76, SF76]. anionized [BW79]. Annulate [SH77b]. Anolis [RR75]. anomalies [BG76b]. Antagonism [Sto75, WJTR77]. antagonist [AKD75b, KAD75]. antagonist-induced [AKD75b]. anterior [Far78, FAE75]. anthocodium [SC75b]. anti [HP79, STdDT79b, SRR79, CFAS78]. anti-actin [HP79]. anti-HRP [CFAS78]. anti-plasma [STdDT79b]. anti-receptor [SRR79]. antibiotic [LLL75, SGCS75]. antibiotics [AV78, DTT75]. Antibodies [SH79a, KHV^{+}78, NT78, NVH^{+}75, PL77, SRR79]. antibody [FP76, FP78b, GLM77, Laz75, MO77, PB79, RFB^{+}76]. antidynein [GOG76]. antigen [dP75, GGGM78]. antigenic [Koe75, MM76c]. antigenicity [PL77]. antigens [AJG78, CGK75, EW77a, EW77b, MB77a, MDRY75, dP78]. antimitotic [DF77]. antimyosins [RMP^{+}78]. antiparallel [MEM79]. antisera [GJ78, WDB79]. antiserum [OAB77, FF79b]. antitubulin [GLM77]. aorta [FSW77]. aortic [FSW77]. ape [BP79]. Aplysia [GKS76]. apoprotein [AHH76]. apparatus [BBC78, BMP76, DR77a, FD78, HO77, LS79b, Sch78b, Ste75, WDBM79]. Apparent [BG76b, WGB78]. Appearance [STWB78]. applicability [SEI75]. application [FRT^{+}76, DVEM79, SMS78]. Applications [MSS75]. approach [Pfe78]. arabinoside [JM77]. arachidonic [HB79]. Arbacia [HMK79, Lon77, SEBL77]. area [PBJ77, WOPB79]. areas [SHWPJ78]. Arginine [WR77a, WR77b, WR77c]. arginine-deprived [WR77c]. arising [Her76]. arm [BBDP79]. arms [AEJL75, WM78]. Arrangement [Amo77, BCS77, VRN76]. arrangements [SS76a]. arrays [FGSM78, HG78, Ros78, WGG77b, Wei77b]. arrest [CB75a, MS75b, WS78a]. artefacts [HH78]. arterial [DR78, NSRL76, WR75c]. arteries [WR75b, WR75c, SSP76]. arthropoda [BD78]. Artifact [WP79]. artificial [HH77b]. aryl [WGC76]. Ascaris [RT76]. ascites [IO79]. ascorbic [SGV75]. Aspects [RW78, EWR77, Edd77, JK77]. Aspergillus [DF77, FSM76]. assay [MB76, Sne76b]. Assays [MST78, SVW79]. assembled [DGR75, PBH78]. Assembly [CM77b, JKZM77, BR78, BR76a, BS79c, BH78, DR77b, DMMW76, FSSZ75, HSGW76, HGW77, HRL77, LLC79, MB75b, MFPO75, NP75, OAB76, RF77, RW75b, Sal75b, Sch76, SO77, Slu76, SK79, TV79, TR79a, TKST78, Til78]. Assessment [MV75a]. associate [BMP76]. associated [Amo77, BRR78, BB79a, BW78, BH78, BG76c, CP79a, DGR75, DDSL77, FGSM78, GW75, GP78, GRW75, KP77a, KAB75, Let79, LP77, NSP76, PS79, SS78a, Sta76, Til76b, UKY^{+}79, WGG77b, WGB78, Gre75]. Association [MCC79, MBBS76, Qui76, Blo77, CLP76, Con79, KBR79, LAS^{+}75, MD75, PH75b, STF75, ST77a, TD75, Til76a]. associations [KS79, LE77]. Aster [HK75, Lon76]. asters [WR75a]. astrocytes [SAPA77]. asymmetrical [Bro79]. Asymmetry [CB76, AEP79, ND77, SWF78]. asynchronous [HBD75]. ATP [BS77a, HJP78, LLP77, SS77a]. ATP-induced [BS77a, SS77a]. ATP-utilizing [HJP78]. ATPase [And77, BTS^{+}78, DJM77, Ern75, EKS^{+}79, KEP76, KKKS76, KCH79, Kyt76a, Kyt76b, SHC^{+}79, vOMW77]. attached [DR77b, EB75, Hop77, Mau76]. attachment [BE79a, BE79b, GTH76, HR79b, MT75, NCPW75, TM76b, TCT79, War78, KAB75]. attachments [McN78]. attending [FOPY77]. attention [GST79]. Atypical [KP75]. aurelia [FKP76, TK77]. auricularia [SS78e, SS78f]. autologous [BCFS75]. autonomy [CM77a]. autophagic [Paa78b, Pfe78]. autophagy [PS75, Pfe78, SA77]. Autoradiographic [MM79, RV78, CB75c, EKS^{+}79, GGS^{+}76, HWAS79, HS79b, JRFS75, NEK78b, NEK78a, SB76, TCPM77, KKKS76, MZ79]. autoradiographs [SMS78]. autoradiography [AHB79, FS76, FB76, HARO79, HG79, PSG^{+}76, SFS77, SRKM78, SMS78]. auxin [BVBM76]. auxin-treated [BVBM76]. available [Slu76]. avellana [VRN76]. avian [BH75a, EM77, SFSB78, TWR^{+}77, VC77]. avidin [HA77]. avidin-biotin [HA77]. axolotls [LMPI76]. axon [BH79, HL75, LGB77, PSGL78, VV76]. Axonal [GKS76, WWLS79, GSB77, HL75, MKK79, SST79]. axonally [Wil77a]. axonemal [SS76a, WPS78]. axoneme [Edd75a]. axonemes [BBDP79]. axons [EB78, GTL77, KKAL79, LP76, LKKA79, MKK79, SMM78, TDHO76]. axoplasm [LKKA79]. axoplasmic [PM75]. axopodia [Edd75a]. axopodium [Sch76]. axostyle [LI79, Tam78]. azurophil [FSB75].

B [JKB^{+}79, AHH76, AL78, BSVG75, BSU79, CH75, HS76a, HLG76, MPO79, OPF77, SS79c, Wei76b]. B-cells [MPO79, OPF77]. B-type [JKB^{+}79]. b5 [FRT^{+}76, RFB^{+}76, Rem78]. BA [CBDY79]. baby [GCY79, GTH75, GTH76, SBGG78]. Bacillus [LLL75]. bacteria [MBBS78]. Bacteriorhodopsin [WAC78]. bacterium [Zaa79]. bacteroids [VKZB78]. BALB [NSP76]. BALB/3T3 [NSP76]. Balbiani [DAF77, WD77]. BALD [GS75]. BALD-2 [GS75]. band [Coh78]. bands [BBDP79]. Barbulanympha [IR78, RIK78b]. barrier [EF75]. Basal [BCB77, CBB77, And77, AC79, CK78, Dip76, GLM77, Gou75, HK75, Nan75, NCM78, NM78, SMM78, SH79a, Ste75, WGG77b]. based [LZK75, Pau75, dSN77]. basement [KF79, MSK^{+}79, SM75a]. basic [EL75, LLLG78, MH76a, MW78]. basis [CM77a, CT77, Edd75b, HT79b, NSBS77a, TRH76]. Basolateral [EM77]. be [TK79]. beads [CKJB77, KNM79, LFR78, LH76]. bean [BTC78]. bearing [PE79]. beat [TSD75]. beating [Bro79]. becomes [Til76a]. behavior [BH75a, Bel77, BTDT76, HA76, MG75b, PPHM75, WS78b, ZM76]. Behaviour [KKH77]. beige [BES75, EH77, OKB76]. benzimidazol [DF77]. benzimidazol- [DF77]. benzo [WGC76]. beta [BGFK79, HS75a, Hir77, SM75b]. beta-1 [SM75b]. beta-bungarotoxin [Hir77]. beta-galactoside-binding [BGFK79]. beta-nerve [HS75a]. betaC [BPN76]. between [AMJR78, All75, AA76, BBW77, BB78, CCTT77, Con78, EG77, FMR79, FDB77, GB79, GW78, GGGM78, HO77, HLR76, Her76, LWS79, LLL75, LP77, LAF^{+}77, MIT76, MEM79, MPO79, MH75, MC78b, MKK79, MCM77, NCM78, NS77, PM79, RR78, RD78, RMP75, RGK78, SOK79, SHWPJ78, ST77b, Tak77, TMGC76, UIOO76, ZVR75]. BHK [GCY79, SBGG78, WG78]. BHK-21 [GCY79, SBGG78, WG78]. biased [WRP75]. bicarbonate [vOMW77]. bicarbonate-stimulated [vOMW77]. bidirectional [WRP75]. bifurcating [MKK79]. bilayer [DVEM79, SPS76]. bind [OLK77, SH79a]. bindin [GL79c]. bindin- [GL79c]. Binding [BSHP79, HK79, KHV^{+}78, LF77, Rem78, SC77, SLJ77, SC78, UT79, AKLL75, AL78, BKS78, BGFK79, CP79a, DF77, DL79, FSB75, FWS75, FS76, FK78, HMC79a, HS76a, HMK79, KH75, KUS78, KFP^{+}78, KAGV79, LMPI76, MV75c, NYY75, NUY^{+}77, PSM^{+}75, PK79, PAW76, Rod76, Rou78, RS75b, RS75c, SMG77, ST78b, SS78b, SBC78, Sne76b, SH76, TSC78a, TSC78b, WM77, WMP79, CC76, CBDY79]. Biochemical [And77, Bou79, DCEO^{+}79, GCM77, JSGR78, Mar77, RR75, SL78, SBGG78, BAGB77, BPL^{+}78, DWM75, DFW^{+}78, EWR77, EP79b, ET79, GLR77, JZR76, KKG76, LWB78, RS75a, RW75b, STF76, TCPM77, VV77, WP78, CBM78, CF76a, Md78b]. Biogenesis [ASD75a, ASD75b, ESAD75, BCB79, KSS76, KL79, MB75a, RW78]. Biological [SPS76, CG78, JB78, AL78]. biology [Pal77]. bioluminescent [SC75b]. Biosynthesis [FKNMT79, GSB77, CB78, HARO79, HB76a, NRH78, NBB77a, NBB77b, OBKP75, TV79]. biosynthetic [HARO79]. biotin [HA77]. biphosphate [ICS75]. birefringence [Sal75a, SEI75, Toy76]. Birefringent [BE79b]. bisphosphate [SDTD^{+}79]. bivalent [OLL76]. black [CGH79a, FCG^{+}76, GHC78]. bladder [Har78a, MM78, MC78b]. blastomeres [MO77]. Blattella [Cav76]. Blattidae [Cav76]. blebbing [ZW77]. blebs [DS77, ZW77]. blisters [HH78]. block [GSJH79]. blockers [KD78]. blood [HLS75, RMP75, SD76, SF76]. bloodstream [FWHB78]. blue [FMR79]. blue-treated [FMR79]. bodies [And77, CK78, Dip76, GLM77, Gou75, HK75, HB78, MG76, NM78, OCO75, OSO76, SV77]. body [BLR76, Lu78, Nan75, NCM78, NLT76, PP79, SP76, SS78e, SS78f, TBZW78, WGG77b, Wil77b]. Bombyx [ST76a, ST76b]. bone [CI78, KH75, Mil78, NCCB77, RA76]. bongkrekic [SS78d]. border [BW78, KCH79, LG75a, Moo76b, MP78]. borders [Moo76b, RNK76]. botulinum [HK79]. bound [BD75a, CBSP76, DJM77, FKNMT79, GWS77, HB79, KFP^{+}78, KS77, LP77, MWS79, MV75a, MV75b, MV75c, OKS77, SH79b, TR79b]. boundary [LG75a]. bounded [All75]. bovine [FFHL78, FFSH77, GWML76, HR79a, KRJ77, MB75b, PJB77, RD78]. brachymorphic [OWC^{+}77]. brain [Amo77, BFT77, BR76a, DGR75, GSN78, HW75, KKG76, KBR79, KR79a, MB75b, SB76, TM76a, ZM76]. branches [MKK79]. brasiliensis [BBW75]. breakdown [HS76b, WS78b]. brevis [LLL75]. bridge [Pol75, SG79a]. bridges [FGSM78, War78]. Bridging [Som79, ZMW79]. bristle [DMB78, FKS76, LTMS79]. bristle-coated [DMB78]. bromodeoxyuridine [BAB79, GPPR76, KSS76]. bronchial [MR75a]. brucei [FWHB78, HLS75]. Brush [Moo76b, BW78, KCH79, MP78, RNK76]. budding [BG76c, DSTW77]. bug [Cav75]. bullfrog [SR78b]. bundles [BC76, BS77c, Fuk78, GCY79, MBBS76, RGE78]. bungarotoxin [CF79, CP79a, FWS75, Hir77, JL79, FS76, SB76]. butyrylcholinesterase [DK78a].

c [MMS^{+}76, GBM78]. C6 [WHC79]. Ca [BSU79, CGH79a, FS77, LFR78, Mat77, MCP^{+}78, RGS^{+}76, Täl78, TKST78, VL78, WJTR77, WGB78]. Ca/Mg [WGB78]. cables [FTBS78]. Caenorhabditis [CT79]. Caerulein [DH75]. caffeine [Mat77, NR76]. caffeine-derived [Mat77]. Calcium [Bro79, CM75, GESG75, HMT79, HT79a, HMK79, JWJ75, PD75, Rou78, TSC78a, TSC78b, TZ77, WS78a, BFT77, BFPU78b, GJRR78, Har78a, HW78, HS76b, Lar75, LS79a, MC79b, NSBS77b, NSBS79, ODS79, PJ76a, RD76, SS78f, WC75]. Calcium-binding [HMK79, Rou78, TSC78a, TSC78b]. Calcium-containing [GESG75]. Calcium-induced [Bro79]. calcium-sequestering [BFT77]. calf [BGFK79, DFW^{+}78, FSW77, VHvdVB78]. californica [GKS76]. callus [VRN76]. calmodulin [WDBM79]. calvaria [KH75]. CAM [RTBE78, RGE78]. cAMP [UGB^{+}79]. cAMP-dependent [UGB^{+}79]. campaniform [MRZV76]. can [BP79]. canine [Con78, GSS77, GSS79, PM79]. cap [SMPGR75]. capable [PK79]. capacity [CW77, FMR79, SSR76]. capillaries [SSP75b, WW75, WW78]. capillary [BCFS75]. Capping [SB79, AA77, BO78, Bou79, Sto75]. caps [Con79, FKS76]. captured [HRD^{+}79]. carbamate [DF77]. Carbamycholine [OKB76]. carbamylcholine [SGV75]. carbohydrate [HWAS79, HS79b]. carbohydrates [CB75b]. carbon [DTT75]. carboxyatractyloside [SS78d]. carboxylase [ICS75, SDTD^{+}79]. carcinogen [RHDL77]. carcinogen-induced [RHDL77]. carcinoma [CMC79, HMC79a, HMC79b]. Cardiac [Bal79, GSS77, GSS79, IW76, LMPI76, VV76]. carlsbergensis [Dam76]. carolinensis [RR75]. cartilage [DST^{+}78, RBC77]. cartilaginous [OWC^{+}77]. casein [FLK^{+}76]. castellanii [RHHB76]. cat [DK78a, GLR77, TP76]. catalyzed [GL79a]. catecholamine [CP77b]. catesbeiana [SR78b]. Catharanthus [MRDC77]. Cation [War78, CW78b, CW78c, FKP76]. Cation-induced [War78]. cationic [COK76, Far78, FYS78, KF79]. cationized [BAGB77, BW79, RCV75]. cations [CG79, DTSM76, Sch76, ZMW79]. caudatum [KH76]. Caulobacter [AEP79]. caupo [GSJH79]. Causal [FJD76]. caused [Cla76]. causing [FCG^{+}76]. Cautionary [DH76]. cecropia [TBZW78]. Cell [GR78, GEB78, Gor75, HWW75, Har78b, KEG77, LGB77, Moo76a, PRZE75, RS75b, SPT78, TR79a, VK77, WB78b, BGRC78, BS79a, BG76a, BF78, Ber75, BRBM77, BSOS76, BCHN75, BR76b, BS77d, BS77e, BP77, CB75a, CPO77, CW76, CBSP76, CCB76, CE77, CKBD76, Cra79, DWB79, DMB78, DR77a, DL79, DCEO^{+}79, DWM75, DUKA77, EWR77, EW77a, EW77b, EGFB78, EKS^{+}79, FBF79, FJD76, FPC76, FF79b, GDR77, GST79, GNB^{+}77, GD77, GLR75, GMDT75, GET^{+}78, HU77, HH77a, HH78, HR79a, HMT79, HD77, Hel79, Her76, HMDH77, HBD75, HPH77, HWAS79, ICS75, Jau75, JKZM77, Jul78, KS79, KEP76, KKKS76, KMS77, KPF75, Kri75, KAGV79, LP76, LGGB79, LFR78, LL79, LN75, LE76, LE77, LZK75, MKKV78, MM76a, MM76b, MM76c, MG75b, MPJ79a, MPJ79b, MGF77]. cell [McG76, MH75, MZ78, Mil77, MHC77, Mil78, MG78a, MC78b, MR75b, NSRL76, NCPW75, NCGC79, NSP76, NSH78, OWS75, OWW78, PT77, Pal77, PDD78, PK79, PR75, PPHM75, PL76, PPL78, PM79, PEA77, PCJG79, Pro75, PB75, QWTI79, QLHTB79, RV78, RR77a, RH77, Rie79, RP77, RCSS79, RBK78, RHDL77, RTBE78, RGE78, SPM79, SR78a, SSD^{+}78, SBC78, SMPGR75, SSP75a, SSP76, SS77b, SF76, SH79b, SG79b, Slu79, Sne76a, Sne76b, SC75a, SVW79, Tak77, TV79, TZ77, UIOO76, WOPB79, WDMR77, War79, Wei76b, Wei77a, WNQK78, WPMW75, WGG77a, WR77a, WP75, WC77, WDB79, YYP77, Yam78, YR75, vVRN76, GW75, TM79, TS77]. cell-substrate [WDB79]. cell-surface [KS79, MGF77, Mil77]. Cell-to-cell [GEB78, KEG77, LGB77, RS75b]. cells [AA77, AB76, AB77, AB79, ASGR78, AA76, AL78, BRS^{+}76, BGRC77, BGRC78, BPN76, BH75b, BK75, Bel77, BCM^{+}77, BDCD75, BHFT79, BC79, BFD78, BO76, BAB79, BK78, BC76, BBMB77, BRS78, BB78, BA79, BS78b, CW78a, CB75b, CCH78, CRD^{+}78, CO76, CW78b, CW78c, CM77a, CAS78, CFAS78, CMC79, COK76, CKJB77, CI78, Con78, CGVR79, CP79b, CH75, CBH^{+}78, CRBH78, DR78, DMB78, DPM76, DB75, DMMW76, DBB^{+}79, DDR75, DTT75, EP79a, EP79b, EG77, Far78, FFHL78, FC77, FK78, FSW77, FSOW79, FLRS77, FDB77, FP76, FP78b, GCH78, GB79, GHA76, GESG75, GTPWD76, GD75, GDC76, GDR77, GR78, GMDT75, GCY79, GBBZ78, GB77, GGGM78, Gra75a, Gra75b, Gre75, Gre78, GTH75, GTH76, GRW75, HMC79a, HMC79b, HMPB76, HWW75]. cells [Han75a, Han75b, HO77, HG78, HT79a, HP79, HSM76, Hit77, HMDH77, HR79b, Hop77, HG77, HWLC78, HP75, HC75a, HC75b, IO79, IP78b, JT79, JJB79, JLA^{+}76, JK77, JSG79, KSS76, KP77a, KMS77, KG78, KW76, KH75, KSP75, Kra77, KRJ77, KG77, LC76, LWS79, LBDL76, LGGB79, LRG^{+}77, Laz75, Laz76, LE78a, LE78b, LMPI76, Lem79, LN75, LH79, Ley75, LZK75, MBPTS79, ML76, MR77, MD77a, MH77, MPJ79a, MPJ79b, MSS75, MGF77, MPS79, MV75a, MV75b, MV75c, MCP^{+}78, MR75a, MHC77, MC78b, Moo76a, MT75, MP78, MS75a, MB77b, MOS^{+}77, MSK^{+}79, MF75, NCCB77, NS75, NSP76, NPM75, OY78, OSM79, OPD76, OH78, OB78, Paa78a, Paa78b, PH75a, Pai75, PSZK78b, PJ76b, PB79, PPHM75, PS79, PR76b, PRZE75, PEA77, Qui76, QLHTB79]. cells [RF77, Rod76, RR77b, RCMI75, RS75a, RW75b, RWLC78, RW79, RR76, RS75c, RTBE78, RGE78, SV76, SOK79, SE75, SGMB76, SAB75, SBU76, SBULB76, SL78, ST76b, SMG77, SC77, SFSB78, SM78b, SSD^{+}78, SHWPJ78, SO77, SBC78, SFFB75, SM75c, SA77, SC78, SBGG78, SH77b, SB79, SLR77, Täl78, THC76, TP76, TWR^{+}77, TR79a, TE79, UT79, VP75, VJGG79, VV77, VMR79, WBE75, WF75, WS78a, WG78, WLW78, WNQK78, WPMW75, WR77a, WR77b, WR77c, WGC76, WHC79, WR75c, WWLS79, WK79, ZJSF76, ZW77, MPO79, OPF77]. Cellular [KS79, SOK79, AJG78, CM77a, CCB76, GD75, GDC76, JJ75, LE78a, NEK78b, NEK78a, PS75, PT79, Roo75, SB78, WPMW75, YK79, RW75a, SPM79]. cellulases [BVBM76]. cellulose [SM75b, Zaa79]. center [SEHW79]. centers [CAS78, SM75c]. Central [LHH76, HK79, KCGN76, LC77, MEM79, MMER79, MOS^{+}77, SL77, WPS78]. centrifugation [WCRdD78]. centriolar [BR77, BRBM77]. centriole [PR76a, MRG77]. centrioles [BBMB77, CK78, MB75b, NM78, Rie79]. centrolobular [WDM^{+}75]. centrosomes [PB79]. cerebellum [HHP76, SHWH^{+}77, TS77]. cerebral [BCS77, BPS78, CBBS77, UGB^{+}79]. cerevisiae [HU77, HB78, RH77]. cervical [DK78a]. chain [GTPWD76, KKG76, MV75c]. chains [BD75a, CRSH75]. chamber [SE75]. chameleon [RR75]. change [AZW^{+}79, MNK79, WAC78]. Changes [BBW75, CO76, Cla76, DM79, DR77a, DWC^{+}77, DUKA77, ES76, JF75, JLA^{+}76, NSBS77a, SWV77, TWC75, AB79, AKD75b, BS77a, CSP79, CCH78, CPO77, CMC79, CGVR79, DN76, Dam76, DPM76, DDR75, FOPY77, GG77, HWW75, HB76a, HGP^{+}76, KSP75, LB78, LGGB79, Mil77, MG78b, MR76b, NCGC79, PD76, Per77, PRZE75, RBB76, RBC77, RJ76, RLJ76, SS78a, SM78a, SEHW79, SPS76, SS77a, SLR77, Swe76, TM79, Toy76, VK77, VL78, NUY^{+}77, ODS79, Sal75a, SBC78]. Changing [BB79a]. channels [SSP75b]. characean [KW76]. characteristic [And76]. Characteristics [SK79, DWM75, FN76, Jul78, KUS78, PJB77, WDMR77, HP75]. Characterization [GHA76, KRM77, MP78, PFK^{+}76, RS79, SHC^{+}79, SAPA77, SSC76a, SSC76b, SCS76, TK77, BCM^{+}77, BC77, BFD78, BCS77, BSPK78, CB75b, CBM78, CB77, CJP75, CCNAN78, DDSL77, DCEO^{+}79, EL75, Edd79, FWHB78, FFHL78, GV77, GJAE75, GSN78, Gou75, GCM77, JZR76, KR79a, LNS76, LLP77, NBB77a, OT78, OBKP75, PSM^{+}75, PAW76, RR75, Ste75, TC75, VKZB78, Wei76a, Yam78, vVSL^{+}76, CBH^{+}78, GBBZ78, KC78, QWTI79, WR77b]. charge [BW79, CB76, RCV75]. charged [BS79c]. chase [NBB77b]. Chediak [OKB76, BES75, FTBS78]. chelating [RW75a]. Chemical [SM78a]. chemically [ML76]. chemotactic [BTS^{+}78, GG77, NSBS77b, NSBS79, OKS^{+}79, Zig77, ZS79a]. Chemotaxis [Zig78, CG79, DR76, MRG77, SGV75]. Chick [KR79a, BN76, BH77, BH75b, CF79, CO76, CP79a, CI78, DN76, Gra75a, Gra75b, GKB78, Hir77, JL79, JM79, KMS77, MB75a, MB76, McG76, MG78b, NVH^{+}75, OSM79, RTBE78, RGE78, TSC78a, TSC78b, VV77]. chicken [BTB79, BGFK79, CI78, GL79a, HL79, LSWB78, Smi77, TR79b, VL78, YBB79, WTE77, WKP^{+}77]. Chinese [BO76, FK78, GB77, Jul78, Ley75, PT77, RS75a, RW75b, SC78, WR77c]. Chironomus [BLR76, DAF77]. Chitinous [HKS77]. Chlamydomonas [WPS78, BGG^{+}75, Blo77, BLB79, BP76, CMB75, CBSP76, CG77, DH76, DR77b, FGG78, GS75, GW75, GW78, GJ78, Gou75, HS76c, HBD75, HPH77, HRL77, ICS75, LNMR78, MG75a, MG78a, SDTD^{+}79, Sne76a, Sne76b, WGG77b, WC77, WGB77]. chloramphenicol [WBE75, WF75]. chloride [KEP76, KKKS76, SPM79]. chlorophyll [CMB75]. chlorophyll-protein [CMB75]. Chloroplast [AMBM76, ASA77, CBSP76, Col79, DH76, JJ78, JHW78, Sta76, WC77, PH75b]. chloroplasts [CS79, GCRB78]. chloroquine [QLHTB79]. Chlorotetracycline [Täl78, CW78b, CW78c, NSBS79]. chlorpohyll [MMM76]. chlorpohyll-protein [MMM76]. Chlortetracycline [SS79a]. CHO [vVRN76]. cholera [DMMW76, JSG79, JSG79]. cholesterol [AC79, DS75, HW75]. cholesteryl [BGK^{+}79]. Choline [GNB^{+}77, PRZE75]. cholinergic [BPRC78, BBW75, HSM76, MIMMF77]. cholines [KP77c, KP77b]. chondriome [Pau75]. chondrocyte [DST^{+}78]. chondrocytes [PT79]. chondrogenesis [VV77]. chondroitin [RD78]. chorioallantoic [TSC78a, TSC78b]. chorionic [AKLL75]. choroid [McN78, vDK79]. chromatin [CSHS77, CH76, DTSM77, HWP78, MWS79, OCO75, OSO76, OO79, RH79, TKK79, SSD^{+}78]. chromatin-bound [MWS79]. chromatin-depleted [HWP78]. chromatoid [PP79, SP76]. Chromosomal [CM77a, BE79a, BE79b, DBB^{+}79]. Chromosome [CW78a, BE79a, BE79b, CM77a, DB76a, Fus75b, MWS79, TMGC76, ZM76, BE79a]. Chromosome-spindle [BE79a]. chromosomes [AJ78, BE79b, BDKS79, KMB78, Leó76, MB75b, Moe79, PH75c]. chronic [Hir77]. chronically [RMP^{+}78, TE79]. chrysophyte [HKS77]. chymotrypsinogen [GST79, RD78]. cilia [DK78b, GLM77, KH76, MS75b, SS76a, WM78]. ciliary [AH77, KP75, Nan75, PL76, TSD75, WS78a, War78, WM78, ZMW79]. ciliated [Ste75]. Ciliogenesis [SBU76, SBULB76]. circadian [Swe76]. circular [FWB78, WF76]. Circulation [THCP75]. cistrons [STF75, SCS76]. clam [WR75a]. class [Bon75b]. classes [KBS79]. Cleavage [HS76b, BCB79, FWB78, FKS76, FP76, FPP78, ODS79]. clonal [Gre78, MH76a]. Cloned [Cra79]. Close [FDB77]. Clostridium [HK79]. clustering [WMP79]. Clusters [CP79a, Blo79, FF79a, VD76]. co [GNB^{+}77]. co-culture [GNB^{+}77]. coat [FKS76, FLK^{+}76, MPDGR76, MR75b, SMPGR75, SM76]. Coated [KKG76, BFT77, CKJB77, DMB78, MSS75, MH76a, RS75c, WPMW75, WMP79]. Coated-vesicle [KKG76]. coats [LTMS79]. cockroach [Cav76, NS75]. code [ST77b]. coenocytotomic [FKS76]. cohesion [SB78]. colcemid [BC76, CH75]. Colchicine [PSM^{+}75, RBHP75, CG78, FMSB75, FK78, LC76, PAW76, PT79, RBM^{+}78, WGE75]. Colchicine-binding [PSM^{+}75, PAW76]. colchicines [PBJ77]. cold [RW75b]. coli [HO75, SHM^{+}77, Wei77b]. collagen [BH77, Bru76, EP79a, HP77, MSK^{+}79, NLT76, OBKP75, VV77]. Collagenous [RA76]. collecting [WOPB79]. colloidal [ST78b]. colony [BAN75, BPN76, CI78, JB78, PM79]. colony-forming [BPN76, PM79]. colony-stimulating [JB78]. colorless [SSC76a, SSC76b, SCS76]. colostrum [PJB77]. Comments [DB75]. Commitment [BPN76]. communication [CKBD76, EG77, FAE75, GEB78, KEG77, TE79]. companion [AA76]. Comparative [TVD78, RS79]. Comparision [KVB75]. Comparison [DJM77, HP79, CLM77, ERSP76, JKB^{+}79, BGFK79, GHC78]. compartment [LG75a]. compartments [GD77, GHC78, RCM77, RMP75]. compensation [CM77a]. compensatory [MKM76, OAR^{+}77, SPS76]. competence [KMS77]. complementary [vDK79]. complex [AMS76, ASA77, BE78, CMB75, Con78, CP79b, DL79, DB76b, GBB78, GST79, GEB78, HA77, HO75, LH76, MMM76, MT75, NT78, Paa78a, Paa78b, RKL^{+}76, ST78a, SS77a, Tam78]. complex-endoplasmic [LH76]. complex-lamina [DB76b, GBB78]. complex-specific [HO75]. complexes [CFAS78, EVMP76, EGFB78, GWS77, HO75, KRM77, Mau77, MD77a, MHP78a, SSF76, TM76a, VV76]. component [HL75, RKL^{+}76]. components [BD75b, FSSZ75, GES79, LLMW77, MPDGR76, MB79, MGF77, PB75, RS79, RS75a, Sne76a, SC78, WPS78, WDB79]. composition [BRV78, CBBS77, DFW^{+}78, HMDH77, LCK75, Md78a, Md78b, WPS78, WH77, SSS79]. Compositional [EB78, EZF76]. compounds [RP78]. comprising [Wei76a]. Computer [BCHN75]. con [SF76]. Concanavalin [HSGW76, MJM76, PPL78, ZS79b, dP75, AA77, CDH^{+}79, CT76, Con79, HB79, LFR78, Let79, LAF^{+}77, OLL76, PELG76, Sto75, WM77, DMKR75, dP78]. Concentration [CSHS77, BAN75, CSP79]. concentrations [GHC78, RD76, SS78a, VV76]. Concerning [NM78]. condensation [DB76a, TMGC76, ZM76]. condensing [NMQY77]. conditional [FGG78]. conditioned [GNB^{+}77]. conditions [EP79a, SC78]. conductance [KD78]. cone [Bur78, RBB76, SR78a, WB78a]. cones [Lan78]. configuration [Bal79]. conformation [WM78]. Conformational [MNK79]. conjugate [HMC79a]. conjugated [NSH78, NVH^{+}75]. conjugates [KOBP75]. Conjugation [OLL76, TM79]. connecting [SKT^{+}76]. connection [ZVR75]. connective [BRS78, BB78]. Conservation [MKM76, MWS79]. constancy [MS75a]. constant [KT77a]. constituents [JBKF77, STdDT79a]. constraints [NCM78]. contact [AL75, Bel77, BB78, GDR77, RW75b]. contacts [RG75, WDMR77]. contain [Lan78]. containing [BDKS79, CCTT77, GESG75, GLR77, MHP78a, MCM77, RW75a, UKY^{+}79]. contains [LSWB78, TR79b]. contaminants [CP78]. contemporary [Pal77]. content [BS78a, CJP75, EP79b, KNL76, MD77a, RR77a, RR78, RWLC78, TZ77, ZJSF76, LEDW76]. contents [BG76b]. continous [GET^{+}78]. Continuation [BR77]. Continued [WR77c]. continuous [BCHN75]. contractile [CH76, CT77, HDS77, HT79b, LS79b, TRH76, WGG77b, WH77]. contractility [Edd75b, TRH76]. contracting [MPS79]. Contraction [RNK76, Bur78, CT77, FWS75, HT79b, MC78b, Moo76b, Pol76, dCdM78]. contrast [BHFT79, KP77c, SE75]. Contrasting [BFPU78b]. Contributions [BW79]. Control [DR76, GBBZ78, ZM76, CCB76, CT77, DWB79, HU77, HMT79, Jau75, LW76, ODS79, SE75, SBULB76, STdDT79b, Slu79, SVW79, TSD75, TM76c, WS78a, SES76]. controlled [DDS79, Von79]. controlling [WS78b]. Conventional [PD75]. Conversion [Wil77b, HARO79, NBB77a, NBB77b, RBM^{+}78]. convoluted [Kyt76a, Kyt76b]. copolymer [Pol75]. copolymerization [HL79]. copper [SSF76]. copper-tolerant [SSF76]. Coprinus [Lu78]. Copurification [HL79]. cord [BF78, FBF79, GNB^{+}77, PSG^{+}76, SF78, THCP75]. core [BCB79, LLLG78]. cornea [BH75a, BH77]. corneal [EVMP76, HH77a, HH78, MH75, NVH^{+}75]. corpus [Paa78a, Paa78b]. corpuscle [OAR^{+}77]. Correction [WP78]. correlated [HRD^{+}79, Zaa79, CGMP77]. Correlation [DBB^{+}79, LAF^{+}77, MB75a, FMR79, GGGM78, SHWPJ78, Tak77, UIOO76, HARO79]. correlations [BAGB77]. correlative [ET79]. cortex [BR79, BCS77, BPS78, CBBS77, Ern75, HD77, ST78a, UGB^{+}79]. cortical [CH79a, DDSL77, DUKA77, HG78, HS76b, KSS76, SS79d, TSD75]. corticosteroid [WOPB79]. corticosterone [HMPB76]. Corylus [VRN76]. cotyledons [BTC78]. couples [SPS76]. coupling [CW78b, CW78c, HT79b, IP78b, KG78, LFR78, LL79, MS76]. course [Lu78]. Covalent [HR79b]. covalently [TR79b]. cowpea [HH76]. crab [JC79]. crayfish [FN76, MCR77, PD76]. creatine [IW76, WTE77, WKP^{+}77]. crest [SBC78]. cristae [MCR77]. criteria [QWTI79]. Crithidia [FMW75, FWB78, MW76]. critical [SFS77, UIOO76]. Cross [MEM79, FGSM78, KNL76, Lon77, PSZK78b, Pol75, RG78, SG79a, War78, ZMW79, KFP^{+}78]. cross-bridge [Pol75, SG79a]. cross-bridges [War78]. cross-fertilization [Lon77]. cross-linked [PSZK78b, RG78]. Cross-linking [KFP^{+}78]. cross-reactions [KNL76]. Cross-sectional [MEM79]. Crosslinking [BSU79]. crustacean [EWR78]. cruzi [MPDGR76]. cryo [SSS77]. cryoultramicrotomy [PB75]. crypt [WNQK78]. Crystallin [VHvdVB78]. Crystallomitin [BDP^{+}77]. crystals [BDP^{+}77, RR75]. cultivated [DMMW76]. culture [BN76, Bel77, CBH^{+}78, DB75, EB75, EG77, FBF79, FP78b, GNB^{+}77, HLS75, JJB79, KSS76, LW76, MPDGR76, Moo76a, MOS^{+}77, NSRL76, Pro75, QLHTB79, RBB76, RWLC78, SV76, SFSB78, SLR77, TWR^{+}77, VMR79, WBE75, WDMR77, WGC76, WR75c]. cultured [AA77, BC79, Blo79, BRS78, BS78b, CP79a, DR78, EP79a, FLRS77, HVW78, HWLC78, HP75, KH75, Lan78, LGGB79, MPS79, OKB76, PH75a, PT79, PRZE75, RG78, RR77b, SMG77, STdDT79b, SHWPJ78, SVW79, THCP75, UT79, VD76, WJB78, WKBL75a, WKBL75b, WP75, WY78, WMP79]. cultures [ASGR78, BF78, BWDP77, BFD78, BCHN75, CRSH75, CI78, CBC75, DST^{+}78, GNB^{+}77, GM75, HKA^{+}79, HBD75, JKZM77, MKKV78, PWSA77, PR75, PPHM75, QWTI79, RHDL77, RGE78, TS77, YGH78]. cumulus [GEB78]. cumulus-oocyte [GEB78]. current [NJ75]. currents [WNJ75]. cyclase [CF76a, CF76b, LJ75, RBC77]. cycle [BGRC78, BLTP^{+}76, CBSP76, FJD76, FPC76, Hel79, HMDH77, HBD75, HPH77, ICS75, Jau75, KPF75, Kri75, KKH77, MG75b, MR77, MZ78, Mil77, MRZV76, PRZE75, PEA77, QLHTB79, RH77, RT76, SG79a, SSD^{+}78, Slu79, TR79a, WR77a, WC77, ZJSF76]. cycle-dependent [TR79a]. cycles [GTPWD76, MC78b]. Cyclic [DDS79, DS79, KP77a, WP75, CCH78, DMMW76, GDR77, GHMD77, HMPB76, MOR78, MC79b, RW75b, SGV75, Sto75, WS78b, vVRN76, BP76]. cycling [AKD75b, MD77a]. cynodontis [DTT75]. Cyprinodon [KEP76]. cyst [RHHB76]. Cytasters [MBV77]. Cytochalasin [BS79b, CW78d, SS79c, Wei76b, BSVG75, CH75, GMDT75, HS76a, HLG76, PCJG79]. Cytochalasin-stimulated [CW78d]. Cytochemical [BKK75, And77, BDKK75, MPDGR76, RdC77, CF76b, Paa78a, Paa78b]. cytochemistry [Ern75]. cytochrome [FRT^{+}76, FKNMT79, IP78a, MRDC77, MFKT78, RFB^{+}76, Rem78, RTC76, MMS^{+}76]. cytochromes [JKB^{+}79]. Cytofluorometric [KPF75, FPC76, Kri75]. cytokinesis [FPP78, IR78, MB77b]. Cytological [NEK78b, NEK78a]. cytolysis [LAF^{+}77]. cytomembrane [KP77c]. cytoplasm [FSSZ75, FP76, MC78b, RGK78, SSS79, ZM76]. Cytoplasmic [EL76, GCH78, JJ75, KAB75, Sch75, WBE75, BC79, Bon75b, BFPU78a, BMP76, BB79b, BP77, CB78, DUKA77, HLS75, JF75, SSC76a, SSC76b, SCS76, SSR76, SH76, TK77, WG78, WS78b, WC77, WP79, Edd75b, ST77b, dP75]. cytosine [JM77]. cytoskeletal [CDH^{+}79, DWB79, ET79, IF79, NSP76]. cytoskeleton [Coo76, Edd79, MB79, MH77]. cytoskeletons [She79]. cytosol [ASD75a]. Cytotoxicity [COK76].

D [BDCD75, GMDT75, MB77b, PCJG79, RBK78, STF75]. D-98S [MB77b]. D-amino [RBK78]. D2O [RW75b]. daily [OY78]. Dallas [Cav75]. Damage [YB75]. darkfield [SK79]. darwiniensis [BD78]. data [FPC76, MCPG77]. Daughter [AB77]. deafferentation [GWN77]. dealkylations [DMS^{+}76]. deaminase [AFB79]. death [Gre78, LP76, PL76]. decanoic [CB75c]. decidual [Her76]. decline [KMS77]. Decoration [GKBM78]. decrease [Gra75a, RBC77, WGB78, ZS79b, YYP77]. decreased [Hil75]. deep [HS79a]. deep-etched [HS79a]. Defective [BES75]. Defects [OWC^{+}77]. defend [AV78]. deferens [BH79, MPS79]. deficiency [KMS77]. deficient [TSS79]. deformability [MW77]. degradation [CC76, Hil75, MC79c, Pfe78, SV76]. degrade [Gra75a]. degranulation [LFR78]. degree [RGK78]. degrees [GBM78]. delay [SC76]. Delineation [BH78]. demembranated [BS77b, Bro79]. demonstrate [KCH79, RMP^{+}78]. demonstrated [KH75]. Demonstration [FSB75, SRR79, SM76, WD77, BCFS75, BKK75, GLM77, LJ75, SVW79, OT76, SEI75]. denaturation [DTSM76, DTSM77]. dendrites [JF75]. dendritic [CAS78, CFAS78]. denervated [BPRC78]. denrites [BPS78]. dense [CCTT77, FBSS78, KRM77]. densities [BCS77, CBBS77, WPMW75]. density [CB75a, CS78, CCB76, CW78d, KC78, KAGV79, LW76, TWR^{+}77, WP78, MW76]. dentritic [HHP76]. denuded [SCRP77]. deoxyglucose [PRZE75]. Deoxyribonucleic [HO75]. Dependence [PR76a, Fus75b, MSRT78]. dependency [AKD75a]. dependent [CMC77, Ern75, IO79, KSS76, MRDC77, MCP^{+}78, OAB76, PRZE75, Pol76, Rod76, She79, SH77a, TR79a, TKK79, TSC78b, UGB^{+}79]. depleted [HWP78, Wei77b]. depletion [FCG^{+}76, LLP77]. depolymerization [Fus75b, Sal75a, Sal75b, SGMB76]. depolymerized [EP79a, PPMSK77]. deposition [DST^{+}78, WB78b]. deposits [Lar75]. deprivation [Ley75, WR77a, WR77b]. deprived [PL76, WR77c]. derivative [SAPA77]. Derivatized [PRGH78]. derived [DR78, KMS77, Mat77, SS79b]. dermal [DH75, RR75]. description [ABL76]. desensitization [OKS^{+}79]. desmin [HL79]. desmosomes [DFW^{+}78]. destined [BBC78]. destruction [BR77, MSP78, WY78]. detachment [Jul78, LNMR78]. detail [SWF78]. details [AZW^{+}79]. detected [Far78]. Detection [Ley75, OWS75, PM79, KNL76]. determinant [MM76c]. determination [CF76a, Kyt76a, Kyt76b, MD77a, HBD75]. determined [HARO79, JRFS75, LWS79, TK79]. determining [MF75]. deuterium [MW76]. developing [BH75a, DF75, GKB78, HH77a, HH78, NJ75]. development [AEP79, CP77b, CP77c, CP77d, GLR75, GCM77, HS75a, HS75b, JL79, MG78b, NR76, Pal77, RPGR77, RBB76, RTBE78, WT79]. Developmental [KMS77, LH79, RJ76, RLJ76, Col79, CP77c]. diamide [EVMP76]. diamidino [Col79, JJ78]. diaphragm [GHC78, KR75, SM75a, SSP78]. diaphragmed [DM79]. diatom [TPH77, TSPH78]. Diatoma [MPHMT77, MEM79, MMER79]. Dibutyryl [GDR77, DMMW76, MC79b, Sto75, vVRN76, RW75b]. dichlorobenzimidazole [THC76]. dichlororibobenzimidazole [Gra75a]. Dictyostelium [CT77, DR77a, DDS79, DS79, EWR77, EL78, Fuk78, Moe76, MJM76, RdC77, WM77]. Differences [RGK78, BP79, MPDGR76, MHC77, PL76]. Different [DTSM77, JSGR78, Moe79, RS75b]. Differential [BH75b, DF77, JM77, SGMB76]. differentiate [VMR79]. differentiated [LW76, RCSS79, SM78b, VJGG79]. differentiating [ASGR78, JKZM77]. Differentiation [KT77b, BLTP^{+}76, BGG^{+}75, CW76, CH79b, CP77b, Cra79, Dec76, EF76, GES79, GW75, Gre78, MG75a, MH75, MS75a, NCGC79, RG78, SBC78, SF76, VHvdVB78, ZT75, SMM78]. differentiations [SSP75a, SSP76, WDMR77, WGG77a]. differentiative [GPPR76]. differently [ST77b]. differing [MD77a]. diffraction [CGMP77, GSS77, KHT^{+}79, MCPG77]. Diffuse [WMP79]. Diffusion [EW77a, EW77b, PPL78]. digested [San78]. digestion [BS77b, Dip76, WKBL75b]. digestive [DR77a, RdC77]. digitonin [EGFB78]. Dihydrocytochalasin [AL78]. dihydrofolate [DBB^{+}79]. diiodofluorescein [GES79]. dimensional [MMER79, Pau75, Sch78b, LLC79]. dimethyl [Fuk78]. diphosphate [MCM77]. diphtheria [DCEO^{+}79, UKY^{+}79]. diploid [BGRC77, BGRC78, LAS^{+}75, WGC76]. Diptera [AJ78, BLR76]. Direct [CLP76, HMC79a, LAS^{+}75, RW75b, SB78, Edd75a, HS76b]. direction [TSD75]. directional [AB79]. directionality [DR77b]. Disappearance [Hir77]. disassembly [ALMS76, KS77, MCP^{+}78, RW75b, Sch78a, SK79]. discharge [HA76, HSGW76]. discoideum [CT77, DR77a, DDS79, DS79, EWR77, Moe76, RdC77, WM77]. disintegration [SKT^{+}76]. disks [AC79, PSZK78a]. Dispersal [Blo79]. Dispersed [Kra77]. dispersion [BP77, GHMD77]. dispersions [DS75]. displaying [AKD75a]. disposed [HC75a, HC75b]. disposition [BSPK78]. disrupted [LF77]. dissociated [FBF79, Hop77, HG77, PR76b, VP75]. Dissociation [RR78, She79]. dissolution [ES79]. distal [HD77, Kyt76a]. distinct [BFPU78a, BFPU78b, CRBH78, KC78]. distinguish [BP79]. distributed [ST77b, Til76a]. Distribution [AKLL75, AH77, BBW77, CFAS78, GL79b, GKB78, HR79a, HW75, KCGN76, MPJ79a, MPJ79b, RR77a, WT79, War79, vOMW77, AMS76, AB79, CLM77, CO76, CM75, CCB76, DMB78, DM79, GBB78, GTH75, HM75, JF75, Koe75, Kyt76a, Kyt76b, MPDGR76, MH77, MC78a, MD77b, ODS79, STWB78, SB76, SF76, SSS77, SWV77, TSPH78, TZ77, UGB^{+}79, WR75b, WP75, dP78, HWAS79, NBB77b, ST76b]. distributions [FPC76]. diuresis [HGP^{+}76]. diurnal [PS75]. Divalent [FKP76, CW78b, CW78c, DTSM76, Sch76, Sch75, ZMW79]. dividing [HP79, JG76, Slu76]. division [BBMB77, DB76a, GWML76, HU77, Har78b, KKH77, Lu78, MO77, MR77, PEA77, ZJSF76]. DNA [BG76b, BS78a, BO76, BDKS79, CBDY79, Cav75, CW76, CW77, Cle75, Col79, DN76, DTSM76, DTSM77, DH76, FWHB78, FMW75, FWB78, GP77, Han75a, Han75b, HG79, JJ78, JT79, JHW78, KMS77, KBS79, LW76, MW76, MD77a, MRB75, MF75, NPM75, OPD76, Rag77, RS79, SSD^{+}78, TC75, WR77b, WR77c, YGH78]. DNA-binding [CBDY79]. DNA-specific [JJ78]. DNase [She79]. DNase-I-dependent [She79]. Do [AB76]. dodecyl [AJG78, LMPI76]. does [HIP78, SC75a]. dog [BRS^{+}76, BCS77, CBBS77, RCSS79, WBY79]. dogfish [EKS^{+}79]. Dolabella [SS78e, SS78f]. Domains [TSS79]. dosage [CM77a]. Double [MF75, BS79a, FP78b]. doublet [BR78, GS75, MNK79]. droplets [TP76]. Drosophila [BC79, BFD78, CM77a, GP77, HSRP79, MIT76, PE79, WF76]. drug [BDCD75]. drugs [CDH^{+}79, MCP^{+}78, SS76b]. dual [NLT76]. during [AEP79, AA77, AP75, BN76, BFTH79, BO78, BJ75, BK78, BB79a, BS77b, BS77d, BP77, CPO77, CGH79a, CGH79b, CO76, CW76, CW77, CW78b, CW78c, CH79a, Che79, CBSP76, Cle75, Con79, CE77, CG79, DB76a, DR77a, DUKA77, FKS76, FPP78, GCH78, GWK77, GGS^{+}76, Han75a, Hil75, HMDH77, HBD75, HS75b, ICS75, JL79, JF75, KT77b, LP76, LRG^{+}77, LE78a, LE77, Lu78, LLMW77, LLP77, MRG77, MG75b, Mau77, MPO79, MRB75, MKM76, Mil77, MB77a, MC78b, MG78b, MR75b, MRZV76, MS75a, NSBS77a, NUY^{+}77, OAR^{+}77, OPF77, Paa78a, Paa78b, PL76, PEA77, QLHTB79, Rag77, RBB76, RG78, RBC77, RJ76, RLJ76, RHHB76, SS79a, SPT78, STF76, STdDT79b, STdDT79a, SBC78, SF76, SP76, SBC76, SWV77, SS78f, Tam79, TDO78, VV77, WT79, WR77b, WF79, WGF79]. during [dSN77, dMdS78]. Dynamic [DPM76, Edd77]. Dynamics [NSH78, BOP^{+}78, SMPGR75, IR78]. dynein [AEJL75, BBDP79, OAB77, SG79a, War78, WM78, ZMW79]. dyskinetoplastic [RS79]. dystrophic [SS79b].

each [AB77]. ear [FC77, Hir77]. Early [FF79a, AP75, BS77d, DUKA77, GM75, GWN77, HBS79, ODS79, WT79]. earthworm [Ros78]. echinoderm [TKST78]. Echinosphaerium [Edd75a, Edd75b]. ECORI [FWB78, WF76]. EcoRI-sensitive [WF76]. Ecto [SRR76, CDH^{+}79]. Ecto-ganglioside-sialidase [SRR76]. Ectogalactosyltransferase [PELG76]. ectoplasm [PH75b, TRH76]. ectoplasm-endoplasm [PH75b]. edges [SS79c]. editor [Pro75]. Effect [BE78, BWDP77, CB75a, DS75, EVMP76, GWML76, HLG76, KAGV79, LC76, MR77, SSG79, BGRC77, BH75b, GOG76, GM75, Gra75b, MO77, OLL76, PBJ77, RF77, RBM^{+}78, RW75b, SFS77, SS76c, SS76d, WP78, ZM76, BAN75, BSVG75, CP77d, DTSM76, EW77b, LH79, MCP^{+}78, NSBS77b, PCJG79, WPMW75, WR77a]. Effects [CDH^{+}79, Dip76, HS75b, JM77, KSS76, NR76, NSP76, OAB76, PW78, PT79, SGV75, STF75, Wei77a, ZMW79, AL78, BFPU78b, BH78, DTT75, GHC78, HBD75, LCK75, MSRT78, NRH78, SS76b, SSS77, TWR^{+}77, BN76, CGH79a, CGH79b, GDR77, GHW75, HS76a, LNMR78, WDBM79, WP75]. efferent [Hir77]. egasyn [LP77]. egg [BR79, ES76, GJRR78, GL79c, HMK79, Kan75, Kan76, Mil77, TM76c]. egg-laying [Mil77]. eggs [BS77c, CH79a, DB76a, DL79, DDSL77, GV77, GSJH79, HK75, MBV77, NR76, NYY75, NJ75, PAW76, SEBL77, Slu76, SS79d, WT79, WR75a]. eggshell [BK78]. egress [Mau76]. Ehrlich [IO79]. ejaculation [NUY^{+}77]. Elasmobranch [EKS^{+}79]. Elastin [NSRL76]. electric [MIMMF77]. Electrical [IP78b, CGH79b, GHC78, WNJ75]. Electrically [GSJH79]. electricus [BPRC78]. Electron [BH78, CB75c, DK78a, GLM77, Pol75, SSS79, WC75, AZW^{+}79, BBW75, BB79b, BW79, BP77, CBM78, CGMP77, FS76, FRT^{+}76, GR78, GWK77, HARO79, HWAS79, HS79b, IF79, JRFS75, KKG76, KP77c, KW76, KRM77, KHT^{+}79, KOBP75, LAF^{+}77, ML76, MSS75, MR75a, MM78, MDRY75, MJM76, MBBS78, MCM77, NEK78b, NEK78a, NCPW75, NT78, OT78, OKS77, OO79, OWW78, PSZK78b, Pau75, PRGH78, PD75, PSG^{+}76, RTC76, SFS77, SRKM78, SMS78, SAB75, ST78a, SMG77, ST78b, STWB78, SS76c, SS76d, dSN77, KT77a, RFB^{+}76, ST76a, SSS77]. electrophoresis [HD77, LMPI76, PPL78]. Electrophoretic [AG75, BBDP79, LLC79, NSHE75, RHHB76, Sne76a, EL75]. Electrophorus [CBM78, BPRC78, KHV^{+}78]. Electrophysiological [MPS79, GD77]. electroplaque [BPRC78]. electroplax [KHV^{+}78]. electrotonic [HKP78]. elegans [CT79]. Element [CSP79]. Elemental [SSS77]. elements [CSHS77, IF79]. elevated [HS76c]. elevation [DBB^{+}79]. Elimination [Mau77]. elongation [BS79b, Bry76, DR77b, JF75, LNMR78, MPHMT77, WB78a]. elution [San78]. embedded [PSZK78b]. embryo [BN76, BH75b, DN76, DUKA77, Gra75a, Gra75b, Hir77, JL79, ML76, RTBE78, RGE78, SRR76, Smi77, TR79b]. embryogenesis [AP75, DUKA77, ODS79, Rag77]. embryogenic [VRN76]. embryonic [BCB77, BFD78, CBB77, FBF79, FF79b, GPPR76, GM76, JM79, JM77, LMPI76, Lem79, MGF77, MB76, McG76, OSM79, OBKP75, PL76, PPL78, RPGR77, RR77b, YBB79]. embryos [HQ78, KMS77, KBS79, MIT76, ODS79]. emphasizing [Paa78b]. employing [FJD76]. enclosing [VKZB78]. encystation [RHHB76]. encystment [dSN77]. end [CG79]. endings [MIMMF77, PJ76a, SR78a]. endochondral [RA76]. endocytic [DMB78, RKL^{+}76, SSR76]. Endocytosis [JSG79, QLHTB79, HBS79, KT77b, PT79, STdDT79b, STdDT79a, TSS79, ZS79b, BK75]. Endogenous [AKD75a, BTB79, GKB78, HR79a]. endomembranes [JKB^{+}79]. endometrial [AKD75b, PS79]. endonucleases [FWB78]. endoplasm [PH75b]. Endoplasmic [HIP78, ALMS76, BN76, CLP76, Dam76, DMS^{+}76, HLR76, JBKF77, LH76, LF77, LWB78, MCC79, OKS77, PFK^{+}76, ST77a, ST77b, WLR79]. endosperm [Fus75a]. Endothelial [YR75, BC76, DR78, EVMP76, FSOW79, GBBZ78, LBDL76, SS79c, VJGG79]. endothelium [BSHP79, DMB78, DM79, SSP75a, SSP76, SSP78, SD76]. ends [CCTT77]. Energy [CMC77, Blo79]. Energy-dependent [CMC77]. engagement [RIK78b]. enhanced [SWF78]. enhancement [DMS^{+}76, KP77c]. enlarged [Ber75]. enriched [LNS76, NSHE75, RCMI75]. Enrichment [BF78, BS78b, LF76]. Entamoeba [KVB75, KSB75, SMPGR75, vVSL^{+}76]. enterocytes [CSP79]. entry [AMJR78, Bon75a, GWS77, MV75b]. enucleated [CKBD76]. envelope [BSOS76, DL79, ES79, FSSZ75, HO75, KRM77, MG75b, OWS75, RG78, ST78a, SKT^{+}76, TP76, VKZB78, WLR79, Zaa79]. envelopes [BC78]. Environmentally [Dam76]. environments [KEP76, KKKS76]. Enzymatic [HC75a, Md78a]. Enzyme [ND77, AKD75a, BN76, BSVG75, BES75, GHW75, HGW77, HR79b, NSBS77a, SS78a]. enzymes [HJP78, HW78, HIP78, RLJ76, RT76, VC77]. Enzymic [PJB77]. eosinophil [LLLG78]. eosinophilic [KS75]. Epidermal [MSK^{+}79, CC76, CMC79, GBBZ78, HMC79a, HMC79b, MKKV78, NS75, RG78]. epidermis [DFW^{+}78, EF75]. epididymal [NUY^{+}77]. epiehtlium [Cra79]. epipodophyllotoxins [KPF75]. epithelia [BCB77, CBB77, VHvdVB78]. epithelial [CRD^{+}78, CP79b, GCH78, HS75b, Kra77, LN75, MC78b, MT75, MP78, PR76b, RCSS79, WNQK78, ZW77]. Epithelioid [QWTI79]. epithelium [BCB77, CH79b, EB75, EM77, GLM77, KP75, Moo76b, MD75, RNK76, RHDL77, SPM79, Ste75, TWC75, WBY79, vDK79, vOMW77]. epoxy [GTH76]. equation [SEI75]. Equilibrium [SS76b]. equiperdum [RS79]. equivalent [LSWB78]. errors [BG76b]. Erythrocyte [AMJR78, BS77a, BB79b, DS75, GCM77, HBS79, HJP78, LLMW77, MM76c, SS77a, She79, STWB78, TD75]. erythrocytes [BAPG79, Coh78, LLP77, SS76b, TSS79, VL78]. erythroid [BAN75, BPN76, CO76, CI78, GLR75]. erythroleukemia [LH79, SH79b]. erythroleukemic [MR77]. erythrophores [BP77]. erythropoietin [BAN75]. Escherichia [HO75, SHM^{+}77, Wei77b]. essential [CH76]. established [GMDT75, RP77, RCSS79]. esters [BGK^{+}79]. estimated [LWB78]. estimates [WP78]. Estradiol [SNP77, KAD75]. estradiol-17beta [KAD75]. Estradiol-induced [SNP77]. Estrogen [AKD75b, PS79, AKD75a, KAD75]. Estrogen-induced [PS79]. etch [OKS77]. etched [HS79a]. etching [LB78, MEB77, SM76]. ethylmaleimide [MC79a]. eucaryotic [HR79b]. Euglena [BRV78]. Evaluation [MIMMF77]. events [BS77d, FF79a, KG79, LRG^{+}77, NEK78b, NEK78a, TDO78, TCT79, dMdS78]. Evidence [All75, BFT77, BB78, CB78, GTL77, GP78, HDS77, JHW78, LTMS79, OPD76, SEHW79, SA77, WRP75, Woo77, And77, AC79, ET79, GLR77, HB76b, JG76, KG78, KT77a, SS76c, Til76a, ZT75, BRR78, FSSZ75, HG79, Hil75, IP78a, MKKV78, MEM79, NLT76, Pol75, SKT^{+}76, STdDT79a, SSP75b, SC75b, SNP77, Tam78, TD75]. evidenced [NSBS79]. evolution [NCM78]. examination [BBC78, MR75b, RFB^{+}76]. exceeds [JHW78]. exchange [HT79a]. excitability [MT78]. excitable [ERSP76]. excitation [Han79, dCdM78]. excitation-contraction [dCdM78]. exclude [Bon75a, Bon75b]. existence [SSP75b, SC75b]. exocrine [BSVG75, CM75, GST79, IP78b, KRJ77, NMQY77, SPT78, SSG79, SG79b]. exocytic [HA76]. Exocytosis [HSM76, CH79a, HRD^{+}79, LRG^{+}77, OPF77]. exogenous [SB79]. exorbital [HO77]. Expansion [WGB78, MC78b]. expansion-contraction [MC78b]. Experimental [AJ78, SKT^{+}76, Slu76]. experiments [MW76, MST78]. explants [FSW77]. exploring [AB76]. export [Zaa79]. exportable [SES76]. exposed [AV78, KFP^{+}78]. Expression [WK79, CGK75, GGGM78, HWW75, JM77, LWS79, MB77a, TSC78b, KP77a]. extension [ABL76]. extensive [FS77]. External [HVW78, PDD78]. Externally [HC75a, HC75b]. extra [Cle75]. extracellular [CBB77, DST^{+}78, HR79a, MH75, SC75a]. extracted [LG75b]. extracts [CT77, IO79, Kan75, Kan76, Pol76, RF77, TRH76, Wei76b, Wei77a]. extrajunctional [FS76]. extraocular [SRKM78]. eye [KP75].

face [FSB75, Swe76]. facilitate [SE75]. factor [BTS^{+}78, BR76a, CC76, CP77b, CP77c, CP77d, CGVR79, DR78, FCG^{+}76, GM75, GWML76, GBBZ78, Gre78, HMC79a, HMC79b, HS75a, JB78, MS76, MOS^{+}77, NSBS77b, VJGG79, WS78b]. factors [CMC79, GG77, NSBS79, OKS^{+}79, RR76, Zig77]. fasciatus [Cav75]. fast [ERSP76, GSJH79, RMP^{+}78, SRKM78]. fasted [SG79b]. faster [CM77a]. fasting [SSG79]. Fat [BLR76, TBZW78, JBKF77, PJB77, dMdS78]. Fate [LP76, STdDT79b, STdDT79a, SH77a, WF79, WGF79, AP75, EL75, HA76, HC75b, MG75b, MF75, OH78]. fates [HS79b]. fatigue [BBW75]. fatty [SFFB75]. feasibility [FB76]. features [KT75, OSO76]. fed [SG79b]. feedback [Jau75]. feeding [SSG79]. female [Lon77]. femoral [LF76]. fenestrae [DM79]. ferritin [BW79, Far78, HM75, HMC79a, KOBP75, NSH78, NVH^{+}75, RFB^{+}76]. ferritin-conjugated [NSH78, NVH^{+}75]. ferritin-labeled [RFB^{+}76]. ferritin-protein [KOBP75]. ferritins [RCV75]. fertilization [DB76a, DL79, EL75, ES76, Lon77, SS79a]. fertilized [BS77c, SS79d]. fetal [HR79a, KSS76, LW76, MFPO75, THCP75, Wil77b]. feulgen [BG76b]. fiber [GL77, GL79b, Gau79, GW78, HG79, MT78, RV78, SMM78, SH79a, TS77]. fibers [BE79a, BE79b, BB78, Coo76, DF75, ERSP76, ET79, FALP75, HWP78, JT79, JJB79, OO79, RMP^{+}78, RS75c, VD76, WG78]. fibrillary [SHWH^{+}77]. Fibrils [Mau76, Bru76, HKS77, SKT^{+}76]. fibrinogen [FMSB75]. Fibroblast [BH77, YK79, BE78, Che79, GM75, GWML76, GBBZ78, Gra75a, Gra75b, GRW75, HH77a, HMT79, HKA^{+}79, JLA^{+}76, SWV77, VJGG79, Yam78]. Fibroblast-collagen [BH77]. fibroblastic [BG76a]. Fibroblasts [FTBS78, ABL76, BH75a, BAGB77, CC76, CW77, CT76, CKBD76, DN76, GM75, HH78, HVW78, HB79, KAGV79, LAS^{+}75, LWS79, MH76a, NVH^{+}75, NPM75, OKB76, OBKP75, PT77, PELG76, Poh76, RR76, SRR76, STdDT79b, Smi77, SVW79, TR79b, WP75, WY78, WMP79, YBB79]. fibrogenic [CH75]. fibroin [ST76a]. Fibronectin [LWS79, TR79b, CBH^{+}78, DST^{+}78, HR79a, HVW78, KS79, SWV77]. fibronectins [YK79]. ficoll [DWM75]. fifth [ST76a]. filament [BRR78, BFTH79, BC76, BS79b, BS79c, DWC^{+}77, EWR78, GP78, Lem79, MSP78, MT75, TM76b, TCT79, MT75]. filament-membrane [MT75, TM76b, TCT79]. filament-microtubule [GP78]. filamentous [Coo76, MT78]. filaments [BRR78, BB79a, BRS78, BS77c, Bur78, BG76c, CM77b, ET79, FC77, FSOW79, GCH78, GSN78, HL79, JJ75, JF75, KW76, KH75, LBDL76, Laz76, LYSS78, LG75b, MCM77, MD75, NP75, OLK77, Pol75, RH75, SO77, SBGG78, Til75, Til78, TK79, WCC79, WRP75]. Filopodia [AB76]. filopodial [Edd77]. filter [SM75a]. Fine [BPS78, LBDL76, HKP78, KEP76, MG75a, SMJ78, SC75b]. firm [All75]. Fischer [SMJ78]. fish [SPM79, SOW78]. fixation [Lar75]. fixed [HH78, SOK79, SAB75]. fixture [SE75]. flagella [BS77b, Bro79, Gol79, MNK79, OAB77, Rik78a, SG79a, SB75, Sne76b]. flagella-binding [Sne76b]. Flagellar [DR77b, FGG78, LNMR78, BJ75, BR78, BR76a, BLB79, BMP76, GJ78, HRL77, MM77, Sne76a, WPS78, BGG^{+}75, BRV78]. flagellate [HKS77, Tam79, TC75]. flagellates [Tam78]. flagellum [Blo77]. Flexibility [Pol75]. Flow [FPC76, CGK75, EL76, HD77, Kri75, LD76, SBC76]. fluctuation [LE78b]. Fluctuations [LE78a]. fluid [CBH^{+}78, LS79b]. fluidity [KT77a, OKS77, Tam79, UIOO76, WGB78]. Fluorescein [CG78]. fluorescence [HA77, JJ78, NSBS79]. Fluorescent [FP76, CW78b, CW78c, CK78, GL79a, HP79, Pai75, SS79a, Täl78]. fluorescent-heavy [HP79]. fluorescently [WT79]. fluorochrome [Col79]. fluoroorotate [CLP76]. focus [HQ78]. fold [LHH76]. folds [And76, BS77e]. follicle [AKLL75, AA76]. follicles [CE77]. follicular [BK78, CAS78, CFAS78]. followed [WMP79]. food [KT77b]. forces [WM78]. forest [HFK77]. Form [CS78, FWHB78, HLS75, SEI75, TD75, Til76b]. formaldehyde [DMS^{+}76].

Formation [Ber75, BJ75, KD78, BH77, Bel77, BK78, CP77a, CI78, Con79, DL79, Edd77, FF79a, GS75, Goo76, GB77, Gra75b, HDS77, HK75, JL79, KSB75, Lem79, MG76, OKB76, Pfe78, PR76a, Pol75, RG78, RIK78b, RGE78, SV77, SMPGR75, SC76, TCT79, TS77, WHC79, Gou75, Gra75a, MBPTS79]. formations [FKS76]. formed [CRD^{+}78, FF79a, GD76, STF75, War79]. forming [BPN76, KS79, NCCB77, PM79, dP78]. forms [Edd79, MPDGR76, PPMSK77, SDTD^{+}79]. Forssman [SB79]. fraction [DB76b, ESAD75, FLRS77, GBB78, Qui76]. Fractionation [San78, FSW77, SPT78, NBB77a]. fractions [BBC78, BPL^{+}78, CF76a, CF76b, DWM75, FRT^{+}76, HIP78, LNS76, LWB78, MV75a, NBB77a, NBB77b, NSHE75, RCMI75, ST77a, ST77b, War79]. fracture [AC79, BS77d, CGH79a, CGH79b, CH79a, CBC75, Con78, DPM76, DJM77, EGFB78, FB76, FDB77, GWN77, HHP76, HH77a, HH78, KHT^{+}79, LWB78, MPDGR76, McN78, MBBS78, OPF77, PWSA77, RG75, Swe76, Tam79, TDO78, dMdS78, vDK79, BLTP^{+}76]. fractured [GKBM78, KR75]. fracturing [GBM78, PJ76b, PR75, SM78c]. Fragilaria [TSPH78]. fragment [UKY^{+}79]. fragments [BS79b]. framework [BC77]. free [AC79, BH78, CBSP76, GJRR78, Gre78, HD77, LLP77, MV75a, MV75b, CGH79a, HH78, SE75]. free-flow [HD77]. Freeze [AC79, CGH79a, CGH79b, EGFB78, FB76, FDB77, GBM78, HH77a, HH78, OPF77, PWSA77, SM78c, Swe76, dMdS78, BS77d, CH79a, CBC75, Con78, DPM76, DJM77, GKBM78, GWN77, HHP76, KR75, KHT^{+}79, LB78, LWB78, MEB77, MPDGR76, McN78, MBBS78, OKS77, PJ76b, PR75, RG75, SM76, Tam79, TDO78, vDK79, BLTP^{+}76]. freeze-etch [OKS77]. freeze-etching [MEB77, SM76]. Freeze-fracture [AC79, CGH79a, CGH79b, EGFB78, FB76, FDB77, HH77a, HH78, OPF77, PWSA77, Swe76, dMdS78, BS77d, CH79a, CBC75, Con78, DPM76, DJM77, GWN77, HHP76, KHT^{+}79, LWB78, MPDGR76, McN78, MBBS78, RG75, Tam79, TDO78, vDK79, BLTP^{+}76]. freeze-fractured [GKBM78, KR75]. Freeze-fracturing [GBM78, SM78c, PJ76b, PR75]. freezing [CH79a, HRD^{+}79]. Friend [MR77, SH79b]. frog [AG75, CCH78, CGH79a, CGH79b, FALP75, Har78a, KBD77, MD77b, MED77, MD75, PSZK78a, PSZK78b, SHH76, SSG79, ZVR75, dCdM78, Bon75a, Bon75b]. frogs [AC79, SG79b]. front [SD76]. frozen [HS79a, TS76]. fructose [MCM77]. fucoid [NJ75]. fucoidin [HQ78]. function [AB76, Bal79, DWB79, EVMP76, ET79, HO75, Hel79, HRL77, HS79c, KKH77, KVB75, LEDW76, Lon76, MRG77, MC79c, TZ77, VC77, WGB77, WPS78, WDB79]. Functional [FFHL78, HS76d, NCGC79, Tak77, BD75b, EB75, LSWB78, PD76, SEHW79, SB75, SS76d, VJGG79, MRZV76]. functionally [CRBH78]. functioning [MKKV78]. Functions [WG78]. Fundulus [KKKS76]. fungal [AV78, DF77]. fungus [HH76]. furrow [FP76, FPP78]. Further [CCNAN78, OBKP75, RKL^{+}76]. Fusion [TP76, BK75, CH79a, DMKR75, Gor75, GWML76, HWLC78, KG79, LRG^{+}77, LZK75, Lu78, MM76b, MG78a, NPM75, OT76, OT78, OWS75, TKST78, TCT79, VL78, WGG77a, dSN77]. fusions [OPF77]. fuzzy [SM76]. g [GWS77]. G1 [Ley75]. G2 [CB75a]. galactose [WM77, NEK78a]. galactose-binding [WM77]. galactoside [BGFK79]. Galactosyl [FLRS77]. gallbladders [FMR79]. Galloylglucoses [SS76c, SS76d]. GAM [FGG78]. GAM-1 [FGG78]. gamete [SS79a]. gametes [GJ78, MG75a]. Gametic [BGG^{+}75, GW75, MG75a, GW78, MG78a]. gamma [FBF79]. gamma-aminobutyric [FBF79]. ganglia [PL76, RGE78]. ganglion [DK78a, LP76, WWLS79]. ganglionic [LP76]. ganglioside [JSG79, SRR76, SFFB75]. Gap [AA76, CGMP77, MCPG77, Per77, Bal79, BA79, Dec76, EF76, GCH78, Goo76, Her76, KD78, KP75, Lar75, LTMS79, MPO79, PM79, SHWPJ78, Som79, YR78, Her76]. gap-junctional [SHWPJ78]. gastric [BRS^{+}76, RCMI75, SHC^{+}79]. gel [IO79, LLC79, LMPI76]. gel-sol [IO79]. gelation [CT77, Kan76, Pol76, Wei76b, Wei77a]. gels [AJG78]. gene [AJ78, BE78, DBB^{+}79, GGGM78, HWW75, JM77, PP79]. General [SSC76b]. generality [HL75]. generated [CCTT77, NJ75, Til75]. generates [Tam78]. Generation [AEP79, WM78]. genes [PH75c, STF76]. Genetic [BLTP^{+}76, Col79, DCEO^{+}79, Har78b]. genetically [JG76]. genome [KT75, WF76]. genomes [SSC76a, SSC76b, SCS76]. genotype [LWS79]. geometrical [HS76d]. GERL [EH77, GKSA77, HO77, JSG79, Paa78a, Paa78b]. GERL-like [JSG79]. germ [DMKR75, MJM76, SC78]. germanica [Cav76]. germinal [CAS78, MH76b]. ghost [HBS79]. ghosts [CCNAN78, HJP78, TRH76, TD75]. giant [CB76, GTL77, KKAL79, LGB77, LKKA79, MST78, MSRT78, MT78, OKB76, PSGL78, VV76]. Giemsa [MDR75]. gill [KEP76, KKKS76, MS75b, Ste75, WS78a]. gills [SPM79]. gland [AJ78, CJP75, EM77, EKS^{+}79, FS77, FAE75, GLR77, HO77, HSM76, HS75b, JBKF77, KG78, Kra77, MR75a, OB78, Pai75, PPHM75, RJ76, RLJ76, ST76a, ST76b, SC77, SNP77, TP76, dMdS78]. glands [DAF77, DH75, PBJ77, RHDL77, Von79]. glass [SE75]. Glaucocystis [WB78b]. glia [GTL77, HHP76, LGB77]. glia- [LGB77]. glia-neuron [GTL77]. glial [GESG75, GSN78, LGB77, SL78, SHWH^{+}77, WHC79]. glioma [MOS^{+}77]. globin [BAN75, BPN76]. globule [JBKF77, PJB77]. Glomerular [SR78b, KF79, KR75, RCV75, SM75a]. glucagon [BRS^{+}76]. glucagon-producing [BRS^{+}76]. glucan [SM75b]. Glucocorticoid [GET^{+}78, Ber75]. Glucocorticoids [RPGR77]. glucose [NRH78, WJTR77, MR75a]. glucuronidase [LP77]. glucuronyltransferase [BN76]. glutamine [JM77, VC77]. glutaraldehyde [HH78, Lar75]. glutaraldehyde-calcium [Lar75]. glutaraldehyde-fixed [HH78]. glutathione [BOP^{+}78, EVMP76, NR76, OAB76]. glutathione-oxidizing [OAB76]. Glycine [PSG^{+}76]. Glycine-specific [PSG^{+}76]. glycogen [MCC79]. glycogenesis [BWDP77]. glycolipid [SB79]. glycopolypeptide [MG78a]. Glycoprotein [KG77, GWS77, KL79, LF77, MG78b, NEK78a, NSHE75, YYP77, Yam78]. glycoproteins [ASD75a, ASD75b, BCB79, BKS78, BSPK78, DWB79, ESAD75, GHA76, Gou77, WNQK78, WLR79, WDB79]. glycosaminoglycan [CCB76, CBB77, CRBH78]. glycosaminoglycans [WR75c]. glycosides [VV76]. glycosylation [BCB79]. glycosyltransferase [WNQK78]. glyoxylate [RT76]. GM1 [JSG79]. GMP [SGV75]. goats [BAN75, BPN76]. gold [ST78b]. golden [Woo77]. goldfish [OY78]. Golgi [AMS76, BBC78, CF76a, CF76b, ESAD75, FD78, FLRS77, GST79, HO77, HIP78, IP78a, LH76, Paa78a, Paa78b, PSZK78b, PBJ77, SGCS75, WNQK78]. gonadotropin [AKLL75]. Gonyaulax [Swe76]. gradient [WCRdD78]. gradients [DWM75, Zig77]. gram [Zaa79]. gram-negative [Zaa79]. granule [CP78, CH79a, HSGW76, HS76b, LLLG78, OKB76, SLJ77]. granule-formation [OKB76]. granules [CJP75, CP78, COK76, DDSL77, FSB75, FS77, FBSS78, GZM75, HO77, MD75, NS77, NSHE75, OLK77, RD78, RP78, SLJ77, TBZW78]. granulocytes [AL75]. granulosa [AA77, CE77, LGGB79]. greatly [Poh76]. green [AMBM76, FSSZ75, SSF76]. Greene [DMMW76]. ground [BP77, WP79]. groups [SS76d, SF76, RW78]. growing [EB78, HS77, JWJ75, JLA^{+}76, WNJ75]. grown [FSW77, SV76]. Growth [FSSZ75, Lan78, LE78b, LW76, SVW79, WPMW75, AKD75a, BAN75, BG76a, CC76, CMC79, CH79b, CP77b, CP77c, CP77d, CCB76, CGVR79, DR78, GM75, GWML76, GBBZ78, Gre78, HMC79a, HMC79b, HMT79, HS75a, HB76b, JC79, JLA^{+}76, KAD75, LE78a, LE77, MR77, MH76a, MOS^{+}77, NPM75, OAR^{+}77, OWC^{+}77, PS79, PRZE75, PEA77, RBB76, SV76, SO75, SM75c, TZ77, VJGG79, vVRN76]. Growth-related [LE78b, BG76a]. guided [AB79]. guinea [CM75, FOPY77, GB79, GL79a, LF76, LLLG78, MPJ79b, MPS79, MCP^{+}78, NSHE75, Paa78a, Paa78b, RP78, SPT78]. guinea-pig [GB79].

H [SHC^{+}79, SC77, BE78, TKST78]. H-2 [BE78]. H1 [JG76, TKK79]. hairs [FC77]. Halobacterium [BSOS76, WAC78]. halobium [BSOS76, WAC78]. hamster [BO76, DCEO^{+}79, FK78, GCY79, GB77, GTH75, GTH76, Jul78, Ley75, PT77, RS75a, RW75b, SRR76, SC78, SBGG78, WR77c, Woo77]. haplo [CM77a]. haplo-x [CM77a]. haploid [PP79]. Hapten [NCPW75]. Hapten-sandwich [NCPW75]. harvest [LF76]. harvesting [ASA77, MMM76]. HCO3 [KCH79]. HCO3-stimulated [KCH79]. healing [GCH78, JJ75]. heart [ET79, GE79, Lem79, SS78d, WKP^{+}77]. Heat [RB77, DTSM77]. Heat-induced [RB77]. Heavy [BRS78, BBC78, CRSH75, HP79, KH75, LMPI76, LE76, MC79a, SFS77, Wei77a, WRP75]. HeLA [SA77, CKJB77, CCNAN78, GRW75, HMDH77, JK77, MZ78, OPD76, PEA77, SGMB76, SFFB75, THC76, TR79a, War79, WLW78, Wei76b, Wei77a, ZJSF76]. Helical [KKAL79, KVB75]. helices [KSB75]. heliozoan [Sch76]. Hemagglutination [MM76a]. heme [CI78, GLR75, SSP75b]. heme-peptides [SSP75b]. Hemiptera [Cav75]. Hemoglobin [BAN75, BPN76, BLR76, GLR75]. hemolysis [MM76a]. hemopoiesis [RA76]. henbane [Rag77]. hepatectomy [MN77]. Hepatic [DMS^{+}76, MCC79, RR78, RBM^{+}78, SMJ78, SSSW77, Svo78]. Hepatocyte [FI77, RR78]. hepatocytes [BSHP79, BWDP77, BBW77, CM77c, DWM75, FMSB75, FKNMT79, LW76, MFKT78, MN77, NY78, PCJG79, RCM77, RBHP75, WDM^{+}75, WDMR77]. hepatocytic [BPL^{+}78, LWB78, WP78]. hepatoma [PRZE75, QLHTB79]. herpes [SRR76]. heteroclitus [KKKS76]. Heterogeneity [BBDW76, FWB78, GP77]. Heterogeneous [HWP78]. heterokaryon [MDR75]. heterokaryons [EW77a, EW77b, Gor75, LWS79]. heterologous [BD75b]. heteroploid [WGC76]. hexagonal [RB77]. Higashi [OKB76, BES75, FTBS78]. Higashi-like [FTBS78]. High [IF79, MBM^{+}79, SS78c, Amo77, AHB79, BR76a, BDKS79, BP77, CGH79b, CKCK78, CW78d, DGR75, GHC78, KEP76, KKKS76, KRM77, Pau75, PRGH78, SAB75, SS78a, ST78a, TM76a]. high-resolution [AHB79]. high-yield [SAB75]. higher [JJ78, MRDC77]. highly [BG76c, RCMI75]. HindIII [FWB78]. histochemical [AFB79, JRFS75, LJ75, MHC77]. histocompatibility [CGK75]. Histogenesis [TS77, HP77]. histolytica [SMPGR75]. Histone [KMB78, JG76, TMGC76, TKK79]. histones [AG75, Bon75a, FSM76]. history [RLRV79]. histotypic [RTBE78]. HL [NCGC79]. HL-60 [NCGC79]. HnRNA [MHP78a]. Hoechst [MDR75]. homogenates [OT76, WR75a]. Homogeneous [HD77]. homogeneously [DBB^{+}79]. homogenization [SPT78]. Hormonal [BA79, Dec76, SBULB76, TWR^{+}77]. Hormonally [BAPG79, LGGB79]. hormone [ES79, HARO79, Hop77, HG77, Mil78, MC79c, NBB77b, SO75, SM75b]. hormone-accelerated [ES79]. hormone-releasing [HG77]. hormones [BH75b, Kah76, NBB77b, RHDL77, TM79, Von79, WOPB79]. horseradish [CFAS78, JL79, OH78, PT79, TDHO76, WW75]. Host [AV78]. Host-pathogen [AV78]. HRP [CFAS78, CFAS78]. Human [MB75b, NP75, AEJL75, AZW^{+}79, AKLL75, BGRC77, BGRC78, BS79a, BAGB77, BP79, BS77a, BDKK75, BB79b, BOP^{+}78, CC76, CMC79, COK76, CKBD76, CG79, EW77a, EW77b, FSOW79, FP76, FP78b, FBSS78, GG77, GSN78, GHW75, GM75, GBBZ78, HMC79a, HMC79b, HMT79, HVW78, HKA^{+}79, HSGW76, HGW77, HW78, JF75, KNM79, KPF75, KAGV79, LAS^{+}75, LLP77, MOR78, MRG77, MR75a, MDR75, NEK78b, NEK78a, NCGC79, NPM75, OAB76, Poh76, RG78, RBK78, SL77, SS76b, SS77a, STWB78, TSS79, VMR79, WBE75]. HVJ [UKY^{+}79]. hyaline [DS77]. hyalocytes [NLT76]. Hyalophora [TBZW78]. hyaluronate [UT79]. hybridization [Leó76]. hybrids [Gor75, WK79]. hydro [MM78]. hydro-osmotic [MM78]. hydrocarbon [WGC76]. hydrogen [BKK75]. hydrostatic [SE75]. hydroxylase [GET^{+}78, WGC76]. hyperplastic [ZW77]. hypertonicity [SSS77]. hypertrophy [Hil75, MKM76, OAR^{+}77]. hypolipidemic [Svo78]. hypothalamus [NZD^{+}79]. hypothesis [GTL77, HT79b, LGB77, LFR78, MCR77, dSN77]. hypoxanthine [PRZE75].

I- [FS76]. I-radioiodine [PL77]. identical [YK79]. Identification [BRS^{+}76, LKKA79, MB79, PH75b, PDD78, RTC76, SEBL77, WM77, EGFB78, ET79, Gau79, KC78, MM79, Wil77a, WDB79, ELD76, HL75, KUS78, PSG^{+}76, WF79]. identified [GKS76, KCGN76]. II [AMBM76, ASA77, ASD75b, BE79b, BGG^{+}75, BHR77, BD75b, BCS77, Bon75b, BRV78, BSPK78, BFPU78b, CGH79b, CW78c, CF76b, CP77c, CH76, DWB79, DS79, Edd75b, EW77b, ERSP76, Gra75b, HH78, HA76, HC75b, HS79b, IR78, JZR76, JM77, KP77b, KKKS76, KUS78, KFP^{+}78, Kyt76b, LEDW76, MM76b, MWS79, MPJ79b, McG76, MV75b, MC78a, MMM76, Md78b, NCPW75, NP75, NBB77b, OT78, Paa78a, PH75a, RLJ76, Sal75b, SBULB76, STdDT79a, ST77b, SS76d, SSC76b, Sne76b, SH76, SS78f, TRH76, Til76a, TSC78b, WKP^{+}77, WP78, WR77b, WR75c, WKBL75b, WGF79, ZM76]. III [CP77d, DR77b, ESAD75, GMDT75, GW75, GM76, HS76a, HQ78, LWB78, MM76c, MV75c, NSRL76, Paa78b, RKL^{+}76, RTBE78, SGMB76, SFS77, SPS76, SCS76, Til76b]. ileum [GB79, JZR76, RKL^{+}76]. images [AB77]. imidodiphosphate [PW78]. immature [AKD75b, NRH78]. immotile [AEJL75]. immune [CFAS78, TC75]. immunoadsorbent [PK79]. immunoadsorption [IP78a]. Immunochemical [MMS^{+}76, STWB78, JM77]. Immunocytochemical [GBB78, GST79, KRJ77, MW78, PSZK78a, PSZK78b]. immunocytochemistry [Gau79]. Immunoelectron [MFKT78, CBDY79, MMS^{+}76, OSM79]. Immunoferritin [Kyt76a, Kyt76b, OBKP75, TS76]. Immunofluorescence [BB79b, JG76, SL77, BTC78, BW78, CLM77, CBDY79, CKCK78, EL78, JJB79, JKZM77, JKM79, KMB78, NM78, OWW78, SOW78]. Immunofluorescent [AFB79]. immunoglobulin [BD75a, BSU79, GTPWD76, TV79]. immunoglobulins [GGS^{+}76, Rod76, STdDT79b, dP75]. Immunohistochemical [TSC78a]. immunolactoperoxidase [Bou79]. immunolatex [MDRY75]. immunologic [QWTI79, SNP77]. Immunological [FF79b, HB76b, Sch77, VV77, Yam78, KNL76, SBGG78, FRT^{+}76, SHC^{+}79]. immunoperoxidase [BCFS75, SHWH^{+}77]. immunospecific [PRGH78, NCPW75]. impaled [MST78]. impermeant [GES79]. implications [HU77, LFR78, MWS79]. importance [SSR76]. Improved [TS76, FBSS78, KOBP75, MB76, OT78, SFSB78, WRP75]. inactivation [PT79]. inactive [JG76]. incisures [PSZK78a]. including [MZ78, PS75, vDK79]. incompatible [VJGG79]. Incorporation [BP76, DVEM79, GD76, Gou77, SP76, ASD75b, FKNMT79, HB79, MZ79, MR75a, NEK78b, NEK78a, WJ76]. Increase [MPO79, RBC77]. Increased [Smi77, GNB^{+}77, Poh76]. incubated [BGK^{+}79]. incubation [NBB77b]. independence [KS79]. independent [Ern75]. Indian [PH75c]. indicator [EB75]. indirect [CLM77, CKCK78]. Individual [OWW78, RMP^{+}78]. induced [AKD75b, Bou79, Bro79, Dam76, EP79a, EF76, Fuk78, GJ78, HA76, HW78, HS76b, KP77a, Kan76, KSB75, LP76, LGGB79, LFR78, LS79a, LG75b, LZK75, MG78a, MR76b, MS75b, PS79, RA76, RB77, RHDL77, RS75b, SV76, Sal75a, Sal75b, SGMB76, SPS76, SFFB75, SH79b, SS78d, SLR77, SNP77, SC76, VL78, VMR79, War78, WGB78, BS77a, SS77a, ZS79b]. induces [HSGW76, WAC78]. Induction [Che79, MN77, AJ78, BAB79, CMC79, FF79a, GG77, GET^{+}78, LLL75, LNMR78, MH75, MGPO76, ZM76, HK75, MM76b]. Inequalities [MKK79]. infected [DDR75, KP77a, Mau76, SH77b, TE79]. infection [HMDH77]. Influence [DMKR75, HWLC78, SMJ78]. inhibited [CW78a]. Inhibition [FMSB75, HV75, Let79, Pfe78, WGE75, AL75, Bel77, BDCD75, BLB79, EP79a, HMPB76, Han75a, HH77b, LW76, LH79, LE77, RBHP75, RBC77, SG79a, SS78a, SS79c, PM75, PB79, SC75a]. inhibitions [LF76]. Inhibitor [HBD75, ALMS76, KSB75]. inhibitor-induced [KSB75]. inhibitors [Blo79, HP77, WDBM79, ZM76]. Inhibitory [PBJ77, SVW79]. inhibits [BS79b, Wei76b]. initates [Til78]. Initial [GTH76, BB79a, SO77]. initiated [Poh76, vVRN76]. Initiation [BR76a, SM75c, ALMS76, CM77a, DS79, Han75a, HG79, LW76, PB79, RHHB76, WR77c]. injected [MF75]. injection [HW75, SB76]. inner [BBDP79, FC77, FGSM78, Hir77, HH77b, SS78d]. innervation [FI77]. insect [BDP^{+}77, BJ75, TM76c]. insertion [CF79, WLR79]. insoluble [SM75b]. instar [ST76a]. insufficiency [EB75]. insulin [BWDP77, MPO79, Pfe78, BSHP79]. intact [GZM75, LF76, MST78]. integral [LLMW77]. Integrated [LWB78, WP78]. integration [BCB79]. integrity [AP75, TM76a]. intense [CGH79a, CGH79b]. Interaction [BK75, CB79, GG77, HP75, PH75a, BB79b, BB78, DMS^{+}76, DS75, GP78, Kan76, Laz76, MH75, MCM77, PM75, SOK79, SD76]. Interactions [HS76a, SH76, AV78, BH77, CCTT77, FKS76, JJB79, MM76a, MM76b, MM76c, MGF77, MEM79, MC79a, NS77, RD78, SS79a, dP75]. Intercellular [FAE75, GM76, Hel79, MB76, McG76, AL75, BE78, Let79, RW75b, WDMR77]. interface [PH75b]. Interference [HP77]. interferon [HV75, MR77]. Intergrated [BPL^{+}78]. interior [GR78]. Intermediate [ET79, FSOW79, LYSS78, BFTH79, BC76, BRS78, HL79]. Intermediate-sized [FSOW79]. Internalization [GKSA77, CC76, GGS^{+}76, HMC79a, JSG79, LTMS79, WGF79]. internodal [KW76]. interphase [DTSM77, Fuk78, RH79, SV76, SSD^{+}78, SWV77, GBB78]. Interrelated [HWLC78]. Interrelationships [GW78, KT77a]. intersegmental [LG75a]. Interspecific [MIT76]. interstitial [BA79]. intestinal [BW78, CB75c, CP79b, KCH79, McN78, MT75, Moo76b, MP78, NT78, RNK76, Rod76, SM76, WNQK78, vOMW77]. intestine [BTB79, QWTI79, RR77a, WC75]. intima [WR75b]. intra [Wil77a]. intra-axonally [Wil77a]. Intracellular [CW78b, CW78c, CKCK78, Har78a, MC78a, OB78, BSU79, CMC77, GTL77, GBE75, HARO79, MC79b, MC79c, Pfe78, SM75b, SSG79, SS78f, TVD78, TV79, TC75, WG78, HS79b]. Intramembrane [RG75, CP79a, HH77b, MPDGR76]. Intramolecular [GP77]. Intranuclear [Fuk78, SSF76]. Intratesticular [RLRV79]. intraventricular [SB76]. intravesicular [RW78]. intrinsic [PSZK78a]. Introduction [MM76c]. intussusceptive [HB76b]. invadens [KVB75, KSB75, vVSL^{+}76]. invasion [AL75, Wei76a]. inversion [VK77]. invertebrate [ELD76, LEDW76]. investigation [GL79c, MBBS78]. investigations [Mar77]. involved [KG79, NEK78b, NEK78a, OWS75, RGS^{+}76, SS76d, Sne76a, TM76b]. Involvement [FKS76, FLK^{+}76, NSBS79, TKK79, WCC79, Edd75a, Gre75, NR76]. iodide [FPC76, Kri75, RW78]. iodide-osmium [RW78]. iodinated [GGS^{+}76, HC75b, Leó76]. iodination [Bou79, HC75a]. Iodine [SFS77]. Iodine-125 [SFS77]. iododeoxyuridine [MF75]. ion [RD76]. Ionic [RD78, CBM78, NSBS77a, VV76]. ionophore [BFPU78b, FS77, HW78, ODS79, Sch76, Sch75]. ionophore-induced [HW78]. ions [ODS79]. Iontophoretic [GHMD77]. iridophores [RR75]. iris [ST78b]. iron [GJAE75]. irradiation [SV76]. islet [NBB77b, Täl78]. islets [LNS76]. isoenzymes [GL79b, WTE77, WKP^{+}77]. isolated [And77, BFT77, BCS77, CBBS77, CM77c, DL79, DFW^{+}78, FFHL78, FRT^{+}76, GCRB78, HMPB76, HK75, HD77, HSM76, HB78, IF79, JZR76, JKB^{+}79, KAB75, KKAL79, LKKA79, MB79, Moo76b, OO79, PFK^{+}76, PSC^{+}76, PCJG79, Qui76, RCM77, RFB^{+}76, Rem78, RNK76, SL78, Sch77, Sch78a, SOW78, SC77, SCRP77, VHvdVB78, WJ76]. Isolation [BFD78, BC78, CB77, Con79, DDSL77, DWM75, Edd79, FWHB78, GV77, GJAE75, GSN78, HKA^{+}79, KNM79, LF76, Md78a, Md78b, MIMMF77, RCMI75, RHHB76, TM76a, TC75, VKZB78, WDM^{+}75, WCRdD78, vVSL^{+}76, BCM^{+}77, BC77, BDKS79, CKJB77, DB76b, FBSS78, GES79, Gou75, Ste75, Wei76a, WH77, CJP75, Kra77, KR79a]. isoleucine [Ley75]. isomaltase [NT78]. isoproterenol [MC79b, OH78]. isoproterenol-stimulated [OH78]. isopycnic [WCRdD78]. isothiocyanate [GES79]. IV [DMS^{+}76, JLA^{+}76, KAB75, LNMR78, MSK^{+}79, RGE78, SMS78, TKST78]. IX [CM77a, LW76].

Japanese [Mil77]. junction [AMJR78, Bal79, CGMP77, Con78, Dec76, EF76, FF79a, Goo76, HS76d, JL79, Lar75, LTMS79, LS79a, LD76, MCPG77, MGPO76, RBB76]. junctional [CP79b, EVMP76, FS76, LHH76, SPM79, SHWPJ78, TM76a]. junctioning [Som79]. junctions [AA76, BA79, CGH79a, CGH79b, Cla76, CE77, ERSP76, FS76, FCG^{+}76, GCH78, HLR76, Her76, HGP^{+}76, KD78, KP75, MPO79, MCP^{+}78, MFPO75, PD76, Per77, PPHM75, PE79, PM79, RW78, SHWPJ78, SSP75a, SSP76, SS77b, SSP78, TWC75, YR75, YR78, vDK79].

K-dependent [TSC78b]. kangaroo [BBMB77]. KB [WR77a, WR77b]. Keith [Pal77]. keratinocyte [MKKV78, RG78]. kidney [Ern75, GCY79, GTH75, GTH76, HD77, KR75, PS75, RCSS79, RW79, SBGG78]. kidneys [BES75]. killifish [KKKS76]. kinase [UGB^{+}79, WTE77, WKP^{+}77]. kinetic [SS76b, MV75b]. Kinetics [HS76c, BK78, Hop77, Pfe78, WKBL75a, WKBL75b]. Kinetochore [Moe79, Moe76, RIK78b]. kinetochores [PB79, Rie79]. kinetoplast [FWHB78, FMW75, FWB78, MW76, RS79]. kinetoplast-mitochondrial [RS79].

L [HC75a, HC75b, SSD^{+}78, THC76, VRN76]. L-929 [THC76]. L. [MT78]. L5178Y [DBB^{+}79]. label [BR76b, HR79b]. labeled [CLP76, CC76, GES79, RFB^{+}76, WT79, NSH78, PSC^{+}76]. labeling [BRS78, GL79a, HPH77, MF75, NCPW75, PRGH78, SHWH^{+}77, STWB78, SRR79, TS76, MPJ79b]. labled [PL77]. Lack [AEJL75, FMR79]. lacking [BBDP79, CMB75, WPS78]. lacrimal [HO77, HSM76]. lactating [JBKF77, SC77, dMdS78]. lactation [RJ76, RLJ76]. Lactose [Von79]. laevis [DH75, HK75, MH76b, ST78a]. lamellae [SH77b]. lamellar [SV77, Wil77b]. lamina [BCB77, CBB77, DB76b, GBB78, KRM77, SMM78, SH79a]. laminae [KF79]. Lanthanum [HMPB76, Con78]. Large [BP79, DAF77, WNJ75, Bon75a, Cav75, DVEM79, PSZK78a]. Large-sized [DAF77]. larvae [RT76]. Laser [MS75b, Tam78, BR77, BRBM77]. Laser-induced [MS75b]. late [GGGM78, KMS77]. later [ST76a]. Lateral [PPL78, EW77b, HH77b]. latipes [GJRR78]. lattice [GSS77, GSS79, WP79]. layer [CE77, GV77, RG75, SHM^{+}77]. laying [Mil77, SBU76]. lead [LJ75]. leading [CG79]. leaves [JHW78]. lecithin [GD76]. Lectin [NUY^{+}77, SBC78, BTB79, BKS78, CO76, GKB78, KCGN76, NYY75, NSH78, RS75c, vVRN76, MPJ79b]. Lectin-binding [NUY^{+}77, BKS78, NYY75]. lectin-coated [RS75c]. lectin-initiated [vVRN76]. Lectin-labeling [MPJ79b]. lectins [BGFK79, GKSA77, PK79, RS75b]. length [DWC^{+}77, Sal75a]. Lens [VHvdVB78]. lenses [VHvdVB78]. lethal [LMPI76]. LETS [MH77]. Letter [DB75, EB75]. Letters [Pro75]. leucine [NEK78b]. leucoplast [SCS76]. leukemic [NCGC79, VMR79]. leukocyte [BTS^{+}78, BB79a, BKK75, NSBS77a, NSBS77b, SGV75]. leukocytes [BO78, BDKK75, BOP^{+}78, FSB75, GHW75, HSGW76, HGW77, HW78, KS75, MOR78, MBBS76, MBBS78, NSHE75, RBK78, WF79, WGF79, Zig77, Zig78, ZS79a]. level [NBB77a, NBB77b]. levels [LP77]. lewisi [FMW75]. Life [RLRV79, MG75b]. lifespan [BGRC77]. ligand [Bou79, NSH78, SD76]. ligand-induced [Bou79]. ligands [BSU79, HWAS79, HS79b]. ligase [FK78]. Ligatin [JM79]. Light [CP77a, ASA77, BBC78, BHR77, BD75a, CRSH75, GTPWD76, MMM76, OT78, OAR^{+}77, SE75, SRKM78, SMG77, SK79, WAC78, CH75]. light-chain [GTPWD76]. light-harvesting [ASA77, MMM76]. light-scattering [WAC78]. like [And76, BBMB77, ES79, FTBS78, JSG79, KH75, MG76, SMG77]. limb [VV77]. limited [MS76]. Limulus [DWC^{+}77, Til75, TCT79]. line [BS79a, BG76a, Ber75, CW76, GET^{+}78, KMS77, NCGC79, NSH78, RCSS79, GL79a]. lines [BCHN75, GMDT75, RP77, YYP77]. lining [BCFS75]. linkages [All75]. linked [OPD76, PSZK78b, RG78]. linking [KFP^{+}78]. Lipid [MM76a, MM76b, MM76c, BK75, DS75, HWLC78, RR77a, UKY^{+}79, WGB78]. lipid-soluble [BK75]. lipids [BW79]. Lipogenesis [TWR^{+}77]. lipolysis [CPO77]. lipoprotein [GBE75, KAGV79, LW76, NY78, RR78, TWR^{+}77]. lipoproteins [AHH76, BAGB77, BGK^{+}79, CW78d, PSC^{+}76]. liposome [EGFB78]. Liposomes [RW75a, DS75]. lipotropin [NZD^{+}79]. lithographic [MBM^{+}79]. liver [AG75, AHH76, ASD75a, ASD75b, BN76, BC77, BBC78, BSHP79, Ber75, BTB79, BBW77, BKS78, BSPK78, CLP76, CF76a, CF76b, CH76, DB76b, ESAD75, FRT^{+}76, GL79a, GJAE75, GBE75, GM76, HWAS79, HS79b, IP78a, JSGR78, JKB^{+}79, KRM77, KS77, LW76, LCK75, MB76, McG76, MHP78a, MHP78b, MFPO75, MGPO76, MR76b, MMS^{+}76, ND77, PSM^{+}75, Pfe78, PB75, RFB^{+}76, Rem78, ST77a, ST77b, TWR^{+}77, VC77, WCRdD78, WGC76, YR78]. living [BHFT79, Pai75, SOK79, SE75, WT79]. loading [GJAE75]. lobular [BCB77]. lobules [HSM76]. Local [EP79a, BFPU78b, NSP76]. Localization [AJG78, BTC78, BTB79, BW78, BDKK75, CB75b, CG79, EL78, JL79, JKM79, KW76, LC77, LLLG78, MED77, NT78, PL77, RBK78, WTE77, WKP^{+}77, WBY79, AHH76, AFB79, BVBM76, BAPG79, BPRC78, BPS78, CBDY79, CW78c, CF76b, CKCK78, CP79b, DK78a, DGR75, ELD76, Ern75, EM77, EKS^{+}79, FRT^{+}76, FP76, FPP78, FP78b, GBB78, GST79, GWK77, HA77, HQ78, HG77, JKZM77, KF79, KKKS76, KRJ77, KR79b, KMB78, Laz75, MRDC77, MFKT78, MW78, MP78, MMS^{+}76, NM78, NYY75, PSZK78a, PSZK78b, PBJ77, RV78, RGS^{+}76, RTC76, RT76, ST76a, SHWH^{+}77, SMG77, TSC78a, VC77, WDBM79, WR75b, BKS78, SS78f, Yam78]. Localized [LFR78]. located [FD78]. Location [NVH^{+}75, MMM76, SMJ78, SM75b, WF76]. Locations [PH75c]. locomotion [AL75]. Locomotory [Bel77]. locust [RW78]. Loligo [MT78]. longitudinal [JC79, SS78e, SS78f]. lorica [HKS77]. Loss [CW77, SSR76, YR78, BHR77, HBR77, KS79, MZ79]. Low [PD76, CBM78, KKKS76, KAGV79, LW76, SHWPJ78, SS76c, SS76d, TWR^{+}77, WPMW75, WGB78]. low-resistance [SHWPJ78]. luciliae [FMW75]. lumenal [CB77, MC78b]. luminal [DPM76, SD76, TP76]. luminally [OH78]. luminelle [SC75b]. lung [PSC^{+}76, Rie79, SV77]. lungs [Wil77b]. luteal [Paa78a, Paa78b]. Luteinizing [ES79, Hop77, HG77]. luteum [Paa78a, Paa78b]. Lygaeidae [Cav75]. lymphatic [LBDL76]. lymphoblasts [KPF75]. lymphocyte [BFPU78a, BFPU78b, SM77, dP75]. lymphocytes [BSU79, GGS^{+}76, MDRY75, PELG76, WGE75]. lymphoid [GHA76]. lymphoma [NSH78]. lysed [BTDT76, CW78a, PB79, SM75c]. lysine [TMGC76]. lysine-rich [TMGC76]. lysis [GW75, TC75]. lysosomal [BES75, GHW75, HGW77, HW78, NSBS77a, SNP77]. lysosome [ES79]. lysosome-like [ES79]. lysosomes [GJAE75, MBBS76, Paa78a, Paa78b, WCRdD78].

M [TV79]. Macroglobulin [WMP79]. macromolecular [BH75b, OWS75]. macromolecules [BFPU78a, BFPU78b, LZK75, PR76b]. macronuclear [BS78a]. Macrophage [MW77, HDS77, JB78, PDD78, PM79]. macrophages [BGK^{+}79, DS77, EH77, FTBS78, GG77, HS76a, HLG76, SH76]. maculosus [HGP^{+}76]. made [BO76, LF76, MV75b]. Magnesium [SS78d, CM75, HMT79, Wei77b]. magnesium-depleted [Wei77b]. Magnesium-induced [SS78d]. maintain [KT77a]. maintained [VJGG79]. maintaining [BCB77]. maize [GCRB78]. major [BDCD75, BCS77, GBB78, HL75, HS76c, LLLG78, MKM76, SWV77, Yam78]. malarial [AMJR78]. male [Lon77]. malhamensis [HKS77]. malignant [SFSB78]. mammalial [HP75]. mammalian [Amo77, AA76, BS78b, CLM77, EF75, ERSP76, GD77, GE79, Han75a, Han75b, HG79, HL75, JT79, Kri75, LG75b, MC78b, NYY75, NMQY77, PH75a, Sch78a, SM75c, THCP75, WF75, WPMW75]. mammary [HS75b, JBKF77, Kra77, PPHM75, RJ76, RLJ76, RHDL77, SC77, Von79, dMdS78]. manipulable [Hop77]. manipulation [Slu76]. map [WF76]. MAP2 [KBR79]. mapping [GD77]. MAPs [SS78c]. marginal [Coh78]. margins [PSZK78a]. marine [GSJH79]. marker [AKD75a, HIP78, RLJ76]. markers [KP77a, MDRY75, NCPW75, PRGH78]. marrow [CI78, LF76, NCCB77]. Mass [Ste75]. mast [BS77d, BS77e, FLRS77, LC76, LRG^{+}77, LFR78]. mastigonemes [BRV78]. mastocytoma [KSP75]. Mastotermes [BD78]. material [BCFS75, CCTT77, FLK^{+}76, GB77, KKG76, MR75b, TR79a]. materials [CBB77, SOK79, SC75a]. mathematical [FPC76]. Mating [KH76, Sne76a, Sne76b, GW75, GW78, GJ78, RH77, TM79]. Mating-reactive [KH76]. mating-structure [GJ78]. matrix [BC77, BB78, DST^{+}78, HR79a, HVW78, HKA^{+}79, HMDH77, MH75, RA76, WH77, WGB78]. matrix-induced [RA76]. maturation [BFTH79, CO76, DB76a, MC79c, NUY^{+}77, RBC77, ZS79b]. mature [JK77, JHW78, MPJ79b, SDTD^{+}79]. maturing [ES79]. maximal [BDCD75, Paa78a]. may [GPPR76]. MDCK [RCSS79]. mean [HIP78]. means [JL79, JC79, JJ78, OWS75, RFB^{+}76]. measles [TE79]. measure [DH76, LL79]. measured [CSHS77]. measurement [MB76, SB78]. measurements [RW75b]. measuring [PPMSK77]. mechanical [BE79a, BE79b, SS78f, SS78e]. Mechanism [FP78a, BS77a, BAB79, DF77, GWS77, KSB75, MPHMT77, MKM76, MRZV76, PH75a, PP79, RW75a, SG79a, SS77a, VK77, WY78, DMS^{+}76]. Mechanisms [AHB79, GHW75, BK75, BFPU78a, BFPU78b]. mechanoreceptor [MRZV76]. medaka [GJRR78]. mediated [GL79c, GSJH79, GP78, JF75, LS79a, Let79]. Mediation [DR78]. mediterranea [AMBM76, FKS76]. medium [GNB^{+}77, Gre78]. medullary [CB75b, FFHL78, Mil78]. megakaryocytes [LF76]. Meiosis [Lu78]. Meiotic [MRB75, BBMB77]. meiotic-like [BBMB77]. melanin [JF75]. melanocytes [JF75, JRFS75]. melanocytic [JF75]. melanogaster [CM77a, GP77, HSRP79, PE79, WF76]. melanoma [DMMW76]. melanophore [GHMD77, SEHW79]. melanophores [Sch78b, SOW78]. melanosomes [GHMD77, JF75]. Membrane [BCB79, CH79a, CKJB77, FOPY77, GCRB78, HHP76, HA76, Kah76, KG79, LS75, MSRT78, MV75a, MV75b, MV75c, NS77, RBC77, SWF78, SR78a, SMPGR75, SH79b, SBC76, Tam79, TCT79, WGG77a, WGG77b, dSN77, All75, AH77, AC79, AMBM76, ASD75a, ASD75b, BBH76, BAPG79, BTS^{+}78, BRR78, BLTP^{+}76, BCFS75, BD75a, BB79a, BB79b, BS77d, BS77e, BG76c, CF79, CPO77, CB77, CJP75, CG77, CCNAN78, DVEM79, DJM77, DDSL77, DDR75, EW77b, ESAD75, EM77, Far78, FYS78, FGG78, FF79a, FSSZ75, FKS76, FGSM78, FB78, FKNMT79, GJ78, GWS77, Han79, HH78, HJP78, HFK77, HS79a, HB79, HK79, HH77b, HR79b, HC75a, HC75b, JZR76, Jul78, KF79, KAB75, KH76, KUS78, KFP^{+}78, LTMS79, LRG^{+}77, LNS76, LP77, LAF^{+}77, LLMW77, MFKT78, McN78, MS76, MMM76, MB77a]. membrane [MC78b, MBBS78, MT75, MR75b, MSK^{+}79, NSBS79, NSP76, NT78, OPF77, PSZK78a, PDD78, PS79, PPL78, Qui76, RV78, RKL^{+}76, RLJ76, RdC77, ST78a, STdDT79b, STdDT79a, SM75a, STWB78, SC75a, SC78, SHM^{+}77, SS78d, SCRP77, Swe76, TP76, THCP75, TDO78, TM76b, TKST78, TSC78a, TSC78b, UIOO76, VKZB78, VV76, VL78, WOPB79, WM76, WDMR77, WS78b, WNQK78, Wei77b, WMP79, WLR79, WGB78, WDB79, ZVR75, BHR77, BPL^{+}78, HBR77, RLJ76, TCT79, dMdS78]. membrane-associated [BRR78, BB79a, BG76c, DDSL77, NSP76, WGB78]. Membrane-bound [MV75a, MV75b, MV75c, SH79b, BD75a, DJM77, FKNMT79, GWS77, LP77]. membrane-bounded [All75]. membrane-denuded [SCRP77]. membrane-enriched [LNS76]. membrane-label [HR79b]. membrane-to-membrane [FGSM78]. membraneless [GZM75]. membranes [ALMS76, AMBM76, ASA77, ASD75b, BGG^{+}75, BS77a, BD75a, BD75b, BBW77, BPL^{+}78, BCB79, BKS78, BSPK78, BP76, BBW75, BW79, CLP76, CP78, CF76b, DF75, DS75, ESAD75, EGFB78, ERSP76, FN76, FRT^{+}76, GGS^{+}76, GKBM78, GRW75, GWN77, HM75, HA76, HLR76, HS76c, IP78a, JZR76, JBKF77, JKB^{+}79, KHV^{+}78, KCGN76, KNM79, KT77b, KT77a, KUS78, KFP^{+}78, Kyt76a, Kyt76b, LAS^{+}75, Lar75, LF77, LWB78, LLMW77, MCC79, MM76c, MV75c, Md78a, Md78b, NSBS77a, NSBS77b, NYY75, NUY^{+}77, ND77, OY78, OKS77, PFK^{+}76, PJB77, RFB^{+}76, Rem78, RKL^{+}76, Ros78, RTC76, SHC^{+}79, SPS76, SS77a, SLJ77, Sta76, Swe76, Til76a, TSS79, WP78, WNQK78, WM77, Wil77b, WGB78, WDB79, vVSL^{+}76]. meroistic [Cav75]. meromyosin [BRS78, CP77a, HP79, KH75, LMPI76, MC79a, Wei77a, WRP75]. mesodermal [VV77]. mesothelial [MBPTS79]. mesothelium [SS77b]. Messenger [Gre75, CLP76, KS77, LAS^{+}75, LE78b, MV75c, SH77a, ST77b]. messenger-dependent [SH77a]. metabolic [DTT75, MST78, MIMMF77, HC75b, MSRT78]. Metabolism [BAGB77, PSC^{+}76, Blo79, BOP^{+}78, Gre75, Hil75, NR76, VC77]. metal [RW75a, SFS77]. metals [OLL76]. metaphase [BDKS79, DTSM77, TMGC76]. metaplasia [EF76]. method [BDKK75, CBC75, FPC76, FBSS78, IP78a, MF75, PJ76b, PPMSK77, WRP75, dCdM78]. methodology [FRT^{+}76]. Methods [PR75, KOBP75, MV75a]. methyl [DH76, DF77]. methyl-3H [DH76]. methylxanthines [NR76]. metrizamide [WCRdD78]. mexican [LMPI76]. mexicanum [LMPI76]. Mg [WGB78]. mice [AC79, BES75, FTBS78, KS79, OKB76, OWC^{+}77]. microanalysis [CSHS77, CSP79]. microbeam [BR77, BRBM77, SV76, Tam78]. microchemical [MBM^{+}79]. micrococcal [San78]. microdensitometric [BG76b]. Microendocytosis [KS75]. microexudate [SMPGR75]. microexudates [WPMW75]. microfilament [AA77, BW78, GCY79, GW75, MRG77, McN78, MBBS76, WDBM79, WP75, PH75b]. microfilament-associated [BW78, GW75]. Microfilament-chloroplast [PH75b]. microfilament-membrane [McN78]. Microfilamentous [BSVG75]. Microfilaments [HW78, CCTT77, Cra79, Edd77, EVMP76, LTMS79, PCJG79, SMG77]. microfluorimetry [CGK75]. microinjected [Bon75a, Bon75b, Pai75, SA77]. microinjection [HS76b]. Micromanipulation [BE79a, BE79b, EL76]. microperoxidase [WW78]. microperoxisomes [LN75]. microplasmodia [Mat77]. Microplicae [And76]. microprobe [GESG75]. microscope [BH78, CB75c, CBM78, CH75, DK78a, FS76, GWK77, GLM77, HARO79, HWAS79, HS79b, JRFS75, KKG76, KW76, KT77a, MFKT78, MR75a, MM78, MBBS78, MMS^{+}76, MCM77, NEK78b, NEK78a, OT78, OSM79, OKS77, OAR^{+}77, PSZK78b, PSG^{+}76, RFB^{+}76, SFS77, SRKM78, SMS78, ST76a, ST78a, SMG77, ST78b, STWB78, dSN77, Gra75b]. microscopy [AZW^{+}79, BTC78, BHFT79, BW78, BB79b, BW79, BP77, CGMP77, CBDY79, FRT^{+}76, GR78, HA77, IF79, KRM77, KHT^{+}79, KOBP75, LAF^{+}77, ML76, MSS75, MDRY75, MJM76, NCPW75, NT78, OO79, OWW78, Pau75, PRGH78, Pol75, PD75, RTC76, SE75, SAB75, SOW78, SS76c, SS76d, SK79, VP75]. Microsomal [SS79b, ASD75a, ASD75b, BCB79, BKS78, CB78, ESAD75, KUS78, KFP^{+}78, LAS^{+}75, LWB78, ND77, OKS77, SH77a]. microsomes [ASD75a, BD75b, BKS78, BSPK78, FRT^{+}76, FKNMT79, KUS78, KS77, MFKT78, MMS^{+}76, RFB^{+}76, SS79b]. microspheres [MJM76]. microsporidian [Wei76a]. Microtrabecular [WP79]. microtubles [CCTT77]. Microtubular [SEI75, Sch78b, LI79]. Microtubule [AA77, BOP^{+}78, DMMW76, PB79, SOW78, WHC79, Bry76, CB79, CKCK78, DR77b, EP79b, Fus75b, GB77, GP78, HG78, HSGW76, HGW77, MRG77, Moe79, OAB76, PM75, RCM77, RR78, RW75b, Sch76, SEHW79, SM77, WDBM79, WP75, WK79, ZMW79]. microtubule-associated [GP78]. microtubule-dependent [OAB76]. microtubule-organizing [SEHW79]. Microtubules [GE79, MOR78, NS75, PM75, TM76c, WB78a, WL76, All75, Amo77, BMP76, BHP75, CT79, DGR75, DR77b, EP79a, FTBS78, GS75, GW78, HW78, HB76b, HB78, KBR79, MEM79, MB75b, MNK79, OWW78, PSM^{+}75, PBH78, RR77a, RB77, Sal75a, Sal75b, SGMB76, ST76b, Sch76, SLJ77, Slu79, SM75c, SK79, TR79a, TSPH78, WCC79, WLW78, WR75a]. microvasculature [SSP75a]. Microvillar [Moo76b]. microvilli [BW78, BS77c, KSP75, MBPTS79, McN78, MT75, SM76]. microvillous [NT78]. microvillus [MB79]. microwell [TS77]. might [TK79]. migrating [MKK79, MD75]. migration [Bon75a, Bon75b, HARO79, MRG77, TS77]. milk [JBKF77, PJB77, dMdS78]. milkweed [Cav75]. miniature [SE75]. minutes [BS79a]. mirror [AB77]. mitochondria [AP75, Ber75, BTDT76, CS79, Dam76, JSGR78, KAB75, MST78, MSRT78, MCR77, PJ76a, RJ76, RLJ76, ST77a, SSS79, SS78d]. Mitochondrial [HLS75, LLC79, PEA77, BDP^{+}77, DN76, GP77, HM75, HH77b, KSS76, KKH77, LB78, RS79, ST77b, TK77, WGB77, WF76]. Mitogen [WJTR77]. Mitogen-stimulated [WJTR77]. Mitogenic [GM75, RHDL77, Smi77, WGE75]. Mitosis [Fus75a, IR78, RIK78b, Roo75, TPH77, BH75b, BR77, BRBM77, BC76, GWK77, HH76, MG75b, MWS79, Mau77, OWS75, SWV77]. Mitotic [DB76a, JRFS75, CSHS77, CLM77, CW78a, FP76, GBB78, Ley75, MEM79, MMER79, PB79, SEI75, SS78c, SM75c, WDBM79]. mitotically [CSP79, JG76]. Mobility [GD75, LE76, OKS77, CO76, GDR77, GTH75, KCGN76, RS75c, SC78, TSS79]. mobilization [BC76, BSU79]. mode [SV77]. model [KP77b, LS79b, NCGC79, RD78]. Modification [RF77]. modified [DB76b, MC79a]. modulate [RPGR77]. modulating [EF76]. Modulation [KAD75, WOPB79, BA79, WGB78]. modulator [Har78a, HMK79]. moiety [SS76d]. mold [Roo75]. molds [SB78]. Molecular [GZM75, JB78, LRG^{+}77, Leó76, Amo77, BR76a, BDKS79, CKCK78, DGR75, Edd75b, FN76, MHP78b, RD78, RCV75, RS79, SS78c, SS76c, SS76d]. molecule [RTBE78, RGE78]. molecules [BK75, FWB78, TV79, TD75]. molluscan [Ste75]. molting [MRZV76]. monitor [DDS79]. monkeys [RG75]. monogenase [MRDC77]. monolayer [CBC75, LW76, PWSA77, TWR^{+}77, WDMR77]. monolayers [CRD^{+}78, PR75]. monophosphate [DMMW76, RW75b, Sto75, WJTR77, ZJSF76]. mordant [SS76c, SS76d]. mordanting [SS76c, SS76d]. mori [ST76a, ST76b]. morphogenesis [IR78, MB75a]. Morphologic [ZT75, BAGB77, QWTI79]. Morphological [BH79, CPO77, HGP^{+}76, RBB76, SS78a, SFFB75, BCM^{+}77, CMC79, ET79, FFHL78, KP77a, KG78, MS75a, STF76, TCPM77, JZR76]. morphologically [SM78b, vVRN76]. morphology [BCB77, CBBS77, CT76, CGVR79, DWB79, DWM75, DTT75, FSW77, FF79a, KAD75, MM78, Moe76, NSP76, WP75, BH75a]. Morphometric [MC78b, MR76b, RdC77, PS75, Pfe78, MH75]. morphometrical [SG79b]. Morphometry [MCR77, OAR^{+}77, SM75a]. motile [TRH76]. Motility [Blo77, BS77b, Edd75a, Edd75b, LI79, AZW^{+}79, BLB79, EWR77, Hit77, LG75b, Moo76b, Tam79]. motion [HH77b, Til75]. motions [Edd75a]. motoneurons [BF78]. motor [Tam78]. Mouse [EL75, MDR75, SS79a, ABL76, BH79, BCM^{+}77, CGK75, CAS78, CFAS78, CH76, DUKA77, EW77a, EW77b, EH77, FS76, GHA76, GHC78, GRW75, HHP76, HT79a, HW75, HLG76, HC75a, HC75b, JB78, KP77a, KG78, KT75, KH75, KRM77, LP77, MBPTS79, ML76, MZ79, MRB75, MB77a, MS75a, RP77, RB76, RLRV79, RCMI75, RJ76, RLJ76, SRKM78, SHWH^{+}77, SSD^{+}78, TS77, TZ77, Von79, WL76]. Mouse-human [MDR75, EW77a, EW77b]. Movement [CCTT77, Rik78a, BH75a, BE79a, BE79b, CW78a, Che79, CT77, GOG76, HT79b, JF75, LAF^{+}77, Pai75, Pro75, SC75a, SH76, TRH76, WCC79]. Movements [LE77, GMDT75, NSBS77a, PP79, SEHW79, Sta76, Tam79, WG78]. moving [AMJR78]. mRNA [ALMS76, Gre75, GWS77, JLA^{+}76, LF77]. mRNA-associated [Gre75]. mucols [NS77]. mucopolysaccharide [SC75a]. mucosa [BRS^{+}76]. Mucosal [MM78, SHC^{+}79]. mucous [EF76, KG77, MR75a, SBU76, SBULB76, TP76]. muller [SL78]. mülleri [SC75b]. multiflagellate [BD78]. Multiple [HSRP79]. multiplication [Smi77]. multiplication-stimulating [Smi77]. mung [BTC78]. muntjac [PH75c]. murine [BD75a, CT76, CRBH78, GLR75, GM75, JSG79, LH79, MR77, SAPA77, SH79b]. murinus [CB76]. Muscle [OLK77, ASGR78, AFB79, BSM79, CRSH75, Coo76, CM77b, DR78, DF75, DWC^{+}77, EWR78, ERSP76, EZF76, FWS75, FSW77, FALP75, FF79b, FDB77, GB79, GL77, GL79b, Gau79, GNB^{+}77, GSS77, GE79, GSS79, HLR76, HL79, JKZM77, JKM79, LMPI76, MPS79, MCR77, NSRL76, OSM79, PPL78, PD75, RGS^{+}76, RR76, SRKM78, SMM78, SH79a, SS79b, SSP75b, SSS77, Som79, SSS79, SS78e, SS78f, Toy76, VD76, WTE77, WKP^{+}77, WR75c, WW75, WW78, YGH78, ZVR75, dCdM78]. muscle-conditioned [GNB^{+}77]. muscles [ELD76, JC79, LEDW76, RMP^{+}78]. musk [CB76]. mussel [WS78a]. mutant [BLTP^{+}76, DF77, FGG78, GW75, LMPI76, SAPA77]. mutants [CMB75, DCEO^{+}79, FJD76, Jul78, WPS78]. mutation [GS75, PE79]. mutations [HRL77, WGB77]. muzzle [DFW^{+}78]. myelin [GD76, Gou77, KHT^{+}79, MW78, SV77, Wil77b, YB75]. Myelinogenesis [TCPM77]. myeloid [DM79, NCGC79]. myeloma [BD75a, MV75a, MV75b, MV75c, SCRP77]. myoblast [FF79b, KG79]. myoblasts [BFD78, DMKR75, GWML76]. myocardial [LMPI76]. myofibers [SMM78]. myofibril [GL79a, SO77]. myofibrillar [RGS^{+}76]. myofibrils [EZF76, WTE77, WKP^{+}77]. myofibroblasts [GCH78]. myogenesis [BFTH79, KEG77]. myogenic [CW76, CH75, HWLC78, SO77]. myoid [WB78a]. myopathy [LS79a]. Myosin [SMG77, YBB79, BFPU78a, CRSH75, Con79, CM77b, FP76, FP78b, GL77, GL79b, HS76a, HA77, KR79a, KR79b, MO77, MP78, NP75, Pol75, SAPA77, SH76, Wei77a, Wil77a]. myosin-phosphorylating [SAPA77]. myosinlike [Bur78]. Myosins [OLK78, KNL76]. myotubes [Blo79, CP79a]. myristate [GHW75]. Mytilus [Lon77, MS75b]. Myxicola [KKAL79, LKKA79].

N [MC79a, MM77, DMS^{+}76]. N-dealkylations [DMS^{+}76]. N-ethylmaleimide-modified [MC79a]. N6 [RW75b]. NADH [BDKK75]. NADP [NR76]. nadph [IP78a, MMS^{+}76]. nadph-cytochrome [IP78a, MMS^{+}76]. nafenopin [SSSW77]. Narcine [BBW75]. Nascent [LSWB78, BD75a, MV75c, OPD76]. native [OAB77]. natriferic [MM78]. naturally [BBDP79]. nature [SRR79, CBB77]. necklace [Gra75b]. necklaces [Gra75a, Gra75b]. neclear [ST78a]. Necturus [HGP^{+}76]. negative [DJM77, Zaa79]. nematode [CT79]. neonatal [Rod76, SSSW77, TSS79, ZT75]. neoplastic [FLRS77, JRFS75, PPHM75]. Nerve [Gre78, JJB79, BBW75, BB78, BSM79, CP77b, CP77c, CP77d, CGVR79, FB78, GD76, Gou77, HS75a, Hir77, MT78, MIMMF77, MOS^{+}77, PJ76a, PR75, RV78, Sch77, SM78a, SF78, ST78b, TCPM77]. nervous [HK79, KCGN76, LYSS78, MOS^{+}77, SL77]. network [MT78]. networks [HS76d, WK79]. neural [JM79, JJB79, JM77, KMS77, MGF77, MG78b, NLT76, RTBE78, RGE78, SBC78]. neurite [RGE78]. neuritic [RBB76]. neuroblastoma [JSG79, MS75a, SM78b, JSG79]. Neurofilament [PSGL78, SF78]. neurofilaments [KKAL79, LKKA79, SL77, Sch77, SM78a, Sch78a]. neurohypophysial [GSB77, TDHO76, TDO78]. neuromuscular [CGH79a, CGH79b, Cla76, ERSP76, FS76, FF79a, FCG^{+}76, JL79, LS79a, PE79, RW78]. neuron [GTL77, GKS76, LGB77, RBB76]. neuronal [BCHN75, CT79, GKSA77, Gre78, GWN77]. neurons [CF79, CP77b, CP77c, CP77d, EB78, FBF79, HLR76, HL75, KCGN76, KR79b, Lan78, THCP75, WJB78]. Neurospora [LLC79]. neurotoxin [HK79]. neurotransmitter [CGH79a, CGH79b]. neurotransmitters [Kah76, SES76]. neurulation [MR75b]. neutrophil [AL75, COK76, MRG77, OKS^{+}79]. neutrophils [CG79, NSBS79, OAB76]. Newly [FD78, CCB76, FF79a, GD76, Hil75, MV75b, STF75, War79]. newt [Rie79]. Nexus [KBD77, CE77]. Nexuses [GB79, FDB77]. NH2 [SDTD^{+}79]. NH2-terminal [SDTD^{+}79]. nidulans [DF77, FSM76]. ninety [MZ78]. Nitella [PH75b]. nitrate [LJ75]. nitrophenylphosphatase [BPS78]. nm [LKKA79, OO79, SBGG78, WCC79]. node [YB75]. nodules [VKZB78]. non [CCTT77, JRFS75, JJB79]. non-actin-containing [CCTT77]. non-neoplastic [JRFS75]. non-neural [JJB79]. nonfilamentous [TD75, Til76a, Til76b]. nonhepatocytes [BBW77]. nonhistone [CH76]. nonhydrolyzable [PW78]. Nonidentity [HS77]. nonimmune [CK78]. nonjunctional [ERSP76]. nonlabeled [NT78]. nonmuscle [Hit77, Laz75, Laz76]. nonpermissive [GGGM78]. nonproliferating [TWR^{+}77]. Nonrandom [DMB78]. nonrandomly [Til76a]. nonreleasable [RMP75]. nonstrain [MHC77]. nonstrain-specific [MHC77]. nontoxic [UKY^{+}79]. Nonuniform [dP78]. normal [BCFS75, BPRC78, BB78, CT76, EP79b, GD75, GDC76, KR75, LP76, LWS79, MH77, MH76a, Moo76a, NY78, OAR^{+}77, Paa78b, PS75, PPHM75, PL76, RR77b, RWLC78, RW79, RHDL77, SS79b, SAPA77, SM78c, SLR77, WK79, DK78a]. note [DH76]. Novel [MG78a, Bou79, LS79b]. Novikoff [PRZE75]. nu [OSO76]. Nuclear [BC77, CH76, HMDH77, Mau77, ST78a, WGB77, BG76b, Bon75b, BC78, EL75, ES79, FP78a, FSSZ75, GBB78, GWK77, HWP78, IR78, JHW78, KRM77, MG75b, Mau76, MD77a, MZ79, MHP78a, MHP78b, Md78a, Md78b, OWS75, PFK^{+}76, RHHB76, SKT^{+}76, SSC76a, SSC76b, SCS76, SC76, WS78b, WH77, WGB78]. nuclease [Dip76, San78]. nuclease-digested [San78]. nucleated [Coh78]. nucleates [GB77]. nucleating [MB75b]. Nucleation [BS79c, HB78, PB79]. nuclei [BC77, Bon75a, Bon75b, CH76, DB76b, HWP78, KRM77, OO79, PB75, San78, SSD^{+}78, SSS79, SCRP77, WCC79, ZM76]. nucleic [SSC76b]. Nucleocytoplasmic [Pai75]. nucleofilum [Edd75a, Edd75b]. nucleoids [KKH77]. Nucleolar [Gra75a, Gra75b]. nucleoli [AHB79, MG76, RHHB76]. Nucleolus [SV76, MG76, RS75a]. nucleolus-like [MG76]. nucleosides [PRZE75]. Nucleosome [RH79, TKK79]. nucleosomes [San78]. nucleotidase [CDH^{+}79, RBC77]. nucleotides [BO76, RMP75]. nucleus [DB76a, Fuk78, JG76, Mau76, MH76b, RGK78, SCS76]. nude [KS79]. number [CO76, PBH78]. numbers [Moe79]. nutritional [HB76a].

O [DS75]. O2’ [RW75b]. O2’-dibutyryl [RW75b]. observation [dSN77]. Observations [Coh78, Sch78a, CH75, DB75, Gra75b, RKL^{+}76, dMdS78]. observed [AZW^{+}79, KHT^{+}79, ST78a, SK79]. obtained [IP78a]. Occluding [PPHM75, MCP^{+}78]. occurrence [MD75]. occurring [Blo77]. occurs [CB78]. Ochromonas [HKS77]. octanoic [CB75c]. off [FB78]. old [SES76]. oligosaccharides [AV78]. omental [MBPTS79]. Oncopeltus [Cav75, LG75a]. one [Swe76]. onto [TR79a]. oocyte [AA76, Bon75a, Bon75b, Cav76, GEB78, ZM76]. oocytes [AHB79, CM78, ES79, FP78a, HS76b, MH76b, MC78a, RB76, ST78a, STF75, Sch75, WJ76, ZM76]. oogenesis [STF76]. Oorthoptera [Cav76]. Opaque [Lar75]. Open [SSP78]. openings [FALP75]. opsin [PSZK78b]. optic [GKB78, TCPM77]. Optical [GSS77, AZW^{+}79, SE75, Toy76]. Optional [SE75]. order [MOR78]. ordered [BG76c]. Orderly [VRN76]. organ [BBW75, MIMMF77]. organelle [BCHN75, DMS^{+}76, GLR77, SSC76b, SGCS75, SC75b, Til78]. organelles [BFT77, BBW77, ES79, Rem78, SSC76a, SSR76]. organism [BBDW76, HB76b]. Organization [BMP76, CT79, HS79a, MT75, RW79, SS77b, BRV78, CBB77, DWB79, DUKA77, GZM75, HS76d, JT79, KR79b, Laz76, MB79, NSP76, RG75, RWLC78, Sch78b, SSD^{+}78, TKK79]. organizing [SEHW79]. organogenesis [BJ75]. orient [Zig77]. orientation [BKS78, BSPK78, MRG77, TSD75, WP75]. Oriented [FYS78, RH75]. origin [AV78, BG76a, MG75b, NLT76, Paa78b, RP77, SEI75]. original [BSM79, SMM78]. Oryzias [GJRR78]. Oscillations [RD76]. osmium [MSP78, RW78]. Osmotic [BTDT76, Cla76, MM78]. osmotically [BTDT76]. ossification [RA76]. Osteoclast [Mil77, Mil78]. other [AB77, CF76a, CF76b, DB75, Gra75a, Kan76, LN75, MBPTS79, MFKT78, MB76, NR76]. Ouabain [SS78b, EM77, Smi77, EKS^{+}79]. ouabain-sensitive [EM77, Smi77, EKS^{+}79]. outer [AC79, BBDP79, BBH76, BHR77, BR78, FSB75, HBR77, KAB75, MNK79, MS76, OY78, PSZK78a, PSZK78b, RG75, RHHB76, SHM^{+}77]. outer-segment [OY78]. ovalbumin [LSWB78]. ovalis [TPH77]. ovarian [AA77, BA79, CE77, Her76, LGGB79]. ovary [AA76, BO76, Cav75, GB77, RS75a, RW75b, SC78, WR77c]. overall [CBBS77]. oviduct [And77, AH77, SBU76, SBULB76]. ovulation [CE77, GEB78]. ox [SO75]. oxidase [BDKK75, RBK78, RTC76]. oxidation [EVMP76]. oxide [MW76]. oxidizing [OAB76]. Oxygen [BGRC78, BGRC77]. Oxygen-sensitive [BGRC78]. oxyntic [CCH78]. oxytocin [SC77].

P [FKNMT79, IP78a, MRDC77, MFKT78]. P-450 [FKNMT79, IP78a, MFKT78]. P-450-dependent [MRDC77]. P815Y [KSP75]. packing [EWR78, RH79, RB77, Til75]. palaestinae [OB78]. Palmivora [dSN77]. pancreas [BSVG75, BTB79, CM75, GPPR76, NMQY77, RPGR77, RP78, SPT78, SSG79, WL76]. pancreatic [CW78b, CW78c, CM75, GST79, KRJ77, LNS76, MPJ79a, MPJ79b, MPO79, MCP^{+}78, OPF77, SS78a, SG79b, Täl78, WL76]. paracellular [FMR79, LD76]. paracrystal [CP77a]. Paracrystalline [FGSM78]. paramecium [HA76, BLTP^{+}76, BS78a, Dip76, DK78b, FKP76, Han79, KH76, SSR76, TK77, TSD75]. parameters [Jau75, SFS77]. Paramyosin [ELD76, LEDW76]. paramyxovirus [DDR75]. parasite [AMJR78]. parasites [AMJR78]. parathormone [MC79b, MC79c]. Parathyroid [HARO79, Mil78]. parenchyma [BBW77]. parental [LWS79]. parietal [RCMI75]. parotid [CJP75, CP78, DPM76, FS77, OH78, PBJ77, SES76]. part [SS78c]. parthenogenetically [MBV77]. partial [BFD78, Gou75, MN77, WF76]. participation [LTMS79]. Particle [Ros78, AC79, Edd75a, FOPY77, KKG76, Sta76, TCT79, WGG77b, Wei77b, HH78]. particle-free [AC79, HH78]. particle/chain [KKG76]. particles [CP79a, GBE75, HH77b, LZK75, MPDGR76, Mau76, MV75b, NY78, Swe76, TM76b]. particular [BA79]. particulate [RKL^{+}76]. Passage [BBC78, GM75]. patches [AC79]. patching [Bou79, HB79]. patent [SSP75b]. pathogen [AV78]. pathway [FMR79]. pattern [Bel77, BP79, JF75, NJ75, SB75]. patterned [HB76b]. Patterns [Nan75, Rag77, RR77b, BB79a, CB76, Han75a, HP79, MPJ79b, NSHE75, RTBE78]. PC12 [CGVR79, Gre78]. pea [BVBM76]. pealei [MT78]. pellucidae [NYY75]. penetration [RW75a]. peptide [NBB77b]. peptides [GSB77, LE78a, SSP75b, ZS79a]. peptidohydrolase [BTC78]. perfused [GLR77]. perfusion [GTL77]. pericellular [HKA^{+}79]. pericentriolar [GB77, TR79a]. perilobular [WDM^{+}75]. perinuclear [BC76]. period [DR77a, HW75]. Periodic [CBSP76, KBR79, SHM^{+}77]. Peripheral [DS77, HJP78, GD76, Gou77, HH77b, SL77, Sch77, SF78, YB75]. peripherally [PL76]. peritoneal [AL75, HLG76, LRG^{+}77, MBPTS79, SS77b]. permanent [DTT75, YYP77]. Permeability [SSP75b, WW78, BTDT76, EF75, Har78a, RCV75, SR78b, WW75]. permeable [BO76, CRD^{+}78]. permissive [GGGM78]. permits [SE75]. Peroxidase [RSM76, AKD75a, CFAS78, JL79, OH78, PL77, PT79, SHH76, TDHO76, WW75, HSM76]. peroxidase-alpha-bungarotoxin [JL79]. peroxide [BKK75]. peroxidomes [SSSW77]. peroxisome [MR76b]. peroxisomes [LCK75, Svo78]. Persistence [WJB78]. persistent [KMS77]. perspective [Har78b]. perturbant [CDH^{+}79]. pH [BR79, BH78, Rod76]. pH-dependent [Rod76]. phagocytizing [BOP^{+}78]. Phagocytosis [MBBS78, BO78, HDS77, MW77, SH76, WF79, WGF79, ZW77]. phagolysosome [HDS77]. phagolysosomes [OT76, OT78]. phagosomes [BKK75]. phalloidin [PCJG79]. phase [FSSZ75, MB77b, RdC77, SE75, WGE75]. phase-contrast [SE75]. phaseoli [HH76]. phases [BB79a, Cle75, HMDH77]. phenobarbital [BN76, WDM^{+}75]. phenol [RHHB76]. phenol-soluble [RHHB76]. phenotype [CGK75]. phenylindole [Col79, JJ78]. pheochromocytoma [CGVR79, GET^{+}78, Gre78]. phloem [SH77b]. phorbol [GHW75]. phosphatase [Ern75, MCC79, VMR79]. phosphate [TR79b]. phosphatidyl [KP77c, KP77b]. phosphatidylcholine [CB78]. phosphatidylethanolamine [CB78]. phospholipid [BK75, CB79, HP75, ND77, PH75a, PT77]. phospholipids [MG75b, RTC76]. phosphoprotein [BAPG79]. phosphorylase [MCC79]. phosphorylated [PSGL78, UGB^{+}79]. phosphorylating [SAPA77]. phosphorylation [TMGC76]. photoreceptor [BHR77, FN76, HBR77, PSZK78b, RSM76, ZT75]. photoreceptors [SHH76]. photosynthetic [MMM76]. photosystem [ASA77, CMB75, MMM76]. Physarum [DR76, KNL76, KKH77, Mat77, RD76]. physical [LCK75]. Physicochemical [FMW75]. Physiological [KG78, PPMK77, SS78e, SS78f]. Phytophthora [dSN77]. pig [CM75, FOPY77, GB79, GL79a, LF76, LLLG78, MPJ79b, MPS79, MCP^{+}78, NSHE75, Paa78a, Paa78b, RP78, SPT78]. pigment [BP77, MD75, SEHW79, LN75]. pigmentation [JF75]. pigmented [Cra79]. pinched [FB78]. pinched-off [FB78]. pinealocytes [ZT75]. pinnate [TPH77]. pinocytosis [DR78, HMC79b, MN77, SBC76, WKBL75a, WKBL75b]. pinosomes [WY78]. pipiens [AG75, HS76b]. Piranhalysis [WY78]. Pitfalls [LJ75]. pits [MBPTS79]. pituitary [BAB79, Far78, FAE75, Hop77, HG77, OLK77, RHDL77, SO75, SLJ77]. placenta [NEK78b, NEK78a]. plane [Woo77]. plant [HG78, MRDC77, PFK^{+}76]. plants [AV78, JJ78, SH77b, AV78]. plasm [MIT76]. Plasma [BS77e, CM77c, TV79, AHH76, AC79, BBH76, BAPG79, BB79a, CF79, CPO77, CB77, CI78, CCNAN78, DDSL77, EM77, FGSM78, GGS^{+}76, HLR76, HB79, HC75a, HC75b, JKB^{+}79, KNM79, Kyt76a, Kyt76b, LNS76, Md78a, Md78b, NYY75, NUY^{+}77, RBHP75, RR76, STdDT79b, STdDT79a, TP76, WM77, WMP79, YK79, ZVR75, CKJB77]. plasmacytoma [GTPWD76]. plasmalemma [And76, DPM76, HH77a, SM78c, TRH76]. plasmalemma-ectoplasm [TRH76]. plasmalemmal [WNQK78]. plasminogen [Qui76, RP77, RR77b]. plasmodia [KNL76]. Platelet [RR76, DR78, KNM79, LF76, NP75]. platelet-derived [DR78]. platelets [AZW^{+}79, FBSS78, RMP75, RAFM75]. plates [OWC^{+}77]. plexiform [RG75]. plexus [McN78, vDK79]. PMNS [AL75]. pneumocytes [KP77b]. podophyllotoxin [KPF75, PM75]. podosomes [DS77]. points [HBD75]. polar [BS79c, CM77b, MIT76, SK79]. Polarity [SO77, BRR78, MT75, TK79]. polarization [Toy76]. Polarized [BS77c, CRD^{+}78, SE75]. polarized-light [SE75]. pole [HB78, Lu78]. poles [BBMB77]. pollen [JWJ75, Rag77, WNJ75]. poly [KS77]. polyacrylamide [AJG78, LMPI76]. polyanions [BC78]. polycationic [HM75, SD76]. polycephalum [DR76, KKH77, Mat77, RD76]. polyedra [Swe76]. polylysine [CKJB77, KNM79, MSS75, OLL76]. polylysine-coated [CKJB77]. polymer [WHC79]. polymerase [FFSH77, YGH78]. polymerases [MWS79]. polymerizability [MC78a]. Polymerization [TKST78, Til78, TK79, BR79, Kan76, MR76a, RGK78, SM78b, Til76a, Til76b, WLW78, WR75a, WRP75]. polymerized [Kan75, PPMK77, PPMSK77]. polymers [MH76a]. polymorphic [Bru76]. Polymorphism [GL77]. polymorphonuclear [BTS^{+}78, BB79a, BKK75, BDKK75, BOP^{+}78, FSB75, GHW75, HSGW76, HGW77, HW78, MBBS76, MBBS78, NSBS77a, NSBS77b, NSHE75, RBK78, WF79, WGF79, Zig77, Zig78]. polyoma [Bel77]. polypeptide [HPH77, KC78, MV75c, HMDH77]. polypeptides [BP76, CG77, FP78a, GBB78, HL75, HS76c, HC75b, MKK79, SH77a, Wil77a]. polyribosome [GWS77]. polysomal [KVB75]. polysomes [ALMS76, DAF77, KS77, WD77]. polyspermy [GSJH79]. Polysphondylium [Roo75]. polytene [KMB78]. Polytoma [SSC76a, SSC76b, SCS76]. Polytomella [BMP76]. polyvinylpyrrolidone [WKBL75a]. pools [CRBH78, MC79b, MC79c]. population [CB75a, EW77a, WPMW75]. populations [BBW75, HD77, PK79]. porcine [RGS^{+}76]. pore [DB76b, GBB78, KRM77, Mau76, Mau77, MD77a, SKT^{+}76, SC76]. pore-connecting [SKT^{+}76]. pores [Zaa79]. Porter [Pal77]. positioning [WCC79]. positively [BS79c]. possible [Gre75, HO75, JC79, LW76, LF76, MT78, PP79, PBJ77, PDD78, RW75a, RD78, DWB79, NSBS77a, WJTR77]. possibly [RGS^{+}76, SS76d]. post [DB76a]. postcapillary [SSP78]. posterior [ST76a, ST76b]. postjunctional [Ros78]. postmitotic [CSP79]. postnatal [OAR^{+}77]. postpartum [Paa78b]. postsynaptic [BCS77, CBBS77, CS78, FF79a, HS79a, KC78]. postulated [HB76b]. potassium [BPS78, CGH79b, Ern75, GHC78, KD78, LL79, NSBS77b]. potassium-dependent [Ern75]. potassium-independent [Ern75]. potassium-stimulated [BPS78]. potency [AP75]. potential [EW77b, GG77, NPM75]. potentials [MSRT78]. Poteriochromonas [HKS77]. preceding [ES79]. precursor [BC79, JK77, SDTD^{+}79]. precursors [HARO79, NBB77b]. predominantly [HVW78, WGC76]. preferential [PJ76a]. preferentially [MSK^{+}79]. prefusion [FF79b]. preimplantation [DUKA77]. preliminary [Ste75]. preovulatory [AKLL75, RB76]. Preparation [Kan75, LNS76, MV75a]. preparations [BP77, LWB78, SS79d, WHC79]. prepared [WRP75]. preparing [KOBP75, SAB75]. prepuberal [BCM^{+}77]. preputial [SNP77]. preribosomes [Gra75a]. Presence [CF76a, CF76b, FSM76, IP78a, AC79, CLP76, DM79, ASD75a, BD75a]. preservation [KP77b]. Pressure [Sal75a, Sal75b, SGMB76, Cla76, SE75]. Pressure-induced [Sal75a, Sal75b, SGMB76]. presynaptic [HK79]. pretreated [JSG79]. prevent [Mau76]. prevention [SS78d]. prevents [Gre78, OKB76]. Primary [SFSB78, BWDP77, BFD78, BCHN75, CW77, LW76, MKKV78, NSHE75, PPHM75, WDMR77]. primate [WR75b, WR75c]. primates [FI77]. principal [CG77]. principle [MCR77]. prior [WGE75]. prismatic [MCR77]. proalbumin [RBM^{+}78]. probable [KP77c]. probasal [Gou75]. probe [CW78b, CW78c, JJ78, MC79a, SS79a, SSS77, SSS79, Täl78, WC75]. probes [KF79, SNP77]. probing [HM75]. problem [LW76]. procaine [MBV77]. procedure [DB76b, NBB77a, SS76c]. procedures [Kra77, Per77, TS76]. process [NMQY77, OLL76, TK79]. processed [BD75a, GCRB78]. processes [FJD76, PCJG79]. processing [Her76, JLA^{+}76, JK77, STdDT79b]. procollagen [NVH^{+}75]. produced [Cle75]. producing [BRS^{+}76]. Production [RP77, CP77b, CP77d, LE78a, RR77b, Zaa79, BCB77, MG75a]. products [PP79]. profile [LHH76]. progeny [PM79]. progesterone [Paa78a]. proglucagon [NBB77a, NBB77b]. program [GPPR76, HPH77, LH79]. programming [WC77]. progress [WR77a]. proinsulin [NBB77a, NBB77b]. prolactin [BAB79, GZM75]. proliferate [Poh76, RR76]. proliferated [LCK75]. Proliferation [BN76, GBBZ78, HWLC78, HS75b, MKKV78, MGPO76, SS79c, SVW79]. proliferative [SSSW77]. proliferator [MR76b]. promote [SM78b]. promoting [Moo76a]. promyelocytic [VMR79]. pronase [SO75]. proparathormone [CMC77]. Properties [BG76a, HS75a, OAB77, PAW76, BE79a, BO78, BGFK79, DTT75, FMW75, Gor75, JB78, LCK75, MV75a, MHC77, MIMMF77, OAB76, RCSS79, Tak77, Toy76, UKY^{+}79, VJGG79, FPC76, McG76]. propidium [FPC76, Kri75]. protease [Dip76, Qui76, RGS^{+}76]. Protein [BDP^{+}77, Bon75a, Bon75b, NSHE75, WJ76, BR76a, BPRC78, CBDY79, CMB75, CBBS77, CKCK78, CM77c, DS75, EL75, FCG^{+}76, GTL77, GCM77, HARO79, Han75a, HS76a, HDS77, HMK79, HVW78, HO75, IW76, KOBP75, LGB77, LNMR78, LH79, LLLG78, LP77, MH77, MW78, MMM76, NRH78, NEK78b, PSGL78, PSZK78a, PSM^{+}75, PM75, PR76a, RBHP75, RBM^{+}78, RGS^{+}76, RG78, RLRV79, SHWH^{+}77, SST79, SES76, SWV77, SHM^{+}77, SH76, TBZW78, TSC78a, TSC78b, UGB^{+}79, Wei76a, WHC79, WKBL75b, ZM76, HWP78, MC79b]. protein-cholesterol [DS75]. proteinase [PS79]. Proteins [KUS78, KFP^{+}78, SM78b, AEP79, Amo77, BBC78, BD75a, BD75b, BCS77, Bon75a, Bon75b, BW78, BW79, BH78, CB79, CS79, COK76, CH76, DVEM79, DGR75, GSB77, GL79a, GCRB78, Gre75, GP78, HS77, HJP78, HH77b, HB76b, HC75a, HC75b, JZR76, JKZM77, JKM79, Kan76, KC78, KRJ77, LLC79, LGB77, LKKA79, LMPI76, Ley75, LS75, MBPTS79, MH76b, MZ78, OB78, PT77, PDD78, Rou78, RHHB76, SPT78, SF78, ST77b, STWB78, SH79b, SSG79, SA77, SNP77, Tak77, TVD78, Til76b, TBZW78, UGB^{+}79, War79, WM77, WF79, WGF79, WGB78]. Proteoglycans [WR75b, WR75c]. proteolytically [BD75a]. proteus [JJ75, MG75b, RH75]. protofilaments [BHP75, PBH78]. protozoa [TC75]. protrusions [ABL76]. proximal [DK78b, HGP^{+}76, HD77, Kyt76b]. psoas [Toy76]. PtK2 [BBMB77, OWW78]. pubescens [AJ78]. pulmonary [HS76a, SH76]. pulse [NBB77b, NJ75]. pulse- [NBB77b]. pump [DJM77, MD77b, MED77]. pumps [WBY79]. punctulata [HMK79, SEBL77]. pupfish [KEP76]. pure [MIMMF77]. Purification [FCG^{+}76, CG78, Kan75, ASD75b, UKY^{+}79]. purified [Amo77, DJM77, HP79, RTC76, SS79b, Smi77, SBGG78, WM77]. Purkinje [ET79]. puromycin [LF77]. purple [FYS78]. Putrescine [Poh76]. pyrene [WGC76]. pyriformis [HB76a, FJD76, OLL76]. Pyrsonympha [LI79].

quadriflagellate [BMP76]. quail [Mil77, SBU76, SBULB76]. quantal [HRD^{+}79]. Quantitation [CE77, FS76, YYP77, OT76, Sne76b, PAW76]. Quantitative [BHFT79, BPRC78, EP79b, GGS^{+}76, GGGM78, MWS79, MD77a, RCM77, SHWPJ78, Toy76, WKBL75a, WKBL75b, ABL76, Bry76, GL79c, MR75a, RR77a, NCM78, OT78]. quick [CH79a, HS79a, HRD^{+}79]. quick-freezing [CH79a]. quick-frozen [HS79a]. quiescent [RR76, Smi77].

rabbit [AL75, BTS^{+}78, CJP75, CP78, CK78, CE77, EZF76, FSB75, FMR79, HS76a, HD77, Koe75, KP75, LS75, Lon76, MM79, NSBS77a, NSBS77b, NSBS79, NUY^{+}77, RMP75, RMP^{+}78, SS78b, SH77a, SH76, Toy76, Wil77a, WWLS79, WF79, WGF79, ZS79b]. rabbits [RG75]. radial [WPS78]. Radiation [SC76]. Radiation-induced [SC76]. radioactive [NBB77b, SOK79, Sne76b]. radioautographic [BSHP79]. radioiodine [PL77]. Rana [AG75, HS76b, SR78b]. random [DM79]. Ranvier [YB75]. Rapid [BDKS79, CMC79, CGVR79, HMC79b, Kri75, Mil78, LB78, SEHW79]. rapidly [ST77a, SBGG78]. rarae [KF79]. Rat [MHP78a, MHP78b, AHH76, AKLL75, ASD75a, ASD75b, BC77, BBC78, BSHP79, BWDP77, BAB79, Blo79, BBW77, BKS78, BSPK78, BBMB77, CLP76, CB75c, CW77, CF76a, CF76b, CP77b, CP77c, CP77d, CM77c, DWM75, DB76b, ESAD75, Ern75, EB78, ES79, FAE75, FRT^{+}76, FKNMT79, GSB77, GZM75, GBE75, GET^{+}78, GLM77, HMPB76, HO77, HR79a, HSM76, HWAS79, HS79b, IP78a, JZR76, JSGR78, JKB^{+}79, JKZM77, JKM79, KSS76, KR75, KCH79, KG77, KS77, LC76, LGGB79, LRG^{+}77, LW76, LCK75, LNS76, MFKT78, MPJ79b, MFPO75, MGPO76, MR76b, MN77, MMS^{+}76, NRH78, NY78, OH78, PL77, PBJ77, Pfe78, PSC^{+}76, PRZE75, PCJG79, PSG^{+}76, QWTI79, RPGR77, RCM77, RBHP75, RFB^{+}76, Rem78, RKL^{+}76, Rod76, RW79, RHDL77, SS78a, SL77, Sch77, SM78a, SF78, SMJ78]. rat [SC77, ST78b, SES76, SM75a, ST77a, ST77b, SB76, SF76, SP76, Svo78, SM76, SNP77, TM76a, WJB78, WDMR77, WCRdD78, WGC76, WKBL75a, WKBL75b, Wil77b, ZT75, dMdS78, vOMW77]. rate [BS79b, MKK79, RR78]. rates [BE78, DH76, EW77a, EW77b]. ratio [CBSP76]. rats [AKD75b, BCFS75, PS75, SSSW77]. ray [CSHS77, CSP79, CGMP77, KHT^{+}79, MCPG77, MBM^{+}79, PD75]. RC [BDCD75]. rDNA [Cav76]. reacting [RW78]. reaction [AZW^{+}79, BGG^{+}75, FOPY77, PAW76, SS79a, Til75, TKST78, TCT79]. reactions [KNL76]. reactivated [GOG76, PW78]. reactive [KH76]. reagent [GES79]. reaggregates [MPS79]. reaggregation [TS77]. reality [WP79]. reappearance [YR78]. rearrangement [BR76b, HBS79]. Rearrangements [LLMW77, AA77]. reassembly [MCP^{+}78]. reassessment [SFS77]. receptacle [MH76b]. Receptor [RS75c, Blo79, BPRC78, Bou79, CDH^{+}79, DF75, FC77, FF79a, HS75a, SEBL77, SF76, SRR79, TSS79, VD76]. receptors [AA77, BSM79, CF79, CO76, FD78, FS76, HS79a, HG77, JL79, Kah76, KHV^{+}78, KCGN76, LC77, Let79, LAF^{+}77, MJM76, NUY^{+}77, NSH78, PPL78, RV78, WM77, dP75, dP78]. recognition [HWAS79, HS79b]. reconstituted [RTC76]. Reconstitution [HFK77, UKY^{+}79, BD75b]. reconstruction [Mau77, OWS75]. recordings [MPS79]. recoveries [BPL^{+}78]. Recovery [Far78, FWS75, ST77a]. rectal [EKS^{+}79]. rectifying [HKP78]. recycling [FB78, SR78a, STdDT79a]. red [SF76]. Redistribution [BFTH79, PR76b, SD76, BFPU78a, BFPU78b, ES79, Let79, SNP77, Yam78, SPT78]. Redox [JBKF77, SH79b]. Reduced [CW76]. reductase [DBB^{+}79, IP78a, MMS^{+}76]. reduction [BBMB77]. reevaluation [HGW77, HS79c]. reference [BA79, CAS78, CFAS78]. reflection [BHFT79]. Reflexive [Her76]. reformation [Blo79]. regenerating [BSM79, SMM78, YR78]. regeneration [FLK^{+}76, LNMR78, LW76]. region [BR77, BRBM77, BC76, DBB^{+}79, DK78b, LH76, WF76]. Regional [GES79]. regions [MS76, WMP79]. regression [Paa78b]. Regular [JZR76, RKL^{+}76]. regulated [BAPG79]. regulates [BR79]. Regulation [BS78a, Han75b, Hit77, MBPTS79, MC79b, RH77, STF76, VMR79, Dec76, HS76c, IW76, LD76, Mat77, PPMK77, Sch76, HB76a]. reinhardi [GW78]. reinhardtii [BGG^{+}75, BP76, CMB75, CBSP76, CG77, DH76, GS75, GW75, Gou75, HS76c, HBD75, HPH77, HRL77, ICS75, MG75a, MG78a, SDTD^{+}79, WC77]. Reinitiation [NPM75, BDCD75]. Reinnervation [SMM78]. related [BG76a, KNL76, KUS78, KFP^{+}78, LE78b, TDHO76, Wei76b, Wei77a]. Relation [RG78, DAF77, JLA^{+}76, LLL75, LEDW76, MH77, PSM^{+}75, RR77a, SPM79, SSD^{+}78]. relations [FJD76]. Relationship [CCB76, HO77, TMGC76, ASA77, BHR77, BS77e, DR77b, FDB77, HBR77, MG76, BFPU78a, RTBE78]. Relationships [LP77]. Relative [LWB78]. relaxed [Toy76]. relay [DDS79, DS79]. releasable [RMP75]. Release [KS77, LLP77, SO75, AZW^{+}79, CGH79a, CGH79b, CW78b, CW78c, FBF79, FS77, GHMD77, GHW75, HSM76, HRD^{+}79, NSBS77a, RBHP75, WL76, ESAD75]. released [Moo76a]. releasing [HG77]. relevant [WDB79]. Remodeling [GWN77]. removal [CP78, DST^{+}78, RW75a, SMM78]. renal [Hil75, Kyt76a, Kyt76b, MKM76, OAR^{+}77, SS78b, Svo78, WOPB79]. Renilla [SC75b]. renin [HD77]. renin-active [HD77]. reorganization [Edd77]. repair [CW76, CW77, KMS77, YB75]. replicas [vDK79]. replicated [HS79a]. Replication [HG79, MW76, BO76, BS78b, CM77a, Han75a, Han75b, JT79, KBS79, MEB77, SSD^{+}78]. reported [DB75]. required [BB78, HQ78]. Requirement [KCH79, MGF77]. Requirements [WLR79, DS79]. resembling [BFT77, Wil77a]. resistance [DF77, PD76, SHWPJ78, WBE75]. resistances [MSRT78]. resistant [DCEO^{+}79, NSH78, WF75]. Resolution [SFS77, SMS78, AHB79, IF79, KRM77, MBM^{+}79, PRGH78, ST78a]. resorption [LNMR78]. respects [Wil77a]. respiratory [AP75]. respond [vVRN76]. Response [CH75, Ber75, CGVR79, DS79, GBBZ78, GWN77, HB76a, MM78, NSBS79, SGCS75, SSSW77, SS78e, TM79, DST^{+}78, KEP76]. responses [DDS79, Svo78]. restacking [Sta76]. resting [JLA^{+}76, OH78, OB78]. Restriction [HB79, FWB78, LE76]. resulting [DM79]. resynthesized [BDCD75]. Retention [ALMS76, FS77, RCSS79]. reticulocyte [SH77a, ZS79b]. reticulocytes [LS75, TSS79, ZS79b]. reticulum [ALMS76, BN76, BFT77, CLP76, Dam76, DMS^{+}76, FDB77, HLR76, HIP78, JBKF77, JKZM77, JKM79, LH76, LF77, LWB78, MCC79, OKS77, PFK^{+}76, PD75, SWF78, ST77a, ST77b, WLR79]. retina [Bur78, JM79, JM77, KMS77, LN75, MM79, MB75a, MGF77, MG78b, NLT76, RG75, RSM76, SL78, SR78a, WB78a]. retinal [AC79, Cra79, MD75, SHH76, WWLS79]. retinas [PSZK78b]. retinotectal [Mar77]. retraction [ABL76]. Retrieval [FB78]. retrograde [ST78b, SST79]. revealed [RW78, SWF78, SOW78]. reveals [SM78c]. reversal [SS78d]. Reversibility [BR76b]. Reversible [BLB79, BGK^{+}79, DTT75, PE79, Sta76, RB77]. Reversibly [WH77]. rhizoid [HQ78]. rho [BPS78]. Rhodopsin [BBH76, DVEM79, MB75a]. Rhythmic [OY78]. Ribonucleoprotein [PB75, Rie79, MHP78a]. ribonucleoproteins [STF75]. ribophorins [KUS78]. riboside [THC76]. Ribosomal [Hil75, Cav75, HS77, LLC79, LE76, LE77, MV75a, MV75b, MKM76, MG76, PH75c, STF75, STF76, War79, WC77, SCS76]. Ribosome [HB76a, JK77, KUS78, KFP^{+}78, KSB75, LLC79]. ribosomes [BD75a, CBSP76, FKNMT79, HLS75, JK77, KAB75, KFP^{+}78, KVB75, MV75a, MV75b, MV75c, OKS77, TK77, VRN76]. ribulose [ICS75, SDTD^{+}79]. ribulose-1 [SDTD^{+}79]. rich [TMGC76, WF76]. ricin [FSB75]. ricin-binding [FSB75]. ridge [And76]. ridge-like [And76]. rigor [PW78, Toy76]. ring [BG76c, DAF77, WD77]. ringlike [SM78c]. Rings [BC76]. RMI [Svo78]. rna [STF76, BP79, BDCD75, CLP76, DAF77, EL76, FFSH77, GWK77, Gra75a, GCM77, Gre75, HWP78, Hil75, JLA^{+}76, KS77, LAS^{+}75, LE78b, Leó76, MWS79, MV75c, MKM76, MG76, MHP78b, OPD76, RB76, RS75a, STF75, ST77b, SCS76, SCRP77, THC76, WD77, WC77, ZM76]. RNA-protein [HWP78]. Roberts [Pal77]. rod [AC79, BBH76, BHR77, HBR77, PSZK78a]. rodents [LN75]. Role [CP77b, CP77c, CP77d, HGW77, KSP75, Lem79, LD76, ODS79, RCV75, SM77, Slu79, VL78, BCB77, BRBM77, BS77a, Cra79, DSTW77, DWB79, HB76b, IO79, JF75, KP77c, MV75c, MT78, MR75b, PDD78, Pol76, Sch76, SS77a, SC75a, TSD75, WJTR77, Yam78, AL75, BG76b, PT77, RGE78, SH76, TKST78]. Roles [IW76]. root [VKZB78]. rootlets [Ste75]. rosette [NCCB77]. rosette-forming [NCCB77]. roseus [MRDC77]. Rotary [MEB77, HS79a, Tam78]. rotary-replicated [HS79a]. rotational [Tam79]. rough [BD75b, BKS78, BSPK78, FKNMT79, JBKF77, KUS78, KFP^{+}78, KS77, OKS77, SH77a, ST77a, ST77b]. round [NRH78]. Rous [DN76, HWW75]. rows [Nan75]. rRNA [JLA^{+}76]. ruffling [SC75a]. rust [HH76].

S [Cle75, WGE75]. S3 [CCNAN78, HMDH77]. sac [WKBL75a, WKBL75b]. saccharides [MPJ79a, MPJ79b]. Saccharomyces [Dam76, HU77, HB78, RH77]. sacs [JB78]. SaH [MR76b]. salamander [Lem79, Leó76]. saline [HGP^{+}76]. salinity [KEP76, KKKS76]. salivary [AJ78, BCB77, CBB77, DAF77, KG78, Pai75]. salt [EM77, San78, SPM79, TKK79]. salt-dependent [TKK79]. same [Her76, OWW78]. sandwich [NCPW75]. sap [CH76]. sarcolemmas [ERSP76]. sarcoma [DN76, HWW75, KP77a]. sarcomere [JC79]. sarcoplasmic [BFT77, FDB77, JKZM77, JKM79, PD75, SWF78]. saturated [KP77b]. Scalar [NCM78]. Scanning [ML76, VP75, GR78, IF79, KW76, KRM77, LAF^{+}77, MDRY75, MJM76, MBBS78, NCPW75, SAB75, ST78a, MM78]. scattering [WAC78]. Schwann [BB78, RV78]. Sciaridae [AJ78]. sciatic [SM78a]. sea [AP75, BR79, BS77b, Bro79, BS77c, DL79, DDSL77, ES76, GOG76, GV77, GL79c, Gol79, HMK79, Kan75, Kan76, KBS79, MNK79, MBV77, NR76, OAB77, PW78, PAW76, Rik78a, SG79a, SEBL77, Slu76, SS79d, WT79]. secondary [BH77, NSHE75]. secretagogues [SS78a]. Secretion [CJP75, CP78, MC79c, MOS^{+}77, SES76, TDHO76, BSVG75, BES75, BS77d, CW78b, CW78c, CH79a, CM75, DH75, EP79a, FLK^{+}76, GTPWD76, HV75, HGW77, HW78, Hop77, LFR78, LS79b, MPO79, MC79b, NS77, Paa78a, PBJ77, RR78, SS78a, TWR^{+}77, TDO78, WGG77b, WR75c, dSN77, dMdS78]. Secretion-related [TDHO76]. Secretory [BK78, BS77e, CP78, EM77, FS77, GLR77, HO77, HSM76, KRJ77, LE76, LE77, MC79b, NMQY77, OB78, OLK77, OLK78, PCJG79, SLJ77, TVD78]. section [CS78, McN78, MBBS78, OKS77, WP78]. sectional [MEM79]. sectioning [DJM77]. sections [BRR78, PL77, Pau75, SSS77, TS76]. sedimentation [RCMI75]. sedimenting [ST77a]. seedlings [BTC78]. seen [CLM77]. segment [BBH76, Kyt76a, Kyt76b, OY78, PSZK78a]. Segmental [SSP75a, SSP76]. segments [AC79, BHR77, HBR77, PSZK78b]. segregated [MG76]. segregation [GGS^{+}76, AEP79, HJP78]. selected [Pau75]. Selection [WF75, FP78a, Ley75]. Selective [OKS^{+}79, ST78b, SST79, BR77, MH76b, PJ76b, CP78]. selectivity [GM76, MB76, McG76]. semitendinosus [ZVR75]. Semliki [HFK77]. Sendai [BB79b, UKY^{+}79, VL78]. senescent [NPM75]. sensilla [MRZV76]. sensitive [BGRC78, EM77, FJD76, HRL77, KP77a, MGF77, PPMSK77, PE79, SSD^{+}78, Smi77, VV76, YYP77, Yam78, EKS^{+}79, WF76]. sensitivity [DTSM77, FWB78]. Sensory [ZS79a, FC77, MRZV76]. separate [MC79b]. separation [Kra77, LF76, PK79, RCMI75, WNQK78]. sepharose [LFR78]. septate [LD76]. sequence [LSWB78]. sequences [Gre75, SDTD^{+}79, WD77]. sequential [CGVR79]. sequestering [BFT77]. sera [CK78]. serial [CS78, Pau75]. series [vVRN76]. serotonin [GKS76, RAFM75, SGV75]. serous [MR75a]. Sertoli [Con78]. serum [BBC78, GOG76, Gre78, HR79a, HS75b, MBPTS79, RBM^{+}78, RR76]. serum-destined [BBC78]. serum-free [Gre78]. sets [GS75]. several [YYP77]. sex [CH79a]. Sexual [TM79, FGG78]. SH [RW78]. SH-groups [RW78]. shape [AZW^{+}79, BS77a, Cra79, FALP75, LGGB79, PCJG79, SPS76, SS77a, TM76c, VK77, WP75, Yam78]. shapes [SS76b]. sheath [YB75]. shedding [OY78]. sheep [BAN75, BPN76, LLMW77]. shells [KKG76]. shibire [PE79]. shift [MW76]. Short [Hop77]. shortening [DR77b, LNMR78]. shrew [CB76]. shuttle [STdDT79a]. sialic [DMB78, DM79]. sialidase [SRR76]. sialyltransferase [SFFB75]. side [CB78, CM77b]. side-polar [CM77b]. signal [LSWB78]. signaling [FGG78]. significance [HS76d, LW76]. sikes [UKY^{+}79]. silica [PRGH78]. silk [ST76a, ST76b]. silkmoth [TBZW78]. silkmoths [BK78]. silkworm [ST76a, ST76b]. simian [GGGM78]. similar [YK79]. Similarity [HLR76, CH76]. simple [CBC75, PJ76b]. simplex [SRR76]. Simultaneous [FP78b, SE75]. sindbis [BCB79, WLR79]. single [BBDW76, DVEM79, GHMD77, MST78]. sinusoidal [DMB78, DM79]. Site [Wei77b, BLR76, CRBH78, PB79, SM75b]. Site-specific [Wei77b, PB79]. Sites [NY78, AMS76, AKLL75, AH77, BKS78, BSM79, CG77, CP79a, EM77, FSB75, FKP76, GTH75, GKBM78, HM75, KF79, Koe75, KFP^{+}78, MB75b, MD77b, MED77, NYY75, NUY^{+}77, SMM78, SH79a, SD76, SF76, WGG77a, WF76, Zaa79]. situ [BH75a, DTSM76, DTSM77, IF79, PH75b, PD75, TD75]. Size [FALP75, KBS79, DWM75]. sized [DAF77, FSOW79, Moe79]. skate [RSM76]. skeletal [AFB79, BFTH79, GL77, GL79b, JKZM77, JKM79, OSM79, RGS^{+}76, SS79b, Som79, WTE77, ZVR75, dCdM78]. skeletin [ET79]. skeleton [MHP78a, MHP78b]. skin [MD77b, MED77]. slices [SS78a, SO75]. slics [SES76]. sliding [LG75b, SS76a, WM78, ZMW79]. slime [Roo75, SB78]. slit [KR75, SM75a]. slow [ERSP76, HL75]. Slowly [MKK79]. Small [GWK77, Bon75a, BS78b, KCH79, LZK75, MHP78b, QWTI79, RR77a, SDTD^{+}79, SSP75b, SFFB75, WC75]. smooth [Coo76, CM77b, DR78, FSW77, FDB77, FKNMT79, GB79, HJP78, HL79, MCC79, MPS79, NSRL76, PD75, RR76, SSS79, WR75c]. snPI [BP79]. Sodium [Han79, AJG78, DJM77, EM77, LMPI76, NSBS77b]. soft [MBM^{+}79]. sol [IO79]. solation [CT77, HT79b]. solation-contraction [HT79b]. soluble [ASD75b, BK75, JZR76, RHHB76]. solutions [CGH79a]. somatic [DK78b]. some [BCS77, Gor75, Tak77, UKY^{+}79, Wil77a]. Sonchus [SH77b]. Sorting [GDC76]. source [TCT79]. sources [DTT75]. soybean [DMKR75, VKZB78]. spanning [Som79]. spasmoneme [Rou78]. Spatial [ASA77, BKS78, BSPK78]. special [CAS78, CFAS78, GST79]. specialization [MKKV78]. specializations [ERSP76, HHP76]. specialized [RG75, WGG77a]. Species [GL79c, BP79, BDCD75, KNL76, MHP78b, NCM78, SB78, GM76]. Species-specific [GL79c, SB78]. Specific [DN76, GL79a, MG78b, AEP79, AKD75a, FSB75, GL79c, GKBM78, HO75, HSGW76, KFP^{+}78, Laz75, LE78a, MEM79, MHC77, PB79, PSG^{+}76, Qui76, RMP^{+}78, Sne76a, Sne76b, SB78, Wei77b, WPS78, LW76, JJ78]. specifically [SH79a]. specificities [PK79]. specificity [EG77, Mar77, WJ76]. Spectrin [HBS79, LLP77, SS77a, TD75, TSS79]. spectrin-free [LLP77]. Spectrum [BDCD75]. Sperm [DL79, Lon76, AP75, BDP^{+}77, BS77b, Bro79, CB76, DB76a, EL75, GES79, GOG76, GL79c, Gol79, LG75b, MNK79, OAB77, PW78, Rik78a, RLRV79, SG79a, SS79a, SEBL77, Sch75, Til75, Til76a, TKST78, Til78, TCT79, Woo77]. spermatid [KT75]. spermatids [NRH78]. Spermatogenesis [MZ79, MRB75, MB77a, SP76]. Spermatogenic [BCM^{+}77]. spermatozoa [AEJL75, FOPY77, FFSH77, Koe75, NUY^{+}77]. spermatozoon [BD78]. spermiation [CB76]. spheres [MDRY75, PRGH78]. spherical [PJ76b]. spider [CGH79a, FCG^{+}76, GHC78]. spinal [BF78, FBF79, GNB^{+}77, PSG^{+}76, RGE78, SF78, THCP75]. Spindle [Moe76, BE79a, BE79b, BBMB77, CLM77, FP76, HB78, Lu78, MPHMT77, MEM79, MMER79, RB77, Sal75a, Sal75b, SGMB76, SEI75, SS78c, Slu76, Slu79, TSPH78, RIK78b, BE79a]. spindles [Rie79, WR75a]. spines [HHP76]. Splayed [SS76a]. spleen [CAS78, CFAS78]. splitting [JC79]. spoke [SS76a]. spokes [WPS78]. spontaneous [NCCB77, WHC79]. spontaneously [MPS79]. spore [Wei76a]. sporulation [LLL75]. spray [LB78]. spray-freeze-etching [LB78]. spreading [AB76, ABL76, Che79, DB75, HDS77, MS75b]. Squalus [EKS^{+}79]. squid [GTL77, KKAL79, LGB77, LKKA79, MT78, PSGL78, RV78, VV76]. stability [GZM75, HO75, SGMB76]. stage [IR78, PRZE75, ST76a, SO77]. stage-dependent [PRZE75]. stages [BGRC78, Con79, MRZV76, RLRV79]. stained [BHP75]. staining [BH78, CK78, DJM77, DBB^{+}79, FP78b, HP79, Kri75, Rie79, SFS77]. starfish [MO77, Sch75]. Starting [Gol79]. starvation [DR77a, LE77, WR77b]. starved [HS77]. state [Toy76]. states [JSGR78, KHT^{+}79]. step [LF76]. Stereo [OO79, BP77]. stereologic [SBC76]. Stereological [SMJ78, BBW77, BPL^{+}78, JSGR78, LWB78, SS79b, SM75a, WP78]. steroid [GLR77]. steroid-containing [GLR77]. steroidogenesis [CW78d]. steroids [NM78]. sterol [EGFB78]. sterol-digitonin [EGFB78]. Sticholonche [CCTT77]. stickiness [EB75]. stimulate [RR76]. stimulated [BH79, BBW75, BPS78, CW78d, DPM76, DMMW76, OH78, PJ76a, RMP^{+}78, WJTR77, vOMW77, HMPB76, KCH79, PELG76]. stimulates [Gre78]. stimulating [JB78, Smi77]. Stimulation [CI78, MH76a, MH75, CGH79b, GHC78, HMC79b, HSM76, MPO79, Smi77, WGE75]. stimuli [DDS79]. stimulus [CW78b, CW78c, LFR78]. stimulus-secretion [CW78b, CW78c, LFR78]. stipitata [HKS77]. stomach [KG77]. stomatitis [GWS77, KL79]. stop [SC75a]. storage [FBSS78, MH76b, RAFM75, Til76b, TBZW78]. strain [MHC77, SE75]. Strain-free [SE75]. strain-specific [MHC77]. strains [BLTP^{+}76, DF77, GW75, RS79]. Stratification [MKKV78]. strength [CBM78]. streptolysin [DS75]. striated [DWC^{+}77, EZF76, GW78, SSS77, Ste75, Toy76]. strives [KT77a]. stroma [BH77]. Structural [DDR75, EWR77, KT75, KHT^{+}79, MRG77, VV76, WM78, BPS78, HS77, HL75, HB76b, JK77, LB78, Laz76, MCR77, SM78a, SEHW79, TM76a, VJGG79, PD76, Per77]. Structure [BSOS76, DFW^{+}78, HG78, LCK75, SSD^{+}78, BC77, BCS77, Bru76, BP77, CGH79a, Cla76, CBBS77, GW75, GW78, GJ78, HKP78, HS79c, KEP76, KKH77, KVB75, LEDW76, Lon76, MM77, MG75a, MWS79, MEM79, MMER79, MBBS78, NP75, OKS77, RIK78b, SMJ78, SMPGR75, SC75b, WB78b, WGB77, WH77, TRH76, WPS78]. structures [BJ75, CGMP77, DR77b, FKS76, GESG75, JZR76, JSG79, MCPG77, MFKT78, RKL^{+}76, SM78c, Som79, dP75]. studied [CH79a, EL76, GES79, Hop77]. Studies [AL75, CKBD76, DWB79, ERSP76, GLR75, GBE75, GWS77, Koe75, KR79b, KKH77, LMPI76, LN75, NBB77a, NBB77b, NMQY77, ST76a, ST76b, WDB79, BH79, BE79a, BE79b, BPL^{+}78, BPRC78, BH78, CGH79a, CGH79b, CP77c, Col79, EZF76, FSW77, FF79b, FDB77, GTL77, GGS^{+}76, HH77a, HH78, JRFS75, JM77, KPF75, LWB78, MZ79, MV75b, MW78, MMS^{+}76, Paa78a, Paa78b, PELG76, RCV75, RdC77, SHC^{+}79, SL78, SPT78, SL77, Sch77, Sch76, SS78e, SS78f, Swe76, TVD78, Toy76, VV77, WP78, WKBL75a, WKBL75b, JK77, PH75a]. Study [BDCD75, SG79a, AP75, BLTP^{+}76, BSHP79, BRBM77, BBW77, BW79, BS77d, BP77, CSP79, CB75c, CS78, Con78, DPM76, EG77, EH77, FAE75, FRT^{+}76, GEB78, GWK77, GWN77, HHP76, HWAS79, HS79b, KKG76, KW76, LB78, LRG^{+}77, LCK75, Lon77, Lu78, LAF^{+}77, MPDGR76, McN78, MR75a, MM78, MCM77, NCCB77, NEK78b, NEK78a, OSM79, OKS77, OAR^{+}77, OBKP75, Pau75, PS75, RG75, RR77a, RFB^{+}76, RS79, SBU76, SHH76, STF76, SMG77, ST78b, SM75a, STWB78, SG79b, Sne76a, Sne76b, SC75b, Tam78, Tam79, TDO78, vDK79]. SU-13 [SSSW77]. Su-13437 [LCK75]. Subaxolemmal [MT78]. Subcellular [AHH76, BVBM76, CBDY79, FSW77, MRDC77, RT76, UGB^{+}79, BPL^{+}78, FRT^{+}76, Gor75, NBB77a, NBB77b, ODS79, RFB^{+}76, Rem78, vVSL^{+}76, CW78c, CM75, RGS^{+}76]. subchromosomal [Han75b]. subclones [vVRN76]. subcultures [CW77, FSW77]. subfractionation [CJP75, DWM75]. subfractions [CJP75]. subfragment [SMG77]. sublobular [SMJ78]. submandibular [TP76]. submembranous [CG79, HB76b]. Submitochondrial [VC77]. submucosal [MR75a]. subsequent [FKNMT79, SST79]. substance [BP77, WP79]. substrata [MH76a]. substrate [AB76, CRBH78, WDB79]. substrate-exploring [AB76]. substrates [Hop77, WPMW75]. substratum [GTH76, Moo76a, WP75]. Substructure [KR75, BH78, KKAL79, OCO75]. subunit [BBDW76, LKKA79, LE76, SDTD^{+}79]. subunits [BFTH79, ICS75, LE77, MV75b, NS75]. Successful [LF76]. successive [CW77, GTPWD76]. succinate [DMS^{+}76]. suckling [HW75, JZR76, RKL^{+}76]. sucrase [NT78]. sucrase-isomaltase [NT78]. suitable [NCPW75]. sulfate [AJG78, LMPI76, RD78]. sulfate-polyacrylamide [AJG78]. Sulfated [RP78]. Sulfation [HQ78]. Sulfhydryls [MR76a]. sulfoxide [Fuk78]. Suncus [CB76]. superior [DK78a]. superstructures [TKK79]. support [CRD^{+}78, Pro75]. suppression [BDCD75]. Supramolecular [Bru76]. surf [WR75a]. Surface [BRV78, CT76, MH77, AMS76, ABL76, BO78, Blo77, BLB79, BFPU78a, BFPU78b, BSU79, BDKK75, BR76b, BS79c, BB79b, BW79, BS77e, CGVR79, CB76, DWB79, DMB78, DM79, DR77a, DL79, DUKA77, EW77a, EW77b, EB75, Far78, FKS76, FLK^{+}76, FGSM78, FF79b, GES79, GHA76, GR78, GMDT75, GGS^{+}76, HWW75, HH77a, HH78, HB76b, HG77, KS79, KSP75, Koe75, KG77, Let79, LWB78, MBPTS79, MPDGR76, MPJ79a, MPJ79b, MGF77, MH75, MS76, Mil77, MHC77, Mil78, MM78, MC78b, MR75b, NCPW75, NUY^{+}77, OAB76, PT77, PR76b, Pro75, RBK78, RTBE78, RGE78, SPM79, ST78a, SEBL77, STWB78, SBC78, SMPGR75, SD76, Sne76a, SM76, Tak77, TCT79, TM79, UT79, WP78, WNQK78, WF79, WGF79, WDB79, YYP77, Yam78, dP75, dP78]. surfaces [DB75, FYS78, GTH75, HM75, MSS75, ST78a]. surirella [TPH77]. Survival [CP77b]. suspension [SAB75]. suum [RT76]. sv40 [SAPA77, GDR77, GR78, HT79a, Mau76]. SV40-3T3 [HT79a]. SV40-virus-transformed [GR78]. switching [BAN75, BPN76]. symbiotes [TC75]. sympathetic [CF79, CP77b, CP77c, CP77d, EB78, Lan78, WJB78]. syn [HKS77]. synapse [HKP78, TS77]. synapses [FF79a, LP76, LC77, PSG^{+}76]. Synaptic [HRD^{+}79, KC78, SHH76, BBW75, BSM79, FB78, FCG^{+}76, KCGN76, RBB76, RW78, SMM78, SH79a, THCP75, TM76a]. synaptosomal [WM76]. synaptosomes [FB78, KKG76]. synchronized [FK78, GTPWD76, HWW75, KSP75, Ley75, ZJSF76]. synchronous [CM77a, WC77]. syncytia [WCC79]. syncytial [NEK78b, NEK78a]. syndrome [FTBS78]. synthase [MCC79]. Synthesis [CF79, CRSH75, DST^{+}78, GTPWD76, ICS75, WNQK78, BAN75, BPN76, BH75b, BLR76, BAB79, CG77, CM77c, CBH^{+}78, DN76, DH76, GBE75, GGGM78, Gra75a, GCM77, HMPB76, Han75a, HP77, HS76c, IW76, LLL75, LNMR78, LW76, LH79, MRB75, MZ78, MG76, MG78a, NSRL76, NEK78b, NEK78a, NPM75, OWS75, PR76a, Rag77, RBM^{+}78, RG78, RB76, RS75a, SM75b, SSG79, SC75a, THC76, TWR^{+}77, VHvdVB78, WR77c, YBB79, ZM76, CG78, GLR75, MG75b, WR75c]. synthesized [CCB76, FD78, FSSZ75, GCRB78, Hil75, LS75, SH77a, WR77b]. synthetase [JM77, SM75b, Von79]. synthetic [Ber75, Pol75]. system [BSVG75, Coh78, EF76, HK79, KCGN76, MKKV78, MOS^{+}77, RdC77, SS76a, SS79b, SL77, SOW78, SFSB78, SAPA77, SM77, SH77a, Sne76a, Sne76b, WJB78, Wil77a]. systematic [BG76b]. systems [HWAS79, HS79b, KP77b]. Sythesis [ASGR78].

T [GGGM78, SS79b]. T-antigen [GGGM78]. TA3 [MHC77]. tails [Woo77]. Tannic [BHP75, BH78, GCY79, KP77c, KP77b, SWF78]. target [RBB76]. technique [Bou79, DDS79, LF76, LZK75, MDR75, SAB75]. techniques [MBM^{+}79, OT78, OBKP75]. tectum [GKB78]. Teleost [KEP76, KKKS76, Bur78, WB78a]. teleostean [SPM79]. telotrophic [Cav75]. Temperature [Fus75b, HRL77, FJD76, HBR77, IO79, KP77a, Kan76, NRH78, Pol76, PE79, SE75, SSD^{+}78, UIOO76]. temperature-dependent [IO79, Pol76]. temperature-induced [Kan76]. Temperature-sensitive [HRL77, FJD76, KP77a, PE79, SSD^{+}78]. temperatures [HS76c]. Temporal [JT79, MB77a, WC77]. temporary [BJ75]. tendon [NVH^{+}75, OBKP75, SFSB78]. tentans [DAF77]. term [Hop77]. Terminal [MB77b, BBW75, BW78, HS79c, RG78, SDTD^{+}79]. terminals [FB78, Hir77, RSM76, ST78b]. termination [DS79]. termite [BD78, Tam78, Tam79]. test [Edd75a, KS79]. testes [NRH78]. testis [AG75, Con78]. testololactone [Sto75]. tetanus [ST78b, SST79]. Tetrahymena [CB75a, Cle75, FJD76, HB76a, HS77, Jau75, KT77b, KT77a, Nan75, NCM78, OLL76, SS76a, WM78]. tetraurelia [BS78a, DK78b]. tetroxide [MSP78, RW78]. thecal [Swe76]. Their [MJM76, ASD75b, BS77e, Con79, DAF77, DR77b, ESAD75, HLS75, HL75, HWLC78, HL79, JJB79, MG75a, MIMMF77, MKK79, PDD78, TC75, VJGG79, WJB78, WHC79, WK79, BC76, TV79]. themselves [AV78]. theproduction [JLA^{+}76]. there [DSTW77]. Thermodynamics [Sal75b]. Thermotropic [HH77b]. these [EP79a, TD75]. theta [dP75]. thick [Bur78, DWC^{+}77, EWR78, Pau75, RH75]. thickness [WP78]. Thin [Bur78, BRR78, DJM77, EWR78, McN78, MBBS78, OKS77, RH75]. thin-section [McN78, MBBS78, OKS77]. things [MOR78]. thiol [EVMP76]. thiol-oxidation [EVMP76]. those [JKB^{+}79]. Three [MMER79, DCEO^{+}79, Pau75, Sch78b]. Three-dimensional [MMER79, Pau75]. threonine [MR75a]. thrombin [GBBZ78, RAFM75]. thrombin-treated [RAFM75]. throughout [KKH77, MZ78]. thummi [BLR76]. thylakoid [BP76, CBSP76, CG77, HS76c, MS76]. thylakoid-bound [CBSP76]. thymidine [DH76, MF75]. thymine [WF76]. thymocytes [Md78a, Md78b, WJTR77, dP78]. Thyone [Til78]. thyroglobulin [PL77]. thyroid [HP77, JSGR78, PL77, TWC75, Von79, ZW77]. Tight [vDK79, EF76, HS76d, HGP^{+}76, MFPO75, MGPO76, TWC75]. Tilia [Fus75a]. time [Bal79, Lu78]. time-course [Lu78]. timing [Slu79]. Tipping [GJ78]. tips [DR77b, JWJ75]. tissue [BRS78, BB78, FP78b, KSS76, NBB77b, PR75, PFK^{+}76, SPT78, VP75, VRN76, WBE75]. tissues [AKD75a, BGFK79, LYSS78, LP77, MB76, OLK78, PPMK77, PPMSK77]. toad [MM78]. tolerant [SSF76]. tonofilaments [DFW^{+}78]. Topochemistry [SMPGR75]. Topographical [HG77, MPDGR76, CO76]. Topography [WM76, AKD75b, ES76, GR78, MBPTS79, AMBM76]. Torpedo [HS79a, KHV^{+}78, MIMMF77]. torque [Tam78]. total [PPMK77]. toxin [DMMW76, DCEO^{+}79, JSG79, NSH78, ST78b, SST79, UKY^{+}79]. toxin-resistant [NSH78]. tracer [Con78, ST78b, RCV75]. tracers [Pai75]. tracheal [GLM77, WBY79]. traffic [Far78]. transcription [AHB79, BDCD75, STF76, WC77]. transcriptional [KT75]. transduction [MRZV76]. transected [SM78a]. transendothelial [SSP75b]. transepithelial [FMR79, Har78a, LD76]. Transfer [BD75a, BD75b, LL79, RMP75, ESAD75, GTL77, JF75, LGB77, SST79, WBE75, SOK79]. transferase [FLRS77]. transfilter [MH75]. transformant [SAPA77]. Transformation [AZW^{+}79, CCB76, IO79, KP77a, RMP^{+}78, SFSB78, TZ77, YYP77, Yam78]. transformation-associated [KP77a]. transformation-sensitive [YYP77, Yam78]. Transformations [BP77]. transformed [Bel77, CT76, DN76, EP79b, GD75, GDC76, GDR77, HB79, LWS79, ML76, MH77, Moo76a, Qui76, RWLC78, RW79, SRR76, WK79, GR78]. transglutaminase [GL79a, RG78]. transients [Gol79]. transition [HBD75, RJ76, RLJ76, Toy76, VV77]. transitional [CH79b, MC78b]. translational [HH77b]. translocation [BCHN75, CMC77, FMSB75, HDS77, MC78b, SS78f]. translucent [CRD^{+}78]. Transmembrane [KL79, TE79, GG77, SRR79, BSPK78]. transmembranous [RTC76]. transmission [KW76]. transmitter [HRD^{+}79, WJB78]. transplantation [MIT76]. transplanted [ZM76]. Transport [CS79, Ern75, NSBS77b, RAFM75, EM77, FMR79, GSB77, GBE75, GKS76, HT79a, HL75, KG77, MKK79, OB78, PP79, PM75, PRZE75, Poh76, RR77a, ST78b, SST79, SSG79, SCRP77, TVD78, TV79, WC75, WJTR77, WWLS79, SCRP77]. transported [Wil77a]. transsynaptic [SST79]. Transverse [JC79, FALP75, ZVR75]. traverse [Hel79, KPF75, WNJ75]. traverses [GJRR78]. treated [BVBM76, Blo79, FMR79, Moo76b, RAFM75, SS79d]. treatment [Bal79, BO76, GDR77, Hir77, KD78, ODS79, SO75, SM75b, SSSW77, WDM^{+}75, YB75]. triacylglycerol [CB78]. triad [Som79]. trichocyst [BLTP^{+}76, HA76]. Trichosia [AJ78]. triggering [DR77a]. triggers [BSU79]. triphosphate [LG75b, PW78]. triphosphate-induced [LG75b]. triplet [GS75]. triton [BS77b, Bro79, Moo76b, SS79d, LG75b]. triton-demembranated [BS77b, Bro79]. triton-treated [Moo76b, SS79d]. Triturus [STF75]. trophic [LCK75]. trophoblast [NEK78b, NEK78a]. Tropomyosin [Laz75, ASGR78, Laz76, Lem79]. Troponin [OSM79]. true [LHH76]. Trypanosoma [FWHB78, FMW75, HLS75, MPDGR76, RS79]. trypanosomatid [LS79b]. trypanosomatids [Pau75]. trypsin [BS77b, MGF77, YB75]. trypsin-sensitive [MGF77]. tube [Wei76a]. tubes [JWJ75, WNJ75]. tubular [SV77, Wil77b]. tubule [FALP75, HD77]. tubules [HGP^{+}76, Kyt76a, Kyt76b, PS75, SS78b, WOPB79, WPS78, ZVR75]. Tubulin [MS75a, WDBM79, BBDW76, BR78, BR76a, BH78, CLM77, DF77, FP78b, GLM77, KKG76, KC78, MB75b, MR76a, PB79, PAW76, PPMK77, PPMSK77, RF77, RCM77, RW79, SM78b, SM77, Slu76]. tubulin-microtubule [SM77]. tumor [BS79a, BAB79, COK76, IO79, KS79, RHDL77]. tumor-forming [KS79]. tumorigenicity [KS79]. turkey [BAPG79]. Turnover [BHR77, HBR77, CF79, DF75, ICS75, LCK75, RW78, RGS^{+}76, RdC77, THCP75]. turtle [SL78, SR78a]. tween [BO76]. tween-80 [BO76]. twisted [Woo77]. twitch [ERSP76, FALP75]. Two [BFPU78a, BFPU78b, CRBH78, KOBP75, ST77a, ST77b, Edd79, GTPWD76, IR78, JZR76, LF76, MKK79, TD75, UGB^{+}79, Wil77a, LLC79]. Two-dimensional [LLC79]. two-stage [IR78]. two-step [LF76]. type [AEP79, BLTP^{+}76, GW75, JKB^{+}79, KP77b, KAB75, MM77, MSK^{+}79, RMP^{+}78, SAPA77, VV77, MM77]. type-specific [AEP79, RMP^{+}78]. types [GL77, GL79b, Gau79, HWAS79, LN75]. tyrosine [GET^{+}78]. tyrosylation [RF77]. Tyrosyltubulin [FK78].

UDP [BN76]. UDP-glucuronyltransferase [BN76]. ultracytochemical [AP75]. ultrahigh [GBM78]. Ultramicroinjection [LZK75]. Ultrastructural [BCFS75, CCH78, DGR75, EH77, FN76, Gau79, GLR77, MHC77, NCCB77, NYY75, OSO76, PL76, SHWH^{+}77, SS78e, SLR77, SS78f, Ern75, GGS^{+}76, Lon77, MB75a, MR76b, MR75b, RR77a, RS75a, SWF78, Sch77, VL78, GESG75, GJAE75, GTH76, SBU76, SSC76a, Sne76a, WDMR77, WR75b]. Ultrastructure [BS79a, DJM77, DK78b, GCY79, HH76, SHM^{+}77, VD76, ZW77, BO78, BWDP77, BBW75, Dip76, FMR79, GDR77, KP77b, Md78a, MCR77, RLJ76, SS78e, Wei76a]. ultrathin [TS76]. ultraviolet [JF75]. ultraviolet-mediated [JF75]. umbilical [GBBZ78]. unable [Pro75]. uncoupling [Bal79, IP78b, PD76, Per77, dCdM78]. undergo [VJGG79]. undergoing [BBMB77, HGP^{+}76]. understanding [SFSB78]. undulations [Woo77]. Unequal [HU77]. unfertilized [HMK79, PAW76, SS79d]. unicellular [HB76b]. uniform [BRR78]. uninucleolate [Cav76]. unit [JZR76, RKL^{+}76]. units [BS78b, Han75b, KBS79]. unlimited [NPM75]. unprocessed [BD75a]. unstacked [MS76]. unstacking [Sta76]. untransformed [YYP77]. untranslated [Gre75]. Unusual [EWR78, Svo78]. upon [HSM76, KS77, NPM75, SEHW79]. upper [DB75, EB75, Pro75]. upsilon [OCO75]. Uptake [FBF79, OH78, BK75, BSHP79, Han79, HP75, KAGV79, LLL75, RSM76, SHH76, ST78b, Smi77, TDHO76, WJB78, WKBL75a, WKBL75b, STdDT79b]. urchin [AP75, BR79, BS77b, Bro79, BS77c, CH79a, DL79, DDSL77, ES76, GOG76, GV77, Gol79, HMK79, Kan75, Kan76, KBS79, MNK79, MBV77, NR76, OAB77, PW78, PAW76, Rik78a, SG79a, SEBL77, Slu76, SS79d, WT79]. urchins [GL79c]. urechis [DB76a, GSJH79]. uridine [SP76]. urinary [Har78a, MM78, MC78b]. Uromyces [HH76]. uropod [dP78]. uropod-forming [dP78]. urothelial [CB77]. Use [Col79, HA77, RMP^{+}78, CGK75, DH76, GCY79, GLM77, LMPI76, LJ75, OBKP75, PK79, SWF78, WJB78, CW78b, WDB79]. using [AL75, Bou79, BP77, FPC76, MW76, MBM^{+}79, NSH78, NT78, PL77, Sch76, ST78b, WRP75]. Ustilago [DTT75]. uterine [KAD75]. utilization [LE78b]. utilizing [HJP78]. uv [SV76]. uv-microbeam [SV76].

V [BAN75, CT77, FRT^{+}76, KKH77, ST76a, Til78, WJ76]. vacuolar [KT77b]. vacuole [LS79b, OT76, OT78, WGG77b]. vacuoles [NMQY77, Paa78b, Pfe78]. vacuum [GBM78, SM78c]. valinomycin [MSRT78]. vanadate [CW78a, SG79a]. var [HH76]. variability [BBDW76]. variant [NSH78]. Variation [WGC76]. Variations [SF76, BP76, CBSP76, Cla76, HS76d, NCM78, PS75, Far78]. varicosities [BH79]. variegatus [KEP76]. various [Con79, DTT75, FRT^{+}76, KEP76, PK79]. vas [BH79, MPS79]. Vascular [VJGG79, BC76, GBBZ78, SSP75a, SSP76, SSS79]. vasopressin [MM78]. vegetative [FSSZ75, RdC77]. vein [GBBZ78]. veins [SSP76]. velocity [Fus75b, RCMI75]. venom [CGH79a, FCG^{+}76, GHC78, OB78]. ventricle [KBD77]. venules [SSP78]. vero [DDR75]. versatile [PK79]. versus [SES76]. Vertebrate [BGFK79, Coo76, GESG75]. very [LW76, TWR^{+}77]. vesicle [BBW75, FLK^{+}76, FB78, HRD^{+}79, KKG76, MM76a, MM76b, MM76c, MH76b, THCP75]. vesicle-cell [MM76a, MM76b, MM76c]. vesicles [All75, BK75, BFT77, CB79, CB78, DVEM79, DMB78, FCG^{+}76, Goo76, HA76, HP75, KH76, Lan78, LLP77, MRDC77, PH75a, PT77, RW78, SH77a, UKY^{+}79]. vesicular [GWS77, KL79]. vesiculated [MBPTS79]. vessel [SD76]. VI [BPN76, HT79b, RFB^{+}76, ST76b, TK79]. viability [MST78]. viable [HR79b, RCMI75]. vicilin [BTC78]. viewed [OWW78]. Vignae [HH76].

VII [LN75]. VIII [Lu78]. villus [WNQK78]. vinblastine [HV75]. violaceum [Roo75]. Vipera [OB78]. viral [GGGM78, LF77]. virally [HB79]. virus [Bel77, BCB79, BB79b, DSTW77, GD75, GDC76, GDR77, GR78, GGGM78, GWS77, HWW75, HFK77, KP77a, KL79, LZK75, Moo76a, SRR76, SH79b, SH77b, TE79, UKY^{+}79, VL78]. virus-induced [LZK75, SH79b, VL78]. virus-infected [SH77b]. virus-transformed [Bel77, GD75, GDC76, GDR77, Moo76a, SRR76]. viruses [DN76]. visual [MDRY75, OY78, Wil77a, HM75]. Visualization [CK78, JJ78, OCO75, PJ76a, Zaa79, BRR78, BS78b, HMC79a, MJM76, GR78, ST78a]. visualized [AHB79, JJB79, PB75]. Vitamin [EF76, BGRC77, TSC78b]. Vitamin-A-induced [EF76]. vitelline [GV77]. vitellogenin [WJ76]. Vitreous [NLT76]. vitro [AL75, BGRC77, BH75a, BAN75, BFTH79, BR78, BCB79, BBMB77, CH79b, CP77b, CP77c, CP77d, DMKR75, DGR75, EF76, ES79, FAE75, FF79a, GZM75, Goo76, HG79, HS75b, HL79, HB78, KD78, KF79, KBR79, LB78, LLMW77, ML76, MKKV78, MR77, MB75b, MR76a, MR75a, NP75, OT76, OT78, OSM79, OLK77, PBH78, PM79, RPGR77, RF77, SS79a, SS79c, SO77, SBC78, Sta76, SK79, TR79a, VV77, WLW78, WR75a, WKBL75a, WKBL75b, WL76]. vitro-assembled [DGR75]. vivo [ALMS76, BSHP79, Bon75a, BBW75, FWS75, FAE75, JRFS75, KF79, LE76, MFPO75, MGPO76, MF75, RBM^{+}78, Sal75b, Sch76, VJGG79, ZW77]. VM [KPF75]. VM-26 [KPF75]. voltage [BP77, Pau75]. volume [WP78]. Volvox [VK77]. vorticellid [Rou78]. VP [KPF75]. VP-16-213 [KPF75]. vulgare [MPHMT77, MEM79, MMER79].

wall [GW75, HQ78, SS78e, SS78f, WB78b]. walls [RHHB76]. water [BK75, Har78a, JZR76]. water-soluble [BK75, JZR76]. wave [GJRR78, PW78]. weaver [HHP76]. web [BW78, HS79c]. weight [Amo77, BR76a, BDKS79, CKCK78, DGR75, MHP78b, SS78c, SS76c, SS76d]. wheat [DMKR75, MJM76, SC78]. where [EP79a]. which [BP79, ST77b, SC78, WJB78, vVRN76]. whole [BP77, FSW77]. whole-cell [BP77]. WI [HMT79]. WI-38 [HMT79]. widow [CGH79a, FCG^{+}76, GHC78]. Wiener [SEI75]. wild [BLTP^{+}76, GW75, RS79, SAPA77]. wild-type [BLTP^{+}76, GW75, SAPA77]. windows [SE75]. within [BBDW76, BCS77, BS77c, CF76b, GHMD77, GD77, JWJ75, NCM78]. worm [GSJH79]. wound [GCH78, JJ75, SS79c].

x [CM77a, CSHS77, CSP79, CGMP77, CM77a, KHT^{+}79, LG75b, Lon77, MCPG77, MBM^{+}79, PD75, SHC^{+}79]. X-100 [LG75b]. X-ray [CSHS77, CSP79, CGMP77, KHT^{+}79, MCPG77, MBM^{+}79, PD75]. X537A [SGCS75]. Xenopus [CM78, DH75, FP78a, HK75, MH76b, MC78a, ODS79, ST78a]. XXX [BD78]. xylinum [Zaa79]. y-1 [HS76c]. yeast [BG76c, Dam76, HB78, KAB75, SM78c, TM79]. yield [SAB75]. yl [DF77]. yolk [JB78, WKBL75a, WKBL75b].

Z [GL79a, GSS77, GSS79]. Z-line [GL79a]. zanclea [CCTT77]. Zeiosis [GMDT75]. zinc [RW78]. ZIO [RW78]. Zn [KCH79]. zonae [NYY75]. zone [CGH79a, EL76]. zones [PSZK78b]. zoospores [dSN77]. zygotes [Lon76]. zymogen [RD78, RP78].

References

Anderson:1976:GJB

[AA76] E. Anderson and D. F. Albertini. Gap junctions between the oocyte and companion follicle cells in the mammalian ovary. Journal of Cell Biology, 71(2):680–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/680.

Albertini:1977:MMR

[AA77] D. F. Albertini and E. Anderson. Microtubule and microfilament rearrangements during capping of concanavalin A receptors on cultured ovarian granulosa cells. Journal of Cell Biology, 73(1):111–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/1/111.

Albrecht-Buehler:1976:FSC

[AB76] G. Albrecht-Buehler. Filopodia of spreading 3T3 cells. Do they have a substrate-exploring function? Journal of Cell Biology, 69(2):275–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/275.

Albrecht-Buehler:1977:DCT

[AB77] G. Albrecht-Buehler. Daughter 3T3 cells. Are they mirror images of each other? Journal of Cell Biology, 72(3):595–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/3/595.

Albrecht-Buehler:1979:ADD

[AB79] G. Albrecht-Buehler. The angular distribution of directional changes of guided 3T3 cells. Journal of Cell Biology, 80(1):53–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
1/53.

Albrecht-Buehler:1976:QDE

[ABL76] G. Albrecht-Buehler and R. M. Lancaster. A quantitative description of the extension and retraction of surface protrusions in spreading 3T3 mouse fibroblasts. Journal of Cell Biology, 71 (2):370–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/2/370.

Andrews:1979:FFE

[AC79] L. D. Andrews and A. I. Cohen. Freeze-fracture evidence for the presence of cholesterol in particle-free patches of basal disks and the plasma membrane of retinal rod outer segments of mice and frogs. Journal of Cell Biology, 81(1):215–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/1/215.

Afzelius:1975:LDA

[AEJL75] B. A. Afzelius, R. Eliasson, O. Johnsen, and C. Lindholmer. Lack of dynein arms in immotile human spermatozoa. Journal of Cell Biology, 66(2):225–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/2/225.

Agabian:1979:GAD

[AEP79] N. Agabian, M. Evinger, and G. Parker. Generation of asymmetry during development. Segregation of type-specific proteins in Caulobacter. Journal of Cell Biology, 81(1):123–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/1/123.

Ashby:1979:IHL

[AFB79] B. Ashby, C. Frieden, and R. Bischoff. Immunofluorescent and histochemical localization of AMP deaminase in skeletal muscle. Journal of Cell Biology, 81(2):361–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/2/361.

Alder:1975:EAL

[AG75] D. Alder and M. A. Gorovsky. Electrophoretic analysis of liver and testis histones of the frog Rana pipiens. Journal of Cell Biology, 64(2):389–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/2/389.

Anderson:1977:DAS

[AH77] R. G. Anderson and C. E. Hein. Distribution of anionic sites on the oviduct ciliary membrane. Journal of Cell Biology, 72 (2):482–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/2/482.

Angelier:1979:MTN

[AHB79] N. Angelier, D. Hemon, and M. Bouteille. Mechanisms of transcription in nucleoli of amphibian oocytes as visualized by highresolution autoradiography. Journal of Cell Biology, 80(2):277–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
2/277.

Alexander:1976:SLB

[AHH76] C. A. Alexander, R. L. Hamilton, and R. J. Havel. Subcellular localization of B apoprotein of plasma lipoproteins in rat liver. Journal of Cell Biology, 69(2):241–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/241.

Amabis:1978:EIG

[AJ78] J. M. Amabis and C. Janczur. Experimental induction of gene activity in the salivary gland chromosomes of Trichosia pubescens (Diptera: Sciaridae). Journal of Cell Biology, 78(1):1–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/1/1.

Adair:1978:LCA

[AJG78] W. S. Adair, D. Jurivich, and U. W. Goodenough. Localization of cellular antigens in sodium dodecyl sulfate-polyacrylamide gels. Journal of Cell Biology, 79(1):281–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/1/281.

Anderson:1975:EPS

[AKD75a] W. A. Anderson, Y. H. Kang, and E. R. DeSombre. Endogenous peroxidase: specific marker enzyme for tissues displaying growth dependency on estrogen. Journal of Cell Biology, 64(3):668–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
3/668.

Anderson:1975:EAI

[AKD75b] W. A. Anderson, Y. H. Kang, and E. R. DeSombre. Estrogen and antagonist-induced changes in endometrial topography of immature and cycling rats. Journal of Cell Biology, 64(3):692–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
3/692.

Amsterdam:1975:DBS

[AKLL75] A. Amsterdam, Y. Koch, M. E. Lieberman, and H. R. Lindner. Distribution of binding sites for human chorionic gonadotropin in the preovulatory follicle of the rat. Journal of Cell Biology, 67 (3):894–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/3/894.

Armstrong:1975:SII

[AL75] P. B. Armstrong and J. M. Lackie. Studies of intercellular invasion in vitro using rabbit peritoneal neutrophil granulocytes (PMNS). I. Role of contact inhibition of locomotion. Journal of Cell Biology, 65(2):439–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/439.

Atlas:1978:DBB

[AL78] S. J. Atlas and S. Lin. Dihydrocytochalasin B. Biological effects and binding to 3T3 cells. Journal of Cell Biology, 76(2):360–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
2/360.

Allen:1975:EFL

[All75] R. D. Allen. Evidence for firm linkages between microtubules and membrane-bounded vesicles. Journal of Cell Biology, 64 (2):497–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/2/497.

Adesnik:1976:RME

[ALMS76] M. Adesnik, M. Lande, T. Martin, and D. D. Sabatini. Retention of mRNA on the endoplasmic reticulum membranes after in vivo disassembly of polysomes by an inhibitor of initiation. Journal of Cell Biology, 71(1):307–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/1/307.

Apel:1976:CMG

[AMBM76] K. Apel, K. R. Miller, L. Bogorad, and G. J. Miller. Chloroplast membranes of the green alga Acetabularia mediterranea. II. Topography of the chloroplast membrane. Journal of Cell Biology, 71(3):876–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/3/876.

Aikawa:1978:EEM

[AMJR78] M. Aikawa, L. H. Miller, J. Johnson, and J. Rabbege. Erythrocyte entry by malarial parasites. A moving junction between erythrocyte and parasite. Journal of Cell Biology, 77(1):72–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/1/72.

Amos:1977:AHM

[Amo77] L. A. Amos. Arrangement of high molecular weight associated proteins on purified mammalian brain microtubules. Journal of Cell Biology, 72(3):642–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/3/642.

Abe:1976:DAS

[AMS76] H. Abe, M. A. Moscarello, and J. M. Sturgess. The distribution of anionic sites on the surface of the Golgi complex. Journal of Cell Biology, 71(3):973–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/3/973.

Andrews:1976:MCR

[And76] P. M. Andrews. Microplicae: characteristic ridge-like folds of the plasmalemma. Journal of Cell Biology, 68(3):420–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/3/420.

Anderson:1977:BCE

[And77] R. G. Anderson. Biochemical and cytochemical evidence for ATPase activity in basal bodies isolated from oviduct. Journal of Cell Biology, 74(2):547–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/547.

Anderson:1975:USR

[AP75] W. A. Anderson and M. E. Perotti. An ultracytochemical study of the respiratory potency, integrity, and fate of the sea urchin sperm mitochondria during early embryogenesis. Journal of Cell Biology, 66(2):367–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/367.

Armond:1977:SRP

[ASA77] P. A. Armond, L. A. Staehelin, and C. J. Arntzen. Spatial relationship of photosystem I, photosystem II, and the lightharvesting complex in chloroplast membranes. Journal of Cell Biology, 73(2):400–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/2/400.

Autuori:1975:BMMa

[ASD75a] F. Autuori, H. Svensson, and G. Dallner. Biogenesis of microsomal membrane glycoproteins in rat liver. I. Presence of glycoproteins in microsomes and cytosol. Journal of Cell Biology, 67(3):687–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/687.

Autuori:1975:BMMb

[ASD75b] F. Autuori, H. Svensson, and G. Dallner. Biogenesis of microsomal membrane glycoproteins in rat liver. II. Purification of soluble glycoproteins and their incorporation into microsomal membranes. Journal of Cell Biology, 67(3):700–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/3/700.

Allen:1978:STC

[ASGR78] R. E. Allen, M. H. Stromer, D. E. Goll, and R. M. Robson. Sythesis of tropomyosin in cultures of differentiating muscle cells. Journal of Cell Biology, 76(1):98–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/98.

Albersheim:1978:HPI

[AV78] P. Albersheim and B. S. Valent. Host-pathogen interactions in plants. Plants, when exposed to oligosaccharides of fungal origin, defend themselves by accumulating antibiotics. Journal of Cell Biology, 78(3):627–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/627.

Allen:1979:TMH

[AZW^{+}79] R. D. Allen, L. R. Zacharski, S. T. Widirstky, R. Rosenstein, L. M. Zaitlin, and D. R. Burgess. Transformation and motility of human platelets: details of the shape change and release reaction observed by optical and electron microscopy. Journal of Cell Biology, 83 (1):126–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/1/126.

Burghardt:1979:HMO

[BA79] R. C. Burghardt and E. Anderson. Hormonal modulation of ovarian interstitial cells with particular reference to gap junctions. Journal of Cell Biology, 81(1):104–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/1/104.

Biswas:1979:MBI

[BAB79] D. K. Biswas, K. T. Abdullah, and B. A. Brennessel. On the mechanism of 5-bromodeoxyuridine induction of prolactin synthesis in rat pituitary tumor cells. Journal of Cell Biology, 81(1):1–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/81/1/1.

Basu:1977:MCL

[BAGB77] S. K. Basu, R. G. W. Anderson, J. L. Goldstein, and M. S. Brown. Metabolism of cationized lipoproteins by human fibroblasts: biochemical and morphologic correlations. Journal of Cell Biology, 74(1):119–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/119.

Baldwin:1979:CGJ

[Bal79] K. M. Baldwin. Cardiac gap junction configuration after an uncoupling treatment as a function of time. Journal of Cell Biology, 82(1):66–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/66.

Barker:1975:HSS

[BAN75] J. E. Barker, W. F. Anderson, and A. W. Nienhuis. Hemoglobin switching in sheep and goats. V. Effect of erythropoietin concentration on in vitro erythroid colony growth and globin synthesis. Journal of Cell Biology, 64(3):515–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/3/515.

Beam:1979:HRP

[BAPG79] K. G. Beam, S. L. Alper, G. E. Palade, and P. Greengard. Hormonally regulated phosphoprotein of turkey erythrocytes: localization to plasma membrane. Journal of Cell Biology, 83(1):1–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/1.

Bunge:1978:ECC

[BB78] R. P. Bunge and M. B. Bunge. Evidence that contact with connective tissue matrix is required for normal interaction between Schwann cells and nerve fibers. Journal of Cell Biology, 78(3): 943–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/943.

Boyles:1979:CPP

[BB79a] J. Boyles and D. F. Bainton. Changing patterns of plasma membrane-associated filaments during the initial phases of polymorphonuclear leukocyte adherence. Journal of Cell Biology, 82 (2):347–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/2/347.

Buechi:1979:IEM

[BB79b] M. Büechi and T. Bächi. Immunofluorescence and electron microscopy of the cytoplasmic surface of the human erythrocyte membrane and its interaction with Sendai virus. Journal of Cell Biology, 83(2):338–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/2/338.

Bergeron:1978:PSD

[BBC78] J. J. Bergeron, D. Borts, and J. Cruz. Passage of serum-destined proteins through the Golgi apparatus of rat liver. An examination of heavy and light Golgi fractions. Journal of Cell Biology, 76(1):87–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/1/87.

Baccetti:1979:DEB

[BBDP79] B. Baccetti, A. G. Burrini, R. Dallai, and V. Pallini. The dynein electrophoretic bands in axonemes naturally lacking the inner or the outer arm. Journal of Cell Biology, 80(2):334–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/2/334.

Bibring:1976:HAS

[BBDW76] T. Bibring, J. Baxandall, S. Denslow, and B. Walker. Heterogeneity of the alpha subunit of tubulin and the variability of tubulin within a single organism. Journal of Cell Biology, 69(2):301–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
2/301.

Basinger:1976:RRO

[BBH76] S. Basinger, D. Bok, and M. Hall. Rhodopsin in the rod outer segment plasma membrane. Journal of Cell Biology, 69(1):29–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/69/1/29.

Brenner:1977:ACS

[BBMB77] S. Brenner, A. Branch, S. Meredith, and M. W. Berns. The absence of centrioles from spindle poles of rat kangaroo (PtK2) cells undergoing meiotic-like reduction division in vitro. Journal of Cell Biology, 72(2):368–??, February 1977. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/2/368.

Boyne:1975:CCS

[BBW75] A. F. Boyne, T. P. Bohan, and T. H. Williams. Changes in cholinergic synaptic vesicle populations and the ultrastructure of the nerve terminal membranes of Narcine brasiliensis electron organ stimulated to fatigue in vivo. Journal of Cell Biology, 67 (3):814–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/3/814.

Blouin:1977:DOM

[BBW77] A. Blouin, R. P. Bolender, and E. R. Weibel. Distribution of organelles and membranes between hepatocytes and nonhepatocytes in the rat liver parenchyma. A stereological study. Journal of Cell Biology, 72(2):441–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/2/441.

Blose:1976:RIF

[BC76] S. H. Blose and S. Chacko. Rings of intermediate (100 A) filament bundles in the perinuclear region of vascular endothelial cells. Their mobilization by colcemid and mitosis. Journal of Cell Biology, 70(2):459–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/2/459.

Berezney:1977:NMI

[BC77] R. Berezney and D. S. Coffey. Nuclear matrix: isolation and characterization of a framework structure from rat liver nuclei. Journal of Cell Biology, 73(3):616–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/3/616.

Bornens:1978:INE

[BC78] M. Bornens and J. C. Courvalin. Isolation of nuclear envelopes with polyanions. Journal of Cell Biology, 76(1):191–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/1/191.

Berger:1979:PCA

[BC79] E. Berger and G. Cox. A precursor of cytoplasmic actin in cultured Drosophila cells. Journal of Cell Biology, 81(3):680–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/3/680.

Banerjee:1977:BLE

[BCB77] S. D. Banerjee, R. H. Cohn, and M. R. Bernfield. Basal lamina of embryonic salivary epithelia. Production by the epithelium and role in maintaining lobular morphology. Journal of Cell Biology, 73(2):445–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/2/445.

Bonatti:1979:MBV

[BCB79] S. Bonatti, R. Cancedda, and G. Blobel. Membrane biogenesis. In vitro cleavage, core glycosylation, and integration into microsomal membranes of sindbis virus glycoproteins. Journal of Cell Biology, 80(1):219–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/219.

Bignon:1975:UID

[BCFS75] J. Bignon, P. Chahinian, G. Feldmann, and C. Sapin. Ultrastructural immunoperoxidase demonstration of autologous albumin in the alveolar capillary membrane and in the alveolar lining material in normal rats. Journal of Cell Biology, 64(2):503–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/2/503.

Breuer:1975:CAO

[BCHN75] A. C. Breuer, C. N. Christian, M. Henkart, and P. G. Nelson. Computer analysis of organelle translocation in primary neuronal cultures and continuous cell lines. Journal of Cell Biology, 65(3):562–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/3/562.

Bellve:1977:SCP

[BCM^{+}77] A. R. Bellve, J. C. Cavicchia, C. F. Millette, D. A. O’Brien, Y. M. Bhatnagar, and M. Dym. Spermatogenic cells of the prepuberal mouse: isolation and morphological characterization. Journal of Cell Biology, 74(1):68–??, July 1977. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/1/68.

Blomberg:1977:SPD

[BCS77] F. Blomberg, R. S. Cohen, and P. Siekevitz. The structure of postsynaptic densities isolated from dog cerebral cortex: II. characterization and arrangement of some of the major proteins within the structure. Journal of Cell Biology, 74(1):204–??, July 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/1/204.

Blobel:1975:TPAa

[BD75a] G. Blobel and B. Dobberstein. Transfer of proteins across membranes. I. Presence of proteolytically processed and unprocessed nascent immunoglobulin light chains on membrane-bound ribosomes of murine myeloma. Journal of Cell Biology, 67(3):835–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/835.

Blobel:1975:TPAb

[BD75b] G. Blobel and B. Dobberstein. Transfer of proteins across membranes. II. Reconstitution of functional rough microsomes from heterologous components. Journal of Cell Biology, 67(3):852–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/852.

Baccetti:1978:SAX

[BD78] B. Baccetti and R. Dallai. The spermatozoon of arthropoda. XXX. The multiflagellate spermatozoon in the termite Mastotermes darwiniensis. Journal of Cell Biology, 76(3):569–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/3/569.

Benedetto:1975:SRT

[BDCD75] A. Benedetto, C. Delfini, G. Carloni, and W. Djaczenko. Study on the reinitiation of transcription in 37 RC cells after actinomycin D inhibition. Spectrum of major RNA species resynthesized after maximal suppression by the drug. Journal of Cell Biology, 67 (3):538–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/3/538.

Briggs:1975:LNO

[BDKK75] R. T. Briggs, D. B. Drath, M. L. Karnovsky, and M. J. Karnovsky. Localization of NADH oxidase on the surface of human polymorphonuclear leukocytes by a new cytochemical method. Journal of Cell Biology, 67(3):566–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/566.

Blumenthal:1979:RIM

[BDKS79] A. B. Blumenthal, J. D. Dieden, L. N. Kapp, and J. W. Sedat. Rapid isolation of metaphase chromosomes containing high molecular weight DNA. Journal of Cell Biology, 81(1):255–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/1/255.

Baccetti:1977:PIS

[BDP^{+}77] B. Baccetti, R. Dallai, V. Pallini, F. Rosati, and B. A. Afzelius. Protein of insect sperm mitochondrial crystals. Crystallomitin. Journal of Cell Biology, 73(3):594–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/3/594.

Bartlett:1978:EHG

[BE78] P. F. Bartlett and M. Edidin. Effect of the H-2 gene complex rates of fibroblast intercellular adhesion. Journal of Cell Biology, 77(2):377–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/377.

Begg:1979:MSCa

[BE79a] D. A. Begg and G. W. Ellis. Micromanipulation studies of chromosome movement. I. Chromosome-spindle attachment and the mechanical properties of chromosomal spindle fibers. Journal of Cell Biology, 82(2):528–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/2/528.

Begg:1979:MSCb

[BE79b] D. A. Begg and G. W. Ellis. Micromanipulation studies of chromosome movement. II. Birefringent chromosomal fibers and the mechanical attachment of chromosomes to the spindle. Journal of Cell Biology, 82(2):542–??, August 1979. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/2/542.

Bell:1977:LBC

[Bel77] P. B. Bell. Locomotory behavior, contact inhibition, and pattern formation of 3T3 and polyoma virus-transformed 3T3 cells in culture. Journal of Cell Biology, 74(3):963–??, September 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/3/963.

Berliner:1975:FEM

[Ber75] J. A. Berliner. Formation of enlarged mitochondria in a liver cell line in response to a synthetic glucocorticoid. Journal of Cell Biology, 64(3):711–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/64/3/711.

Brandt:1975:DLE

[BES75] E. J. Brandt, R. W. Elliott, and R. T. Swank. Defective lysosomal enzyme secretion in kidneys of Chediak–Higashi (beige) mice. Journal of Cell Biology, 67(3):774–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/3/774.

Berg:1978:ESC

[BF78] D. K. Berg and G. D. Fischbach. Enrichment of spinal cord cell cultures with motoneurons. Journal of Cell Biology, 77(1):83–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/77/1/83.

Bernstein:1978:IPC

[BFD78] S. I. Bernstein, E. A. Fyrberg, and J. J. Donady. Isolation and partial characterization of Drosophila myoblasts from primary cultures of embryonic cells. Journal of Cell Biology, 78 (3):856–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/856.

Braun:1978:TDMa

[BFPU78a] J. Braun, K. Fujiwara, T. D. Pollard, and E. R. Unanue. Two distinct mechanisms for redistribution of lymphocyte surface macromolecules. I. Relationship to cytoplasmic myosin. Journal of Cell Biology, 79(2):409–??, November 1978. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/2/409.

Braun:1978:TDMb

[BFPU78b] J. Braun, K. Fujiwara, T. D. Pollard, and E. R. Unanue. Two distinct mechanisms for redistribution of lymphocyte surface macromolecules. II. Contrasting effects of local anesthetics and a calcium ionophore. Journal of Cell Biology, 79(2):419–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/2/419.

Blitz:1977:ECV

[BFT77] A. L. Blitz, R. E. Fine, and P. A. Toselli. Evidence that coated vesicles isolated from brain are calcium-sequestering organelles resembling sarcoplasmic reticulum. Journal of Cell Biology, 75 (1):135–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/135.

Bennett:1979:RIF

[BFTH79] G. S. Bennett, S. A. Fellini, Y. Toyama, and H. Holtzer. Redistribution of intermediate filament subunits during skeletal myogenesis and maturation in vitro. Journal of Cell Biology, 82(2):577–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
2/577.

Bartos:1976:PGR

[BG76a] E. M. Bartos and A. D. Glinos. Properties of growth-related acetylcholinesterase in a cell line of fibroblastic origin. Journal of Cell Biology, 69(3):638–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/3/638.

Bedi:1976:AAN

[BG76b] K. S. Bedi and D. J. Goldstein. Apparent anomalies in nuclear feulgen–DNA contents. Role of systematic microdensitometric errors. Journal of Cell Biology, 71(1):68–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/68.

Byers:1976:HOR

[BG76c] B. Byers and L. Goetsch. A highly ordered ring of membraneassociated filaments in budding yeast. Journal of Cell Biology, 69(3):717–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/717.

Briles:1979:VLC

[BGFK79] E. B. Briles, W. Gregory, P. Fletcher, and S. Kornfeld. Vertebrate lectins, Comparison of properties of beta-galactoside-binding lectins from tissues of calf and chicken. Journal of Cell Biology, 81(3):528–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/3/528.

Bergman:1975:GDC

[BGG^{+}75] K. Bergman, U. W. Goodenough, D. A. Goodenough, J. Jawitz, and H. Martin. Gametic differentiation in Chlamydomonas reinhardtii. II. Flagellar membranes and the agglutination reaction. Journal of Cell Biology, 67(3):606–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/3/606.

Brown:1979:RAC

[BGK^{+}79] M. S. Brown, J. L. Goldstein, M. Krieger, Y. K. Ho, and R. G. Anderson. Reversible accumulation of cholesteryl esters in macrophages incubated with acetylated lipoproteins. Journal of Cell Biology, 82(3):597–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/3/597.

Balin:1977:EOV

[BGRC77] A. K. Balin, D. B. Goodman, H. Rasmussen, and V. J. Cristofalo. The effect of oxygen and vitamin E on the lifespan of human diploid cells in vitro. Journal of Cell Biology, 74(1):58–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/1/58.

Balin:1978:OSS

[BGRC78] A. K. Balin, D. B. Goodman, H. Rasmussen, and V. J. Cristofalo. Oxygen-sensitive stages of the cell cycle of human diploid cells. Journal of Cell Biology, 78(2):390–??, August 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/78/2/390.

Bard:1975:BFD

[BH75a] J. B. Bard and E. D. Hay. The behavior of fibroblasts from the developing avian cornea. Morphology and movement in situ and in vitro. Journal of Cell Biology, 67(2):400–??, November 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/2/400.

Baseman:1975:DEH

[BH75b] J. B. Baseman and N. S. Hayes. Differential effect of hormones on macromolecular synthesis and mitosis in chick embryo cells. Journal of Cell Biology, 67(2):492–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/2/492.

Bard:1977:FCI

[BH77] J. B. Bard and K. Higginson. Fibroblast-collagen interactions in the formation of the secondary stroma of the chick cornea. Journal of Cell Biology, 74(3):816–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/3/816.

Burton:1978:EMS

[BH78] P. R. Burton and R. H. Himes. Electron microscope studies of pH effects on assembly of tubulin free of associated proteins. Delineation of substructure by tannic acid staining. Journal of Cell Biology, 77(1):120–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/1/120.

Basbaum:1979:MSS

[BH79] C. B. Basbaum and J. E. Heuser. Morphological studies of stimulated adrenergic axon varicosities in the mouse vas deferens. Journal of Cell Biology, 80(2):310–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/80/2/310.

Bereiter-Hahn:1979:QRC

[BHFT79] J. Bereiter-Hahn, C. H. Fox, and B. Thorell. Quantitative reflection contrast microscopy of living cells. Journal of Cell Biology, 82 (3):767–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/767.

Burton:1975:TAS

[BHP75] P. R. Burton, R. E. Hinkley, and G. B. Pierson. Tannic acidstained microtubules with 12, 13, and 15 protofilaments. Journal of Cell Biology, 65(1):227–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/1/227.

Besharse:1977:TRP

[BHR77] J. C. Besharse, J. G. Hollyfield, and M. E. Rayborn. Turnover of rod photoreceptor outer segments. II. Membrane addition and loss in relationship to light. Journal of Cell Biology, 75(2):507–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
2/507.

Berry:1975:FTF

[BJ75] S. J. Berry and E. Johnson. Formation of temporary flagellar structures during insect organogenesis. Journal of Cell Biology, 65(2):489–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/489.

Batzri:1975:IPV

[BK75] S. Batzri and E. D. Korn. Interaction of phospholipid vesicles with cells. Endocytosis and fusion as alternate mechanisms for the uptake of lipid-soluble and water-soluble molecules. Journal of Cell Biology, 66(3):621–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/3/621.

Blau:1978:SKF

[BK78] H. M. Blau and F. C. Kafatos. Secretory kinetics in the follicular cells of silkmoths during eggshell formation. Journal of Cell Biology, 78(1):131–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/131.

Briggs:1975:CDH

[BKK75] R. T. Briggs, M. L. Karnovsky, and M. J. Karnovsky. Cytochemical demonstration of hydrogen peroxide in polymorphonuclear leukocyte phagosomes. Journal of Cell Biology, 64(1):254–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
1/254.

Boulan:1978:SOGa

[BKS78] E. Rodriguez Boulan, G. Kreibich, and D. D. Sabatini. Spatial orientation of glycoproteins in membranes of rat liver rough microsomes. I. Localization of lectin-binding sites in microsomal membranes. Journal of Cell Biology, 78(3):874–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/874.

Bloodgood:1979:RIC

[BLB79] R. A. Bloodgood, E. M. Leffler, and A. T. Bojczuk. Reversible inhibition of Chlamydomonas flagellar surface motility. Journal of Cell Biology, 82(3):664–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/3/664.

Bloodgood:1977:MOA

[Blo77] R. A. Bloodgood. Motility occurring in association with the surface of the Chlamydomonas flagellum. Journal of Cell Biology, 75 (3):983–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/983.

Bloch:1979:DRA

[Blo79] R. J. Bloch. Dispersal and reformation of acetylcholine receptor clusters of cultured rat myotubes treated with inhibitors of energy metabolism. Journal of Cell Biology, 82(3):626–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/3/626.

Bergtrom:1976:FBS

[BLR76] G. Bergtrom, H. Laufer, and R. Rogers. Fat body: a site of hemoglobin synthesis in Chironomus thummi (diptera). Journal of Cell Biology, 69(2):264–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/264.

Beisson:1976:GAM

[BLTP^{+}76] J. Beisson, M. Lefort-Tran, M. Pouphile, M. Rossignol, and B. Satir. Genetic analysis of membrane differentiation in Paramecium. Freeze-fracture study of the trichocyst cycle in wild-type and mutant strains. Journal of Cell Biology, 69(1):126–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/1/126.

Brown:1976:OFA

[BMP76] D. L. Brown, A. Massalski, and R. Patenaude. Organization of the flagellar apparatus and associate cytoplasmic microtubules in the quadriflagellate alga Polytomella agilis. Journal of Cell Biology, 69(1):106–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/106.

Banjo:1976:PER

[BN76] A. O. Banjo and A. M. Nemeth. Proliferation of endoplasmic reticulum with its enzyme, UDP-glucuronyltransferase, in chick embryo liver during culture. Effects of phenobarbital. Journal of Cell Biology, 70(2):319–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/2/319.

Billen:1976:DRC

[BO76] D. Billen and A. C. Olson. DNA replication in Chinese hamster ovary cells made permeable to nucleotides by tween-80 treatment. Journal of Cell Biology, 69(3):732–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/3/732.

Berlin:1978:AUS

[BO78] R. D. Berlin and J. M. Oliver. Analogous ultrastructure and surface properties during capping and phagocytosis in leukocytes. Journal of Cell Biology, 77(3):789–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/789.

Bonner:1975:PMNa

[Bon75a] W. M. Bonner. Protein migration into nuclei. I. Frog oocyte nuclei in vivo accumulate microinjected histones, allow entry to small proteins, and exclude large proteins. Journal of Cell Biology, 64 (2):421–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/2/421.

Bonner:1975:PMNb

[Bon75b] W. M. Bonner. Protein migration into nuclei. II. Frog oocyte nuclei accumulate a class of microinjected oocyte nuclear proteins and exclude a class of microinjected oocyte cytoplasmic proteins. Journal of Cell Biology, 64(2):431–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/2/431.

Burchill:1978:MDG

[BOP^{+}78] B. R. Burchill, J. M. Oliver, C. B. Pearson, E. D. Leinbach, and R. D. Berlin. Microtubule dynamics and glutathione metabolism in phagocytizing human polymorphonuclear leukocytes. Journal of Cell Biology, 76(2):439–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/2/439.

Bourguignon:1979:BAL

[Bou79] L. Y. Bourguignon. Biochemical analysis of ligand-induced receptor patching and capping using a novel immunolactoperoxidase iodination technique. Journal of Cell Biology, 83(3):649–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
3/649.

Bourguignon:1976:IPT

[BP76] L. Y. Bourguignon and G. E. Palade. Incorporation of polypeptides into thylakoid membranes of Chlamydomonas reinhardtii. Cyclic variations. Journal of Cell Biology, 69(2):327–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/2/327.

Byers:1977:TSC

[BP77] H. R. Byers and K. R. Porter. Transformations in the structure of the cytoplasmic ground substance in erythrophores during pigment aggregation and dispersion. I. A study using whole-cell preparations in stereo high voltage electron microscopy. Journal of Cell Biology, 75(2):541–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/2/541.

Benecke:1979:LSD

[BP79] B. J. Benecke and S. Penman. Large species differences in the pattern of snPI RNA which can distinguish ape from human. Journal of Cell Biology, 80(3):778–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/3/778.

Bolender:1978:ISB

[BPL^{+}78] R. P. Bolender, D. Paumgartner, G. Losa, D. Muellener, and E. R. Weibel. Intergrated stereological and biochemical studies of hepatocytic membranes. I. Membrane recoveries in subcellular fractions. Journal of Cell Biology, 77(2):565–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/565.

Barker:1976:HSS

[BPN76] J. E. Barker, J. E. Pierce, and A. W. Nienhuis. Hemoglobin switching in sheep and goats. VI. Commitment of erythroid colony-forming cells to the synthesis of betaC globin. Journal of Cell Biology, 71(3):715–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/3/715.

Bourgeois:1978:QSL

[BPRC78] J. P. Bourgeois, J. L. Popot, A. Ryter, and J. P. Changeux. Quantitative studies on the localization of the cholinergic receptor protein in the normal and denervated electroplaque from Electrophorus electricus. Journal of Cell Biology, 79(1):200–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/200.

Broderson:1978:FSL

[BPS78] S. H. Broderson, D. L. Patton, and W. L. Stahl. Fine structural localization of potassium-stimulated rho nitrophenylphosphatase activity in denrites of the cerebral cortex. Journal of Cell Biology, 77(2):??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/R13.

Bloodgood:1976:IBT

[BR76a] R. A. Bloodgood and J. L. Rosenbaum. Initiation of brain tubulin assembly by a high molecular weight flagellar protein factor.

Journal of Cell Biology, 71(1):322–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/322.

Brown:1976:RCS

[BR76b] S. S. Brown and J. P. Revel. Reversibility of cell surface label rearrangement. Journal of Cell Biology, 68(3):629–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/3/629.

Berns:1977:CMA

[BR77] M. W. Berns and S. M. Richardson. Continuation of mitosis after selective laser microbeam destruction of the centriolar region. Journal of Cell Biology, 75(3):977–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/3/977.

Binder:1978:VAF

[BR78] L. I. Binder and J. L. Rosenbaum. The in vitro assembly of flagellar outer doublet tubulin. Journal of Cell Biology, 79(2):500–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
2/500.

Begg:1979:PRP

[BR79] D. A. Begg and L. I. Rebhun. pH regulates the polymerization of actin in the sea urchin egg cortex. Journal of Cell Biology, 83(1):241–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/1/241.

Berns:1977:RCR

[BRBM77] M. W. Berns, J. B. Rattner, S. Brenner, and S. Meredith. The role of the centriolar region in animal cell mitosis. A laser microbeam study. Journal of Cell Biology, 72(2):351–??, February 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/72/2/351.

Brokaw:1979:CIA

[Bro79] C. J. Brokaw. Calcium-induced asymmetrical beating of tritondemembranated sea urchin sperm flagella. Journal of Cell Biology, 82(2):401–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/2/401.

Begg:1978:VAF

[BRR78] D. A. Begg, R. Rodewald, and L. I. Rebhun. The visualization of actin filament polarity in thin sections. Evidence for the uniform polarity of membrane-associated filaments. Journal of Cell Biology, 79(3):846–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/3/846.

Baetens:1976:IGP

[BRS^{+}76] D. Baetens, C. Rufener, B. C. Srikant, R. Dobbs, R. Unger, and L. Orci. Identification of glucagon-producing cells (A cells) in dog gastric mucosa. Journal of Cell Biology, 69(2):455–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/2/455.

Buckley:1978:HML

[BRS78] I. K. Buckley, T. R. Raju, and M. Stewart. Heavy meromyosin labeling of intermediate filaments in cultured connective tissue cells. Journal of Cell Biology, 78(3):644–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/78/3/644.

Bruns:1976:SSP

[Bru76] R. R. Bruns. Supramolecular structure of polymorphic collagen fibrils. Journal of Cell Biology, 68(3):521–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/521.

Bouck:1978:SOC

[BRV78] G. B. Bouck, A. Rogalski, and A. Valaitis. Surface organization and composition of Euglena. II. Flagellar mastigonemes. Journal of Cell Biology, 77(3):805–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/3/805.

Bryan:1976:QAM

[Bry76] J. Bryan. A quantitative analysis of microtubule elongation. Journal of Cell Biology, 71(3):749–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/749.

Birchmeier:1977:MAI

[BS77a] W. Birchmeier and S. J. Singer. On the mechanism of ATPinduced shape changes in the human erythrocyte membranes: the role of ATP. Journal of Cell Biology, 73(3):647–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/3/647.

Brokaw:1977:MTD

[BS77b] C. J. Brokaw and T. F. Simonick. Motility of tritondemembranated sea urchin sperm flagella during digestion by trypsin. Journal of Cell Biology, 75(3):650–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/3/650.

Burgess:1977:PBA

[BS77c] D. R. Burgess and T. E. Schroeder. Polarized bundles of actin filaments within microvilli of fertilized sea urchin eggs. Journal of Cell Biology, 74(3):1032–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/3/1032.

Burwen:1977:FFS

[BS77d] S. J. Burwen and B. H. Satir. A freeze-fracture study of early membrane events during mast cell secretion. Journal of Cell Biology, 73(3):660–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/3/660.

Burwen:1977:PMF

[BS77e] S. J. Burwen and B. H. Satir. Plasma membrane folds on the mast cell surface and their relationship to secretory activity. Journal of Cell Biology, 74(3):690–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/3/690.

Berger:1978:RMD

[BS78a] J. D. Berger and H. J. Schmidt. Regulation of macronuclear DNA content in Paramecium tetraurelia. Journal of Cell Biology, 76 (1):116–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/1/116.

Burks:1978:EVS

[BS78b] D. J. Burks and P. J. Stambrook. Enrichment and visualization of small replication units from cultured mammalian cells. Journal of Cell Biology, 77(3):762–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/3/762.

Barker:1979:UDM

[BS79a] P. E. Barker and E. Stubblefield. Ultrastructure of double minutes from a human tumor cell line. Journal of Cell Biology, 83 (3):663–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/3/663.

Brown:1979:CIR

[BS79b] S. S. Brown and J. A. Spudich. Cytochalasin inhibits the rate of elongation of actin filament fragments. Journal of Cell Biology, 83 (3):657–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/3/657.

Brown:1979:NPA

[BS79c] S. S. Brown and J. A. Spudich. Nucleation of polar actin filament assembly by a positively charged surface. Journal of Cell Biology, 80(2):499–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/499.

Bergeron:1979:BUI

[BSHP79] J. J. Bergeron, R. Sikstrom, A. R. Hand, and B. I. Posner. Binding and uptake of 125I-insulin into rat liver hepatocytes and endothelium. An in vivo radioautographic study. Journal of Cell Biology, 80(2):427–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/427.

Burden:1979:ARR

[BSM79] S. J. Burden, P. B. Sargent, and U. J. McMahan. Acetylcholine receptors in regenerating muscle accumulate at original synaptic sites in the absence of the nerve. Journal of Cell Biology, 82 (2):412–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/2/412.

Blaurock:1976:SCE

[BSOS76] A. E. Blaurock, W. Stoeckenius, D. Oesterhelt, and G. L. Scherfhof. Structure of the cell envelope of Halobacterium halobium. Journal of Cell Biology, 71(1):1–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/1.

Boulan:1978:SOGb

[BSPK78] E. Rodriguez Boulan, D. D. Sabatini, B. N. Pereyra, and G. Kreibich. Spatial orientation of glycoproteins in membranes of rat liver rough microsomes. II. Transmembrane disposition and characterization of glycoproteins. Journal of Cell Biology, 78(3): 894–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/894.

Braun:1979:CLS

[BSU79] J. Braun, R. I. Sha’afi, and E. R. Unanue. Crosslinking by ligands to surface immunoglobulin triggers mobilization of intracellular ^{45}Ca^{2+} in B lymphocytes. Journal of Cell Biology, 82 (3):755–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/755.

Bauduin:1975:MSS

[BSVG75] H. Bauduin, C. Stock, D. Vincent, and J. F. Grenier. Microfilamentous system and secretion of enzyme in the exocrine pancreas. Effect of cytochalasin B. Journal of Cell Biology, 66(1):165–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/66/
1/165.

Beyer:1979:LEL

[BTB79] E. C. Beyer, K. T. Tokuyasu, and S. H. Barondes. Localization of an endogenous lectin in chicken liver, intestine, and pancreas. Journal of Cell Biology, 82(2):565–??, August 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/82/2/565.

Baumgartner:1978:LVP

[BTC78] B. Baumgartner, K. T. Tokuyasu, and M. J. Chrispeels. Localization of vicilin peptidohydrolase in the cotyledons of mung bean seedlings by immunofluorescence microscopy. Journal of Cell Biology, 79(1):10–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/1/10.

Bowman:1976:OBP

[BTDT76] C. L. Bowman, H. Tedeschi, B. J. DiDomenico, and F. D. Tung. Osmotic behavior and permeability of osmotically lysed mitochondria. Journal of Cell Biology, 70(2):348–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/70/2/348.

Becker:1978:ARP

[BTS^{+}78] E. L. Becker, V. Talley, H. J. Showell, P. H. Naccache, and R. I. Sha’afi. Activation of the rabbit polymorphonuclear leukocyte membrane “Na+, K+”-ATPase by chemotactic factor. Journal of Cell Biology, 77(2):329–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/329.

Burnside:1978:TAT

[Bur78] B. Burnside. Thin (actin) and thick (myosinlike) filaments in cone contraction in the teleost retina. Journal of Cell Biology, 78(1):227–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/227.

Bal:1976:SLC

[BVBM76] A. K. Bal, D. P. Verma, H. Byrne, and G. A. Maclachlan. Subcellular localization of cellulases in auxin-treated pea. Journal of Cell Biology, 69(1):97–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/1/97.

Bretscher:1978:LAM

[BW78] A. Bretscher and K. Weber. Localization of actin and microfilament-associated proteins in the microvilli and terminal web of the intestinal brush border by immunofluorescence microscopy. Journal of Cell Biology, 79(3):839–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/3/839.

Burry:1979:CLP

[BW79] R. W. Burry and J. G. Wood. Contributions of lipids and proteins to the surface charge of membranes. An electron microscopy study with cationized and anionized ferritin. Journal of Cell Biology, 82 (3):726–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/726.

Bernaert:1977:EIU

[BWDP77] D. Bernaert, J. C. Wanson, P. Drochmans, and A. Popowski. Effect of insulin on ultrastructure and glycogenesis in primary cultures of adult rat hepatocytes. Journal of Cell Biology, 74 (3):878–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/878.

Chen:1978:AGC

[CAS78] L. L. Chen, J. C. Adams, and R. M. Steinman. Anatomy of germinal centers in mouse spleen, with special reference to “follicular dendritic cells”. Journal of Cell Biology, 77(1):148–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/1/148.

Cave:1975:ARD

[Cav75] M. D. Cave. Absence of ribosomal DNA amplification in the meroistic (telotrophic) ovary of the large milkweed bug Oncopeltus fasciatus (Dallas) (Hemiptera: Lygaeidae). Journal of Cell Biology, 66(3):461–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/3/461.

Cave:1976:ARA

[Cav76] M. D. Cave. Absence of rDNA amplification in the uninucleolate oocyte of the cockroach Blattella germanica (Oorthoptera: Blattidae). Journal of Cell Biology, 71(1):49–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/49.

Cameron:1975:ECP

[CB75a] I. L. Cameron and N. C. Bols. Effect of cell population density on G2 arrest in Tetrahymena. Journal of Cell Biology, 67(3):518–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/518.

Cantin:1975:LCC

[CB75b] M. Cantin and S. Benchimol. Localization and characterization of carbohydrates in adrenal medullary cells. Journal of Cell Biology, 65(2):463–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/463.

Carlier:1975:EMA

[CB75c] H. Carlier and J. Bezard. Electron microscope autoradiographic study of intestinal absorption of decanoic and octanoic acids in the rat. Journal of Cell Biology, 65(2):383–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/2/383.

Cooper:1976:ASS

[CB76] G. W. Cooper and J. M. Bedford. Asymmetry of spermiation and sperm surface charge patterns over the giant acrosome in the musk shrew Suncus murinus. Journal of Cell Biology, 69(2):415– ??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
2/415.

Caruthers:1977:ICU

[CB77] J. S. Caruthers and M. A. Bonneville. Isolation and characterization of the urothelial lumenal plasma membrane. Journal of Cell Biology, 73(2):382–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/2/382.

Coleman:1978:EBP

[CB78] R. Coleman and R. M. Bell. Evidence that biosynthesis of phosphatidylethanolamine, phosphatidylcholine, and triacylglycerol occurs on the cytoplasmic side of microsomal vesicles. Journal of Cell Biology, 76(1):245–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/245.

Caron:1979:IMP

[CB79] J. M. Caron and R. D. Berlin. Interaction of microtubule proteins with phospholipid vesicles. Journal of Cell Biology, 81(3):665–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/665.

Cohn:1977:BLE

[CBB77] R. H. Cohn, S. D. Banerjee, and M. R. Bernfield. Basal lamina of embryonic salivary epithelia. Nature of glycosaminoglycan and organization of extracellular materials. Journal of Cell Biology, 73(2):464–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/2/464.

Cohen:1977:SPD

[CBBS77] R. S. Cohen, F. Blomberg, K. Berzins, and P. Siekevitz. The structure of postsynaptic densities isolated from dog cerebral cortex: I. overall morphology and protein composition. Journal of Cell Biology, 74(1):181–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/181.

Collins:1975:SMF

[CBC75] T. R. Collins, J. C. Bartholomew, and M. Calvin. A simple method for freeze-fracture of monolayer cultures. Journal of Cell Biology, 67(3):904–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/904.

Catino:1979:SLD

[CBDY79] J. J. Catino, H. Busch, Y. Daskal, and L. C. Yeoman. Subcellular localization of DNA-binding protein BA by immunofluorescence and immunoelectron microscopy. Journal of Cell Biology, 83(2):462–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/2/462.

Crouch:1978:AFF

[CBH^{+}78] E. Crouch, G. Balian, K. Holbrook, D. Duksin, and P. Bornstein. Amniotic fluid fibronectin. Characterization and synthesis by cells in culture. Journal of Cell Biology, 78(3):701–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/701.

Cartaud:1978:EAB

[CBM78] J. Cartaud, S. Bon, and J. Massoulié. Electrophorus acetylcholinesterase. Biochemical and electron microscope characterization of low ionic strength aggregates. Journal of Cell Biology, 77(2):315–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/315.

Chua:1976:PVR

[CBSP76] N. H. Chua, G. Blobel, P. Siekevitz, and G. E. Palade. Periodic variations in the ratio of free to thylakoid-bound chloroplast ribosomes during the cell cycle of Chlamydomonas reinhardtii. Journal of Cell Biology, 71(2):497–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/497.

Carpenter:1976:LHE

[CC76] G. Carpenter and S. Cohen. 125i-labeled human epidermal growth factor. Binding, internalization, and degradation in human fibroblasts. Journal of Cell Biology, 71(1):159–??, October 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/1/159.

Cohn:1976:RTC

[CCB76] R. H. Cohn, J. J. Cassiman, and M. R. Bernfield. Relationship of transformation, cell density, and growth control to the cellular distribution of newly synthesized glycosaminoglycan. Journal of Cell Biology, 71(1):280–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/1/280.

Carlisle:1978:UCC

[CCH78] K. S. Carlisle, C. S. Chew, and S. J. Hersey. Ultrastructural changes and cyclic AMP in frog oxyntic cells. Journal of Cell Biology, 76(1):31–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/76/1/31.

Costantino-Ceccarini:1978:FCH

[CCNAN78] E. Costantino-Ceccarini, P. M. Novikoff, P. H. Atkinson, and A. B. Novikoff. Further characterization of HeLa S3 plasma membrane ghosts. Journal of Cell Biology, 77(2):448–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/448.

Cachon:1977:MGI

[CCTT77] J. Cachon, M. Cachon, L. G. Tilney, and M. S. Tilney. Movement generated by interactions between the dense material at the ends of microtubles and non-actin-containing microfilaments in Sticholonche zanclea. Journal of Cell Biology, 72(2):314–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
2/314.

Carraway:1979:ECP

[CDH^{+}79] K. L. Carraway, R. C. Doss, J. W. Huggins, R. W. Chesnut, and C. A. Carraway. Effects of cytoskeletal perturbant drugs on ecto 5’-nucleotidase, a concanavalin A receptor. Journal of Cell Biology, 83(3):529–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/83/3/529.

Coons:1977:QNJ

[CE77] L. W. Coons and L. L. Espey. Quantitation of nexus junctions in the granulosa cell layer of rabbit ovarian follicles during ovulation. Journal of Cell Biology, 74(1):321–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/1/321.

Cheng:1976:PACa

[CF76a] H. Cheng and M. G. Farquhar. Presence of adenylate cyclase activity in Golgi and other fractions from rat liver. I. Biochemical determination. Journal of Cell Biology, 70(3):660–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/3/660.

Cheng:1976:PACb

[CF76b] H. Cheng and M. G. Farquhar. Presence of adenylate cyclase activity in Golgi and other fractions from rat liver. II. Cytochemical localization within Golgi and ER membranes. Journal of Cell Biology, 70(3):671–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/671.

Carbonetto:1979:SIP

[CF79] S. Carbonetto and D. M. Fambrough. Synthesis, insertion into the plasma membrane, and turnover of alpha-bungarotoxin receptors in chick sympathetic neurons. Journal of Cell Biology, 81(3):555– ??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/555.

Chen:1978:DHP

[CFAS78] L. L. Chen, A. M. Frank, J. C. Adams, and R. M. Steinman. Distribution of horseradish peroxidase (HRP)–anti-HRP immune complexes in mouse spleen with special reference to follicular dendritic cells. Journal of Cell Biology, 79(1):184–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/184.

Chua:1977:SSP

[CG77] N. H. Chua and N. W. Gillham. The sites of synthesis of the principal thylakoid membrane polypeptides in Chlamydomonas reinhardtii. Journal of Cell Biology, 74(2):441–??, August 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/2/441.

Clark:1978:FCS

[CG78] J. I. Clark and D. Garland. Fluorescein colchicine. Synthesis, purification, and biological activity. Journal of Cell Biology, 76 (3):619–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/3/619.

Cramer:1979:LSC

[CG79] E. B. Cramer and J. I. Gallin. Localization of submembranous cations to the leading end of human neutrophils during chemotaxis. Journal of Cell Biology, 82(2):369–??, August 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/82/2/369.

Ceccarelli:1979:FFSa

[CGH79a] B. Ceccarelli, F. Grohovaz, and W. P. Hurlbut. Freeze-fracture studies of frog neuromuscular junctions during intense release of neurotransmitter. I. Effects of black widow spider venom and Ca^{2+}-free solutions on the structure of the active zone. Journal of Cell Biology, 81(1):163–??, April 1979. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/1/163.

Ceccarelli:1979:FFSb

[CGH79b] B. Ceccarelli, F. Grohovaz, and W. P. Hurlbut. Freeze-fracture studies of frog neuromuscular junctions during intense release of neurotransmitter. II. Effects of electrical stimulation and high potassium. Journal of Cell Biology, 81(1):178–??, April 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/81/1/178.

Campbell:1975:UFM

[CGK75] G. L. Campbell, L. T. Goldstein, and B. B. Knowles. The use of flow microfluorimetry in the analysis of the phenotype expression of mouse histocompatibility antigens. Journal of Cell Biology, 64(3):719–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/719.

Caspar:1977:GJS

[CGMP77] D. L. Caspar, D. A. Goodenough, L. Makowski, and W. C. Phillips. Gap junction structures. I. Correlated electron microscopy and X-ray diffraction. Journal of Cell Biology, 74 (2):605–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/2/605.

Connolly:1979:RSC

[CGVR79] J. L. Connolly, L. A. Greene, R. R. Viscarello, and W. D. Riley. Rapid, sequential changes in surface morphology of PC12 pheochromocytoma cells in response to nerve growth factor. Journal of Cell Biology, 82(3):820–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/3/820.

Croop:1975:RMF

[CH75] J. Croop and H. Holtzer. Response of myogenic and fibrogenic cells to cytochalasin B and to colcemid. I. Light microscope observations. Journal of Cell Biology, 65(2):271–??, May 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/65/2/271.

Comings:1976:NPI

[CH76] D. E. Comings and D. C. Harris. Nuclear proteins. II. Similarity of nonhistone proteins in nuclear sap and chromatin, and essential absence of contractile proteins from mouse liver nuclei. Journal of Cell Biology, 70(2):440–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/2/440.

Chandler:1979:MFD

[CH79a] D. E. Chandler and J. Heuser. Membrane fusion during secretion: cortical granule exocytosis in sex urchin eggs as studied by quickfreezing and freeze-fracture. Journal of Cell Biology, 83(1):91–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/91.

Chlapowski:1979:GDT

[CH79b] F. J. Chlapowski and L. Haynes. The growth and differentiation of transitional epithelium in vitro. Journal of Cell Biology, 83 (3):605–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/3/605.

Chen:1979:ISD

[Che79] W. T. Chen. Induction of spreading during fibroblast movement. Journal of Cell Biology, 81(3):684–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/3/684.

Coll:1978:SHA

[CI78] J. Coll and V. M. Ingram. Stimulation of heme accumulation and erythroid colony formation in cultures of chick bone marrow cells by chicken plasma. Journal of Cell Biology, 76(1):184–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
1/184.

Castle:1975:SGR

[CJP75] J. D. Castle, J. D. Jamieson, and G. E. Palade. Secretion granules of the rabbit parotid gland. Isolation, subfractionation, and characterization of the membrane and content subfractions. Journal of Cell Biology, 64(1):182–??, January 1975. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/182.

Connolly:1978:VCB

[CK78] J. A. Connolly and V. I. Kalnins. Visualization of centrioles and basal bodies by fluorescent staining with nonimmune rabbit sera. Journal of Cell Biology, 79(2):526–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/2/526.

Cox:1976:SCC

[CKBD76] R. P. Cox, M. R. Krauss, M. E. Balis, and J. Dancis. Studies on cell communication with enucleated human fibroblasts. Journal of Cell Biology, 71(3):693–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/3/693.

Connolly:1978:ILH

[CKCK78] J. A. Connolly, V. I. Kalnins, D. W. Cleveland, and M. W. Kirschner. Intracellular localization of the high molecular weight microtubule accessory protein by indirect immunofluorescence. Journal of Cell Biology, 76(3):781–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/76/3/781.

Cohen:1977:MIP

[CKJB77] C. M. Cohen, D. I. Kalish, B. S. Jacobson, and D. Branton. Membrane isolation on polylysine-coated beads. Plasma membrane from HeLa cells. Journal of Cell Biology, 75(1):119–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/119.

Clark:1976:CSN

[Cla76] A. W. Clark. Changes in the structure of neuromuscular junctions caused by variations in osmotic pressure. Journal of Cell Biology, 69(3):521–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/521.

Cleffmann:1975:ADP

[Cle75] G. Cleffmann. Amount of DNA produced during extra S phases in Tetrahymena. Journal of Cell Biology, 66(1):204–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/1/204.

Cande:1977:CDA

[CLM77] W. Z. Cande, E. Lazarides, and J. R. McIntosh. A comparison of the distribution of actin and tubulin in the mammalian mitotic spindle as seen by indirect immunofluorescence. Journal of Cell Biology, 72(3):552–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/72/3/552.

Cardelli:1976:DAM

[CLP76] J. Cardelli, B. Long, and H. C. Pitot. Direct association of messenger RNA labeled in the presence of fluoroorotate with membranes of the endoplasmic reticulum in rat liver. Journal of Cell Biology, 70(1):47–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/1/47.

Clemente:1975:CPS

[CM75] F. Clemente and J. Meldolesi. Calcium and pancreatic secretion. I. Subcellular distribution of calcium and magnesium in the exocrine pancreas of the guinea pig. Journal of Cell Biology, 65(1):88–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/65/1/88.

Chatterjee:1977:CBD

[CM77a] R. N. Chatterjee and A. S. Mukherjee. Chromosomal basis of dosage compensation in Drosophila: IX. cellular autonomy of the faster replication of the X chromosome in haplo-x cells of Drosophila melanogaster and synchronous initiation. Journal of Cell Biology, 74(1):168–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/74/1/168.

Craig:1977:ASM

[CM77b] R. Craig and J. Megerman. Assembly of smooth muscle myosin into side-polar filaments. Journal of Cell Biology, 75(3):990–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/990.

Crane:1977:PPS

[CM77c] L. J. Crane and D. L. Miller. Plasma protein synthesis by isolated rat hepatocytes. Journal of Cell Biology, 72(1):11–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/72/1/11.

Clark:1978:AXO

[CM78] T. G. Clark and R. W. Merriam. Actin in Xenopus oocytes. Journal of Cell Biology, 77(2):427–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/427.

Chua:1975:CPC

[CMB75] N. H. Chua, K. Matlin, and P. Bennoun. A chlorophyll-protein complex lacking in photosystem I mutants of Chlamydomonas reinhardtii. Journal of Cell Biology, 67(2):361–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/2/361.

Chu:1977:EDI

[CMC77] L. L. Chu, R. R. MacGregor, and D. V. Cohn. Energy-dependent intracellular translocation of proparathormone. Journal of Cell Biology, 72(1):1–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/72/1/1.

Chinkers:1979:RIM

[CMC79] M. Chinkers, J. A. McKanna, and S. Cohen. Rapid induction of morphological changes in human carcinoma cells A-431 by epidermal growth factors. Journal of Cell Biology, 83(1):260–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/260.

Chan:1976:CNM

[CO76] L. N. Chan and J. M. Oliver. Changes in number, mobility, and topographical distribution of lectin receptors during maturation of chick erythroid cells. Journal of Cell Biology, 69(3):647–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
3/647.

Cohen:1978:OMB

[Coh78] W. D. Cohen. Observations of the marginal band system of nucleated erythrocytes. Journal of Cell Biology, 78(1):260–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/1/260.

Clark:1976:CTC

[COK76] R. A. Clark, I. Olsson, and S. J. Klebanoff. Cytotoxicity for tumor cells of cationic proteins from human neutrophil granules. Journal of Cell Biology, 70(3):719–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/719.

Coleman:1979:UFD

[Col79] A. W. Coleman. Use of the fluorochrome 4’6-diamidino-2-phenylindole in genetic and developmental studies of chloroplast DNA. Journal of Cell Biology, 82(1):299–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/1/299.

Connell:1978:FFL

[Con78] C. J. Connell. A freeze-fracture and lanthanum tracer study of the complex junction between Sertoli cells of the canine testis. Journal of Cell Biology, 76(1):57–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/57.

Condeelis:1979:ICC

[Con79] J. Condeelis. Isolation of concanavalin a caps during various stages of formation and their association with actin and myosin. Journal of Cell Biology, 80(3):751–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/3/751.

Cooke:1976:FCV

[Coo76] P. Cooke. A filamentous cytoskeleton in vertebrate smooth muscle fibers. Journal of Cell Biology, 68(3):539–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/539.

Chowrashi:1977:LMP

[CP77a] P. K. Chowrashi and F. A. Pepe. Light meromyosin paracrystal formation. Journal of Cell Biology, 74(1):136–??, July 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/1/136.

Chun:1977:RNGa

[CP77b] L. L. Chun and P. H. Patterson. Role of nerve growth factor in the development of rat sympathetic neurons in vitro. I. Survival, growth, and differentiation of catecholamine production. Journal of Cell Biology, 75(3):694–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/3/694.

Chun:1977:RNGb

[CP77c] L. L. Chun and P. H. Patterson. Role of nerve growth factor in the development of rat sympathetic neurons in vitro. II. Developmental studies. Journal of Cell Biology, 75(3):705–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/3/705.

Chun:1977:RNGc

[CP77d] L. L. Chun and P. H. Patterson. Role of nerve growth factor in the development of rat sympathetic neurons in vitro. III. Effect on acetylcholine production. Journal of Cell Biology, 75(3):712–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/712.

Castle:1978:SGR

[CP78] J. D. Castle and G. E. Palade. Secretion granules of the rabbit parotid. Selective removal of secretory contaminants from granule membranes. Journal of Cell Biology, 76(2):323–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/323.

Cohen:1979:CIP

[CP79a] S. A. Cohen and D. W. Pumplin. Clusters of intramembrane particles associated with binding sites for alpha-bungarotoxin in cultured chick myotubes. Journal of Cell Biology, 82(2):494–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
2/494.

Craig:1979:AAL

[CP79b] S. W. Craig and J. V. Pardo. alpha-actinin localization in the junctional complex of intestinal epithelial cells. Journal of Cell Biology, 80(1):203–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/1/203.

Carpentier:1977:MCA

[CPO77] J. Carpentier, A. Perrelet, and L. Orci. Morphological changes of the adipose cell plasma membrane during lipolysis. Journal of Cell Biology, 72(1):104–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/1/104.

Crawford:1979:CPR

[Cra79] B. Crawford. Cloned pigmented retinal epiehtlium. The role of microfilaments in the differentiation of cell shape. Journal of Cell Biology, 81(2):301–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/2/301.

Culp:1978:TFD

[CRBH78] L. A. Culp, B. J. Rollins, J. Buniel, and S. Hitri. Two functionally distinct pools of glycosaminoglycan in the substrate adhesion site of murine cells. Journal of Cell Biology, 79(3):788–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/3/788.

Cereijido:1978:PMF

[CRD^{+}78] M. Cereijido, E. S. Robbins, W. J. Dolan, C. A. Rotunno, and D. D. Sabatini. Polarized monolayers formed by epithelial cells on a permeable and translucent support. Journal of Cell Biology, 77(3):853–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/3/853.

Chi:1975:SMH

[CRSH75] J. C. Chi, N. Rubinstein, K. Strahs, and H. Holtzer. Synthesis of myosin heavy and light chains in muscle cultures. Journal of Cell Biology, 67(3):523–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/523.

Cohen:1978:FPD

[CS78] R. S. Cohen and P. Siekevitz. Form of the postsynaptic density. A serial section study. Journal of Cell Biology, 78(1):36–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/1/36.

Chua:1979:TPM

[CS79] N. H. Chua and G. W. Schmidt. Transport of proteins into mitochondria and chloroplasts. Journal of Cell Biology, 81(3):461–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/461.

Cameron:1977:CEM

[CSHS77] I. L. Cameron, R. L. Sparks, K. L. Horn, and N. R. Smith. Concentration of elements in mitotic chromatin as measured by Xray microanalysis. Journal of Cell Biology, 73(1):193–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/1/193.

Cameron:1979:ECC

[CSP79] I. L. Cameron, N. K. Smith, and T. B. Pool. Element concentration changes in mitotically active and postmitotic enterocytes. An X-ray microanalysis study. Journal of Cell Biology, 80(2):444–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
2/444.

Collard:1976:SMA

[CT76] J. G. Collard and J. H. Temmink. Surface morphology and agglutinability with concanavalin A in normal and transformed murine fibroblasts. Journal of Cell Biology, 68(1):101–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/1/101.

Condeelis:1977:CBA

[CT77] J. S. Condeelis and D. L. Taylor. The contractile basis of amoeboid movement: V. The control of gelation, solation, and contraction in extracts from Dictyostelium discoideum. Journal of Cell Biology, 74(3):901–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/3/901.

Chalfie:1979:ONM

[CT79] M. Chalfie and J. N. Thomson. Organization of neuronal microtubules in the nematode Caenorhabditis elegans. Journal of Cell Biology, 82(1):278–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/82/1/278.

Chan:1976:RDR

[CW76] A. C. Chan and I. G. Walker. Reduced DNA repair during differentiation of a myogenic cell line. Journal of Cell Biology, 70 (3):685–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/685.

Chan:1977:LDR

[CW77] A. C. Chan and I. G. Walker. Loss of DNA repair capacity during successive subcultures of primary rat fibroblasts. Journal of Cell Biology, 74(2):365–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/365.

Cande:1978:CML

[CW78a] W. Z. Cande and S. M. Wolniak. Chromosome movement in lysed mitotic cells is inhibited by vanadate. Journal of Cell Biology, 79 (2):573–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/573.

Chandler:1978:IDCa

[CW78b] D. E. Chandler and J. A. Williams. Intracellular divalent cation release in pancreatic acinar cells during stimulus-secretion coupling. I. Use of chlorotetracycline as fluorescent probe. Journal of Cell Biology, 76(2):371–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/2/371.

Chandler:1978:IDCb

[CW78c] D. E. Chandler and J. A. Williams. Intracellular divalent cation release in pancreatic acinar cells during stimulus-secretion coupling. II. Subcellular localization of the fluorescent probe chlorotetracycline. Journal of Cell Biology, 76(2):386–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/386.

Cortese:1978:CSS

[CW78d] F. Cortese and J. Wolf. Cytochalasin-stimulated steroidogenesis from high density lipoproteins. Journal of Cell Biology, 77(2):507– ??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/77/
2/507.

Daneholt:1977:LSP

[DAF77] B. Daneholt, K. Anderson, and M. Fagerlind. Large-sized polysomes in Chironomus tentans salivary glands and their relation to Balbiani ring 75S RNA. Journal of Cell Biology, 73(1):149– ??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
1/149.

Damsky:1976:EIC

[Dam76] C. H. Damsky. Environmentally induced changes in mitochondria and endoplasmic reticulum of Saccharomyces carlsbergensis yeast. Journal of Cell Biology, 71(1):123–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/123.

DiPasquale:1975:LCR

[DB75] A. DiPasquale and P. B. Bell. Letter: Comments on reported observations of cells spreading on the upper surfaces of other cells in culture. Journal of Cell Biology, 66(1):216–??, July 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/1/216.

Das:1976:MCC

[DB76a] N. K. Das and C. Barker. Mitotic chromosome condensation in the sperm nucleus during post fertilization maturation division in urechis eggs. Journal of Cell Biology, 68(1):155–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/1/155.

Dwyer:1976:MPI

[DB76b] N. Dwyer and G. Blobel. A modified procedure for the isolation of a pore complex-lamina fraction from rat liver nuclei. Journal of Cell Biology, 70(3):581–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/581.

Dolnick:1979:CDR

[DBB^{+}79] B. J. Dolnick, R. J. Berenson, J. R. Bertino, R. J. Kaufman, J. H. Nunberg, and R. T. Schimke. Correlation of dihydrofolate reductase elevation with gene amplification in a homogeneously staining chromosomal region in L5178Y cells. Journal of Cell Biology, 83 (2):394–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/2/394.

delCastillo:1978:NME

[dCdM78] J. del Castillo and G. Escalona de Motta. A new method for excitation-contraction uncoupling in frog skeletal muscle. Journal of Cell Biology, 78(3):782–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/782.

Draper:1979:BGC

[DCEO^{+}79] R. K. Draper, D. Chin, D. Eurey-Owens, I. E. Scheffler, and M. I. Simon. Biochemical and genetic characterization of three hamster cell mutants resistant to diphtheria toxin. Journal of Cell Biology, 83(1):116–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/1/116.

Dubois-Dalcq:1975:SCM

[DDR75] M. Dubois-Dalcq and T. S. Reese. Structural changes in the membrane of vero cells infected with a paramyxovirus. Journal of Cell Biology, 67(3):551–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/551.

Devreotes:1979:CARa

[DDS79] P. N. Devreotes, P. L. Derstine, and T. L. Steck. Cyclic 3’,5’ AMP relay in Dictyostelium discoideum. I. A technique to monitor responses to controlled stimuli. Journal of Cell Biology, 80 (2):291–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/291.

Detering:1977:ICP

[DDSL77] N. K. Detering, G. L. Decker, E. D. Schmell, and W. J. Lennarz. Isolation and characterization of plasma membrane-associated cortical granules from sea urchin eggs. Journal of Cell Biology, 75 (3):899–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/899.

Decker:1976:HRG

[Dec76] R. S. Decker. Hormonal regulation of gap junction differentiation. Journal of Cell Biology, 69(3):669–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/3/669.

Devreotes:1975:ART

[DF75] P. N. Devreotes and D. M. Fambrough. Acetylcholine receptor turnover in membranes of developing muscle fibers. Journal of Cell Biology, 65(2):335–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/2/335.

Davidse:1977:DBM

[DF77] L. C. Davidse and W. Flach. Differential binding of methyl benzimidazol-2-yl carbamate to fungal tubulin as a mechanism of resistance to this antimitotic agent in mutant strains of Aspergillus nidulans. Journal of Cell Biology, 72(1):174–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/1/174.

Drochmans:1978:SBC

[DFW^{+}78] P. Drochmans, C. Freudenstein, J. C. Wanson, L. Laurent, T. W. Keenan, J. Stadler, R. Leloup, and W. W. Franke. Structure and biochemical composition of desmosomes and tonofilaments isolated from calf muzzle epidermis. Journal of Cell Biology, 79 (2):427–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/427.

Dentler:1975:ULH

[DGR75] W. L. Dentler, S. Granett, and J. L. Rosenbaum. Ultrastructural localization of the high molecular weight proteins associated with in vitro-assembled brain microtubules. Journal of Cell Biology, 65(1):237–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/1/237.

Dockray:1975:CSD

[DH75] G. J. Dockray and C. R. Hopkins. Caerulein secretion by dermal glands in Xenopus laevis. Journal of Cell Biology, 64(3):724–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
3/724.

Dashe:1976:CNU

[DH76] S. W. Dashe and S. H. Howell. Cautionary note on the use of [methyl-3H]thymidine to measure rates of chloroplast DNA synthesis in Chlamydomonas reinhardtii. Journal of Cell Biology, 69(1):215–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/215.

Dippell:1976:ENP

[Dip76] R. V. Dippell. Effects of nuclease and protease digestion on the ultrastructure of Paramecium basal bodies. Journal of Cell Biology, 69(3):622–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/622.

Deguchi:1977:USP

[DJM77] N. Deguchi, P. L. Jorgensen, and A. B. Maunsbach. Ultrastructure of the sodium pump: Comparison of thin sectioning, negative staining, and freeze-fracture of purified, membrane-bound (Na^{+}, K^{+})-ATPase. Journal of Cell Biology, 75(3):619–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/3/619.

Davis:1978:EML

[DK78a] R. Davis and G. B. Koelle. Electron microscope localization of acetylcholinesterase and butyrylcholinesterase in the superior cervical ganglion of the cat. I. Normal ganglion. Journal of Cell Biology, 78(3):785–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/3/785.

Dute:1978:UPR

[DK78b] R. Dute and C. Kung. Ultrastructure of the proximal region of somatic cilia in Paramecium tetraurelia. Journal of Cell Biology, 78(2):451–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/2/451.

Decker:1979:SBF

[DL79] G. L. Decker and W. J. Lennarz. Sperm binding and fertilization envelope formation in a cell surface complex isolated from sea urchin eggs. Journal of Cell Biology, 81(1):92–??, April 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/81/1/92.

DeBruyn:1979:CRD

[DM79] P. P. De Bruyn and S. Michelson. Changes in the random distribution of sialic acid at the surface of the myeloid sinusoidal endothelium resulting from the presence of diaphragmed fenestrae. Journal of Cell Biology, 82(3):708–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/3/708.

DeBruyn:1978:NDS

[DMB78] P. P. De Bruyn, S. Michelson, and R. P. Becker. Nonrandom distribution of sialic acid over the cell surface of bristle-coated endocytic vesicles of the sinusoidal endothelium cells. Journal of Cell Biology, 78(2):379–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/379.

deMenezes:1978:FFO

[dMdS78] A. Peixoto de Menezes and P. Pinto da Silva. Freeze-fracture observations of the lactating rat mammary gland. Membrane events during milk fat secretion. Journal of Cell Biology, 76(3):767–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
3/767.

Den:1975:ICW

[DMKR75] H. Den, D. A. Malinzak, H. J. Keating, and A. Rosenberg. Influence of Concanavalin A, wheat germ agglutinin, and soybean agglutinin on the fusion of myoblasts in vitro. Journal of Cell Biology, 67(3):826–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/67/3/826.

DiPasquale:1976:MAC

[DMMW76] A. M. DiPasquale, J. McGuire, G. Moellmann, and S. J. Wasserman. Microtubule assembly in cultivated Greene melanoma cells is stimulated by dibutyryl adenosine 3’: 5’-cyclic monophosphate or cholera toxin. Journal of Cell Biology, 71(3):735–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/3/735.

Denk:1976:HOI

[DMS^{+}76] H. Denk, P. W. Moldeus, R. A. Schulz, J. B. Schenkman, S. R. Keyes, and D. L. Cinti. Hepatic organelle interaction. IV. Mechanism of succinate enhancement of formaldehyde accumulation from endoplasmic reticulum N-dealkylations. Journal of Cell Biology, 69(3):589–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/589.

DAgostino:1976:SCS

[DN76] M. A. D’Agostino and M. M. Nass. Specific changes in the synthesis of mitochondrial DNA in chick embryo fibroblasts transformed by Rous sarcoma viruses. Journal of Cell Biology, 71(3):781–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
3/781.

dePetris:1975:CRI

[dP75] S. de Petris. Concanavalin A receptors, immunoglobulins, and theta antigen of the lymphocyte surface. Interactions with concanavalin A and with Cytoplasmic structures. Journal of Cell Biology, 65(1):123–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/65/1/123.

dePetris:1978:NDC

[dP78] S. de Petris. Nonuniform distribution of concanavalin–A receptors and surface antigens on uropod-forming thymocytes. Journal of Cell Biology, 79(1):235–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/1/235.

DeCamilli:1976:DCL

[DPM76] P. De Camilli, D. Peluchetti, and J. Meldolesi. Dynamic changes of the luminal plasmalemma in stimulated parotid acinar cells. A freeze-fracture study. Journal of Cell Biology, 70(1):59–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/1/59.

Durham:1976:CCP

[DR76] A. C. Durham and E. B. Ridgway. Control of chemotaxis in Physarum polycephalum. Journal of Cell Biology, 69(1):218–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
1/218.

DeChastellier:1977:CCS

[DR77a] C. De Chastellier and A. Ryter. Changes of the cell surface and of the digestive apparatus of Dictyostelium discoideum during the starvation period triggering aggregation. Journal of Cell Biology, 75(1):218–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/218.

Dentler:1977:FES

[DR77b] W. L. Dentler and J. L. Rosenbaum. Flagellar elongation and shortening in Chlamydomonas. III. structures attached to the tips of flagellar microtubules and their relationship to the directionality of flagellar microtubule assembly. Journal of Cell Biology, 74 (3):747–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/747.

Davies:1978:MPC

[DR78] P. F. Davies and R. Ross. Mediation of pinocytosis in cultured arterial smooth muscle and endothelial cells by platelet-derived growth factor. Journal of Cell Biology, 79(3):663–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/3/663.

Duncan:1975:ESE

[DS75] J. L. Duncan and R. Schlegel. Effect of streptolysin O on erythrocyte membranes, liposomes, and lipid dispersions. A proteincholesterol interaction. Journal of Cell Biology, 67(1):160–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
1/160.

Davies:1977:PHB

[DS77] W. A. Davies and T. P. Stossel. Peripheral hyaline blebs (podosomes) of macrophages. Journal of Cell Biology, 75(3):941–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/941.

Devreotes:1979:CARb

[DS79] P. N. Devreotes and T. L. Steck. Cyclic 3’,5’ AMP relay in Dictyostelium discoideum. II. Requirements for the initiation and termination of the response. Journal of Cell Biology, 80(2):300–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
2/300.

daSilva:1977:MFD

[dSN77] P. Pinto da Silva and M. L. Nogueira. Membrane fusion during secretion. A hypothesis based on electron microscope observation of Phytophthora Palmivora zoospores during encystment. Journal of Cell Biology, 73(1):161–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/1/161.

Dessau:1978:SED

[DST^{+}78] W. Dessau, J. Sasse, R. Timpl, F. Jilek, and K. von der Mark. Synthesis and extracellular deposition of fibronectin in chondrocyte cultures. Response to the removal of extracellular cartilage matrix. Journal of Cell Biology, 79(2):342–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/2/342.

Damsky:1977:TRA

[DSTW77] C. H. Damsky, J. B. Sheffield, G. P. Tuszynski, and L. Warren. Is there a role for actin in virus budding? Journal of Cell Biology, 75 (2):593–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/593.

Darzynkiewicz:1976:DDS

[DTSM76] Z. Darzynkiewicz, F. Traganos, T. Sharpless, and M. R. Melamed. DNA denaturation in situ. Effect of divalent cations and alcohols. Journal of Cell Biology, 68(1):1–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/68/1/1.

Darzynkiewicz:1977:DSD

[DTSM77] Z. Darzynkiewicz, F. Traganos, T. Sharpless, and M. R. Melamed. Different sensitivity of DNA in situ in interphase and metaphase chromatin to heat denaturation. Journal of Cell Biology, 73 (1):128–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/1/128.

Durieu-Trautmann:1975:RPE

[DTT75] O. Durieu-Trautmann and J. Tavlitzki. Reversible and permanent effects of the carbon sources and various antibiotics on the morphology and metabolic properties of Ustilago cynodontis cells. Journal of Cell Biology, 66(1):102–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/1/102.

Ducibella:1977:CCS

[DUKA77] T. Ducibella, T. Ukena, M. Karnovsky, and E. Anderson. Changes in cell surface and cortical cytoplasmic organization during early embryogenesis in the preimplantation mouse embryo. Journal of Cell Biology, 74(1):153–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/74/1/153.

Darszon:1979:IMP

[DVEM79] A. Darszon, C. A. Vandenberg, M. H. Ellisman, and M. Montal. Incorporation of membrane proteins into large single bilayer vesicles. Application to rhodopsin. Journal of Cell Biology, 81(2):446– ??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
2/446.

Damsky:1979:SFC

[DWB79] C. H. Damsky, D. E. Wylie, and C. A. Buck. Studies on the function of cell surface glycoproteins. II. Possible role of surface glycoproteins in the control of cytoskeletal organization and surface morphology. Journal of Cell Biology, 80(2):403–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/2/403.

Dewey:1977:CTF

[DWC^{+}77] M. M. Dewey, B. Walcott, D. E. Colflesh, H. Terry, and R. J. Levine. Changes in thick filament length in Limulus striated muscle. Journal of Cell Biology, 75(2):366–??, November 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/75/2/366.

Drochmans:1975:ISF

[DWM75] P. Drochmans, J. C. Wanson, and R. Mosselmans. Isolation and subfractionation on ficoll gradients of adult rat hepatocytes. Size, morphology, and biochemical characteristics of cell fractions. Journal of Cell Biology, 66(1):1–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/1.

Elsdale:1975:LSU

[EB75] T. Elsdale and J. Bard. Letter: Is stickiness of the upper surface of an attached epithelium in culture an indicator of functional insufficiency? Journal of Cell Biology, 66(1):218–??, July 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/1/218.

Estridge:1978:CAG

[EB78] M. Estridge and R. Bunge. Compositional analysis of growing axons from rat sympathetic neurons. Journal of Cell Biology, 79 (1):138–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/1/138.

Edds:1975:MENa

[Edd75a] K. T. Edds. Motility in Echinosphaerium nucleofilum. I. An analysis of particle motions in the axopodia and a direct test of the involvement of the axoneme. Journal of Cell Biology, 66(1):145– ??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/66/
1/145.

Edds:1975:MENb

[Edd75b] K. T. Edds. Motility in Echinosphaerium nucleofilum. II. Cytoplasmic contractility and its molecular basis. Journal of Cell Biology, 66(1):156–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/1/156.

Edds:1977:DAF

[Edd77] K. T. Edds. Dynamic aspects of filopodial formation by reorganization of microfilaments. Journal of Cell Biology, 73(2):479–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
2/479.

Edds:1979:ICT

[Edd79] K. T. Edds. Isolation and characterization of two forms of a cytoskeleton. Journal of Cell Biology, 83(1):109–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/1/109.

Elias:1975:PBM

[EF75] P. M. Elias and D. S. Friend. The permeability barrier in mammalian epidermis. Journal of Cell Biology, 65(1):180–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/1/180.

Elias:1976:VIM

[EF76] P. M. Elias and D. S. Friend. Vitamin-a-induced mucous metaplasia. An in vitro system for modulating tight and gap junction differentiation. Journal of Cell Biology, 68(2):173–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/2/173.

Epstein:1977:SCS

[EG77] M. L. Epstein and N. B. Gilula. A study of communication specificity between cells in culture. Journal of Cell Biology, 75 (3):769–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/769.

Elias:1978:FFI

[EGFB78] P. M. Elias, J. Goerke, D. S. Friend, and B. E. Brown. Freezefracture identification of sterol-digitonin complexes in cell and liposome membranes. Journal of Cell Biology, 78(2):577–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/78/
2/577.

Essner:1977:USG

[EH77] E. Essner and H. Haimes. Ultrastructural study of GERL in beige mouse alveolar macrophages. Journal of Cell Biology, 75 (2):381–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/381.

Eveloff:1979:ERG

[EKS^{+}79] J. Eveloff, K. J. Karnaky, P. Silva, F. H. Epstein, and W. B. Kinter. Elasmobranch rectal gland cell: autoradiographic localization of [3H]ouabain-sensitive Na, K–ATPase in rectal gland of dogfish, Squalus acanthias. Journal of Cell Biology, 83(1):16–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/83/1/16.

Ecklund:1975:MSB

[EL75] P. S. Ecklund and L. Levine. Mouse sperm basic nuclear protein. Electrophoretic characterization and fate after fertilization. Journal of Cell Biology, 66(2):251–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/2/251.

Edstrom:1976:CZA

[EL76] J. E. Edström and U. Lönn. Cytoplasmic zone analysis. RNA flow studied by micromanipulation. Journal of Cell Biology, 70 (3):562–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/562.

Eckert:1978:LAD

[EL78] B. S. Eckert and E. Lazarides. Localization of actin in Dictyostelium amebas by immunofluorescence. Journal of Cell Biology, 77(3):714–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/3/714.

Elfvin:1976:PIM

[ELD76] M. Elfvin, R. J. Levine, and M. M. Dewey. Paramyosin in invertebrate muscles. I. Identification and localization. Journal of Cell Biology, 71(1):261–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/1/261.

Ernst:1977:BPM

[EM77] S. A. Ernst and J. W. Mills. Basolateral plasma membrane localization of ouabain-sensitive sodium transport sites in the secretory epithelium of the avian salt gland. Journal of Cell Biology, 75(1):74–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/74.

Eichhorn:1979:LAI

[EP79a] J. H. Eichhorn and B. Peterkofsky. Local anesthetic-induced inhibition of collagen secretion in cultured cells under conditions where microtubules are not depolymerized by these agents. Journal of Cell Biology, 81(1):26–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/1/26.

Eichhorn:1979:QBA

[EP79b] J. H. Eichhorn and B. Peterkofsky. Quantitative biochemical analysis of microtubule content in normal and transformed 3T3 cells. Journal of Cell Biology, 82(2):572–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/2/572.

Ernst:1975:TAC

[Ern75] S. A. Ernst. Transport ATPase cytochemistry: ultrastructural localization of potassium-dependent and potassium-independent phosphatase activities in rat kidney cortex. Journal of Cell Biology, 66(3):586–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/3/586.

Ellisman:1976:SEM

[ERSP76] M. H. Ellisman, J. E. Rash, L. A. Staehelin, and K. R. Porter. Studies of excitable membranes. II. A comparison of specializations at neuromuscular junctions and nonjunctional sarcolemmas of mammalian fast and slow twitch muscle fibers. Journal of Cell Biology, 68(3):752–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/3/752.

Eddy:1976:CTS

[ES76] E. M. Eddy and B. M. Shapiro. Changes in the topography of the sea urchin egg after fertilization. Journal of Cell Biology, 71(1):35–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/1/35.

Ezzell:1979:LHA

[ES79] R. M. Ezzell and C. M. Szego. Luteinizing hormone-accelerated redistribution of lysosome-like organelles preceding dissolution of the nuclear envelope in rat oocytes maturing in vitro. Journal of Cell Biology, 82(1):264–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/1/264.

Elhammer:1975:BMM

[ESAD75] A. Elhammer, H. Svensson, F. Autuori, and G. Dallner. Biogenesis of microsomal membrane glycoproteins in rat liver. III. Release of glycoproteins from the Golgi fraction and their transfer to microsomal membranes. Journal of Cell Biology, 67(3):715–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/715.

Eriksson:1979:ISF

[ET79] A. Eriksson and L. E. Thornell. Intermediate (skeletin) filaments in heart Purkinje fibers. A correlative morphological and biochemical identification with evidence of a cytoskeletal function. Journal of Cell Biology, 80(2):231–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/2/231.

Edelhauser:1976:ETO

[EVMP76] H. F. Edelhauser, D. L. Van Horn, P. Miller, and H. J. Pederson. Effect of thiol-oxidation of glutathione with diamide on corneal endothelial function, junctional complexes, and microfilaments. Journal of Cell Biology, 68(3):567–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/567.

Edidin:1977:DRCa

[EW77a] M. Edidin and T. Y. Wei. Diffusion rates of cell surface antigens of mouse-human heterokaryons. I. Analysis of the population. Journal of Cell Biology, 75(2):475–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/2/475.

Edidin:1977:DRCb

[EW77b] M. Edidin and T. Y. Wei. Diffusion rates of cell surface antigens of mouse-human heterokaryons. II. Effect of membrane potential on lateral diffusion. Journal of Cell Biology, 75(2):483–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/2/483.

Eckert:1977:SBA

[EWR77] B. S. Eckert, R. H. Warren, and R. W. Rubin. Structural and biochemical aspects of cell motility in amebas of Dictyostelium discoideum. Journal of Cell Biology, 72(2):339–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/2/339.

Eastwood:1978:UTT

[EWR78] A. B. Eastwood, D. S. Wood, and J. P. Reuben. Unusual thick and thin filament packing in a crustacean muscle. Journal of Cell Biology, 77(1):48–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/1/48.

Etlinger:1976:CSM

[EZF76] J. D. Etlinger, R. Zak, and D. A. Fischman. Compositional studies of myofibrils from rabbit striated muscle. Journal of Cell Biology, 68(1):123–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/1/123.

Fletcher:1975:ICR

[FAE75] W. H. Fletcher, N. C. Anderson, and J. W. Everett. Intercellular communication in the rat anterior pituitary gland: an in vivo and in vitro study. Journal of Cell Biology, 67(2):469–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/2/469.

Franzini-Armstrong:1975:SST

[FALP75] C. Franzini-Armstrong, L. Landmesser, and G. Pilar. Size and shape of transverse tubule openings in frog twitch muscle fibers. Journal of Cell Biology, 64(2):493–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/2/493.

Farquhar:1978:RSM

[Far78] M. G. Farquhar. Recovery of surface membrane in anterior pituitary cells. Variations in traffic detected with anionic and cationic ferritin. Journal of Cell Biology, 77(3):??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/R35.

Fisher:1976:FFA

[FB76] K. A. Fisher and D. Branton. Freeze-fracture autoradiography: feasibility. Journal of Cell Biology, 70(2):453–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/2/453.

Fried:1978:RRS

[FB78] R. C. Fried and M. P. Blaustein. Retrieval and recycling of synaptic vesicle membrane in pinched-off nerve terminals (synaptosomes). Journal of Cell Biology, 78(3):685–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/685.

Farb:1979:URA

[FBF79] D. H. Farb, D. K. Berg, and G. D. Fischbach. Uptake and release of [3H]gamma-aminobutyric acid by embryonic spinal cord neurons in dissociated cell culture. Journal of Cell Biology, 80 (3):651–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/3/651.

Fukami:1978:IMI

[FBSS78] M. H. Fukami, J. S. Bauer, G. J. Stewart, and L. Salganicoff. An improved method for the isolation of dense storage granules from human platelets. Journal of Cell Biology, 77(2):389–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/2/389.

Flock:1977:AFS

[FC77] A. Flock and H. C. Cheung. Actin filaments in sensory hairs of inner ear receptor cells. Journal of Cell Biology, 75(2):339–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
2/339.

Frontali:1976:PBW

[FCG^{+}76] N. Frontali, B. Ceccarelli, A. Gorio, A. Mauro, P. Siekevitz, M. C. Tzeng, and W. P. Hurlbut. Purification from black widow spider venom of a protein factor causing the depletion of synaptic vesicles at neuromuscular junctions. Journal of Cell Biology, 68(3):462–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/462.

Fambrough:1978:NSA

[FD78] D. M. Fambrough and P. N. Devreotes. Newly synthesized acetylcholine receptors are located in the Golgi apparatus. Journal of Cell Biology, 76(1):237–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/237.

Fry:1977:FFS

[FDB77] G. N. Fry, C. E. Devine, and G. Burnstock. Freeze-fracture studies of nexuses between smooth muscle cells. Close relationship to sarcoplasmic reticulum. Journal of Cell Biology, 72(1):26–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/72/1/26.

Frank:1979:EEN

[FF79a] E. Frank and G. D. Fischbach. Early events in neuromuscular junction formation in vitro: induction of acetylcholine receptor clusters in the postsynaptic membrane and morphology of newly formed synapses. Journal of Cell Biology, 83(1):143–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/1/143.

Friedlander:1979:ISE

[FF79b] M. Friedlander and D. A. Fischman. Immunological studies of the embryonic muscle cell surface. Antiserum to the prefusion myoblast. Journal of Cell Biology, 81(1):193–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/1/193.

Fenwick:1978:FMC

[FFHL78] E. M. Fenwick, P. B. Fajdiga, N. B. Howe, and B. G. Livett. Functional and morphological characterization of isolated bovine adrenal medullary cells. Journal of Cell Biology, 76(1):12–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/76/1/12.

Fuster:1977:RPA

[FFSH77] C. D. Fuster, D. Farrell, F. A. Stern, and N. B. Hecht. RNA polymerase activity in bovine spermatozoa. Journal of Cell Biology, 74 (3):698–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/698.

Forest:1978:FMA

[FGG78] C. L. Forest, D. A. Goodenough, and U. W. Goodenough. Flagellar membrane agglutination and sexual signaling in the conditional GAM-1 mutant of Chlamydomonas. Journal of Cell Biology, 79(1):74–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/1/74.

Franke:1978:PAM

[FGSM78] W. W. Franke, C. Grund, E. Schmid, and E. Mandelkow. Paracrystalline arrays of membrane-to-membrane cross bridges associated with the inner surface of plasma membrane. Journal of Cell Biology, 77(2):323–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/323.

Forssmann:1977:HIP

[FI77] W. G. Forssmann and S. Ito. Hepatocyte innervation in primates. Journal of Cell Biology, 74(1):299–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/1/299.

Frankel:1976:CRA

[FJD76] J. Frankel, L. M. Jenkins, and L. E. DeBault. Causal relations among cell cycle processes in Tetrahymena pyriformis. An analysis employing temperature-sensitive mutants. Journal of Cell Biology, 71(1):242–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/71/1/242.

Forrest:1978:TLC

[FK78] G. L. Forrest and R. R. Klevecz. Tyrosyltubulin ligase and colchicine binding activity in synchronized Chinese hamster cells. Journal of Cell Biology, 78(2):441–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/78/2/441.

Fujii-Kuriyama:1979:BCP

[FKNMT79] Y. Fujii-Kuriyama, M. Negishi, R. Mikawa, and Y. Tashiro. Biosynthesis of cytochrome P-450 on membrane-bound ribosomes and its subsequent incorporation into rough and smooth microsomes in rat hepatocytes. Journal of Cell Biology, 81(3):510–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/510.

Fisher:1976:DCA

[FKP76] G. Fisher, E. S. Kaneshiro, and P. D. Peters. Divalent cation affinity sites in Paramecium aurelia. Journal of Cell Biology, 69(2):429–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/429.

Franke:1976:IBC

[FKS76] W. W. Franke, J. Kartenbeck, and H. Spring. Involvement of bristle coat structures in surface membrane formations and membrane interactions during coenocytotomic cleavage in caps of Acetabularia mediterranea. Journal of Cell Biology, 71(1):196–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/1/196.

Franke:1976:IVC

[FLK^{+}76] W. W. Franke, M. R. Lüder, J. Kartenbeck, H. Zerban, and T. W. Keenan. Involvement of vesicle coat material in casein secretion and surface regeneration. Journal of Cell Biology, 69(1):173–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
1/173.

Freilich:1977:GTG

[FLRS77] L. S. Freilich, R. G. Lewis, A. C. Reppucci, and J. E. Silbert. Galactosyl transferase of a Golgi fraction from cultured neoplastic mast cells. Journal of Cell Biology, 72(3):655–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/3/655.

Frederiksen:1979:LCB

[FMR79] O. Frederiksen, K. Møllgård, and J. Rostgaard. Lack of correlation between transepithelial transport capacity and paracellular pathway ultrastructure in Alcian blue-treated rabbit gallbladders. Journal of Cell Biology, 83(2):383–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/2/383.

Feldmann:1975:ICF

[FMSB75] G. Feldmann, M. Maurice, C. Sapin, and J. P. Benhamou. Inhibition by colchicine of fibrinogen translocation in hepatocytes. Journal of Cell Biology, 67(1):237–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/1/237.

Fouts:1975:PPK

[FMW75] D. L. Fouts, J. E. Manning, and D. R. Wolstenholme. Physicochemical properties of kinetoplast DNA from Crithidia acanthocephali. Crithidia luciliae, and Trypanosoma lewisi. Journal of Cell Biology, 67(2):378–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/2/378.

Fernandez:1976:UMC

[FN76] H. R. Fernandez and E. E. Nickel. Ultrastructural and molecular characteristics of crayfish photoreceptor membranes. Journal of Cell Biology, 69(3):721–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/3/721.

Friend:1977:MPC

[FOPY77] D. S. Friend, L. Orci, A. Perrelet, and R. Yanagimachi. Membrane particle changes attending the acrosome reaction in guinea pig spermatozoa. Journal of Cell Biology, 74(2):561–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/561.

Fujiwara:1976:FAL

[FP76] K. Fujiwara and T. D. Pollard. Fluorescent antibody localization of myosin in the cytoplasm, cleavage furrow, and mitotic spindle of human cells. Journal of Cell Biology, 71(3):848–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/3/848.

Feldherr:1978:MSN

[FP78a] C. M. Feldherr and J. Pomerantz. Mechanism for the selection of nuclear polypeptides in Xenopus oocytes. Journal of Cell Biology, 78(1):168–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/168.

Fujiwara:1978:SLM

[FP78b] K. Fujiwara and T. D. Pollard. Simultaneous localization of myosin and tubulin in human tissue culture cells by double antibody staining. Journal of Cell Biology, 77(1):182–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/1/182.

Fried:1976:FCA

[FPC76] J. Fried, A. G. Perez, and B. D. Clarkson. Flow cytofluorometric analysis of cell cycle distributions using propidium iodide. Properties of the method and mathematical analysis of the data. Journal of Cell Biology, 71(1):172–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/1/172.

Fujiwara:1978:AAL

[FPP78] K. Fujiwara, M. E. Porter, and T. D. Pollard. Alpha-actinin localization in the cleavage furrow during cytokinesis. Journal of Cell Biology, 79(1):268–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/1/268.

Fowler:1976:ASM

[FRT^{+}76] S. Fowler, J. Remacle, A. Trouet, H. Beaufay, J. Berthet, M. Wibo, and P. Hauser. Analytical study of microsomes and isolated subcellular membranes from rat liver. V. Immunological localization of cytochrome b5 by electron microscopy: methodology and application to various subcellular fractions. Journal of Cell Biology, 71(2):535–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/2/535.

Fertuck:1976:QJE

[FS76] H. C. Fertuck and M. M. Salpeter. Quantitation of junctional and extrajunctional acetylcholine receptors by electron microscope autoradiography after (125)I-α-bungarotoxin binding at mouse neuromuscular junctions. Journal of Cell Biology, 69(1):144–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/1/144.

Flashner:1977:RAS

[FS77] Y. Flashner and M. Schramm. Retention of amylase in the secretory granules of parotid gland after extensive release of Ca++ by ionophore A-23187. Journal of Cell Biology, 74(3): 789–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/789.

Feigenson:1975:DRB

[FSB75] M. E. Feigenson, H. P. Schnebli, and M. Baggiolini. Demonstration of ricin-binding sites on the outer face of azurophil and specific granules of rabbit polymorphonuclear leukocytes. Journal of Cell Biology, 66(1):183–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/183.

Felden:1976:PHA

[FSM76] R. A. Felden, M. M. Sanders, and N. R. Morris. Presence of histones in Aspergillus nidulans. Journal of Cell Biology, 68(3):430– ??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/430.

Franke:1979:ISF

[FSOW79] W. W. Franke, E. Schmid, M. Osborn, and K. Weber. Intermediate-sized filaments of human endothelial cells. Journal of Cell Biology, 81(3):570–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/3/570.

Franke:1975:GNE

[FSSZ75] W. W. Franke, H. Spring, U. Scheer, and H. Zerban. Growth of the nuclear envelope in the vegetative phase of the green alga Acetabularia. Evidence for assembly from membrane components synthesized in the cytoplasm. Journal of Cell Biology, 66(3): 681–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/681.

Fowler:1977:SFM

[FSW77] S. Fowler, H. Shio, and H. Wolinsky. Subcellular fractionation and morphology of calf aortic smooth muscle cells: studies on whole aorta, aortic explants, and subcultures grown under. Journal of Cell Biology, 75(1):166–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/1/166.

Frankel:1978:FMM

[FTBS78] F. R. Frankel, R. W. Tucker, J. Bruce, and R. Stenberg. Fibroblasts and macrophages of mice with the Chediak–Higashilike syndrome have microtubules and actin cables. Journal of Cell Biology, 79(2):401–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/2/401.

Fukui:1978:IAB

[Fuk78] Y. Fukui. Intranuclear actin bundles induced by dimethyl sulfoxide in interphase nucleus of Dictyostelium. Journal of Cell Biology, 76(1):146–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/1/146.

Fuseler:1975:MTA

[Fus75a] J. W. Fuseler. Mitosis in Tilia americana endosperm. Journal of Cell Biology, 64(1):159–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/159.

Fuseler:1975:TDA

[Fus75b] J. W. Fuseler. Temperature dependence of anaphase chromosome velocity and microtubule depolymerization. Journal of Cell Biology, 67(3):789–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/67/3/789.

Fouts:1978:HSC

[FWB78] D. L. Fouts, D. R. Wolstenholme, and H. W. Boyer. Heterogeneity in sensitivity to cleavage by the restriction endonucleases ECORI and HindIII of circular kinetoplast DNA molecules of Crithidia acanthocephali. Journal of Cell Biology, 79(2):329–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/2/329.

Fairlamb:1978:ICK

[FWHB78] A. H. Fairlamb, P. O. Weislogel, J. H. Hoeijmakers, and P. Borst. Isolation and characterization of kinetoplast DNA from bloodstream form of Trypanosoma brucei. Journal of Cell Biology, 76 (2):293–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/2/293.

Fertuck:1975:VRM

[FWS75] H. C. Fertuck, W. Woodward, and M. M. Salpeter. In vivo recovery of muscle contraction after alpha-bungarotoxin binding. Journal of Cell Biology, 66(1):209–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/209.

Fisher:1978:OAP

[FYS78] K. A. Fisher, K. Yanagimoto, and W. Stoeckenius. Oriented adsorption of purple membrane to cationic surfaces. Journal of Cell Biology, 77(2):611–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/2/611.

Gauthier:1979:UIM

[Gau79] G. F. Gauthier. Ultrastructural identification of muscle fiber types by immunocytochemistry. Journal of Cell Biology, 82(2):391–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
2/391.

Gould:1977:PMC

[GB77] R. R. Gould and G. G. Borisy. The pericentriolar material in Chinese hamster ovary cells nucleates microtubule formation. Journal of Cell Biology, 73(3):601–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/3/601.

Gabella:1979:NBS

[GB79] G. Gabella and D. Blundell. Nexuses between the smooth muscle cells of the guinea-pig ileum. Journal of Cell Biology, 82(1):239– ??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
1/239.

Gerace:1978:ILM

[GBB78] L. Gerace, A. Blum, and G. Blobel. Immunocytochemical localization of the major polypeptides of the nuclear pore complexlamina fraction. Interphase and mitotic distribution. Journal of Cell Biology, 79(2):546–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/2/546.

Gospodarowicz:1978:CPH

[GBBZ78] D. Gospodarowicz, K. D. Brown, C. R. Birdwell, and B. R. Zetter. Control of proliferation of human vascular endothelial cells. Characterization of the response of human umbilical vein endothelial cells to fibroblast growth factor, epidermal growth factor, and thrombin. Journal of Cell Biology, 77(3):774–??, June 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/3/774.

Glaumann:1975:SSI

[GBE75] H. Glaumann, A. Bergstrand, and J. L. Ericsson. Studies on the synthesis and intracellular transport of lipoprotein particles in rat liver. Journal of Cell Biology, 64(2):356–??, February 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/2/356.

Gross:1978:FFU

[GBM78] H. Gross, E. Bas, and H. Moor. Freeze-fracturing in ultrahigh vacuum at −196 degrees C. Journal of Cell Biology, 76(3):712–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
3/712.

Gabbiani:1978:CFG

[GCH78] G. Gabbiani, C. Chaponnier, and I. Hüttner. Cytoplasmic filaments and gap junctions in epithelial cells and myofibroblasts during wound healing. Journal of Cell Biology, 76(3):561–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/3/561.

Grasso:1977:BCR

[GCM77] J. A. Grasso, N. C. Chromey, and C. F. Moxey. Biochemical characterization of RNA and protein synthesis in erythrocyte development. Journal of Cell Biology, 73(1):206–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/1/206.

Grebanier:1978:MPS

[GCRB78] A. E. Grebanier, D. M. Coen, A. Rich, and L. Bogorad. Membrane proteins synthesized but not processed by isolated maize chloroplasts. Journal of Cell Biology, 78(3):734–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/734.

Goldman:1979:UMB

[GCY79] R. D. Goldman, B. Chojnacki, and M. J. Yerna. Ultrastructure of microfilament bundles in baby hamster kidney (BHK-21) cells. The use of tannic acid. Journal of Cell Biology, 80(3):759–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/759.

Gershman:1975:MNV

[GD75] H. Gershman and J. Drumm. Mobility of normal and virustransformed cells in cellular aggregates. Journal of Cell Biology, 67(2):419–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/2/419.

Gould:1976:INF

[GD76] R. M. Gould and R. M. Dawson. Incorporation of newly formed lecithin into peripheral nerve myelin. Journal of Cell Biology, 68(3):480–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/3/480.

Giulian:1977:EMC

[GD77] D. Giulian and E. G. Diacumakos. The electrophysiological mapping of compartments within a mammalian cell. Journal of Cell Biology, 72(1):86–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/72/1/86.

Gershman:1976:SNV

[GDC76] H. Gershman, J. Drumm, and L. Culp. Sorting out of normal and virus-transformed cells in cellular aggregates. Journal of Cell Biology, 68(2):276–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/2/276.

Gershman:1977:DCA

[GDR77] H. Gershman, J. Drumm, and J. J. Rosen. Dibutyryl cyclic AMP treatment of 3T3 and SV40 virus-transformed 3T3 cells in aggregates. Effects on mobility and cell contact ultrastructure. Journal of Cell Biology, 72(2):424–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/2/424.

Goldstein:1979:MMH

[GE79] M. A. Goldstein and M. L. Entman. Microtubules in mammalian heart muscle. Journal of Cell Biology, 80(1):183–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/1/183.

Gilula:1978:CCC

[GEB78] N. B. Gilula, M. L. Epstein, and W. H. Beers. Cell-to-cell communication and ovulation. A study of the cumulus-oocyte complex. Journal of Cell Biology, 78(1):58–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/78/1/58.

Gabel:1979:RDS

[GES79] C. A. Gabel, E. M. Eddy, and B. M. Shapiro. Regional differentiation of the sperm surface as studied with 125I-diiodofluorescein isothiocyanate, an impermeant reagent that allows isolation of the labeled components. Journal of Cell Biology, 82(3):742– ??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/742.

Gambetti:1975:CCS

[GESG75] P. Gambetti, S. E. Erulkar, A. P. Somlyo, and N. K. Gonatas. Calcium-containing structures in vertebrate glial cells. Ultrastructural and microprobe analysis. Journal of Cell Biology, 64(2): 322–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/2/322.

Goodman:1978:GIT

[GET^{+}78] R. Goodman, D. Edgar, H. Thoenen, W. Wechsler, and H. Herschman. Glucocorticoid induction of tyrosine hydroxylase in a continous cell line of rat pheochromocytoma. Journal of Cell Biology, 78(1):??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/R1.

Gallin:1977:ICF

[GG77] E. K. Gallin and J. I. Gallin. Interaction of chemotactic factors with human macrophages: induction of transmembrane potential changes. Journal of Cell Biology, 75(1):277–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/277.

Graessmann:1978:QCB

[GGGM78] A. Graessmann, M. Graessmann, E. Guhl, and C. Mueller. Quantitative correlation between simian virus 40 T-antigen synthesis and late viral gene expression in permissive and nonpermissive cells. Journal of Cell Biology, 77(1):??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/1/R1.

Gonatas:1976:QUA

[GGS^{+}76] N. K. Gonatas, J. O. Gonatas, A. Stieber, J. C. Antoine, and S. Avrameas. Quantitative ultrastructural autoradiographic studies of iodinated plasma membranes of lymphocytes during segregation and internalization of surface immunoglobulins. Journal of Cell Biology, 70(3):477–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/477.

Gahmberg:1976:CSG

[GHA76] C. G. Gahmberg, P. Häyry, and L. C. Andersson. Characterization of surface glycoproteins of mouse lymphoid cells. Journal of Cell Biology, 68(3):642–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/3/642.

Gorio:1978:ACM

[GHC78] A. Gorio, W. P. Hurlbut, and B. Ceccarelli. Acetylcholine compartments in mouse diaphragm. Comparison of the effects of black widow spider venom, electrical stimulation, and high concentrations of potassium. Journal of Cell Biology, 78(3):716–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/78/
3/716.

Geschwind:1977:IRC

[GHMD77] I. I. Geschwind, J. M. Horowitz, G. M. Mikuckis, and R. D. Dewey. Iontophoretic release of cyclic AMP and dispersion of melanosomes within a single melanophore. Journal of Cell Biology, 74(3):928–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/3/928.

Goldstein:1975:MLE

[GHW75] I. M. Goldstein, S. T. Hoffstein, and G. Weissmann. Mechanisms of lysosomal enzyme release from human polymorphonuclear leukocytes. Effects of phorbol myristate acetate. Journal of Cell Biology, 66(3):647–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/3/647.

Goodenough:1978:TMS

[GJ78] U. W. Goodenough and D. Jurivich. Tipping and matingstructure activation induced in Chlamydomonas gametes by flagellar membrane antisera. Journal of Cell Biology, 79(3):680–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
3/680.

Glaumann:1975:ILL

[GJAE75] H. Glaumann, H. Jansson, B. Arborgh, and J. L. Ericsson. Isolation of liver lysosomes by iron loading. Ultrastructural characterization. Journal of Cell Biology, 67(3):887–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/3/887.

Gilkey:1978:FCW

[GJRR78] J. C. Gilkey, L. F. Jaffe, E. B. Ridgway, and G. T. Reynolds. A free calcium wave traverses the activating egg of the medaka, Oryzias latipes. Journal of Cell Biology, 76(2):448–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/448.

Gremo:1978:DEL

[GKB78] F. Gremo, D. Kobiler, and S. H. Barondes. Distribution of an endogenous lectin in the developing chick optic tectum. Journal of Cell Biology, 79(2):491–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/2/491.

Gross:1978:DSS

[GKBM78] H. Gross, O. Kuebler, E. Bas, and H. Moor. Decoration of specific sites on freeze-fractured membranes. Journal of Cell Biology, 79 (3):646–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/646.

Goldman:1976:ATI

[GKS76] J. E. Goldman, K. S. Kim, and J. H. Schwartz. Axonal transport of [3H]serotonin in an identified neuron of Aplysia californica. Journal of Cell Biology, 70(2):304–??, August 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/70/2/304.

Gonatas:1977:ILN

[GKSA77] N. K. Gonatas, S. U. Kim, A. Stieber, and S. Avrameas. Internalization of lectins in neuronal GERL. Journal of Cell Biology, 73(1):1–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/1/1.

Gauthier:1977:PMA

[GL77] G. F. Gauthier and S. Lowey. Polymorphism of myosin among skeletal muscle fiber types. Journal of Cell Biology, 74(3): 760–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/760.

Gard:1979:SFL

[GL79a] D. L. Gard and E. Lazarides. Specific fluorescent labeling of chicken myofibril Z-line proteins catalyzed by guinea pig liver transglutaminase. Journal of Cell Biology, 81(2):336–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/336.

Gauthier:1979:DMI

[GL79b] G. F. Gauthier and S. Lowey. Distribution of myosin isoenzymes among skeletal muscle fiber types. Journal of Cell Biology, 81(1):10–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/1/10.

Glabe:1979:SSS

[GL79c] C. G. Glabe and W. J. Lennarz. Species-specific sperm adhesion in sea urchins. A quantitative investigation of bindin- mediated egg agglutination. Journal of Cell Biology, 83(3):595–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/3/595.

Gordon:1977:EMD

[GLM77] R. E. Gordon, B. P. Lane, and F. Miller. Electron microscope demonstration of tubulin in cilia and basal bodies of rat tracheal epithelium by the use of an antitubulin antibody. Journal of Cell Biology, 75(2):586–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/2/586.

Glass:1975:SME

[GLR75] J. Glass, L. M. Lavidor, and S. H. Robinson. Studies of murine erythroid cell development. Synthesis of heme and hemoglobin. Journal of Cell Biology, 65(2):298–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/2/298.

Gemmell:1977:UBE

[GLR77] R. T. Gemmell, S. G. Laychock, and R. P. Rubin. Ultrastructural and biochemical evidence for a steroid-containing secretory organelle in the perfused cat adrenal gland. Journal of Cell Biology, 72(1):209–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/1/209.

Gospodarowicz:1975:MEF

[GM75] D. Gospodarowicz and J. S. Moran. Mitogenic effect of fibroblast growth factor on early passage cultures of human and murine fibroblasts. Journal of Cell Biology, 66(2):451–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/2/451.

Grady:1976:IAS

[GM76] S. R. Grady and E. J. McGuire. Intercellular adhesive selectivity. III. Species selectivity of embryonic liver intercellular adhesion. Journal of Cell Biology, 71(1):96–??, October 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/1/96.

Godman:1975:ACD

[GMDT75] G. C. Godman, A. F. Miranda, A. D. Deitch, and S. W. Tanenbaum. Action of cytochalasin D on cells of established lines. III. Zeiosis and movements at the cell surface. Journal of Cell Biology, 64(3):644–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/644.

Giller:1977:CAA

[GNB^{+}77] E. L. Giller, J. H. Neale, P. N. Bullock, B. K. Schrier, and P. G. Nelson. Choline acetyltransferase activity of spinal cord cell cultures increased by co-culture with muscle and by muscleconditioned medium. Journal of Cell Biology, 74(1):16–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/1/16.

Gibbons:1976:EAS

[GOG76] B. H. Gibbons, K. Ogawa, and I. R. Gibbons. The effect of antidynein 1 serum on the movement of reactivated sea urchin sperm. Journal of Cell Biology, 71(3):823–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/823.

Goldstein:1979:STS

[Gol79] S. F. Goldstein. Starting transients in sea urchin sperm flagella. Journal of Cell Biology, 80(1):61–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/80/1/61.

Goodenough:1976:VFG

[Goo76] D. A. Goodenough. In vitro formation of gap junction vesicles. Journal of Cell Biology, 68(2):220–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/68/2/220.

Gordon:1975:CFS

[Gor75] S. Gordon. Cell fusion and some subcellular properties of heterokaryons and hybrids. Journal of Cell Biology, 67(2):257–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
2/257.

Gould:1975:BBC

[Gou75] R. R. Gould. The basal bodies of Chlamydomonas reinhardtii. Formation from probasal bodies, isolation, and partial characterization. Journal of Cell Biology, 65(1):65–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/1/65.

Gould:1977:IGP

[Gou77] R. M. Gould. Incorporation of glycoproteins into peripheral nerve myelin. Journal of Cell Biology, 75(2):326–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/2/326.

Goldring:1977:IHM

[GP77] E. S. Goldring and W. J. Peacock. Intramolecular heterogeneity of mitochondrial DNA of Drosophila melanogaster. Journal of Cell Biology, 73(2):279–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/2/279.

Griffith:1978:EAF

[GP78] L. M. Griffith and T. D. Pollard. Evidence for actin filamentmicrotubule interaction mediated by microtubule-associated proteins. Journal of Cell Biology, 78(3):958–??, September 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/78/3/958.

Githens:1976:BMA

[GPPR76] S. Githens, R. Pictet, P. Phelps, and W. J. Rutter. 5bromodeoxyuridine may alter the differentiative program of the embryonic pancreas. Journal of Cell Biology, 71(2):341–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
2/341.

Gershman:1978:CAC

[GR78] H. Gershman and J. J. Rosen. Cell adhesion and cell surface topography in aggregates of 3T3 and SV40-virus-transformed 3T3 cells. Visualization of interior cells by scanning electron microscopy. Journal of Cell Biology, 76(3):639–??, March 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/3/639.

Granick:1975:NNCa

[Gra75a] D. Granick. Nucleolar necklaces in chick embryo fibroblast cells. I. Formation of necklaces by dichlororibobenzimidazole and other adenosine analogues that decrease RNA synthesis and degrade preribosomes. Journal of Cell Biology, 65(2):398–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/2/398.

Granick:1975:NNCb

[Gra75b] D. Granick. Nucleolar necklaces in chick embryo fibroblast cells. II. Microscope observations of the effect of adenosine analogues on nucleolar necklace formation. Journal of Cell Biology, 65(2):418– ??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/65/
2/418.

Greenberg:1975:MRM

[Gre75] Greenberg. Messenger RNA metabolism of animal cells: possible involvement of untranslated sequences and mRNA-associated proteins. Journal of Cell Biology, 64(2):269–??, February 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/2/269.

Greene:1978:NGF

[Gre78] L. A. Greene. Nerve growth factor prevents the death and stimulates the neuronal differentiation of clonal PC12 pheochromocytoma cells in serum-free medium. Journal of Cell Biology, 78 (3):747–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/747.

Gruenstein:1975:AAM

[GRW75] E. Gruenstein, A. Rich, and R. R. Weihing. Actin associated with membranes from 3T3 mouse fibroblast and HeLa cells. Journal of Cell Biology, 64(1):223–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/223.

Goodenough:1975:BMA

[GS75] U. W. Goodenough and H. S. StClair. BALD-2: a mutation affecting the formation of doublet and triplet sets of microtubules in Chlamydomonas reinhardtii. Journal of Cell Biology, 66(3): 480–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/480.

Gainer:1977:BAT

[GSB77] H. Gainer, Y. Sarne, and M. J. Brownstein. Biosynthesis and axonal transport of rat neurohypophysial proteins and peptides. Journal of Cell Biology, 73(2):366–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/2/366.

Gould-Somero:1979:EMF

[GSJH79] M. Gould-Somero, L. A. Jaffe, and L. Z. Holland. Electrically mediated fast polyspermy block in eggs of the marine worm, Urechis caupo. Journal of Cell Biology, 82(2):426–??, August 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/82/2/426.

Goldman:1978:ICG

[GSN78] J. E. Goldman, H. H. Schaumburg, and W. T. Norton. Isolation and characterization of glial filaments from human brain. Journal of Cell Biology, 78(2):426–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/426.

Goldstein:1977:ODZ

[GSS77] M. A. Goldstein, J. P. Schroeter, and R. L. Sass. Optical diffraction of the Z lattice in canine cardiac muscle. Journal of Cell Biology, 75(3):818–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/75/3/818.

Goldstein:1979:ZLC

[GSS79] M. A. Goldstein, J. P. Schroeter, and R. L. Sass. The Z lattice in canine cardiac muscle. Journal of Cell Biology, 83(1):187–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/187.

Geuze:1979:ILA

[GST79] J. J. Geuze, J. W. Slot, and K. T. Tokuyasu. Immunocytochemical localization of amylase and chymotrypsinogen in the exocrine pancreatic cell with special attention to the Golgi complex. Journal of Cell Biology, 82(3):697–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/3/697.

Grinnell:1975:DMA

[GTH75] F. Grinnell, M. Q. Tobleman, and C. R. Hackenbrock. The distribution and mobility of anionic sites on the surfaces of baby hamster kidney cells. Journal of Cell Biology, 66(3):470–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/3/470.

Grinnell:1976:IAB

[GTH76] F. Grinnell, M. Q. Tobleman, and C. R. Hackenbrock. Initial attachment of baby hamster kidney cells to an epoxy substratum. Ultrastructural analysis. Journal of Cell Biology, 70(3): 707–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/707.

Gainer:1977:EGN

[GTL77] H. Gainer, I. Tasaki, and R. J. Lasek. Evidence for the glia-neuron protein transfer hypothesis from intracellular perfusion studies of squid giant axons. Journal of Cell Biology, 74(2):524–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/524.

Garatun-Tjeldsto:1976:SSL

[GTPWD76] O. Garatun-Tjeldsto, I. F. Pryme, J. K. Weltman, and R. M. Dowben. Synthesis and secretion of light-chain immunoglobulin in two successive cycles of synchronized plasmacytoma cells. Journal of Cell Biology, 68(2):232–??, February 1976. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/2/232.

Glabe:1977:ICV

[GV77] C. G. Glabe and V. D. Vacquier. Isolation and characterization of the vitelline layer of sea urchin eggs. Journal of Cell Biology, 75 (2):410–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/410.

Goodenough:1975:GDC

[GW75] U. W. Goodenough and R. L. Weiss. Gametic differentiation in Chlamydomonas reinhardtii. III. Cell wall lysis and microfilamentassociated mating structure activation in wild-type and mutant strains. Journal of Cell Biology, 67(3):623–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/3/623.

Goodenough:1978:IBM

[GW78] U. W. Goodenough and R. L. Weiss. Interrelationships between microtubules, a striated fiber, and the gametic mating structure of Chlamydomonas reinhardi. Journal of Cell Biology, 76(2):430–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
2/430.

Goldstein:1977:SNR

[GWK77] L. Goldstein, G. E. Wise, and C. Ko. Small nuclear RNA localization during mitosis. An electron microscope study. Journal of Cell Biology, 73(2):322–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/2/322.

Gospodarowicz:1976:EFG

[GWML76] D. Gospodarowicz, J. Weseman, J. S. Moran, and J. Lindstrom. Effect of fibroblast growth factor on the division and fusion of bovine myoblasts. Journal of Cell Biology, 70(2):395–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/2/395.

Gulley:1977:RNM

[GWN77] R. L. Gulley, R. J. Wenthold, and G. R. Neises. Remodeling of neuronal membranes as an early response to deafferentation.

A freeze-fracture study. Journal of Cell Biology, 75(3):837–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/837.

Grubman:1977:SME

[GWS77] M. J. Grubman, J. A. Weinstein, and D. A. Shafritz. Studies on the mechanism for entry of vesicular stomatitis virus glycoprotein g mRNA into membrane-bound polyribosome complexes. Journal of Cell Biology, 74(1):43–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/1/43.

Giannattasio:1975:MOR

[GZM75] G. Giannattasio, A. Zanini, and J. Meldolesi. Molecular organization of rat prolactin granules: in vitro stability of intact and “membraneless” granules. Journal of Cell Biology, 64(1):246–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
1/246.

Hausmann:1976:MBE

[HA76] K. Hausmann and R. D. Allen. Membrane behavior of exocytic vesicles: II in fate of the trichocyst membranes in paramecium after induced trichocyst discharge. Journal of Cell Biology, 69(2):313–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/313.

Heggeness:1977:UAB

[HA77] M. H. Heggeness and J. F. Ash. Use of the avidin-biotin complex for the localization of actin and myosin with fluorescence microscopy. Journal of Cell Biology, 73(3):783–??, June 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/73/3/783.

Hand:1975:DRM

[Han75a] R. Hand. DNA replication in mammalian cells. Altered patterns of initiation during inhibition of protein synthesis. Journal of Cell Biology, 67(3):761–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/761.

Hand:1975:RDR

[Han75b] R. Hand. Regulation of DNA replication on subchromosomal units of mammalian cells. Journal of Cell Biology, 64(1):89–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/1/89.

Hansma:1979:SUM

[Han79] H. G. Hansma. Sodium uptake and membrane excitation in Paramecium. Journal of Cell Biology, 81(2):374–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/374.

Hardy:1978:ICM

[Har78a] M. A. Hardy. Intracellular calcium as a modulator of transepithelial permeability to water in frog urinary bladder. Journal of Cell Biology, 76(3):787–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/3/787.

Hartwell:1978:CDG

[Har78b] L. H. Hartwell. Cell division from a genetic perspective. Journal of Cell Biology, 77(3):627–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/3/627.

Habener:1979:PHB

[HARO79] J. F. Habener, M. Amherdt, M. Ravazzola, and L. Orci. Parathyroid hormone biosynthesis. Correlation of conversion of biosynthetic precursors with intracellular protein migration as determined by electron microscope autoradiography. Journal of Cell Biology, 80(3):715–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/3/715.

Hallberg:1976:RBT

[HB76a] R. L. Hallberg and P. J. Bruns. Ribosome biosynthesis in Tetrahymena pyriformis. Regulation in response to nutritional changes. Journal of Cell Biology, 71(2):383–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/383.

Hofmann:1976:ISE

[HB76b] C. Hofmann and G. B. Bouck. Immunological and structural evidence for patterned intussusceptive surface growth in a unicellular organism. A postulated role for submembranous proteins and microtubules. Journal of Cell Biology, 69(3):693–??, June 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/69/3/693.

Hyams:1978:NMV

[HB78] J. S. Hyams and G. G. Borisy. Nucleation of microtubules in vitro by isolated spindle pole bodies of the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 78(2):401–??, August 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/78/2/401.

Hill:1979:RPB

[HB79] D. J. Hill and J. Z. Borysenko. Restriction of patching of bound concanavalin A after incorporation of arachidonic acid into the plasma membrane of virally transformed fibroblasts. Journal of Cell Biology, 83(1):218–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/1/218.

Howell:1975:IED

[HBD75] S. H. Howell, W. J. Blaschko, and C. M. Drew. Inhibitor effects during the cell cycle in Chlamydomonas reinhardtii. Determination of transition points in asynchronous cultures. Journal of Cell Biology, 67(1):126–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/126.

Hollyfield:1977:TRP

[HBR77] J. G. Hollyfield, J. C. Besharse, and M. E. Rayborn. Turnover of rod photoreceptor outer segments. I. Membrane addition and loss in relationship to temperature. Journal of Cell Biology, 75 (2):490–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/490.

Hardy:1979:SRE

[HBS79] B. Hardy, K. G. Bensch, and S. L. Schrier. Spectrin rearrangement early in erythrocyte ghost endocytosis. Journal of Cell Biology, 82 (3):654–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/654.

Hubbard:1975:EDPa

[HC75a] A. L. Hubbard and Z. A. Cohn. Externally disposed plasma membrane proteins. I. Enzymatic iodination of mouse L cells. Journal of Cell Biology, 64(2):438–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/2/438.

Hubbard:1975:EDPb

[HC75b] A. L. Hubbard and Z. A. Cohn. Externally disposed plasma membrane proteins. II. Metabolic fate of iodinated polypeptides of mouse L cells. Journal of Cell Biology, 64(2):461–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/2/461.

Heidrich:1977:HCP

[HD77] H. G. Heidrich and M. E. Dew. Homogeneous cell populations from rabbit kidney cortex. Proximal, distal tubule, and reninactive cell isolated by free-flow electrophoresis. Journal of Cell Biology, 74(3):780–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/74/3/780.

Hartwig:1977:ECP

[HDS77] J. H. Hartwig, W. A. Davies, and T. P. Stossel. Evidence for contractile protein translocation in macrophage spreading, phagocytosis, and phagolysosome formation. Journal of Cell Biology, 75 (3):956–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/956.

Hellerqvist:1979:IAF

[Hel79] C. G. Hellerqvist. Intercellular adhesion as a function of the cell cycle traverse. Journal of Cell Biology, 82(3):682–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/3/682.

Herr:1976:RGJ

[Her76] J. C. Herr. Reflexive gap junctions. Gap junctions between processing arising from the same ovarian decidual cell. Journal of Cell Biology, 69(2):495–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/2/495.

Helenius:1977:RSF

[HFK77] A. Helenius, E. Fries, and J. Kartenbeck. Reconstitution of Semliki forest virus membrane. Journal of Cell Biology, 75(3):866–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/866.

Hopkins:1977:TLR

[HG77] C. R. Hopkins and H. Gregory. Topographical localization of the receptors for luteinizing hormone-releasing hormone on the surface of dissociated pituitary cells. Journal of Cell Biology, 75 (2):528–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/528.

Hardham:1978:SCM

[HG78] A. R. Hardham and B. E. Gunning. Structure of cortical microtubule arrays in plant cells. Journal of Cell Biology, 77(1):14–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/77/1/14.

Hand:1979:RMD

[HG79] R. Hand and J. R. Gautschi. Replication of mammalian DNA in vitro. Evidence for initiation from fiber autoradiography. Journal of Cell Biology, 82(2):485–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/2/485.

Humbert:1976:MCT

[HGP^{+}76] F. Humbert, A. Grandchamp, C. Pricam, A. Perrelet, and L. Orci. Morphological changes in tight junctions of Necturus maculosus proximal tubules undergoing saline diuresis. Journal of Cell Biology, 69(1):90–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/90.

Hoffstein:1977:RMA

[HGW77] S. Hoffstein, I. M. Goldstein, and G. Weissmann. Role of microtubule assembly in lysosomal enzyme secretion from human polymorphonuclear leukocytes. A reevaluation. Journal of Cell Biology, 73(1):242–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/1/242.

Heath:1976:UMC

[HH76] I. B. Heath and M. C. Heath. Ultrastructure of mitosis in the cowpea rust fungus Uromyces phaseoli var. Vignae. Journal of Cell Biology, 70(3):592–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/592.

Hasty:1977:FFS

[HH77a] D. L. Hasty and E. D. Hay. Freeze-fracture studies of the developing cell surface. I. The plasmalemma of the corneal fibroblast. Journal of Cell Biology, 72(3):667–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/3/667.

Hochli:1977:TLT

[HH77b] M. Höchli and C. R. Hackenbrock. Thermotropic lateral translational motion of intramembrane particles in the inner mitochondrial membrane and its inhibition by artificial peripheral proteins. Journal of Cell Biology, 72(2):278–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/2/278.

Hasty:1978:FFS

[HH78] D. L. Hasty and E. D. Hay. Freeze-fracture studies of the developing cell surface. II. Particle-free membrane blisters on glutaraldehyde-fixed corneal fibroblasts are artefacts. Journal of Cell Biology, 78(3):756–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/756.

Hanna:1976:MSD

[HHP76] R. B. Hanna, A. Hirano, and G. D. Pappas. Membrane specializations of dentritic spines and glia in the weaver mouse cerebellum: a freeze-fracture study. Journal of Cell Biology, 68(3):403–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/403.

Hill:1975:RRM

[Hil75] J. M. Hill. Ribosomal RNA metabolism during renal hypertrophy. Evidence of decreased degradation of newly synthesized ribosomal RNA. Journal of Cell Biology, 64(1):260–??, January 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/1/260.

Howell:1978:ERM

[HIP78] K. E. Howell, A. Ito, and G. E. Palade. Endoplasmic reticulum marker enzymes in Golgi fractions–what does this mean? Journal of Cell Biology, 79(2):581–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/2/581.

Hirokawa:1977:DAE

[Hir77] N. Hirokawa. Disappearance of afferent and efferent nerve terminals in the inner ear of the chick embryo after chronic treatment with beta-bungarotoxin. Journal of Cell Biology, 73(1):27–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/73/1/27.

Hitchcock:1977:RMN

[Hit77] S. E. Hitchcock. Regulation of motility in nonmuscle cells. Journal of Cell Biology, 74(1):1–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/1/1.

Hayashi:1978:PPS

[HJP78] H. Hayashi, H. W. Jarrett, and J. T. Penniston. Peripheral proteins and smooth membrane from erythrocyte ghosts. Segregation of ATP-utilizing enzymes into smooth membrane. Journal of Cell Biology, 76(1):105–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/105.

Heidemann:1975:AFE

[HK75] S. R. Heidemann and M. W. Kirschner. Aster formation in eggs of Xenopus laevis. Induction by isolated basal bodies. Journal of Cell Biology, 67(1):105–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/105.

Hirokawa:1979:BCB

[HK79] N. Hirokawa and M. Kitamura. Binding of Clostridium botulinum neurotoxin to the presynaptic membrane in the central nervous system. Journal of Cell Biology, 81(1):43–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/1/43.

Hedman:1979:IPM

[HKA^{+}79] K. Hedman, M. Kurkinen, K. Alitalo, A. Vaheri, S. Johansson, and M. Höök. Isolation of the pericellular matrix of human fibroblast cultures. Journal of Cell Biology, 81(1):83–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/1/83.

Hanna:1978:FSR

[HKP78] R. B. Hanna, J. S. Keeter, and G. D. Pappas. The fine structure of a rectifying electrotonic synapse. Journal of Cell Biology, 79 (3):764–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/764.

Herth:1977:CFL

[HKS77] W. Herth, A. Kuppel, and E. Schnepf. Chitinous fibrils in the lorica of the flagellate chrysophyte Poteriochromonas stipitata (syn. Ochromonas malhamensis). Journal of Cell Biology, 73(2):311– ??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
2/311.

Hoffman:1975:SCA

[HL75] P. N. Hoffman and R. J. Lasek. The slow component of axonal transport. Identification of major structural polypeptides of the axon and their generality among mammalian neurons. Journal of Cell Biology, 66(2):351–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/2/351.

Hubbard:1979:CAD

[HL79] B. D. Hubbard and E. Lazarides. Copurification of actin and desmin from chicken smooth muscle and their copolymerization in vitro to intermediate filaments. Journal of Cell Biology, 80 (1):166–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/166.

Helantjaris:1976:ECB

[HLG76] T. G. Helantjaris, P. S. Lombardi, and L. A. Glasgow. Effect of cytochalasin B on the adhesion of mouse peritoneal macrophages. Journal of Cell Biology, 69(2):407–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/2/407.

Henkart:1976:SJB

[HLR76] M. Henkart, D. M. Landis, and T. S. Reese. Similarity of junctions between plasma membranes and endoplasmic reticulum in muscle and neurons. Journal of Cell Biology, 70(2):338–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/2/338.

Hanas:1975:MCR

[HLS75] J. Hanas, G. Linden, and K. Stuart. Mitochondrial and cytoplasmic ribosomes and their activity in blood and culture form Trypanosoma brucei. Journal of Cell Biology, 65(1):103–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/1/103.

Hackenbrock:1975:DAS

[HM75] C. R. Hackenbrock and K. J. Miller. The distribution of anionic sites on the surfaces of mitochondrial membranes. Visual probing with polycationic ferritin. Journal of Cell Biology, 65(3):615–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/65/
3/615.

Haigler:1979:DVB

[HMC79a] H. T. Haigler, J. A. McKanna, and S. Cohen. Direct visualization of the binding and internalization of a ferritin conjugate of epidermal growth factor in human carcinoma cells A-431. Journal of Cell Biology, 81(2):382–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/2/382.

Haigler:1979:RSP

[HMC79b] H. T. Haigler, J. A. McKanna, and S. Cohen. Rapid stimulation of pinocytosis in human carcinoma cells A-431 by epidermal growth factor. Journal of Cell Biology, 83(1):82–??, October 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/83/1/82.

Hodge:1977:NMH

[HMDH77] L. D. Hodge, P. Mancini, F. M. Davis, and P. Heywood. Nuclear matrix of HeLa S3 cells. Polypeptide composition during adenovirus infection and in phases of the cell cycle. Journal of Cell Biology, 72(1):194–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/72/1/194.

Head:1979:CBM

[HMK79] J. F. Head, S. Mader, and B. Kaminer. Calcium-binding modulator protein from the unfertilized egg of the sea urchin Arbacia punctulata. Journal of Cell Biology, 80(1):211–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/1/211.

Haksar:1976:LIA

[HMPB76] A. Haksar, D. V. Maudsley, F. G. Péron, and E. Bedigian. Lanthanum: inhibition of ACTH-stimulated cyclic AMP and corticosterone synthesis in isolated rat adrenocortical cells. Journal of Cell Biology, 68(1):142–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/1/142.

Hazelton:1979:CMG

[HMT79] B. Hazelton, B. Mitchell, and J. Tupper. Calcium, magnesium, and growth control in the WI-38 human fibroblast cell. Journal of Cell Biology, 83(2):487–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/2/487.

Heidrich:1975:DAE

[HO75] H. G. Heidrich and W. L. Olsen. Deoxyribonucleic acid-envelope complexes from Escherichia coli. A complex-specific protein and its possible function for the stability of the complex. Journal of Cell Biology, 67(2):444–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/2/444.

Hand:1977:RBG

[HO77] A. R. Hand and C. Oliver. Relationship between the Golgi apparatus, GERL, and secretory granules in acinar cells of the rat exorbital lacrimal gland. Journal of Cell Biology, 74(2):399–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/74/
2/399.

Hopkins:1977:STK

[Hop77] C. R. Hopkins. Short term kinetics of luteinizing hormone secretion studied in dissociated pituitary cells attached to manipulable substrates. Journal of Cell Biology, 73(3):685–??, June 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/73/3/685.

Huang:1975:IPV

[HP75] L. Huang and R. E. Pagano. Interaction of phospholipid vesicles with cultured mammalial cells. I. Characteristics of uptake. Journal of Cell Biology, 67(1):38–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/38.

Hilfer:1977:ITH

[HP77] S. R. Hilfer and G. L. Pakstis. Interference with thyroid histogenesis by inhibitors of collagen synthesis. Journal of Cell Biology, 75 (2):446–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/446.

Herman:1979:CPA

[HP79] I. M. Herman and T. D. Pollard. Comparison of purified anti-actin and fluorescent-heavy meromyosin staining patterns in dividing cells. Journal of Cell Biology, 80(3):509–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/80/3/509.

Howell:1977:CCP

[HPH77] S. H. Howell, J. W. Posakony, and K. R. Hill. The cell cycle program of polypeptide labeling in Chlamydomonas reinhardtii. Journal of Cell Biology, 72(2):223–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/2/223.

Hogsett:1978:SFF

[HQ78] W. E. Hogsett and R. S. Quatrano. Sulfation of fucoidin in focus embryos: III. Required for localization in the rhizoid wall. Journal of Cell Biology, 78(3):866–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/866.

Hayman:1979:DFB

[HR79a] E. G. Hayman and E. Ruoslahti. Distribution of fetal bovine serum fibronectin and endogenous rat cell fibronectin in extracellular matrix. Journal of Cell Biology, 83(1):255–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/1/255.

Hogg:1979:CAE

[HR79b] N. M. Hogg and B. Rotman. Covalent attachment of enzyme as a membrane-label for viable eucaryotic cells. Journal of Cell Biology, 83(2):511–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/83/2/511.

Heuser:1979:SVE

[HRD^{+}79] J. E. Heuser, T. S. Reese, M. J. Dennis, Y. Jan, L. Jan, and L. Evans. Synaptic vesicle exocytosis captured by quick freezing and correlated with quantal transmitter release. Journal of Cell Biology, 81(2):275–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/2/275.

Huang:1977:TSM

[HRL77] B. Huang, M. R. Rifkin, and D. J. Luck. Temperature-sensitive mutations affecting flagellar assembly and function in Chlamydomonas reinhardtii. Journal of Cell Biology, 72(1):67–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/72/1/67.

Herrup:1975:PBN

[HS75a] K. Herrup and E. M. Shooter. Properties of the beta-nerve growth factor receptor in development. Journal of Cell Biology, 67(1):118–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/1/118.

Hsueh:1975:ESM

[HS75b] H. W. Hsueh and F. E. Stockdale. Effects of serum on mammary epithelial proliferation in vitro during mammary gland development. Journal of Cell Biology, 66(2):243–??, August 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/2/243.

Hartwig:1976:IAM

[HS76a] J. H. Hartwig and T. P. Stossel. Interactions of actin, myosin, and an actin-binding protein of rabbit pulmonary macrophages. III. Effects of cytochalasin B. Journal of Cell Biology, 71(1):295–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
1/295.

Hollinger:1976:CCG

[HS76b] T. G. Hollinger and A. W. Schuetz. ”cleavage” and cortical granule breakdown in Rana pipiens oocytes induced by direct microinjection of calcium. Journal of Cell Biology, 71(2):395–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
2/395.

Hoober:1976:KRS

[HS76c] J. K. Hoober and W. J. Stegeman. Kinetics and regulation of synthesis of the major polypeptides of thylakoid membranes in Chlamydomonas reinhardtii y-1 at elevated temperatures. Journal of Cell Biology, 70(2):326–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/2/326.

Hull:1976:FSV

[HS76d] B. E. Hull and L. A. Staehelin. Functional significance of the variations in the geometrical organization of tight junction networks. Journal of Cell Biology, 68(3):688–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/688.

Hallberg:1977:NRS

[HS77] R. L. Hallberg and C. A. Sutton. Nonidentity of ribosomal structural proteins in growing and starved Tetrahymena. Journal of Cell Biology, 75(1):268–??, October 1977. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/1/268.

Heuser:1979:OAR

[HS79a] J. E. Heuser and S. R. Salpeter. Organization of acetylcholine receptors in quick-frozen, deep-etched, and rotary-replicated Torpedo postsynaptic membrane. Journal of Cell Biology, 82(1):150– ??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
1/150.

Hubbard:1979:EMAb

[HS79b] A. L. Hubbard and H. Stukenbrok. An electron microscope autoradiographic study of the carbohydrate recognition systems in rat liver. II. Intracellular fates of the 125I-ligands. Journal of Cell Biology, 83(1):65–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/83/1/65.

Hull:1979:TWR

[HS79c] B. E. Hull and L. A. Staehelin. The terminal web. A reevaluation of its structure and function. Journal of Cell Biology, 81(1):67–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/81/1/67.

Hoffstein:1976:CIM

[HSGW76] S. Hoffstein, R. Soberman, I. Goldstein, and G. Weissmann. Concanavalin A induces microtubule assembly and specific granule discharge in human polymorphonuclear leukocytes. Journal of Cell Biology, 68(3):781–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/68/3/781.

Herzog:1976:ESC

[HSM76] V. Herzog, H. Sies, and F. Miller. Exocytosis in secretory cells of rat lacrimal gland. Peroxidase release from lobules and isolated cells upon cholinergic stimulation. Journal of Cell Biology, 70 (3):692–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/692.

Horovitch:1979:MAD

[HSRP79] S. J. Horovitch, R. V. Storti, A. Rich, and M. L. Pardue. Multiple actins in Drosophila melanogaster. Journal of Cell Biology, 82(1):86–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/86.

Hazelton:1979:CTE

[HT79a] B. J. Hazelton and J. T. Tupper. Calcium transport and exchange in mouse 3T3 and SV40-3T3 cells. Journal of Cell Biology, 81(3):538–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/538.

Hellewell:1979:CBA

[HT79b] S. B. Hellewell and D. L. Taylor. The contractile basis of ameboid movement. VI. The solation-contraction coupling hypothesis. Journal of Cell Biology, 83(3):633–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/3/633.

Hartwell:1977:UDS

[HU77] L. H. Hartwell and M. W. Unger. Unequal division in Saccharomyces cerevisiae and its implications for the control of cell division. Journal of Cell Biology, 75(2):422–??, November 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/75/2/422.

Havell:1975:IIS

[HV75] E. A. Havell and J. Vilcek. Inhibition of interferon secretion by vinblastine. Journal of Cell Biology, 64(3):716–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/3/716.

Hedman:1978:EFC

[HVW78] K. Hedman, A. Vaheri, and J. Wartiovaara. External fibronectin of cultured human fibroblasts is predominantly a matrix protein. Journal of Cell Biology, 76(3):748–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/76/3/748.

Hedley-Whyte:1975:DMB

[HW75] E. T. Hedley-Whyte. Distribution of [1,2-3H]cholesterol in mouse brain after injection in the suckling period. Journal of Cell Biology, 66(2):333–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/2/333.

Hoffstein:1978:MMC

[HW78] S. Hoffstein and G. Weissmann. Microfilaments and microtubules in calcium ionophore-induced secretion of lysosomal enzymes from human polymorphonuclear leukocytes. Journal of Cell Biology, 78 (3):769–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/769.

Hubbard:1979:EMAa

[HWAS79] A. L. Hubbard, G. Wilson, G. Ashwell, and H. Stukenbrok. An electron microscope autoradiographic study of the carbohydrate recognition systems in rat liver. I. Distribution of 125I-ligands among the liver cell types. Journal of Cell Biology, 83(1):47–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/47.

Horwitz:1978:ILA

[HWLC78] A. F. Horwitz, A. Wight, P. Ludwig, and R. Cornell. Interrelated lipid alterations and their influence on the proliferation and fusion of cultured myogenic cells. Journal of Cell Biology, 77(2):334–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/77/
2/334.

Herman:1978:HNR

[HWP78] R. Herman, L. Weymouth, and S. Penman. Heterogeneous nuclear RNA-protein fibers in chromatin-depleted nuclei. Journal of Cell Biology, 78(3):663–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/663.

Hale:1975:CSC

[HWW75] A. H. Hale, J. L. Winkelhake, and M. J. Weber. Cell surface changes and Rous sarcoma virus gene expression in synchronized cells. Journal of Cell Biology, 64(2):398–??, February 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/2/398.

Iwanij:1975:STR

[ICS75] V. Iwanij, N. H. Chua, and P. Siekevitz. Synthesis and turnover of ribulose biphosphate carboxylase and of its subunits during the cell cycle of Chlamydomonas reinhardtii. Journal of Cell Biology, 64(3):572–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/572.

Ip:1979:HRS

[IF79] W. Ip and D. A. Fischman. High resolution scanning electron microscopy of isolated and in situ cytoskeletal elements. Journal of Cell Biology, 83(1):249–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/1/249.

Ishiura:1979:RAT

[IO79] M. Ishiura and Y. Okada. The role of actin in temperaturedependent gel-sol transformation of extracts of Ehrlich ascites tumor cells. Journal of Cell Biology, 80(2):465–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/2/465.

Ito:1978:PNC

[IP78a] A. Ito and G. E. Palade. Presence of nadph-cytochrome P-450 reductase in rat liver Golgi membranes: Evidence obtained by immunoadsorption method. Journal of Cell Biology, 79(2):590–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
2/590.

Iwatsuki:1978:ECU

[IP78b] N. Iwatsuki and O. H. Petersen. Electrical coupling and uncoupling of exocrine acinar cells. Journal of Cell Biology, 79 (2):533–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/533.

Inoue:1978:MBI

[IR78] S. Inoué and H. Ritter. Mitosis in Barbulanympha. II. Dynamics of a two-stage anaphase, nuclear morphogenesis, and cytokinesis. Journal of Cell Biology, 77(3):655–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/655.

Ingwall:1976:RCR

[IW76] J. S. Ingwall and K. Wildenthal. Roles of creatine in the regulation of cardiac protein synthesis. Journal of Cell Biology, 68(1):159–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
1/159.

Jauker:1975:FCC

[Jau75] F. Jauker. A feedback control of cell cycle parameters in Tetrahymena. Journal of Cell Biology, 67(3):901–??, December 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/3/901.

Johnson:1978:MBP

[JB78] G. R. Johnson and A. W. Burgess. Molecular and biological properties of a macrophage colony-stimulating factor from mouse yolk sacs. Journal of Cell Biology, 77(1):35–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/1/35.

Jarasch:1977:RCM

[JBKF77] E. D. Jarasch, G. Bruder, T. W. Keenan, and W. W. Franke. Redox constituents in milk fat globule membranes and rough endoplasmic reticulum from lactating mammary gland. Journal of Cell Biology, 73(1):223–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/1/223.

Jahromi:1979:TSS

[JC79] S. S. Jahromi and M. P. Charlton. Transverse sarcomere splitting. A possible means of longitudinal growth in crab muscles. Journal of Cell Biology, 80(3):736–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/3/736.

Jimbow:1975:CDP

[JF75] K. Jimbow and T. B. Fitzpatrick. Changes in distribution pattern of cytoplasmic filaments in human melanocytes during ultravioletmediated melanin pigmentation. The role of the 100-Angstrom filaments in the elongation of melanocytic dendrites and in the movement and transfer of melanosomes. Journal of Cell Biology, 65(2):481–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/481.

Johmann:1976:IEA

[JG76] C. A. Johmann and M. A. Gorovsky. Immunofluorescence evidence for the absence of histone H1 in a mitotically dividing, genetically inactive nucleus. Journal of Cell Biology, 71(1):89–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
1/89.

Jope:1978:EAC

[JHW78] C. A. Jope, A. Hirai, and S. G. Wildman. Evidence that the amount of chloroplast DNA exceeds that of nuclear DNA in mature leaves. Journal of Cell Biology, 79(3):631–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/3/631.

Jeon:1975:CFC

[JJ75] K. W. Jeon and M. S. Jeon. Cytoplasmic filaments and cellular wound healing in amoeba proteus. Journal of Cell Biology, 67 (1):243–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/1/243.

James:1978:VFC

[JJ78] T. W. James and C. Jope. Visualization by fluorescence of chloroplast DNA in higher plants by means of the DNA-specific probe 4’6-diamidino-2-phenylindole. Journal of Cell Biology, 79 (3):623–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/623.

Jockusch:1979:NFC

[JJB79] H. Jockusch, B. M. Jockusch, and M. M. Burger. Nerve fibers in culture and their interactions with non-neural cells visualized by immunofluorescence. Journal of Cell Biology, 80(3):629–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/629.

Johnson:1977:RPH

[JK77] T. R. Johnson and A. Kumar. Ribosome processing in HeLa cells. Studies on structural aspects of precursor and mature ribosomes. Journal of Cell Biology, 73(2):419–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/2/419.

Jarasch:1979:BTC

[JKB^{+}79] E. D. Jarasch, J. Kartenbeck, G. Bruder, A. Fink, D. J. Morré, and W. W. Franke. B-type cytochromes in plasma membranes isolated from rat liver, in comparison with those of endomembranes. Journal of Cell Biology, 80(1):37–??, January 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/1/37.

Jorgensen:1979:LSR

[JKM79] A. O. Jorgensen, V. Kalnins, and D. H. MacLennan. Localization of sarcoplasmic reticulum proteins in rat skeletal muscle by immunofluorescence. Journal of Cell Biology, 80(2):372–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
2/372.

Jorgensen:1977:ASR

[JKZM77] A. O. Jorgensen, V. I. Kalnins, E. Zubrzyca, and D. H. MacLennan. Assembly of the sarcoplasmic reticulum proteins in differentiating rat skeletal muscle cell cultures: localization by immunofluorescence of sarcoplasmic reticulum proteins in differentiating rat skeletal muscle cell cultures. Journal of Cell Biology, 74(1):287– ??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/74/
1/287.

Jacob:1979:LAR

[JL79] M. Jacob and T. L. Lentz. Localization of acetylcholine receptors by means of horseradish peroxidase-alpha-bungarotoxin during formation and development of the neuromuscular junction in the chick embryo. Journal of Cell Biology, 82(1):195–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/1/195.

Johnson:1976:CRR

[JLA^{+}76] L. F. Johnson, R. Levis, H. T. Abelson, H. Green, and S. Penman. Changes in RNA in relation to growth of the fibroblast. IV. Alterations in theproduction and processing of mRNA and rRNA in resting and growing cells. Journal of Cell Biology, 71 (3):933–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/3/933.

Jones:1977:ECA

[JM77] R. E. Jones and A. A. Moscona. Effects of cytosine arabinoside on differential gene expression in embryonic neural retina. II. Immunochemical studies on the accumulation of glutamine synthetase. Journal of Cell Biology, 74(1):30–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/1/30.

Jakoi:1979:LEC

[JM79] E. R. Jakoi and R. B. Marchase. Ligatin from embryonic chick neural retina. Journal of Cell Biology, 80(3):642–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/3/642.

Jimbow:1975:MAN

[JRFS75] K. Jimbow, S. I. Roth, T. B. Fitzpatrick, and G. Szabo. Mitotic activity in non-neoplastic melanocytes in vivo as determined by histochemical, autoradiographic, and electron microscope studies. Journal of Cell Biology, 66(3):663–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/3/663.

Joseph:1979:ECT

[JSG79] K. C. Joseph, A. Stieber, and N. K. Gonatas. Endocytosis of cholera toxin in GERL-like structures of murine neuroblastoma cells pretreated with GM1 ganglioside. Cholera toxin internalization into Neuroblastoma GERL. Journal of Cell Biology, 81(3):543–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/543.

Jakovcic:1978:BSA

[JSGR78] S. Jakovcic, H. H. Swift, N. J. Gross, and M. Rabinowitz. Biochemical and stereological analysis of rat liver mitochondria in different thyroid states. Journal of Cell Biology, 77(3):887–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/3/887.

Jasny:1979:TOR

[JT79] B. R. Jasny and I. Tamm. Temporal organization of replication in DNA fibers of mammalian cells. Journal of Cell Biology, 81(3):692–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/692.

Juliano:1978:ADC

[Jul78] R. L. Juliano. Adhesion and detachment characteristics of Chinese hamster cell membrane mutants. Journal of Cell Biology, 76(1):43–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/1/43.

Jaffe:1975:CAW

[JWJ75] L. A. Jaffe, M. H. Weisenseel, and L. F. Jaffe. Calcium accumulations within the growing tips of pollen tubes. Journal of Cell Biology, 67(2):488–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/2/488.

Jakoi:1976:RSU

[JZR76] E. R. Jakoi, G. Zampighi, and J. D. Robertson. Regular structures in unit membranes. II. Morphological and biochemical characterization of two water-soluble membrane proteins isolated from the suckling rat ileum. Journal of Cell Biology, 70(1):97–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/1/97.

Kellems:1975:CTR

[KAB75] R. E. Kellems, V. F. Allison, and R. A. Butow. Cytoplasmic type 80S ribosomes associated with yeast mitochondria. IV. Attachment of ribosomes to the outer membrane of isolated mitochondria. Journal of Cell Biology, 65(1):1–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/1/1.

Kang:1975:MUM

[KAD75] Y. H. Kang, W. A. Anderson, and E. R. DeSombre. Modulation of uterine morphology and growth by estradiol-17beta and an estrogen antagonist. Journal of Cell Biology, 64(3):682–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/3/682.

Kruth:1979:ECD

[KAGV79] H. S. Kruth, J. Avigan, W. Gamble, and M. Vaughan. Effect of cell density on binding and uptake of low density lipoprotein by human fibroblasts. Journal of Cell Biology, 83(3):588–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
3/588.

Kahn:1976:MRH

[Kah76] C. R. Kahn. Membrane receptors for hormones and neurotransmitters. Journal of Cell Biology, 70(2):261–??, August 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/70/2/261.

Kane:1975:PPP

[Kan75] R. E. Kane. Preparation and purification of polymerized actin from sea urchin egg extracts. Journal of Cell Biology, 66(2): 305–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/305.

Kane:1976:API

[Kan76] R. E. Kane. Actin polymerization and interaction with other proteins in temperature-induced gelation of sea urchin egg extracts. Journal of Cell Biology, 71(3):704–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/704.

Kensler:1977:NFV

[KBD77] R. W. Kensler, P. Brink, and M. M. Dewey. Nexus of frog ventricle. Journal of Cell Biology, 73(3):768–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/3/768.

Kim:1979:PAM

[KBR79] H. Kim, L. I. Binder, and J. L. Rosenbaum. The periodic association of MAP2 with brain microtubules in vitro. Journal of Cell Biology, 80(2):266–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/2/266.

Kurek:1979:SCR

[KBS79] M. P. Kurek, D. Billig, and P. Stambrook. Size classes of replication units in DNA from sea urchin embryos. Journal of Cell Biology, 81(3):698–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/698.

Kelly:1978:SPC

[KC78] P. T. Kelly and C. W. Cotman. Synaptic proteins. Characterization of tubulin and actin and identification of a distinct postsynaptic density polypeptide. Journal of Cell Biology, 79(1):173–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
1/173.

Kelly:1976:DML

[KCGN76] P. Kelly, C. W. Cotman, C. Gentry, and G. L. Nicolson. Distribution and mobility of lectin receptors on synaptic membranes of identified neurons in the central nervous system. Journal of Cell Biology, 71(2):487–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/2/487.

Kaysen:1979:RZD

[KCH79] G. A. Kaysen, L. Y. Chou, and M. H. Humphreys. Requirement of Zn to demonstrate HCO3-stimulated ATPase activity of rat small intestinal brush border. Journal of Cell Biology, 82 (3):780–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/780.

Kannan:1978:FGJ

[KD78] M. S. Kannan and E. E. Daniel. Formation of gap junctions by treatment in vitro with potassium conductance blockers. Journal of Cell Biology, 78(2):338–??, August 1978. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/338.

Kalderon:1977:CCC

[KEG77] N. Kalderon, M. L. Epstein, and N. B. Gilula. Cell-to-cell communication and myogenesis. Journal of Cell Biology, 75(3):788–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/788.

Karnaky:1976:TCCa

[KEP76] K. J. Karnaky, S. A. Ernst, and C. W. Philpott. Teleost chloride cell. I. Response of pupfish Cyprinodon variegatus gill Na,K– ATPase and chloride cell fine structure to various high salinity environments. Journal of Cell Biology, 70(1):144–??, July 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/70/1/144.

Kanwar:1979:ASG

[KF79] Y. S. Kanwar and M. G. Farquhar. Anionic sites in the glomerular basement membrane. In vivo and in vitro localization to the laminae rarae by cationic probes. Journal of Cell Biology, 81(1):137– ??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
1/137.

Kreibich:1978:PRMb

[KFP^{+}78] G. Kreibich, C. M. Freienstein, B. N. Pereyra, B. L. Ulrich, and D. D. Sabatini. Proteins of rough microsomal membranes related to ribosome binding. II. Cross-linking of bound ribosomes to specific membrane proteins exposed at the binding sites. Journal of Cell Biology, 77(2):488–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/488.

Kramer:1977:GTS

[KG77] M. F. Kramer and J. J. Geuze. Glycoprotein transport in the surface mucous cells of the rat stomach. Journal of Cell Biology, 73(3):533–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/3/533.

Kater:1978:PME

[KG78] S. B. Kater and N. J. Galvin. Physiological and morphological evidence for coupling in mouse salivary gland acinar cells. Journal of Cell Biology, 79(1):20–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/1/20.

Kalderon:1979:MEI

[KG79] N. Kalderon and N. B. Gilula. Membrane events involved in myoblast fusion. Journal of Cell Biology, 81(2):411–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/411.

King:1975:ALF

[KH75] G. J. King and M. E. Holtrop. Actin-like filaments in bone cells of cultured mouse calvaria as demonstrated by binding to heavy meromyosin. Journal of Cell Biology, 66(2):445–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/2/445.

Kitamura:1976:MRM

[KH76] A. Kitamura and K. Hiwatashi. Mating-reactive membrane vesicles from cilia of Paramecium caudatum. Journal of Cell Biology, 69(3):736–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/736.

Kirschner:1979:SSM

[KHT^{+}79] D. A. Kirschner, C. J. Hollingshead, C. Thaxton, D. L. Caspar, and D. A. Goodenough. Structural states of myelin observed by X-ray diffraction and freeze-fracture electron microscopy. Journal of Cell Biology, 82(1):140–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/1/140.

Karlin:1978:BAA

[KHV^{+}78] A. Karlin, E. Holtzman, R. Valderrama, V. Damle, K. Hsu, and F. Reyes. Binding of antibodies to acetylcholine receptors in Electrophorus and Torpedo electroplax membranes. Journal of Cell Biology, 76(3):577–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/76/3/577.

Krishnan:1979:HSN

[KKAL79] N. Krishnan, I. R. Kaiserman-Abramof, and R. J. Lasek. Helical substructure of neurofilaments isolated from Myxicola and squid giant axons. Journal of Cell Biology, 82(2):323–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/2/323.

Kadota:1976:CVS

[KKG76] T. Kadota, K. Kadota, and E. G. Gray. Coated-vesicle shells, particle/chain material, and tubulin in brain synaptosomes. An electron microscope and biochemical study. Journal of Cell Biology, 69(3):608–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/608.

Kuroiwa:1977:SMS

[KKH77] T. Kuroiwa, S. Kawano, and M. Hizume. Studies on mitochondrial structure and function in Physarum polycephalum. V. Behaviour of mitochondrial nucleoids throughout mitochondrial division cycle. Journal of Cell Biology, 72(3):687–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/3/687.

Karnaky:1976:TCCb

[KKKS76] K. J. Karnaky, L. B. Kinter, W. B. Kinter, and C. E. Stirling. Teleost chloride cell. II. Autoradiographic localization of gill Na,K–ATPase in killifish Fundulus heteroclitus adapted to low and high salinity environments. Journal of Cell Biology, 70(1):157–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/1/157.

Katz:1979:TBV

[KL79] F. N. Katz and H. F. Lodish. Transmembrane biogenesis of the vesicular stomatitis virus glycoprotein. Journal of Cell Biology, 80(2):416–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/416.

Kurth:1978:HLP

[KMB78] P. D. Kurth, E. N. Moudrianakis, and M. Bustin. Histone localization in polytene chromosomes by immunofluorescence. Journal of Cell Biology, 78(3):910–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/910.

Karran:1977:DDD

[KMS77] P. Karran, A. Moscona, and B. Strauss. Developmental decline in DNA repair in neural retina cells of chick embryos: persistent deficiency of repair competence in a cell line derived from late. Journal of Cell Biology, 74(1):274–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/1/274.

Kessler:1976:ACP

[KNL76] D. Kessler, V. T. Nachmias, and A. G. Loewy. Actomyosin content of Physarum plasmodia and detection of immunological crossreactions with myosins from related species. Journal of Cell Biology, 69(2):393–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/393.

Kinoshita:1979:IHP

[KNM79] T. Kinoshita, R. L. Nachman, and R. Minick. Isolation of human platelet plasma membranes with polylysine beads. Journal of Cell Biology, 82(3):688–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/3/688.

Kishida:1975:TIM

[KOBP75] Y. Kishida, B. R. Olsen, R. A. Berg, and D. J. Prockop. Two improved methods for preparing ferritin-protein conjugates for electron microscopy. Journal of Cell Biology, 64(2):331–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/2/331.

Koehler:1975:SDA

[Koe75] J. K. Koehler. Studies on the distribution of antigenic sites on the surface of rabbit spermatozoa. Journal of Cell Biology, 67 (3):647–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/3/647.

Kogon:1975:AGJ

[KP75] M. Kogon and G. D. Pappas. Atypical gap junctions in the ciliary epithelium of the albino rabbit eye. Journal of Cell Biology, 66 (3):671–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/671.

Kalina:1977:CAI

[KP77a] M. Kalina and D. C. Pease. Cyclic amp-induced morphological transformation of cells infected by temperature-sensitive mouse sarcoma virus: Expression of transformation-associated markers. Journal of Cell Biology, 74(3):707–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/74/3/707.

Kalina:1977:PUS

[KP77b] M. Kalina and D. C. Pease. The preservation of ultrastructure in saturated phosphatidyl cholines by tannic acid in model systems and type II pneumocytes. Journal of Cell Biology, 74(3): 726–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/726.

Kalina:1977:PRP

[KP77c] M. Kalina and D. C. Pease. The probable role of phosphatidyl cholines in the tannic acid enhancement of cytomembrane electron contrast. Journal of Cell Biology, 74(3):742–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/3/742.

Krishan:1975:CSA

[KPF75] A. Krishan, K. Paika, and E. Frei. Cytofluorometric studies on the action of podophyllotoxin and epipodophyllotoxins (VM-26, VP-16-213) on the cell cycle traverse of human lymphoblasts. Journal of Cell Biology, 66(3):521–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/3/521.

Karnovsky:1975:SGS

[KR75] M. J. Karnovsky and G. B. Ryan. Substructure of the glomerular slit diaphragm in freeze-fractured normal rat kidney. Journal of Cell Biology, 65(1):233–??, April 1975. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/1/233.

Kuczmarski:1979:CBA

[KR79a] E. R. Kuczmarski and J. L. Rosenbaum. Chick brain actin and myosin. Isolation and characterization. Journal of Cell Biology, 80(2):341–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/341.

Kuczmarski:1979:SOL

[KR79b] E. R. Kuczmarski and J. L. Rosenbaum. Studies on the organization and localization of actin and myosin in neurons. Journal of Cell Biology, 80(2):356–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/2/356.

Kraehenbuhl:1977:DMG

[Kra77] J. P. Kraehenbuhl. Dispersed mammary gland epithelial cells. I. Isolation and separation procedures. Journal of Cell Biology, 72 (2):390–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/2/390.

Krishan:1975:RFC

[Kri75] A. Krishan. Rapid flow cytofluorometric analysis of mammalian cell cycle by propidium iodide staining. Journal of Cell Biology, 66(1):188–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/1/188.

Kraehenbuhl:1977:ILS

[KRJ77] J. P. Kraehenbuhl, L. Racine, and J. D. Jamieson. Immunocytochemical localization of secretory proteins in bovine pancreatic exocrine cells. Journal of Cell Biology, 72(2):406–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/2/406.

Kirschner:1977:CNE

[KRM77] R. H. Kirschner, M. Rusli, and T. E. Martin. Characterization of the nuclear envelope, pore complexes, and dense lamina of mouse liver nuclei by high resolution scanning electron microscopy. Journal of Cell Biology, 72(1):118–??, January 1977.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/1/118.

Komiyama:1975:MEL

[KS75] A. Komiyama and S. S. Spicer. Microendocytosis in eosinophilic leukocytes. Journal of Cell Biology, 64(3):622–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/3/622.

Kruppa:1977:RPM

[KS77] J. Kruppa and D. D. Sabatini. Release of poly a(+) messenger RNA from rat liver rough microsomes upon disassembly of bound polysomes. Journal of Cell Biology, 74(2):414–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/414.

Kahn:1979:CTN

[KS79] P. Kahn and S. I. Shin. Cellular tumorigenicity in nude mice. Test of associations among loss of cell-surface fibronectin, anchorage independence, and tumor-forming ability. Journal of Cell Biology, 82(1):1–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/1.

Kusamrarn:1975:MFI

[KSB75] T. Kusamrarn, P. Sobhon, and G. B. Bailey. The mechanism of formation of inhibitor-induced ribosome helices in Entamoeba invadens. Journal of Cell Biology, 65(3):529–??, June 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/65/3/529.

Knutton:1975:RMS

[KSP75] S. Knutton, M. C. Sumner, and C. A. Pasternak. Role of microvilli in surface changes of synchronized P815Y mastocytoma cells. Journal of Cell Biology, 66(3):568–??, September 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/3/568.

Kahri:1976:EBA

[KSS76] A. I. Kahri, M. Salmenperä, and A. Saure. Effects of 5bromodeoxyuridine on the ACTH-dependent mitochondrial biogenesis in cortical cells of fetal rat adrenals in tissue culture. Journal of Cell Biology, 71(3):951–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/951.

Kierszenbaum:1975:STF

[KT75] A. L. Kierszenbaum and L. L. Tres. Structural and transcriptional features of the mouse spermatid genome. Journal of Cell Biology, 65(2):258–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/258.

Kitajima:1977:TSM

[KT77a] Y. Kitajima and G. A. Thompson. Tetrahymena strives to maintain the fluidity interrelationships of all its membranes constant. Electron microscope evidence. Journal of Cell Biology, 72(3):744– ??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
3/744.

Kitajima:1977:DFV

[KT77b] Y. Kitajima and G. A. Thompson. Differentiation of food vacuolar membranes during endocytosis in Tetrahymena. Journal of Cell Biology, 75(2):436–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/2/436.

Kreibich:1978:PRMa

[KUS78] G. Kreibich, B. L. Ulrich, and D. D. Sabatini. Proteins of rough microsomal membranes related to ribosome binding. I. Identification of ribophorins I and II, membrane proteins characteristics of rough microsomes. Journal of Cell Biology, 77(2):464–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/2/464.

Kusamrarn:1975:CSF

[KVB75] T. Kusamrarn, K. Vinijchaikul, and G. B. Bailey. Comparision of the structure and function of polysomal and helical ribosomes from Entamoeba invadens. Journal of Cell Biology, 65(3):540–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/65/
3/540.

Kersey:1976:LAF

[KW76] Y. M. Kersey and N. K. Wessells. Localization of actin filaments in internodal cells of characean algae. A scanning and transmission electron microscope study. Journal of Cell Biology, 68(2):264–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
2/264.

Kyte:1976:IDDa

[Kyt76a] J. Kyte. Immunoferritin determination of the distribution of (Na^{+} + K^{+}) ATPase over the plasma membranes of renal convoluted tubules. I. Distal segment. Journal of Cell Biology, 68(2):287–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
2/287.

Kyte:1976:IDDb

[Kyt76b] J. Kyte. Immunoferritin determination of the distribution of (Na^{+} + K^{+}) ATPase over the plasma membranes of renal convoluted tubules. II. Proximal segment. Journal of Cell Biology, 68 (2):304–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/304.

Lustig:1977:CBM

[LAF^{+}77] S. Lustig, O. Ascher, P. Fishman, M. Djaldetti, and D. H. Pluznik. Correlation between movement of concanavalin A membrane receptors and cytolysis. A scanning electron microscopy study. Journal of Cell Biology, 75(2):388–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/2/388.

Landis:1978:GCC

[Lan78] S. C. Landis. Growth cones of cultured sympathetic neurons contain adrenergic vesicles. Journal of Cell Biology, 78(1):??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/1/R8.

Larsen:1975:ODG

[Lar75] W. J. Larsen. Opaque deposits on gap junction membranes after glutaraldehyde-calcium fixation. Journal of Cell Biology, 67 (3):801–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/3/801.

Lande:1975:DAM

[LAS^{+}75] M. A. Lande, M. Adesnik, M. Sumida, Y. Tashiro, and D. D. Sabatini. Direct association of messenger RNA with microsomal membranes in human diploid fibroblasts. Journal of Cell Biology, 65(3):513–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/3/513.

Lazarides:1975:TAS

[Laz75] E. Lazarides. Tropomyosin antibody: the specific localization of tropomyosin in nonmuscle cells. Journal of Cell Biology, 65(3):549–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/3/549.

Lazarides:1976:AAA

[Laz76] E. Lazarides. Actin, alpha-actinin, and tropomyosin interaction in the structural organization of actin filaments in nonmuscle cells. Journal of Cell Biology, 68(2):202–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/68/2/202.

Lang:1978:SRM

[LB78] R. D. Lang and J. R. Bronk. A study of rapid mitochondrial structural changes in vitro by spray-freeze-etching. Journal of Cell Biology, 77(1):134–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/1/134.

Lauweryns:1976:FFL

[LBDL76] J. M. Lauweryns, J. Baert, and W. De-Loecker. Fine filaments in lymphatic endothelial cells. Journal of Cell Biology, 68(1):163–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
1/163.

Lagunoff:1976:ECR

[LC76] D. Lagunoff and E. Y. Chi. Effect of colchicine on rat mast cells. Journal of Cell Biology, 71(1):182–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/182.

Lentz:1977:LAR

[LC77] T. L. Lentz and J. Chester. Localization of acetylcholine receptors in central synapses. Journal of Cell Biology, 75(1):258–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
1/258.

Leighton:1975:SCP

[LCK75] F. Leighton, L. Coloma, and C. Koenig. Structure, composition, physical properties, and turnover of proliferated peroxisomes. A study of the trophic effects of Su-13437 on rat liver. Journal of Cell Biology, 67(2):281–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/2/281.

Lord:1976:RSJ

[LD76] B. A. Lord and D. R. DiBona. Role of the septate junction in the regulation of paracellular transepithelial flow. Journal of Cell Biology, 71(3):967–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/3/967.

Lonn:1976:MRV

[LE76] U. Lönn and J. E. Edström. Mobility restriction in vivo of the heavy ribosomal subunit in a secretory cell. Journal of Cell Biology, 70(3):573–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/3/573.

Lonn:1977:MAR

[LE77] U. Lonn and J. E. Edstrom. Movements and associations of ribosomal subunits in a secretory cell during growth inhibition by starvation. Journal of Cell Biology, 73(3):696–??, June 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/73/3/696.

Lee:1978:FPS

[LE78a] G. T. Lee and D. L. Engelhardt. Fluctuations in the production of specific cellular peptides during the growth of animal cells. Journal of Cell Biology, 78(2):??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/R28.

Lee:1978:GRF

[LE78b] G. T. Lee and D. L. Engelhardt. Growth-related fluctuation in messenger RNA utilization in animal cells. Journal of Cell Biology, 79(1):85–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/1/85.

Levine:1976:PIM

[LEDW76] R. J. Levine, M. Elfvin, M. M. Dewey, and B. Walcott. Paramyosin in invertebrate muscles. II. Content in relation to structure and function. Journal of Cell Biology, 71(1):273–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
1/273.

Lemanski:1979:RTA

[Lem79] L. F. Lemanski. Role of tropomyosin in actin filament formation in embryonic salamander heart cells. Journal of Cell Biology, 82(1):227–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/227.

Leon:1976:MHI

[Leó76] P. E. León. Molecular hybridization of iodinated 4S, 5S, and 18S + 28S RNA to salamander chromosomes. Journal of Cell Biology, 69(2):287–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/2/287.

Letourneau:1979:IIA

[Let79] P. C. Letourneau. Inhibition of intercellular adhesion by concanavalin A is associated with concanavalin A-mediated redistribution of surface receptors. Journal of Cell Biology, 80(1):128–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
1/128.

Ley:1975:DGP

[Ley75] K. D. Ley. Detection of G1 proteins in Chinese hamster cells synchronized by isoleucine deprivation or mitotic selection. Journal of Cell Biology, 66(1):95–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/95.

Levine:1976:IIM

[LF76] R. F. Levine and M. E. Fedorko. Isolation of intact megakaryocytes from guinea pig femoral marrow. Successful harvest made possible with inhibitions of platelet aggregation; enrichment achieved with a two-step separation technique. Journal of Cell Biology, 69(1):159–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/1/159.

Lodish:1977:BVG

[LF77] H. F. Lodish and S. Froshauer. Binding of viral glycoprotein mRNA to endoplasmic reticulum membranes is disrupted by puromycin. Journal of Cell Biology, 74(2):358–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/358.

Lawson:1978:LMC

[LFR78] D. Lawson, C. Fewtrell, and M. C. Raff. Localized mast cell degranulation induced by concanavalin A-sepharose beads. Implications for the Ca^{2+} hypothesis of stimulus-secretion coupling. Journal of Cell Biology, 79(2):394–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/2/394.

Lawerence:1975:ACB

[LG75a] P. A. Lawerence and S. M. Green. The anatomy of a compartment border. The intersegmental boundary in Oncopeltus. Journal of Cell Biology, 65(2):373–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/2/373.

Lindemann:1975:ATI

[LG75b] C. B. Lindemann and I. R. Gibbons. Adenosine triphosphateinduced motility and sliding of filaments in mammalian sperm extracted with Triton X-100. Journal of Cell Biology, 65(1):147– ??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/65/
1/147.

Lasek:1977:CCT

[LGB77] R. J. Lasek, H. Gainer, and J. L. Barker. Cell-to-cell transfer of glial proteins to the squid giant axon: The glia- neuron protein transfer hypothesis. Journal of Cell Biology, 74(2):501–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/501.

Lawrence:1979:HIC

[LGGB79] T. S. Lawrence, R. D. Ginzberg, N. B. Gilula, and W. H. Beers. Hormonally induced cell shape changes in cultured rat ovarian granulosa cells. Journal of Cell Biology, 80(1):21–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/1/21.

Locke:1976:BGC

[LH76] M. Locke and P. Huie. The beads in the Golgi complexendoplasmic reticulum region. Journal of Cell Biology, 70(2): 384–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/2/384.

Levenson:1979:DPM

[LH79] R. Levenson and D. Housman. Developmental program of murine erythroleukemia cells. Effect of the inhibition of protein synthesis. Journal of Cell Biology, 82(3):715–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/3/715.

Larsen:1976:CFT

[LHH76] W. J. Larsen, P. M. Heidger, and J. C. Herr. ”central fold” or true junctional profile. Journal of Cell Biology, 71(1):333–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/1/333.

Langford:1979:MMA

[LI79] G. M. Langford and S. Inoué. Motility of the microtubular axostyle in Pyrsonympha. Journal of Cell Biology, 80(3):521–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/521.

Lemay:1975:PUL

[LJ75] A. Lemay and L. Jarett. Pitfalls in the use of lead nitrate for the histochemical demonstration of adenylate cyclase activity. Journal of Cell Biology, 65(1):39–??, April 1975. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/1/39.

Lasek:1979:ISP

[LKKA79] R. J. Lasek, N. Krishnan, and I. R. Kaiserman-Abramof. Identification of the subunit proteins of 10-nm neurofilaments isolated from axoplasm of squid and Myxicola giant axons. Journal of Cell Biology, 82(2):336–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/2/336.

Ledbetter:1979:TPN

[LL79] M. L. Ledbetter and M. Lubin. Transfer of potassium. A new measure of cell–cell coupling. Journal of Cell Biology, 80(1):150–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
1/150.

Lambowitz:1979:MRA

[LLC79] A. M. Lambowitz, R. J. LaPolla, and R. A. Collins. Mitochondrial ribosome assembly in Neurospora. Two-dimensional gel electrophoretic analysis of mitochondrial ribosomal proteins. Journal of Cell Biology, 82(1):17–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/1/17.

Lee:1975:RBS

[LLL75] S. G. Lee, V. Littau, and F. Lipmann. The relation between sporulation and the induction of antibiotic synthesis and of amino acid uptake in Bacillus brevis. Journal of Cell Biology, 66(2):233–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/66/
2/233.

Lewis:1978:LGP

[LLLG78] D. M. Lewis, J. C. Lewis, D. A. Loegering, and G. J. Gleich. Localization of the guinea pig eosinophil major basic protein to the core of the granule. Journal of Cell Biology, 77(3):702–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/77/
3/702.

Lutz:1977:RIM

[LLMW77] H. U. Lutz, A. J. Lomant, P. McMillan, and E. Wehrli. Rearrangements of integral membrane components during in vitro aging of sheep erythrocyte membranes. Journal of Cell Biology, 74(2):389–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/2/389.

Lutz:1977:RSF

[LLP77] H. U. Lutz, S. C. Liu, and J. Palek. Release of spectrin-free vesicles from human erythrocytes during ATP depletion: 1. characterization of spectrin-free vesicles. Journal of Cell Biology, 73(3):548– ??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
3/548.

Lemanski:1976:SMP

[LMPI76] L. F. Lemanski, M. S. Mooseker, L. D. Peachey, and M. R. Iyengar. Studies of muscle proteins in embryonic myocardial cells of cardiac lethal mutant mexican axolotls (Ambystoma mexicanum) by use of heavy meromyosin binding and sodium dodecyl sulfate polyacrylamide gel electrophoresis. Journal of Cell Biology, 68 (2):375–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/375.

Leuenberger:1975:SMV

[LN75] P. M. Leuenberger and A. B. Novikoff. Studies on microperoxisomes. VII. Pigment epithelial cells and other cell types in the retina of rodents. Journal of Cell Biology, 65(2):324–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/2/324.

Lefebvre:1978:FES

[LNMR78] P. A. Lefebvre, S. A. Nordstrom, J. E. Moulder, and J. L. Rosenbaum. Flagellar elongation and shortening in Chlamydomonas. IV. Effects of flagellar detachment, regeneration, and resorption on the induction of flagellar protein synthesis. Journal of Cell Biology, 78(1):8–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/8.

Lernmark:1976:PCP

[LNS76] A. Lernmark, A. Nathans, and D. F. Steiner. Preparation and characterization of plasma membrane-enriched fractions from rat pancreatic islets. Journal of Cell Biology, 71(2):606–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/2/606.

Longo:1976:SAR

[Lon76] F. J. Longo. Sperm aster in rabbit zygotes: its structure and function. Journal of Cell Biology, 69(3):539–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/3/539.

Longo:1977:USC

[Lon77] F. J. Longo. An ultrastructural study of cross-fertilization (Arbacia female X Mytilus male). Journal of Cell Biology, 73(1):14–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/73/1/14.

Landmesser:1976:FGS

[LP76] L. Landmesser and G. Pilar. Fate of ganglionic synapses and ganglion cell axons during normal and induced cell death. Journal of Cell Biology, 68(2):357–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/2/357.

Lusis:1977:RBL

[LP77] A. J. Lusis and K. Paigen. Relationships between levels of membrane-bound glucuronidase and the associated protein egasyn in mouse tissues. Journal of Cell Biology, 73(3):728–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/3/728.

Lawson:1977:MED

[LRG^{+}77] D. Lawson, M. C. Raff, B. Gomperts, C. Fewtrell, and N. B. Gilula. Molecular events during membrane fusion. A study of exocytosis in rat peritoneal mast cells. Journal of Cell Biology, 72(2):242–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/2/242.

Lodish:1975:MPS

[LS75] H. F. Lodish and B. Small. Membrane proteins synthesized by rabbit reticulocytes. Journal of Cell Biology, 65(1):51–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/1/51.

Leonard:1979:AIM

[LS79a] J. P. Leonard and M. M. Salpeter. Agonist-induced myopathy at the neuromuscular junction is mediated by calcium. Journal of Cell Biology, 82(3):811–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/3/811.

Linder:1979:NMF

[LS79b] J. C. Linder and L. A. Staehelin. A novel model for fluid secretion by the trypanosomatid contractile vacuole apparatus. Journal of Cell Biology, 83(2):371–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/2/371.

Lingappa:1978:NCO

[LSWB78] V. R. Lingappa, D. Shields, S. L. Woo, and G. Blobel. Nascent chicken ovalbumin contains the functional equivalent of a signal sequence. Journal of Cell Biology, 79(2):567–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/2/567.

Larsen:1979:EPA

[LTMS79] W. J. Larsen, H. N. Tung, S. A. Murray, and C. A. Swenson. Evidence for the participation of actin microfilaments and bristle coats in the internalization of gap junction membrane. Journal of Cell Biology, 83(3):576–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/3/576.

Lu:1978:MCV

[Lu78] B. C. Lu. Meiosis in Coprinus. VIII. A time-course study of the fusion and division of the spindle pole body during meiosis. Journal of Cell Biology, 76(3):761–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/3/761.

Leffert:1976:GCD

[LW76] H. L. Leffert and D. B. Weinstein. Growth control of differentiated fetal rat hepatocytes in primary monolayer culture. IX. Specific inhibition of DNA synthesis initiation by very low density lipoprotein and possible significance to the problem of liver regeneration. Journal of Cell Biology, 70(1):20–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/70/1/20.

Losa:1978:ISB

[LWB78] G. A. Losa, E. R. Weibel, and R. P. Bolender. Integrated stereological and biochemical studies on hepatocytic membranes. III. Relative surface of endoplasmic reticulum membranes in microsomal fractions estimated on freeze-fracture preparations. Journal of Cell Biology, 78(2):289–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/289.

Laurila:1979:FED

[LWS79] P. Laurila, J. Wartiovaara, and S. Stenman. Fibronectin expression is determined by the genotype of the transformed parental cells in heterokaryons between normal and transformed fibroblasts. Journal of Cell Biology, 80(1):118–??, January 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/1/118.

Liem:1978:IFN

[LYSS78] R. K. H. Liem, S. Yen, G. D. Salomon, and M. L. Shelanski. Intermediate filaments in nervous tissues. Journal of Cell Biology, 79(3):637–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/637.

Loyter:1975:UMS

[LZK75] A. Loyter, N. Zakai, and R. G. Kulka. ”ultramicroinjection” of macromolecules or small particles into animal cells. A new technique based on virus-induced cell fusion. Journal of Cell Biology, 66(2):292–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/292.

Marchase:1977:BIR

[Mar77] R. B. Marchase. Biochemical investigations of retinotectal adhesive specificity. Journal of Cell Biology, 75(1):237–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/237.

Matthews:1977:CRC

[Mat77] L. M. Matthews. Ca++ regulation in caffeine-derived microplasmodia of Physarum polycephalum. Journal of Cell Biology, 72 (2):502–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/2/502.

Maul:1976:FAN

[Mau76] G. G. Maul. Fibrils attached to the nuclear pore prevent egress of SV40 particles from the infected nucleus. Journal of Cell Biology, 70(3):714–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/3/714.

Maul:1977:NPC

[Mau77] G. G. Maul. Nuclear pore complexes. Elimination and reconstruction during mitosis. Journal of Cell Biology, 74(2):492–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/492.

Mason:1975:CRB

[MB75a] W. T. Mason and K. J. Bighouse. Correlation of rhodopsin biogenesis with ultrastructural morphogenesis in the chick retina. Journal of Cell Biology, 64(1):235–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/1/235.

McGill:1975:HCC

[MB75b] M. McGill and B. R. Brinkley. Human chromosomes and centrioles as nucleating sites for the in vitro assembly of microtubules from bovine brain tubulin. Journal of Cell Biology, 67(1):189–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
1/189.

McGuire:1976:IASa

[MB76] E. J. McGuire and C. L. Burdick. Intercellular adhesive selectivity. I. An improved assay for the measurement of embryonic chick intercellular adhesion (liver and other tissues). Journal of Cell Biology, 68(1):80–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/1/80.

Millete:1977:TEM

[MB77a] C. F. Millete and A. R. Bellve. Temporal expression of membrane antigens during mouse spermatogenesis. Journal of Cell Biology, 74(1):86–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/86.

Mullins:1977:TPC

[MB77b] J. Mullins and J. J. Biesele. Terminal phase of cytokinesis in D-98S cells. Journal of Cell Biology, 73(3):672–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/3/672.

Matsudaira:1979:IOC

[MB79] P. T. Matsudaira and D. R. Burgess. Identification and organization of the components in the isolated microvillus cytoskeleton. Journal of Cell Biology, 83(3):667–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/3/667.

Moore:1976:AMB

[MBBS76] P. L. Moore, H. L. Bank, N. T. Brissie, and S. S. Spicer. Association of microfilament bundles with lysosomes in polymorphonuclear leukocytes. Journal of Cell Biology, 71(2):659–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/2/659.

Moore:1978:PBP

[MBBS78] P. L. Moore, H. L. Bank, N. T. Brissie, and S. S. Spicer. Phagocytosis of bacteria by polymorphonuclear leukocytes: a freezefracture, scanning electron microscope, and thin-section investigation of membrane structure. Journal of Cell Biology, 76(1):158–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
1/158.

McGowan:1979:HRM

[MBM^{+}79] J. W. McGowan, B. Borwein, J. A. Medeiros, T. Beveridge, J. D. Brown, E. Spiller, R. Feder, J. Topalian, and W. Gudat. High resolution microchemical analysis using soft X-ray lithographic techniques. Journal of Cell Biology, 80(3):732–??, March 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/3/732.

Madison:1979:RST

[MBPTS79] L. D. Madison, B. Bergstrom-Porter, A. R. Torres, and E. Shelton. Regulation of surface topography of mouse peritoneal cells. Formation of microvilli and vesiculated pits on omental mesothelial cells by serum and other proteins. Journal of Cell Biology, 82 (3):783–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/783.

Moy:1977:CSU

[MBV77] G. W. Moy, B. Brandriff, and V. D. Vacquier. Cytasters from sea urchin eggs parthenogenetically activated by procaine. Journal of Cell Biology, 73(3):788–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/3/788.

Merriam:1978:AXO

[MC78a] R. W. Merriam and T. G. Clark. Actin in Xenopus oocytes: II. intracellular distribution and polymerizability. Journal of Cell Biology, 77(2):439–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/2/439.

Minsky:1978:MAT

[MC78b] B. D. Minsky and F. J. Chlapowski. Morphometric analysis of the translocation of lumenal membrane between cytoplasm and cell surface of transitional epithelial cells during the expansioncontraction cycles of mammalian urinary bladder. Journal of Cell Biology, 77(3):685–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/77/3/685.

Meeusen:1979:EMH

[MC79a] R. L. Meeusen and W. Z. Cande. N-ethylmaleimide-modified heavy meromyosin. A probe for actomyosin interactions. Journal of Cell Biology, 82(1):57–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/1/57.

Morrissey:1979:RSP

[MC79b] J. J. Morrissey and D. V. Cohn. Regulation of secretion of parathormone and secretory protein–I from separate intracellular pools by calcium, dibutyryl cyclic AMP, and (1)-isoproterenol. Journal of Cell Biology, 82(1):93–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/1/93.

Morrissey:1979:SDP

[MC79c] J. J. Morrissey and D. V. Cohn. Secretion and degradation of parathormone as a function of intracellular maturation of hormone pools. Journal of Cell Biology, 83(3):521–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/3/521.

Margolis:1979:AGS

[MCC79] R. N. Margolis, R. R. Cardell, and R. T. Curnow. Association of glycogen synthase phosphatase and phosphorylase phosphatase activities with membranes of hepatic smooth endoplasmic reticulum. Journal of Cell Biology, 83(2):348–??, November 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/83/2/348.

McGuire:1976:IASb

[McG76] E. J. McGuire. Intercellular adhesive selectivity. II. Properties of embryonic chick liver cell–cell adhesion. Journal of Cell Biology, 68(1):90–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/1/90.

Morton:1977:EMS

[MCM77] D. J. Morton, F. M. Clarke, and C. J. Masters. An electron microscope study of the interaction between fructose diphosphate aldolase and actin-containing filaments. Journal of Cell Biology, 74(3):1016–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/3/1016.

McNutt:1978:TSF

[McN78] N. S. McNutt. A thin-section and freeze-fracture study of microfilament-membrane attachments in choroid plexus and intestinal microvilli. Journal of Cell Biology, 79(3):774–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
3/774.

Meldolesi:1978:CDD

[MCP^{+}78] J. Meldolesi, G. Castiglioni, R. Parma, N. Nassivera, and P. De Camilli. Ca++-dependent disassembly and reassembly of occluding junctions in guinea pig pancreatic acinar cells. Effect of drugs. Journal of Cell Biology, 79(1):156–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/1/156.

Makowski:1977:GJS

[MCPG77] L. Makowski, D. L. D. Caspar, W. C. Phillips, and D. A. Goodenough. Gap junction structures: Analysis of the X-ray diffraction data. Journal of Cell Biology, 74(2):629–??, August 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/2/629.

Murdock:1977:MUP

[MCR77] L. L. Murdock, M. A. Cahill, and A. Reith. Morphometry and ultrastructure of prismatic cristae in mitochondria of a crayfish muscle. A hypothesis of the structural principle. Journal of Cell Biology, 74(1):326–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/1/326.

Murray:1975:OAF

[MD75] R. L. Murray and M. W. Dubin. The occurrence of actinlike filaments in association with migrating pigment granules in frog retinal pigment epithelium. Journal of Cell Biology, 64(3):705–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
3/705.

Maul:1977:QDN

[MD77a] G. G. Maul and L. Deaven. Quantitative determination of nuclear pore complexes in cycling cells with differing DNA content. Journal of Cell Biology, 73(3):748–??, June 1977. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/3/748.

Mills:1977:DPS

[MD77b] J. W. Mills and D. R. DiBona. On the distribution of Na+ pump sites in the frog skin. Journal of Cell Biology, 75(3):968–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/968.

Monneron:1978:IPNa

[Md78a] A. Monneron and J. d’Alayer. Isolation of plasma and nuclear membranes of thymocytes. I. Enzymatic composition and ultrastructure. Journal of Cell Biology, 77(1):211–??, April 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/1/211.

Monneron:1978:IPNb

[Md78b] A. Monneron and J. d’Alayer. Isolation of plasma and nuclear membranes of thymocytes. II. Biochemical composition. Journal of Cell Biology, 77(1):232–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/1/232.

Moser:1975:MHH

[MDR75] F. G. Moser, B. P. Dorman, and F. H. Ruddle. Mouse-human heterokaryon analysis with a 33258 Hoechst–Giemsa technique. Journal of Cell Biology, 66(3):676–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/3/676.

Molday:1975:NIS

[MDRY75] R. S. Molday, W. J. Dreyer, A. Rembaum, and S. P. Yen. New immunolatex spheres: visual markers of antigens on lymphocytes for scanning electron microscopy. Journal of Cell Biology, 64 (1):75–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/1/75.

Margaritis:1977:RRF

[MEB77] L. H. Margaritis, A. Elgsaeter, and D. Branton. Rotary replication for freeze-etching. Journal of Cell Biology, 72(1):47–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/72/1/47.

Mills:1977:LPS

[MED77] J. W. Mills, S. A. Ernst, and D. R. DiBona. Localization of Na+- pump sites in frog skin. Journal of Cell Biology, 73(1):88–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/73/1/88.

McDonald:1979:CSS

[MEM79] K. L. McDonald, M. K. Edwards, and J. R. McIntosh. Crosssectional structure of the central mitotic spindle of Diatoma vulgare. Evidence for specific interactions between antiparallel microtubules. Journal of Cell Biology, 83(2):443–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/2/443.

Myers:1975:DLM

[MF75] D. K. Myers and L. E. Feinendegen. Double labeling with [3H]thymidine and [125I]iododeoxyuridine as a method for determining the fate of injected DNA and cells in vivo. Journal of Cell Biology, 67(2):484–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/2/484.

Matsuura:1978:IML

[MFKT78] S. Matsuura, Y. Fujii-Kuriyama, and Y. Tashiro. Immunoelectron microscope localization of cytochrome P-450 on microsomes and other membrane structures of rat hepatocytes. Journal of Cell Biology, 78(2):503–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/2/503.

Montesano:1975:VAT

[MFPO75] R. Montesano, D. S. Friend, A. Perrelet, and L. Orci. In vivo assembly of tight junctions in fetal rat liver. Journal of Cell Biology, 67(2):310–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/2/310.

Martin:1975:GDC

[MG75a] N. C. Martin and U. W. Goodenough. Gametic differentiation in Chlamydomonas reinhardtii. I. Production of gametes and their fine structure. Journal of Cell Biology, 67(3):587–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/3/587.

Maruta:1975:FON

[MG75b] H. Maruta and L. Goldstein. The fate and origin of the nuclear envelope during and after mitosis in Amoeba proteus. I. Synthesis and behavior of phospholipids of the nuclear envelope during the cell life cycle. Journal of Cell Biology, 65(3):631–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/3/631.

Miller:1976:RRR

[MG76] L. Miller and F. Gonzales. The relationship of ribosomal RNA synthesis to the formation of segregated nucleoli and nucleoluslike bodies. Journal of Cell Biology, 71(3):939–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/3/939.

Minami:1978:NGS

[MG78a] S. A. Minami and U. W. Goodenough. Novel glycopolypeptide synthesis induced by gametic cell fusion in Chlamydomonas reinhardtii. Journal of Cell Biology, 77(1):165–??, April 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/1/165.

Mintz:1978:SGC

[MG78b] G. Mintz and L. Glaser. Specific glycoprotein changes during development of the chick neural retina. Journal of Cell Biology, 79(1):132–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/1/132.

McClay:1977:RTS

[MGF77] D. R. McClay, L. R. Godding, and M. E. Fransen. A requirement for trypsin-sensitive cell-surface components for cell–cell interactions of embryonic neural retina cells. Journal of Cell Biology, 75(1):56–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/56.

Montesano:1976:VIT

[MGPO76] R. Montesano, G. Gabbiani, A. Perrelet, and L. Orci. In vivo induction of tight junction proliferation in rat liver. Journal of Cell Biology, 68(3):793–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/3/793.

Meier:1975:SCD

[MH75] L. Meier and E. D. Hay. Stimulation of corneal differentiation by interaction between cell surface and extracellular matrix. I. Morphometric analysis of transfilter “induction”. Journal of Cell Biology, 66(2):275–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/2/275.

McKeehan:1976:SCG

[MH76a] W. L. McKeehan and R. G. Ham. Stimulation of clonal growth of normal fibroblasts with substrata coated with basic polymers. Journal of Cell Biology, 71(3):727–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/727.

Merriam:1976:GVN

[MH76b] R. W. Merriam and R. J. Hill. The germinal vesicle nucleus of Xenopus laevis oocytes as a selective storage receptacle for proteins. Journal of Cell Biology, 69(3):659–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/3/659.

Mautner:1977:SDL

[MH77] V. Mautner and R. O. Hynes. Surface distribution of LETS protein in relation to the cytoskeleton of normal and transformed cells. Journal of Cell Biology, 75(3):743–??, December 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/75/3/743.

Miller:1977:UHD

[MHC77] S. C. Miller, E. D. Hay, and J. F. Codington. Ultrastructural and histochemical differences in cell surface properties of strainspecific and nonstrain-specific TA3 adenocarcinoma cells. Journal of Cell Biology, 72(3):511–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/3/511.

Miller:1978:RLNa

[MHP78a] T. E. Miller, C. Y. Huang, and A. O. Pogo. Rat liver nuclear skeleton and ribonucleoprotein complexes containing HnRNA. Journal of Cell Biology, 76(3):675–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/3/675.

Miller:1978:RLNb

[MHP78b] T. E. Miller, C. Y. Huang, and A. O. Pogo. Rat liver nuclear skeleton and small molecular weight RNA species. Journal of Cell Biology, 76(3):692–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/76/3/692.

Miller:1977:OCS

[Mil77] S. C. Miller. Osteoclast cell-surface changes during the egg-laying cycle in Japanese quail. Journal of Cell Biology, 75(1):104–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
1/104.

Miller:1978:RAM

[Mil78] S. C. Miller. Rapid activation of the medullary bone osteoclast cell surface by parathyroid hormone. Journal of Cell Biology, 76(3):615–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/3/615.

Morel:1977:IPC

[MIMMF77] N. Morel, M. Israel, R. Manaranche, and P. Mastour-Frachon. Isolation of pure cholinergic nerve endings from Torpedo electric organ. Evaluation of their metabolic properties. Journal of Cell Biology, 75(1):43–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/75/1/43.

Mahowald:1976:ITP

[MIT76] A. P. Mahowald, K. Illmensee, and F. R. Turner. Interspecific transplantation of polar plasm between Drosophila embryos. Journal of Cell Biology, 70(2):358–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/2/358.

Molday:1976:CWG

[MJM76] R. Molday, R. Jaffe, and D. McMahon. Concanavalin A and wheat germ agglutinin receptors on Dictyostelium: Their visualization by scanning electron microscopy with microspheres. Journal of Cell Biology, 71(1):314–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/1/314.

Mori:1979:SMA

[MKK79] H. Mori, Y. Komiya, and M. Kurokawa. Slowly migrating axonal polypeptides. Inequalities in their rate and amount of transport between two branches of bifurcating axons. Journal of Cell Biology, 82(1):174–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/174.

Marcelo:1978:SSP

[MKKV78] C. L. Marcelo, Y. G. Kim, J. L. Kaine, and J. J. Voorhees. Stratification, specialization, and proliferation of primary keratinocyte cultures. Evidence of a functioning in vitro epidermal cell system. Journal of Cell Biology, 79(2):356–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/2/356.

Melvin:1976:CRR

[MKM76] W. T. Melvin, A. Kumar, and R. A. Malt. Conservation of ribosomal RNA during compensatory renal hypertrophy. A major mechanism in RNA accretion. Journal of Cell Biology, 69(3):548– ??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
3/548.

Malick:1976:SEM

[ML76] L. E. Malick and R. Langenbach. Scanning electron microscopy of in vitro chemically transformed mouse embryo cells. Journal of Cell Biology, 68(3):654–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/3/654.

Martin:1976:LVCa

[MM76a] F. J. Martin and R. C. MacDonald. Lipid vesicle-cell interactions. I. Hemagglutination and hemolysis. Journal of Cell Biology, 70 (3):494–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/494.

Martin:1976:LVCb

[MM76b] F. J. Martin and R. C. MacDonald. Lipid vesicle-cell interactions. II. Induction of cell fusion. Journal of Cell Biology, 70 (3):506–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/506.

Martin:1976:LVCc

[MM76c] F. J. Martin and R. C. MacDonald. Lipid vesicle-cell interactions. III. Introduction of a new antigenic determinant into erythrocyte membranes. Journal of Cell Biology, 70(3):515–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/3/515.

Marchand:1977:NTF

[MM77] B. Marchand and X. Mattei. [a new type of flagellar structure. Type 9+n]. Journal of Cell Biology, 72(3):707–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/3/707.

Mills:1978:MSM

[MM78] J. W. Mills and L. E. Malick. Mucosal surface morphology of the toad urinary bladder. Scanning electron microscope study of the natriferic and hydro-osmotic response to vasopressin. Journal of Cell Biology, 77(2):598–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/598.

Masland:1979:AIA

[MM79] R. H. Masland and J. W. Mills. Autoradiographic identification of acetylcholine in the rabbit retina. Journal of Cell Biology, 83 (1):159–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/1/159.

McIntosh:1979:TDS

[MMER79] J. R. McIntosh, K. L. McDonald, M. K. Edwards, and B. M. Ross. Three-dimensional structure of the central mitotic spindle of Diatoma vulgare. Journal of Cell Biology, 83(2):428–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/2/428.

Miller:1976:LHC

[MMM76] K. R. Miller, G. J. Miller, and K. R. McIntyre. The lightharvesting chlorpohyll-protein complex of photosystem II. Its location in the photosynthetic membrane. Journal of Cell Biology, 71(2):624–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/2/624.

Morimoto:1976:IIM

[MMS^{+}76] T. Morimoto, S. Matsuura, S. Sasaki, Y. Yashiro, and T. Omura. Immunochemical and immunoelectron microscope studies on localization of NADPH-cytochrome c reductase on rat liver microsomes. Journal of Cell Biology, 68(2):189–??, February 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/2/189.

Mori:1977:IPR

[MN77] M. Mori and A. B. Novikoff. Induction of pinocytosis in rat hepatocytes by partial hepatectomy. Journal of Cell Biology, 72(3):695–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/3/695.

Miki-Noumura:1979:CCO

[MNK79] T. Miki-Noumura and R. Kamiya. Conformational change in the outer doublet microtubules from sea urchin sperm flagella. Journal of Cell Biology, 81(2):355–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/2/355.

Mabuchi:1977:EMA

[MO77] I. Mabuchi and M. Okuno. The effect of myosin antibody on the division of starfish blastomeres. Journal of Cell Biology, 74(1):251–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/251.

Moens:1976:SKM

[Moe76] P. B. Moens. Spindle and kinetochore morphology of Dictyostelium discoideum. Journal of Cell Biology, 68(1):113–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
1/113.

Moens:1979:KMN

[Moe79] P. B. Moens. Kinetochore microtubule numbers of different sized chromosomes. Journal of Cell Biology, 83(3):556–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/3/556.

Moore:1976:CSA

[Moo76a] E. G. Moore. Cell to substratum adhesion-promoting activity released by normal and virus-transformed cells in culture. Journal of Cell Biology, 70(3):634–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/634.

Mooseker:1976:BBM

[Moo76b] M. S. Mooseker. Brush border motility. Microvillar contraction in triton-treated brush borders isolated from intestinal epithelium. Journal of Cell Biology, 71(2):417–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/417.

Malawista:1978:MCA

[MOR78] S. E. Malawista, J. M. Oliver, and S. A. Rudolph. Microtubules and cyclic AMP in human leukocytes: on the order of things. Journal of Cell Biology, 77(3):881–??, June 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/3/881.

Murphy:1977:SNG

[MOS^{+}77] R. A. Murphy, J. Oger, J. D. Saide, M. H. Blanchard, B. G. Arnason, C. Hogan, and N. J. Pantazis. Secretion of nerve growth factor by central nervous system glioma cells in culture. Journal of Cell Biology, 72(3):769–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/3/769.

Mooseker:1978:CLM

[MP78] M. S. Mooseker and T. D. Pollard. Characterization and localization of myosin in the brush border of intestinal epithelial cells. Journal of Cell Biology, 79(2):444–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/2/444.

Martinez-Palomo:1976:TDD

[MPDGR76] A. Martínez-Palomo, W. DeSouza, and A. Gonzalez-Robles. Topographical differences in the distribution of surface coat components and intramembrane particles. A cytochemical and freezefracture study in culture forms of Trypanosoma cruzi. Journal of Cell Biology, 69(2):507–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/507.

McDonald:1977:MAS

[MPHMT77] K. McDonald, J. D. Pickett-Heaps, J. R. McIntosh, and D. H. Tippit. On the mechanism of anaphase spindle elongation in Diatoma vulgare. Journal of Cell Biology, 74(2):377–??, August 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/74/2/377.

Maylie-Pfenninger:1979:DCSa

[MPJ79a] M. Maylie-Pfenninger and J. D. Jamieson. Distribution of cell surface saccharides on pancreatic cells. Journal of Cell Biology, 80(1):69–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/69.

Maylie-Pfenninger:1979:DCSb

[MPJ79b] M. F. Maylié-Pfenninger and J. D. Jamieson. Distribution of cell surface saccharides on pancreatic cells. II. Lectin-labeling patterns on mature guinea pig and rat pancreatic cells. Journal of Cell Biology, 80(1):77–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/1/77.

Meda:1979:IGJ

[MPO79] P. Meda, A. Perrelet, and L. Orci. Increase of gap junctions between pancreatic B-cells during stimulation of insulin secretion. Journal of Cell Biology, 82(2):441–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/2/441.

McLean:1979:ERS

[MPS79] M. J. McLean, A. Pelleg, and N. Sperelakis. Electrophysiological recordings from spontaneously contracting reaggregates of cultured smooth muscle cells from guinea pig vas deferens. Journal of Cell Biology, 80(3):539–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/3/539.

Meyrick:1975:VIT

[MR75a] B. Meyrick and L. Reid. In vitro incorporation of (3H)threonine and (3H)glucose by the mucous and serous cells of the human bronchial submucosal gland. A quantitative electron microscope study. Journal of Cell Biology, 67(2):320–??, November 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/2/320.

Moran:1975:UER

[MR75b] D. Moran and R. W. Rice. An ultrastructural examination of the role of cell membrane surface coat material during neurulation. Journal of Cell Biology, 64(1):172–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/1/172.

Mellon:1976:SVP

[MR76a] M. G. Mellon and L. I. Rebhun. Sulfhydryls and the in vitro polymerization of tubulin. Journal of Cell Biology, 70(1):226–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/70/
1/226.

Moody:1976:MAU

[MR76b] D. E. Moody and J. K. Reddy. Morphometric analysis of the ultrastructural changes in rat liver induced by the peroxisome proliferator SaH 42-348. Journal of Cell Biology, 71(3):768–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
3/768.

Matarese:1977:EIG

[MR77] G. P. Matarese and G. B. Rossi. Effect of interferon on growth and division cycle of Friend erythroleukemic murine cells in vitro. Journal of Cell Biology, 75(2):344–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/2/344.

Meistrich:1975:MDS

[MRB75] M. L. Meistrich, B. O. Reid, and W. J. Barcellona. Meiotic DNA synthesis during mouse spermatogenesis. Journal of Cell Biology, 64(1):211–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/1/211.

Madyastha:1977:SLC

[MRDC77] K. M. Madyastha, J. E. Ridgway, J. G. Dwyer, and C. J. Coscia. Subcellular localization of a cytochrome P-450-dependent monogenase in vesicles of the higher plant Catharanthus roseus. Journal of Cell Biology, 72(2):302–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/2/302.

Malech:1977:SAH

[MRG77] H. L. Malech, R. K. Root, and J. I. Gallin. Structural analysis of human neutrophil migration: Centriole, microtubule, and microfilament orientation and function during chemotaxis. Journal of Cell Biology, 75(3):666–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/3/666.

Moran:1976:MST

[MRZV76] D. T. Moran, J. C. Rowley, S. N. Zill, and F. G. Varela. The mechanism of sensory transduction in a mechanoreceptor. Functional stages in campaniform sensilla during the molting cycle. Journal of Cell Biology, 71(3):832–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/832.

Morgan:1975:TCD

[MS75a] J. L. Morgan and N. W. Seeds. Tubulin constancy during morphological differentiation of mouse neuroblastoma cells. Journal of Cell Biology, 67(1):136–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/136.

Motokawa:1975:LIS

[MS75b] T. Motokawa and P. Satir. Laser-induced spreading arrest of Mytilus gill cilia. Journal of Cell Biology, 66(2):377–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/2/377.

Miller:1976:ATO

[MS76] K. R. Miller and L. A. Staehelin. Analysis of the thylakoid outer surface. Coupling factor is limited to unstacked membrane regions. Journal of Cell Biology, 68(1):30–??, January 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/1/30.

Murray:1979:ECA

[MSK^{+}79] J. C. Murray, G. Stingl, H. K. Kleinman, G. R. Martin, and S. I. Katz. Epidermal cells adhere preferentially to type IV (basement membrane) collagen. Journal of Cell Biology, 80(1):197–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
1/197.

Maupin-Szamier:1978:AFD

[MSP78] P. Maupin-Szamier and T. D. Pollard. Actin filament destruction by osmium tetroxide. Journal of Cell Biology, 77(3):837–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/3/837.

Maloff:1978:MPR

[MSRT78] B. L. Maloff, S. P. Scordilis, C. Reynolds, and H. Tedeschi. Membrane potentials and resistances of giant mitochondria. Metabolic dependence and the effects of valinomycin. Journal of Cell Biology, 78(1):199–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/199.

Mazia:1975:ACS

[MSS75] D. Mazia, G. Schatten, and W. Sale. Adhesion of cells to surfaces coated with polylysine. Applications to electron microscopy. Journal of Cell Biology, 66(1):198–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/198.

Maloff:1978:AMV

[MST78] B. L. Maloff, S. P. Scordilis, and H. Tedeschi. Assays of the metabolic viability of single giant mitochondria. Experiments with intact and impaled mitochondria. Journal of Cell Biology, 78(1):214–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/214.

Mooseker:1975:OAF

[MT75] M. S. Mooseker and L. G. Tilney. Organization of an actin filament-membrane complex. Filament polarity and membrane attachment in the microvilli of intestinal epithelial cells. Journal of Cell Biology, 67(3):725–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/725.

Metuzals:1978:SFN

[MT78] J. Metuzals and I. Tasaki. Subaxolemmal filamentous network in the giant nerve fiber of the squid (Loligo pealei L.) and its possible role in excitability. Journal of Cell Biology, 78(2):597–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/78/
2/597.

Mechler:1975:MBRa

[MV75a] B. Mechler and P. Vassalli. Membrane-bound ribosomes of myeloma cells. I. Preparation of free and membrane-bound ribosomal fractions. Assessment of the methods and properties of the ribosomes. Journal of Cell Biology, 67(1):1–??, October 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/1/1.

Mechler:1975:MBRb

[MV75b] B. Mechler and P. Vassalli. Membrane-bound ribosomes of myeloma cells. II. Kinetic studies on the entry of newly made ribosomal subunits into the free and the membrane-bound ribosomal particles. Journal of Cell Biology, 67(1):16–??, October 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/1/16.

Mechler:1975:MBRc

[MV75c] B. Mechler and P. Vassalli. Membrane-bound ribosomes of myeloma cells. III. The role of the messenger RNA and the nascent polypeptide chain in the binding of ribosomes to membranes. Journal of Cell Biology, 67(1):25–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/1/25.

Manning:1976:RKD

[MW76] J. E. Manning and D. R. Wolstenholme. Replication of kinetoplast DNA of Crithidia acanthocephali. I. Density shift experiments using deuterium oxide. Journal of Cell Biology, 70(2):406–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/70/
2/406.

Mazur:1977:MDP

[MW77] M. T. Mazur and Williamson. Macrophage deformability and phagocytosis. Journal of Cell Biology, 75(1):185–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/185.

Mendell:1978:ILS

[MW78] J. R. Mendell and J. N. Whitaker. Immunocytochemical localization studies of myelin basic protein. Journal of Cell Biology, 76 (2):502–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/2/502.

Matsui:1979:QCC

[MWS79] S. I. Matsui, H. Weinfeld, and A. A. Sandberg. Quantitative conservation of chromatin-bound RNA polymerases I and II in mitosis. Implications for chromosome structure. Journal of Cell Biology, 80(2):451–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/2/451.

Milcarek:1978:SNP

[MZ78] C. Milcarek and K. Zahn. The synthesis of ninety proteins including actin throughout the HeLa cell cycle. Journal of Cell Biology, 79(3):833–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/833.

Mayer:1979:SMA

[MZ79] J. F. Mayer and B. R. Zirkin. Spermatogenesis in the mouse. I. Autoradiographic studies of nuclear incorporation and loss of 3H-amino acids. Journal of Cell Biology, 81(2):403–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/403.

Nanney:1975:PBB

[Nan75] D. L. Nanney. Patterns of basal body addition in ciliary rows in Tetrahymena. Journal of Cell Biology, 65(3):503–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/3/503.

Noe:1977:SPPa

[NBB77a] B. D. Noe, C. A. Baste, and G. E. Bauer. Studies on proinsulin and proglucagon biosynthesis and conversion at the subcellular level: I. Fractionation procedure and characterization of the subcellular fractions. Journal of Cell Biology, 74(2):578–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/578.

Noe:1977:SPPb

[NBB77b] B. D. Noe, C. A. Baste, and G. E. Bauer. Studies on proinsulin and proglucagon biosynthesis and conversion at the subcellular level: II. Distribution of radioactive peptide hormones and hormone precursors in subcellular fractions after pulse and pulse- chase incubation of islet tissue. Journal of Cell Biology, 74(2):589–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/2/589.

Nabarra:1977:USS

[NCCB77] B. Nabarra, J. Charreire, J. F. Cavellier, and J. F. Bach. Ultrastructural study of spontaneous bone marrow rosette-forming cells. Journal of Cell Biology, 72(3):773–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/3/773.

Newburger:1979:FCH

[NCGC79] P. E. Newburger, M. E. Chovaniec, J. S. Greenberger, and H. J. Cohen. Functional changes in human leukemic cell line HL-60. A model for myeloid differentiation. Journal of Cell Biology, 82 (2):315–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/2/315.

Nanney:1978:SCT

[NCM78] D. L. Nanney, S. S. Chen, and E. B. Meyer. Scalar constraints in Tetrahymena evolution. Quantitative basal body variations within and between species. Journal of Cell Biology, 79(3):727–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
3/727.

Nemanic:1975:HSL

[NCPW75] M. K. Nemanic, D. P. Carter, D. R. Pitelka, and L. Wofsy. Hapten-sandwich labeling. II. Immunospecific attachment of cell surface markers suitable for scanning electron microscopy. Journal of Cell Biology, 64(2):311–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/2/311.

Nilsson:1977:EPA

[ND77] O. S. Nilsson and G. Dallner. Enzyme and phospholipid asymmetry in liver microsomal membranes. Journal of Cell Biology, 72(3):568–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/3/568.

Nelson:1978:CEIb

[NEK78a] D. M. Nelson, A. C. Enders, and B. F. King. Cytological events involved in glycoprotein synthesis in cellular and syncytial trophoblast of human placenta. An electron microscope autoradiographic study of [3H]galactose incorporation. Journal of Cell Biology, 76(2):418–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/76/2/418.

Nelson:1978:CEIa

[NEK78b] D. M. Nelson, A. C. Enders, and B. F. King. Cytological events involved in protein synthesis in cellular and syncytial trophoblast of human placenta. An electron microscope autoradiographic study of [3H]leucine incorporation. Journal of Cell Biology, 76(2): 400–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/2/400.

Nuccitelli:1975:PCP

[NJ75] R. Nuccitelli and L. F. Jaffe. The pulse current pattern generated by developing fucoid eggs. Journal of Cell Biology, 64(3):636–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
3/636.

Newsome:1976:VBC

[NLT76] D. A. Newsome, T. F. Linsenmayer, and R. L. Trelstad. Vitreous body collagen. Evidence for a dual origin from the neural retina and hyalocytes. Journal of Cell Biology, 71(1):59–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/71/1/59.

Nenci:1978:CLS

[NM78] I. Nenci and E. Marchetti. Concerning the localization of steroids in centrioles and basal bodies by immunofluorescence. Journal of Cell Biology, 76(2):255–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/2/255.

Novikoff:1977:SSP

[NMQY77] A. B. Novikoff, M. Mori, N. Quintana, and A. Yam. Studies of the secretory process in the mammalian exocrine pancreas. I. The condensing vacuoles. Journal of Cell Biology, 75(1):148–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
1/148.

Niederman:1975:HPM

[NP75] R. Niederman and T. D. Pollard. Human platelet myosin. II. In vitro assembly and structure of myosin filaments. Journal of Cell Biology, 67(1):72–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/72.

Norwood:1975:RDS

[NPM75] T. H. Norwood, W. R. Pendergrass, and G. M. Martin. Reinitiation of DNA synthesis in senescent human fibroblasts upon fusion with cells of unlimited growth potential. Journal of Cell Biology, 64(3):551–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/551.

Nath:1976:ECO

[NR76] J. Nath and J. I. Rebhun. Effects of caffeine and other methylxanthines on the development and metabolism of sea urchin eggs. Involvement of NADP and glutathione. Journal of Cell Biology, 68(3):440–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/3/440.

Nakamura:1978:ETG

[NRH78] M. Nakamura, L. J. Romrell, and P. F. Hall. The effects of temperature and glucose on protein biosynthesis by immature (round) spermatids from rat testes. Journal of Cell Biology, 79(1):1–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
1/1.

Nagano:1975:MSC

[NS75] T. Nagano and F. Suzuki. Microtubules with 15 subunits in cockroach epidermal cells. Journal of Cell Biology, 64(1):242–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
1/242.

Neutra:1977:MIB

[NS77] M. R. Neutra and S. F. Schaeffer. Membrane interactions between adjacent mucols secretion granules. Journal of Cell Biology, 74 (3):983–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/983.

Naccache:1977:CIM

[NSBS77a] P. H. Naccache, H. J. Showell, E. L. Becker, and R. I. Sha’afi. Changes in ionic movements across rabbit polymorphonuclear leukocyte membranes during lysosomal enzyme release. Possible ionic basis for lysosomal enzyme release. Journal of Cell Biology, 75(3):635–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/635.

Naccache:1977:TSP

[NSBS77b] P. H. Naccache, H. J. Showell, E. L. Becker, and R. I. Sha’afi. Transport of sodium, potassium, and calcium across rabbit polymorphonuclear leukocyte membranes. Effect of chemotactic factor. Journal of Cell Biology, 73(2):428–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/2/428.

Naccache:1979:IMC

[NSBS79] P. H. Naccache, H. J. Showell, E. L. Becker, and R. I. Sha’afi. Involvement of membrane calcium in the response of rabbit neutrophils to chemotactic factors as evidenced by the fluorescence of chlorotetracycline. Journal of Cell Biology, 83(1):179–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/1/179.

Nicolson:1978:DTL

[NSH78] G. L. Nicolson, J. R. Smith, and R. Hyman. Dynamics of toxin and lectin receptors on a lymphoma cell line and its toxin-resistant variant using ferritin-conjugated, 125I-labeled ligand. Journal of Cell Biology, 78(2):565–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/565.

Noseworthy:1975:PGE

[NSHE75] J. Noseworthy, G. H. Smith, S. R. Himmelhoch, and W. H. Evans. Protein and glycoprotein electrophoretic patterns of enriched fractions of primary and secondary granules from guinea pig polymorphonuclear leukocytes. Journal of Cell Biology, 65(3):577–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/65/
3/577.

Nicolson:1976:ELA

[NSP76] G. L. Nicolson, J. R. Smith, and G. Poste. Effects of local anesthetics on cell morphology and membrane-associated cytoskeletal organization in BALB/3T3 cells. Journal of Cell Biology, 68 (2):395–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/395.

Narayanan:1976:SMC

[NSRL76] A. S. Narayanan, L. B. Sandberg, R. Ross, and D. L. Layman. The smooth muscle cell. III. Elastin synthesis in arterial smooth muscle cell culture. Journal of Cell Biology, 68(3):411–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/3/411.

Nishi:1978:LIS

[NT78] Y. Nishi and Y. Takesue. Localization of intestinal sucraseisomaltase complex on the microvillous membrane by electron microscopy using nonlabeled antibodies. Journal of Cell Biology, 79 (2):516–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/516.

Nicolson:1977:LBS

[NUY^{+}77] G. L. Nicolson, N. Usui, R. Yanagimachi, H. Yanagimachi, and Smith. Lectin-binding sites on the plasma membranes of rabbit spermatozoa: Changes in surface receptors during epididymal maturation and after ejaculation. Journal of Cell Biology, 74 (3):950–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/950.

Nist:1975:LPC

[NVH^{+}75] C. Nist, K. Von Der Mark, E. D. Hay, B. R. Olsen, P. Bornstein, R. Ross, and P. Dehm. Location of procollagen in chick corneal and tendon fibroblasts with ferritin-conjugated antibodies. Journal of Cell Biology, 65(1):75–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/1/75.

Novikoff:1978:SLP

[NY78] P. M. Novikoff and A. Yam. Sites of lipoprotein particles in normal rat hepatocytes. Journal of Cell Biology, 76(1):1–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/1/1.

Nicolson:1975:ULL

[NYY75] G. L. Nicolson, R. Yanagimachi, and H. Yanagimachi. Ultrastructural localization of lectin-binding sites on the zonae pellucidae and plasma membranes of mammalian eggs. Journal of Cell Biology, 66(2):263–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/2/263.

Nivaler:1979:ALH

[NZD^{+}79] G. Nivaler, E. A. Zimmerman, R. Defendini, A. S. Liotta, D. T. Kreiger, and M. J. Brownstein. Andrenocorticotropin and βlipotropin in the hypothalamus. Journal of Cell Biology, 81(1):50– ??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
1/50.

Oliver:1976:EGO

[OAB76] J. M. Oliver, D. F. Albertini, and R. D. Berlin. Effects of glutathione-oxidizing agents on microtubule assembly and microtubule-dependent surface properties of human neutrophils. Journal of Cell Biology, 71(3):921–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/921.

Ogawa:1977:PAA

[OAB77] K. Ogawa, D. J. Asai, and C. J. Brokaw. Properties of an antiserum against native dynein 1 from sea urchin sperm flagella. Journal of Cell Biology, 73(1):182–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/1/182.

Olivetti:1977:MRC

[OAR^{+}77] G. Olivetti, P. Anversa, W. Rigamonti, L. Vitali-Mazza, and A. V. Loud. Morphometry of the renal corpuscle during normal postnatal growth and compensatory hypertrophy. A light microscope study. Journal of Cell Biology, 75(2):573–??, November 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/75/2/573.

Oron:1978:ITP

[OB78] U. Oron and A. Bdolah. Intracellular transport of proteins in active and resting secretory cells of the venom gland of Vipera palaestinae. Journal of Cell Biology, 78(2):488–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/2/488.

Olsen:1975:FCE

[OBKP75] B. R. Olsen, R. A. Berg, Y. Kishida, and D. J. Prockop. Further characterization of embryonic tendon fibroblasts and the use of immunoferritin techniques to study collagen biosynthesis. Journal of Cell Biology, 64(2):340–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/2/340.

Olins:1975:VCS

[OCO75] A. L. Olins, R. D. Carlson, and D. E. Olins. Visualization of chromatin substructure: upsilon bodies. Journal of Cell Biology, 64(3):528–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/528.

Osborn:1979:RCI

[ODS79] J. C. Osborn, C. J. Duncan, and J. L. Smith. Role of calcium ions in the control of embryogenesis of Xenopus. Changes in the subcellular distribution of calcium in early cleavage embryos after treatment with the ionophore A23187. Journal of Cell Biology, 80(3):589–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/3/589.

Oliver:1978:UFL

[OH78] C. Oliver and A. R. Hand. Uptake and fate of luminally administered horseradish peroxidase in resting and isoproterenolstimulated rat parotid acinar cells. Journal of Cell Biology, 76 (1):207–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/1/207.

Oliver:1976:CPG

[OKB76] J. M. Oliver, J. A. Krawiec, and R. D. Berlin. Carbamycholine prevents giant granule-formation in cultured fibroblasts from beige (Chediak–Higashi) mice. Journal of Cell Biology, 69(1):205–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/205.

Ojakian:1977:MRB

[OKS77] G. K. Ojakian, G. Kreibich, and D. D. Sabatini. Mobility of ribosomes bound to microsomal membranes. A freeze-etch and thin-section electron microscope study of the structure and fluidity of the rough endoplasmic reticulum. Journal of Cell Biology, 72(3):530–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/3/530.

OFlaherty:1979:SND

[OKS^{+}79] J. T. O’Flaherty, D. L. Kreutzer, H. J. Showell, G. Vitkauskas, E. L. Becker, and P. A. Ward. Selective neutrophil desensitization to chemotactic factors. Journal of Cell Biology, 80(3):564–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/564.

Ostlund:1977:MAF

[OLK77] R. E. Ostlund, J. T. Leung, and D. M. Kipnis. Muscle actin filaments bind pituitary secretory granules in vitro. Journal of Cell Biology, 73(1):78–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/1/78.

Ostlund:1978:MST

[OLK78] R. E. Ostlund, J. T. Leung, and D. M. Kipnis. Myosins of secretory tissues. Journal of Cell Biology, 77(3):827–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/3/827.

Ofer:1976:CTP

[OLL76] L. Ofer, H. Levkovitz, and A. Loyter. Conjugation in Tetrahymena pyriformis. The effect of polylysine, concanavalin A, and bivalent metals on the conjugation process. Journal of Cell Biology, 70 (2):287–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/2/287.

Olins:1979:SEM

[OO79] A. L. Olins and D. E. Olins. Stereo electron microscopy of the 25nm chromatin fibers in isolated nuclei. Journal of Cell Biology, 81(1):260–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/1/260.

Olgiati:1976:ERL

[OPD76] D. D. Olgiati, B. G. Pogo, and S. Dales. Evidence for RNA linked to nascent DNA in HeLa cells. Journal of Cell Biology, 68(3):557– ??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/557.

Orci:1977:FFM

[OPF77] L. Orci, A. Perrelet, and D. S. Friend. Freeze-fracture of membrane fusions during exocytosis in pancreatic B-cells. Journal of Cell Biology, 75(1):23–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/1/23.

Obinata:1979:TEC

[OSM79] T. Obinata, Y. Shimada, and R. Matsuda. Troponin in embryonic chick skeletal muscle cells in vitro. An immunoelectron microscope study. Journal of Cell Biology, 81(1):59–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/1/59.

Olins:1976:UFC

[OSO76] A. L. Olins, M. B. Senior, and D. E. Olins. Ultrastructural features of chromatin nu bodies. Journal of Cell Biology, 68(3):787–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/787.

Oates:1976:VFA

[OT76] P. J. Oates and O. Touster. In vitro fusion of Acanthamoeba phagolysosomes. I. Demonstration and quantitation of vacuole fusion in Acanthamoeba homogenates. Journal of Cell Biology, 68 (2):319–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/319.

Oates:1978:VFA

[OT78] P. J. Oates and O. Touster. In vitro fusion of Acanthamoeba phagolysosomes. II Quantitative characterization of in vitro vacuole fusion by improved electron microscope and new light microscope techniques. Journal of Cell Biology, 79(1):217–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/217.

Orkin:1977:DCG

[OWC^{+}77] R. W. Orkin, B. R. Williams, R. E. Cranley, D. C. Poppke, and K. S. Brown. Defects in the cartilaginous growth plates of brachymorphic mice. Journal of Cell Biology, 73(2):287–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/2/287.

Obara:1975:DMC

[OWS75] Y. Obara, H. Weinfeld, and A. A. Sandberg. Detection by means of cell fusion of macromolecular synthesis involved in the reconstruction of the nuclear envelope in mitosis. Journal of Cell Biology, 64(2):378–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/64/2/378.

Osborn:1978:IMV

[OWW78] M. Osborn, R. E. Webster, and K. Weber. Individual microtubules viewed by immunofluorescence and electron microscopy in the same PtK2 cell. Journal of Cell Biology, 77(3):??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/3/R27.

ODay:1978:RDS

[OY78] W. T. O’Day and R. W. Young. Rhythmic daily shedding of outer-segment membranes by visual cells in the goldfish. Journal of Cell Biology, 76(3):593–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/3/593.

Paavola:1978:CLGa

[Paa78a] L. G. Paavola. The corpus luteum of the guinea pig. II. Cytochemical studies on the Golgi complex, GERL, and lysosomes in luteal cells during maximal progesterone secretion. Journal of Cell Biology, 79(1):45–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/1/45.

Paavola:1978:CLGb

[Paa78b] L. G. Paavola. The corpus luteum of the guinea pig. III. Cytochemical studies on the Golgi complex and GERL during normal postpartum regression of luteal cells, emphasizing the origin of lysosomes and autophagic vacuoles. Journal of Cell Biology, 79(1):59–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/1/59.

Paine:1975:NMF

[Pai75] P. L. Paine. Nucleocytoplasmic movement of fluorescent tracers microinjected into living salivary gland cells. Journal of Cell Biology, 66(3):652–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/3/652.

Palade:1977:KRP

[Pal77] G. E. Palade. Keith Roberts Porter and the development of contemporary cell biology. Journal of Cell Biology, 75(1):??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/D1.

Paulin:1975:CST

[Pau75] J. J. Paulin. The chondriome of selected trypanosomatids. A three-dimensional study based on serial thick sections and high voltage electron microscopy. Journal of Cell Biology, 66 (2):404–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/404.

Pfeffer:1976:PTU

[PAW76] T. A. Pfeffer, C. F. Asnes, and L. Wilson. Properties of tubulin in unfertilized sea urchin eggs. Quantitation and characterization by the colchicine-binding reaction. Journal of Cell Biology, 69(3):599–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/3/599.

Puvion:1975:RCL

[PB75] E. Puvion and W. Bernhard. Ribonucleoprotein components in liver cell nuclei as visualized by cryoultramicrotomy. Journal of Cell Biology, 67(1):200–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/200.

Pepper:1979:MIK

[PB79] D. A. Pepper and B. R. Brinkley. Microtubule initiation at kinetochores and centrosomes in lysed mitotic cells. Inhibition of sitespecific nucleation by tubulin antibody. Journal of Cell Biology, 82(2):585–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/2/585.

Pierson:1978:ANP

[PBH78] G. B. Pierson, P. R. Burton, and R. H. Himes. Alterations in number of protofilaments in microtubules assembled in vitro. Journal of Cell Biology, 76(1):223–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/1/223.

Patzelt:1977:IEC

[PBJ77] C. Patzelt, D. Brown, and B. Jeanrenaud. Inhibitory effect of colchicines on amylase secretion by rat parotid glands: possible localization in the Golgi area. Journal of Cell Biology, 73(3):578– ??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
3/578.

Prentki:1979:AMC

[PCJG79] M. Prentki, C. Chaponnier, B. Jeanrenaud, and G. Gabbiani. Actin microfilaments, cell shape, and secretory processes in isolated rat hepatocytes. Effect of phalloidin and cytochalasin D. Journal of Cell Biology, 81(3):592–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/3/592.

Popescu:1975:CSM

[PD75] L. M. Popescu and I. Diculescu. Calcium in smooth muscle sarcoplasmic reticulum in situ. Conventional and X-ray analytical electron microscopy. Journal of Cell Biology, 67(3):911–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/911.

Peracchia:1976:LRJ

[PD76] C. Peracchia and A. F. Dulhunty. Low resistance junctions in crayfish. Structural changes with functional uncoupling. Journal of Cell Biology, 70(2):419–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/2/419.

Pearlstein:1978:IME

[PDD78] E. Pearlstein, S. R. Dienstman, and V. Defendi. Identification of macrophage external membrane proteins and their possible role in cell adhesion. Journal of Cell Biology, 79(1):263–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/263.

Poodry:1979:RAN

[PE79] C. A. Poodry and L. Edgar. Reversible alteration in the neuromuscular junctions of Drosophila melanogaster bearing a temperature-sensitive mutation, shibire. Journal of Cell Biology, 81(3):520–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/520.

Posakony:1977:MGD

[PEA77] J. W. Posakony, J. M. England, and G. Attardi. Mitochondrial growth and division during the cell cycle in HeLa cells. Journal of Cell Biology, 74(2):468–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/468.

Patt:1976:ESF

[PELG76] L. M. Patt, R. O. Endres, D. O. Lucas, and W. J. Grimes. Ectogalactosyltransferase studies in fibroblasts and concanavalin Astimulated lymphocytes. Journal of Cell Biology, 68(3):799–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/799.

Peracchia:1977:GJS

[Per77] C. Peracchia. Gap junctions. Structural changes after uncoupling procedures. Journal of Cell Biology, 72(3):628–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/72/3/628.

Pfeifer:1978:IIF

[Pfe78] U. Pfeifer. Inhibition by insulin of the formation of autophagic vacuoles in rat liver. A morphometric approach to the kinetics of intracellular degradation by autophagy. Journal of Cell Biology, 78(1):152–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/1/152.

Philipp:1976:CNM

[PFK^{+}76] E. I. Philipp, W. W. Franke, T. W. Keenan, J. Stadler, and E. D. Jarasch. Characterization of nuclear membranes and endoplasmic reticulum isolated from plant tissue. Journal of Cell Biology, 68(1):11–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/1/11.

Pagano:1975:IPV

[PH75a] R. E. Pagano and L. Huang. Interaction of phospholipid vesicles with cultured mammalian cells. II. Studies of mechanism. Journal of Cell Biology, 67(1):49–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/49.

Palevitz:1975:IAS

[PH75b] B. A. Palevitz and P. K. Hepler. Identification of actin in situ at the ectoplasm-endoplasm interface of Nitella. Microfilamentchloroplast association. Journal of Cell Biology, 65(1):29–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/65/1/29.

Pardue:1975:LRG

[PH75c] M. L. Pardue and T. C. Hsu. Locations of 18S and 28S ribosomal genes on the chromosomes of the Indian muntjac. Journal of Cell Biology, 64(1):251–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/251.

Parducz:1976:VSN

[PJ76a] A. Parducz and F. Joó. Visualization of stimulated nerve endings by preferential calcium accumulation of mitochondria. Journal of Cell Biology, 69(2):513–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/513.

Peng:1976:SSM

[PJ76b] H. B. Peng and L. F. Jaffe. A simple, selective method for freezefracturing spherical cells. Journal of Cell Biology, 71(2):674–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
2/674.

Powell:1977:ECF

[PJB77] J. T. Powell, U. Järlfors, and K. Brew. Enzymic characteristics of fat globule membranes from bovine colostrum and bovine milk. Journal of Cell Biology, 72(3):617–??, March 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/72/3/617.

Pereira:1979:VIC

[PK79] M. E. Pereira and E. A. Kabat. A versatile immunoadsorbent capable of binding lectins of various specificities and its use for the separation of cell populations. Journal of Cell Biology, 82(1):185– ??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
1/185.

Pilar:1976:UDD

[PL76] G. Pilar and L. Landmesser. Ultrastructural differences during embryonic cell death in normal and peripherally deprived ciliary ganglia. Journal of Cell Biology, 68(2):339–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/2/339.

Paiement:1977:LTA

[PL77] J. Paiement and C. P. Leblond. Localization of thyroglobulin antigenicity in rat thyroid sections using antibodies labled with peroxidase or (125)I-radioiodine. Journal of Cell Biology, 74(3): 992–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/992.

Paulson:1975:MAT

[PM75] J. C. Paulson and W. O. McClure. Microtubules and axoplasmic transport. Inhibition of transport by podophyllotoxin: an interaction with microtubule protein. Journal of Cell Biology, 67 (2):461–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/2/461.

Porvaznik:1979:DGJ

[PM79] M. Porvaznik and T. J. MacVittie. Detection of gap junctions between the progeny of a canine macrophage colony-forming cell in vitro. Journal of Cell Biology, 82(2):555–??, August 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/82/2/555.

Pohjanpelto:1976:PTG

[Poh76] P. Pohjanpelto. Putrescine transport is greatly increased in human fibroblasts initiated to proliferate. Journal of Cell Biology, 68(3):512–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/3/512.

Pollard:1975:EMS

[Pol75] T. D. Pollard. Electron microscopy of synthetic myosin filaments. Evidence for cross-bridge. Flexibility and copolymer formation. Journal of Cell Biology, 67(1):93–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/67/1/93.

Pollard:1976:RAT

[Pol76] T. D. Pollard. The role of actin in the temperature-dependent gelation and contraction of extracts of Acanthamoeba. Journal of Cell Biology, 68(3):579–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/3/579.

Parvinen:1979:AMC

[PP79] M. Parvinen and L. M. Parvinen. Active movements of the chromatoid body. A possible transport mechanism for haploid gene products. Journal of Cell Biology, 80(3):621–??, March 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/3/621.

Pickett:1975:OJC

[PPHM75] P. B. Pickett, D. R. Pitelka, S. T. Hamamoto, and D. S. Misfeldt. Occluding junctions and cell behavior in primary cultures of normal and neoplastic mammary gland cells. Journal of Cell Biology, 66(2):316–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/316.

Poo:1978:LED

[PPL78] M. M. Poo, W. J. Poo, and J. W. Lam. Lateral electrophoresis and diffusion of concanavalin A receptors in the membrane of embryonic muscle cell. Journal of Cell Biology, 76(2):483–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/483.

Pipeleers:1977:PRT

[PPMK77] D. G. Pipeleers, M. A. Pipeleers-Marichal, and D. M. Kipnis. Physiological regulation of total tubulin and polymerized tubulin in tissues. Journal of Cell Biology, 74(2):351–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/351.

Pipeleers:1977:SMM

[PPMSK77] D. G. Pipeleers, M. A. Pipeleers-Marichal, P. Sherline, and D. M. Kipnis. A sensitive method for measuring polymerized and depolymerized forms of tubulin in tissues. Journal of Cell Biology, 74(2):341–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/2/341.

Pfenninger:1975:MFF

[PR75] K. H. Pfenninger and E. R. Rinderer. Methods for the freezefracturing of nerve tissue cultures and cell monolayers. Journal of Cell Biology, 65(1):15–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/1/15.

Phillips:1976:DCF

[PR76a] S. G. Phillips and J. B. Rattner. Dependence of centriole formation on protein synthesis. Journal of Cell Biology, 70(1):9–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/70/1/9.

Pisam:1976:RSM

[PR76b] M. Pisam and P. Ripoche. Redistribution of surface macromolecules in dissociated epithelial cells. Journal of Cell Biology, 71(3):907–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/3/907.

Peters:1978:DSS

[PRGH78] K. R. Peters, G. Rutter, H. H. Gschwender, and W. Haller. Derivatized silica spheres as immunospecific markers for high resolution labeling in electron microscopy. Journal of Cell Biology, 78(2):309–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/2/309.

Prop:1975:LEU

[Pro75] F. J. Prop. Letters to the editor: Is the upper cell surface unable to support active cell movement in culture? Journal of Cell Biology, 66(1):215–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/1/215.

Plagemenn:1975:CCG

[PRZE75] P. G. Plagemenn, D. P. Richey, J. M. Zylka, and J. Erbe. Cell cycle and growth stage-dependent changes in the transport of nucleosides, hypoxanthine, choline, and deoxyglucose in cultured Novikoff rat hepatoma cells. Journal of Cell Biology, 64(1):29–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
1/29.

Pfeifer:1975:MSC

[PS75] U. Pfeifer and H. Scheller. A morphometric study of cellular autophagy including diurnal variations in kidney tubules of normal rats. Journal of Cell Biology, 64(3):608–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/3/608.

Pietras:1979:EIM

[PS79] R. J. Pietras and C. M. Szego. Estrogen-induced membrane alterations and growth associated with proteinase activity in endometrial cells. Journal of Cell Biology, 81(3):649–??, June 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/81/3/649.

Pietra:1976:MLL

[PSC^{+}76] G. G. Pietra, L. G. Spagnoli, D. M. Capuzzi, C. E. Sparks, A. P. Fishman, and J. B. Marsh. Metabolism of 125I-labeled lipoproteins by the isolated rat lung. Journal of Cell Biology, 70(1):33–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/70/1/33.

Price:1976:GSS

[PSG^{+}76] D. L. Price, A. Stocks, J. W. Griffin, A. Young, and K. Peck. Glycine-specific synapses in rat spinal cord. Identification by electron microscope autoradiography. Journal of Cell Biology, 68 (2):389–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/389.

Pant:1978:NPP

[PSGL78] H. C. Pant, G. Shecket, H. Gainer, and R. J. Lasek. Neurofilament protein is phosphorylated in the squid giant axon. Journal of Cell Biology, 78(2):??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/2/R23.

Patzelt:1975:CBP

[PSM^{+}75] C. Patzelt, A. Singh, Y. L. Marchand, L. Orci, and B. Jeanrenaud. Colchicine-binding protein of the liver. Its characterization and relation to microtubules. Journal of Cell Biology, 66 (3):609–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/609.

Papermaster:1978:ILL

[PSZK78a] D. S. Papermaster, B. G. Schneider, M. A. Zorn, and J. P. Kraehenbuhl. Immunocytochemical localization of a large intrinsic membrane protein to the incisures and margins of frog rod outer segment disks. Journal of Cell Biology, 78(2):415–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/2/415.

Papermaster:1978:ILO

[PSZK78b] D. S. Papermaster, B. G. Schneider, M. A. Zorn, and J. P. Kraehenbuhl. Immunocytochemical localization of opsin in outer segments and Golgi zones of frog photoreceptor cells. An electron microscope analysis of cross-linked albumin-embedded retinas. Journal of Cell Biology, 77(1):196–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/1/196.

Pagano:1977:APV

[PT77] R. E. Pagano and M. Takeichi. Adhesion of phospholipid vesicles to Chinese hamster fibroblasts: Role of cell surface proteins. Journal of Cell Biology, 74(2):531–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/531.

Piasek:1979:ECE

[PT79] A. Piasek and J. Thyberg. Effects of colchicine on endocytosis and cellular inactivation of horseradish peroxidase in cultured chondrocytes. Journal of Cell Biology, 81(2):426–??, May 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/81/2/426.

Penningroth:1978:EAI

[PW78] S. M. Penningroth and G. B. Witman. Effects of adenylyl imidodiphosphate, a nonhydrolyzable adenosine triphosphate analog, on reactivated and rigor wave sea urchin sperm. Journal of Cell Biology, 79(3):827–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/3/827.

Pauli:1977:FFM

[PWSA77] B. Pauli, R. S. Weinstein, L. W. Soble, and J. Alroy. Freezefracture of monolayer cultures. Journal of Cell Biology, 72(3):763– ??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
3/763.

Quintart:1979:ECA

[QLHTB79] J. Quintart, M. A. Leroy-Houyet, A. Trouet, and P. Baudhuin. Endocytosis and chloroquine accumulation during the cell cycle of hepatoma cells in culture. Journal of Cell Biology, 82 (3):644–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/644.

Quigley:1976:APP

[Qui76] J. P. Quigley. Association of a protease (plasminogen activator) with a specific membrane fraction isolated from transformed cells. Journal of Cell Biology, 71(2):472–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/472.

Quaroni:1979:ECC

[QWTI79] A. Quaroni, J. Wands, R. L. Trelstad, and K. J. Isselbacher. Epithelioid cell cultures from rat small intestine. Characterization by morphologic and immunologic criteria. Journal of Cell Biology, 80(2):248–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/248.

Reddi:1976:CBM

[RA76] A. H. Reddi and W. A. Anderson. Collagenous bone matrixinduced endochondral ossification hemopoiesis. Journal of Cell Biology, 69(3):557–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/3/557.

Reimers:1975:TSS

[RAFM75] H. J. Reimers, D. J. Allen, I. A. Feuerstein, and J. F. Mustard. Transport and storage of serotonin by thrombin-treated platelets. Journal of Cell Biology, 65(2):359–??, May 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/65/2/359.

Raghavan:1977:PDS

[Rag77] V. Raghavan. Patterns of DNA synthesis during pollen embryogenesis in henbane. Journal of Cell Biology, 73(2):521–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/2/521.

Rodman:1976:RSP

[RB76] T. C. Rodman and R. Bachvarova. RNA synthesis in preovulatory mouse oocytes. Journal of Cell Biology, 70(1):251–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/1/251.

Rieder:1977:HIR

[RB77] C. Rieder and A. S. Bajer. Heat-induced reversible hexagonal packing of spindle microtubules. Journal of Cell Biology, 74(3): 717–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/717.

Rees:1976:MCN

[RBB76] R. P. Rees, M. B. Bunge, and R. P. Bunge. Morphological changes in the neuritic growth cone and target neuron during synaptic junction development in culture. Journal of Cell Biology, 68 (2):240–??, February 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/2/240.

Rodan:1977:MCD

[RBC77] G. A. Rodan, L. A. Bourret, and L. S. Cutler. Membrane changes during cartilage maturation. Increase in 5’-nucleotidase and decrease in adenosine inhibition of adenylate cyclase. Journal of Cell Biology, 72(2):493–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/2/493.

Redman:1975:CIP

[RBHP75] C. M. Redman, D. Banerjee, K. Howell, and G. E. Palade. Colchicine inhibition of plasma protein release from rat hepatocytes. Journal of Cell Biology, 66(1):42–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/1/42.

Robinson:1978:LDA

[RBK78] J. M. Robinson, R. T. Briggs, and M. J. Karnovsky. Localization of D-amino acid oxidase on the cell surface of human polymorphonuclear leukocytes. Journal of Cell Biology, 77(1):59–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/77/1/59.

Redman:1978:VEC

[RBM^{+}78] C. M. Redman, D. Banerjee, C. Manning, C. Y. Huang, and K. Green. In vivo effect of colchicine on hepatic protein synthesis and on the conversion of proalbumin to serum albumin. Journal of Cell Biology, 77(2):400–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/2/400.

Reaven:1977:QAT

[RCM77] E. P. Reaven, Y. Cheng, and M. D. Miller. Quantitative analysis of tubulin and microtubule compartments in isolated rat hepatocytes. Journal of Cell Biology, 75(3):731–??, December 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/75/3/731.

Romrell:1975:ISH

[RCMI75] L. J. Romrell, M. R. Coppe, D. R. Munro, and S. Ito. Isolation and separation of highly enriched fractions of viable mouse gastric parietal cells by velocity sedimentation. Journal of Cell Biology, 65(2):428–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/428.

Rindler:1979:RDP

[RCSS79] M. J. Rindler, L. M. Chuman, L. Shaffer, and M. H. Saier. Retention of differentiated properties in an established dog kidney epithelial cell line (MDCK). Journal of Cell Biology, 81(3):635– ??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/635.

Rennke:1975:RMC

[RCV75] H. G. Rennke, R. S. Cotran, and M. A. Venkatachalam. Role of molecular charge in glomerular permeability. Tracer studies with cationized ferritins. Journal of Cell Biology, 67(3):638–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/638.

Ridgway:1976:OCI

[RD76] E. B. Ridgway and A. C. Durham. Oscillations of calcium ion concentrations in Physarum polycephalum. Journal of Cell Biology, 69(1):223–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/223.

Reggio:1978:IIB

[RD78] H. Reggio and J. C. Dagorn. Ionic interactions between bovine chymotrypsinogen A and chondroitin sulfate A.B.C.. A possible model for molecular aggregation in zymogen granules. Journal of Cell Biology, 78(3):951–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/951.

Ryter:1977:MCS

[RdC77] A. Ryter and C. de Chastellier. Morphometric and cytochemical studies of Dictyostelium discoideum in vegetative phase. Digestive system and membrane turnover. Journal of Cell Biology, 75 (1):200–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/200.

Remacle:1978:BCB

[Rem78] J. Remacle. Binding of cytochrome b5 to membranes of isolated subcellular organelles from rat liver. Journal of Cell Biology, 79 (2):291–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/291.

Raybin:1977:MTT

[RF77] D. Raybin and M. Flavin. Modification of tubulin by tyrosylation in cells and extracts and its effect on assembly in vitro. Journal of Cell Biology, 73(2):492–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/2/492.

Remacle:1976:ASM

[RFB^{+}76] J. Remacle, S. Fowler, H. Beaufay, A. Amarcostesec, and J. Berthet. Analytical study of microsomes and isolated subcellular membranes from rat liver. VI. Electron microscope examination of microsomes for cytochrome b5 by means of a ferritinlabeled antibody. Journal of Cell Biology, 71(2):551–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
2/551.

Raviola:1975:IOS

[RG75] E. Raviola and N. B. Gilula. Intramembrane organization of specialized contacts in the outer plexiform layer of the retina. A freeze-fracture study in monkeys and rabbits. Journal of Cell Biology, 65(1):192–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/1/192.

Rice:1978:RPS

[RG78] Rh Rice and H. Green. Relation of protein synthesis and transglutaminase activity to formation of the cross-linked envelope during terminal differentiation of the cultured human epidermal keratinocyte. Journal of Cell Biology, 76(3):705–??, March 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/3/705.

Rutishauser:1978:AANb

[RGE78] U. Rutishauser, W. E. Gall, and G. M. Edelman. Adhesion among neural cells of the chick embryo. IV. Role of the cell surface molecule CAM in the formation of neurite bundles in cultures of spinal ganglia. Journal of Cell Biology, 79(2):382–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/2/382.

Rubin:1978:DBN

[RGK78] R. W. Rubin, L. Goldstein, and C. Ko. Differences between nucleus and cytoplasm in the degree of actin polymerization. Journal of Cell Biology, 77(3):698–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/3/698.

Reville:1976:CAP

[RGS^{+}76] W. J. Reville, D. E. Goll, M. H. Stromer, R. M. Robson, and W. R. Dayton. A Ca^{2+}-activated protease possibly involved in myofibrillar protein turnover. Subcellular localization of the protease in porcine skeletal muscle. Journal of Cell Biology, 70(1):1–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/1/1.

Rinaldi:1975:OTT

[RH75] R. A. Rinaldi and B. Hrebenda. Oriented thick and thin filaments in amoeba proteus. Journal of Cell Biology, 66(1):193–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/1/193.

Reid:1977:RMC

[RH77] B. J. Reid and L. H. Hartwell. Regulation of mating in the cell cycle of Saccharomyces cerevisiae. Journal of Cell Biology, 75 (2):355–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/355.

Rattner:1979:NPI

[RH79] J. B. Rattner and B. A. Hamkalo. Nucleosome packing in interphase chromatin. Journal of Cell Biology, 81(2):453–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/453.

Rudland:1977:MAP

[RHDL77] P. S. Rudland, R. C. Hallowes, H. Durbin, and D. Lewis. Mitogenic activity of pituitary hormones on cell cultures of normal and carcinogen-induced tumor epithelium from rat mammary glands. Journal of Cell Biology, 73(3):561–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/3/561.

Rubin:1976:IEA

[RHHB76] R. W. Rubin, M. C. Hill, P. Hepworth, and J. Boehmer. Isolation and electrophoretic analysis of nucleoli, phenol-soluble nuclear proteins, and outer cyst walls from Acanthamoeba castellanii during encystation initiation. Journal of Cell Biology, 68(3):740– ??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/740.

Rieder:1979:RSC

[Rie79] C. L. Rieder. Ribonucleoprotein staining of centrioles and kinetochores in newt lung cell spindles. Journal of Cell Biology, 80(1):1–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/1.

Rikmenspoel:1978:MSU

[Rik78a] R. Rikmenspoel. Movement of sea urchin sperm flagella. Journal of Cell Biology, 76(2):310–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/2/310.

Ritter:1978:MBS

[RIK78b] H. Ritter, S. Inoué, and D. Kubai. Mitosis in Barbulanympha. I. Spindle structure, formation, and kinetochore engagement. Journal of Cell Biology, 77(3):638–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/77/3/638.

Rosano:1976:DCMa

[RJ76] T. G. Rosano and D. H. Jones. Developmental changes in mitochondria during the transition into lactation in the mouse mammary gland. Journal of Cell Biology, 69(3):573–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/3/573.

Robertson:1976:RSU

[RKL^{+}76] J. D. Robertson, S. Knutton, A. R. Limbrick, E. R. Jakoi, and G. Zampighi. Regular structures in unit membranes. III. Further observations on the particulate component of the suckling rat ileum endocytic membrane complex. Journal of Cell Biology, 70(1):112–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/1/112.

Rosano:1976:DCMb

[RLJ76] T. G. Rosano, S. K. Lee, and D. H. Jones. Developmental changes in mitochondria during the transition into lactation in the mouse mammary gland. II. Membrane marker enzymes and membrane ultrastructure. Journal of Cell Biology, 69(3):581–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/69/3/581.

Rodman:1979:LHM

[RLRV79] T. C. Rodman, S. D. Litwin, M. Romani, and G. Vidali. Life history of mouse sperm protein. Intratesticular stages. Journal of Cell Biology, 80(3):605–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/3/605.

Reimers:1975:TAN

[RMP75] H. J. Reimers, J. F. Mustard, and M. A. Packham. Transfer of adenine nucleotides between the releasable and nonreleasable compartments of rabbit blood platelets. Journal of Cell Biology, 67(1):61–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/1/61.

Rubinstein:1978:UTS

[RMP^{+}78] N. Rubinstein, K. Mabuchi, F. Pepe, S. Salmons, J. Gergely, and F. Sreter. Use of type-specific antimyosins to demonstrate the transformation of individual fibers in chronically stimulated rabbit fast muscles. Journal of Cell Biology, 79(1):252–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/252.

Rodewald:1976:CIB

[RNK76] R. Rodewald, S. B. Newman, and M. J. Karnovsky. Contraction of isolated brush borders from the intestinal epithelium. Journal of Cell Biology, 70(3):541–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/3/541.

Rodewald:1976:PDB

[Rod76] R. Rodewald. pH-dependent binding of immunoglobulins to intestinal cells of the neonatal rat. Journal of Cell Biology, 71 (2):666–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/2/666.

Roos:1975:MCS

[Roo75] U. P. Roos. Mitosis in the cellular slime mold Polysphondylium violaceum. Journal of Cell Biology, 64(2):480–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/2/480.

Rosenbluth:1978:PAE

[Ros78] J. Rosenbluth. Particle arrays in earthworm postjunctional membranes. Journal of Cell Biology, 76(1):76–??, January 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/1/76.

Routledge:1978:CBP

[Rou78] L. M. Routledge. Calcium-binding proteins in the vorticellid spasmoneme. Journal of Cell Biology, 77(2):358–??, May 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/2/358.

Rifkin:1977:PPA

[RP77] D. B. Rifkin and R. Pollack. Production of plasminogen activator by established cell lines of mouse origin. Journal of Cell Biology, 73(1):47–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/1/47.

Reggio:1978:SCZ

[RP78] H. A. Reggio and G. E. Palade. Sulfated compounds in the zymogen granules of the guinea pig pancreas. Journal of Cell Biology, 77(2):288–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/288.

Rall:1977:GMV

[RPGR77] L. Rall, R. Pictet, S. Githens, and W. J. Rutter. Glucocorticoids modulate the in vitro development of the embryonic rat pancreas.

Journal of Cell Biology, 75(2):398–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/2/398.

Rohrlich:1975:BCC

[RR75] S. T. Rohrlich and R. W. Rubin. Biochemical characterization of crystals from the dermal iridophores of a chameleon Anolis carolinensis. Journal of Cell Biology, 66(3):635–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/3/635.

Rutherford:1976:PFS

[RR76] R. B. Rutherford and R. Ross. Platelet factors stimulate fibroblasts and smooth muscle cells quiescent in plasma serum to proliferate. Journal of Cell Biology, 69(1):196–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/1/196.

Reaven:1977:DCM

[RR77a] E. P. Reaven and G. M. Reaven. Distribution and content of microtubules in relation to the transport of lipid. An ultrastructural quantitative study of the absorptive cell of the small intestine. Journal of Cell Biology, 75(2):559–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/2/559.

Rohrlich:1977:PPA

[RR77b] S. T. Rohrlich and D. B. Rifkin. Patterns of plasminogen activator production in cultured normal embryonic cells. Journal of Cell Biology, 75(1):31–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/75/1/31.

Reaven:1978:DBR

[RR78] E. P. Reaven and G. M. Reaven. Dissociation between rate of hepatic lipoprotein secretion and hepatocyte microtubule content. Journal of Cell Biology, 77(3):735–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/735.

Royal:1975:RSU

[RS75a] A. Royal and R. Simard. RNA synthesis in the ultrastructural and biochemical components of the nucleolus of Chinese hamster ovary cells. Journal of Cell Biology, 66(3):577–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/66/3/577.

Rutishauser:1975:CCB

[RS75b] U. Rutishauser and L. Sachs. Cell-to-cell binding induced by different lectins. Journal of Cell Biology, 65(2):247–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/2/247.

Rutishauser:1975:RMB

[RS75c] U. Rutishauser and L. Sachs. Receptor mobility and the binding of cells to lectin-coated fibers. Journal of Cell Biology, 66(1):76–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/66/1/76.

Riou:1979:CMC

[RS79] G. F. Riou and J. M. Saucier. Characterization of the molecular components in kinetoplast-mitochondrial DNA of Trypanosoma equiperdum. Comparative study of the dyskinetoplastic and wild strains. Journal of Cell Biology, 82(1):248–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/1/248.

Ripps:1976:PUP

[RSM76] H. Ripps, M. Shakib, and E. D. MacDonald. Peroxidase uptake by photoreceptor terminals of the skate retina. Journal of Cell Biology, 70(1):86–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/1/86.

Rubin:1976:SLG

[RT76] H. Rubin and R. N. Trelease. Subcellular localization of glyoxylate cycle enzymes in Ascaris suum larvae. Journal of Cell Biology, 70 (2):374–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/2/374.

Rutishauser:1978:AANa

[RTBE78] U. Rutishauser, J. P. Thiery, R. Brackenbury, and G. M. Edelman. Adhesion among neural cells of the chick embryo. III. Relationship of the surface molecule CAM to cell adhesion and the development of histotypic patterns. Journal of Cell Biology, 79 (2):371–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/371.

Ruben:1976:ITL

[RTC76] G. C. Ruben, J. N. Telford, and R. C. Carroll. Identification and transmembranous localization of active cytochrome oxidase in reconstituted membranes of purified phospholipids by electron microscopy. Journal of Cell Biology, 68(3):724–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/724.

Rawlins:1978:ALA

[RV78] F. A. Rawlins and J. Villegas. Autoradiographic localization of acetylcholine receptors in the Schwann cell membrane of the squid nerve fiber. Journal of Cell Biology, 77(2):371–??, May 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/2/371.

Rahman:1975:LCC

[RW75a] Y. E. Rahman and B. J. Wright. Liposomes containing chelating agents. Cellular penetration and a possible mechanism of metal removal. Journal of Cell Biology, 65(1):112–??, April 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/65/1/112.

Rubin:1975:DBM

[RW75b] R. W. Rubin and G. D. Weiss. Direct biochemical measurements of microtubule assembly and disassembly in Chinese hamster ovary cells. The effect of intercellular contact, cold, D2O, and N6,O2’-dibutyryl cyclic adenosine monophosphate. Journal of Cell Biology, 64(1):42–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/42.

Reinecke:1978:ATB

[RW78] M. Reinecke and C. Walther. Aspects of turnover and biogenesis of synaptic vesicles at locust neuromuscular junctions as revealed by zinc iodide-osmium tetroxide (ZIO) reacting with intravesicular SH-groups. Journal of Cell Biology, 78(3):839–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/839.

Rubin:1979:OTN

[RW79] R. W. Rubin and R. H. Warren. Organization of tubulin in normal and transformed rat kidney cells. Journal of Cell Biology, 82(1):103–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/103.

Rubin:1978:ACO

[RWLC78] R. W. Rubin, R. H. Warren, D. S. Lukeman, and E. Clements. Actin content and organization in normal and transformed cells in culture. Journal of Cell Biology, 78(1):28–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/78/1/28.

Stacey:1977:EAM

[SA77] D. W. Stacey and V. G. Allfrey. Evidence for the autophagy of microinjected proteins in HeLA cells. Journal of Cell Biology, 75 (3):807–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/807.

Sanders:1975:HYT

[SAB75] S. K. Sanders, E. L. Alexander, and R. C. Braylan. A high-yield technique for preparing cells fixed in suspension for scanning electron microscopy. Journal of Cell Biology, 67(2):476–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/2/476.

Salmon:1975:PIDa

[Sal75a] E. D. Salmon. Pressure-induced depolymerization of spindle microtubules. I. Changes in birefringence and spindle length. Journal of Cell Biology, 65(3):603–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/3/603.

Salmon:1975:PIDb

[Sal75b] E. D. Salmon. Pressure-induced depolymerization of spindle microtubules. II. Thermodynamics of in vivo spindle assembly. Journal of Cell Biology, 66(1):114–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/114.

Sanders:1978:FNS

[San78] M. M. Sanders. Fractionation of nucleosomes by salt elution from micrococcal nuclease-digested nuclei. Journal of Cell Biology, 79 (1):97–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/1/97.

Scrodilis:1977:CMP

[SAPA77] S. P. Scrodilis, J. L. Anderson, R. Pollack, and R. S. Adelstein. Characterization of the myosin-phosphorylating system in normal murine astrocytes and derivative sv40 wild-type and A-mutant transformant. Journal of Cell Biology, 74(3):940–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/3/940.

Schrevel:1975:FFP

[SB75] J. Schrevel and C. Besse. [a functional flagella with a 6 + 0 pattern]. Journal of Cell Biology, 66(3):492–??, September 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/3/492.

Silver:1976:AAB

[SB76] J. Silver and R. B. Billiar. An autoradiographic analysis of [3H]alpha-bungarotoxin distribution in the rat brain after intraventricular injection. Journal of Cell Biology, 71(3):956–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
3/956.

Springer:1978:DMS

[SB78] W. R. Springer and S. H. Barondes. Direct measurement of species-specific cohesion in cellular slime molds. Journal of Cell Biology, 78(3):937–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/78/3/937.

Stern:1979:CEF

[SB79] P. L. Stern and M. S. Bretscher. Capping of exogenous Forssman glycolipid on cells. Journal of Cell Biology, 82(3):829–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
3/829.

Steinman:1976:MFD

[SBC76] R. M. Steinman, S. E. Brodie, and Z. A. Cohn. Membrane flow during pinocytosis. A stereologic analysis. Journal of Cell Biology, 68(3):665–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/3/665.

Sieber-Blum:1978:LBN

[SBC78] M. Sieber-Blum and A. M. Cohen. Lectin binding to neural crest cells. Changes of the cell surface during differentiation in vitro. Journal of Cell Biology, 76(3):628–??, March 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/3/628.

Starger:1978:BIA

[SBGG78] J. M. Starger, W. E. Brown, A. E. Goldman, and R. D. Goldman. Biochemical and immunological analysis of rapidly purified 10-nm filaments from baby hamster kidney (BHK-21) cells. Journal of Cell Biology, 78(1):93–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/1/93.

Sandoz:1976:CMCa

[SBU76] D. Sandoz and E. Biosvieux-Ulrich. [ciliogenesis in the mucous cells of the quail oviduct. I. Ultrastructural study in the laying quail]. Journal of Cell Biology, 71(2):449–??, November 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/2/449.

Sandoz:1976:CMCb

[SBULB76] D. Sandoz, E. Biosvieux-Ulrich, C. Laugier, and E. Brard. [ciliogenesis in the mucous cells of the quail oviduct. II. Hormonal control]. Journal of Cell Biology, 71(2):460–??, November 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/2/460.

Spooner:1975:REM

[SC75a] B. S. Spooner and G. W. Conrad. The role of extracellular materials in cell movement. I. Inhibition of mucopolysaccharide synthesis does not stop ruffling membrane activity or cell movement. Journal of Cell Biology, 65(2):286–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/2/286.

Spurlock:1975:FSS

[SC75b] B. O. Spurlock and M. J. Cormier. A fine structure study of the anthocodium in Renilla mülleri. Evidence for the existence of a bioluminescent organelle, the luminelle. Journal of Cell Biology, 64(1):15–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/1/15.

Szekely:1976:RID

[SC76] J. G. Szekely and T. P. Copps. Radiation-induced delay of nuclear pore formation. Journal of Cell Biology, 70(2):466–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/2/466.

Schroeder:1977:BHC

[SC77] B. T. Schroeder and J. Chakraborty. Binding of [(3)H]oxytocin to cells isolated from the mammary gland of the lactating rat. Journal of Cell Biology, 74(2):428–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/428.

Stanley:1978:BWG

[SC78] P. Stanley and J. P. Carver. Binding of [125I] wheat germ agglutinin to Chinese hamster ovary cells under conditions which affect the mobility of membrane components. Journal of Cell Biology, 79(3):617–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/617.

Schuetz:1975:CAS

[Sch75] A. W. Schuetz. Cytoplasmic activation of starfish oocytes by sperm and divalent ionophore A-23187. Journal of Cell Biology, 66(1):86–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/1/86.

Schliwa:1976:RDC

[Sch76] M. Schliwa. The role of divalent cations in the regulation of microtubule assembly. In vivo studies on microtubules of the heliozoan axopodium using the ionophore A23187. Journal of Cell Biology, 70(3):527–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/3/527.

Schlaepfer:1977:IUS

[Sch77] W. W. Schlaepfer. Immunological and ultrastructural studies of neurofilaments isolated from rat peripheral nerve. Journal of Cell Biology, 74(1):226–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/1/226.

Schlaepfer:1978:ODI

[Sch78a] W. W. Schlaepfer. Observations on the disassembly of isolated mammalian neurofilaments. Journal of Cell Biology, 76(1):50–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
1/50.

Schliwa:1978:MAM

[Sch78b] M. Schliwa. Microtubular apparatus of melanophores: three dimensional organization. Journal of Cell Biology, 76(3):605–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
3/605.

Stuart:1977:RTI

[SCRP77] S. E. Stuart, G. A. Clawson, F. M. Rottman, and R. J. Patterson. RNA transport in isolated myeloma nuclei. Transport from membrane-denuded nuclei. Journal of Cell Biology, 72(1):57–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
1/57.

Siu:1976:CCNc

[SCS76] C. Siu, K. Chiang, and H. Swift. Characterization of cytoplasmic and nuclear genomes in the colorless alga Polytoma. III. Ribosomal RNA cistrons of the nucleus and leucoplast. Journal of Cell Biology, 69(2):383–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/2/383.

Skutelsky:1976:RSA

[SD76] E. Skutelsky and D. Danon. Redistribution of surface anionic sites on the luminal front of blood vessel endothelium after interaction with polycationic ligand. Journal of Cell Biology, 71(1):232–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
1/232.

Schmidt:1979:NTA

[SDTD^{+}79] G. W. Schmidt, A. Devillers-Thiery, H. Desruisseaux, G. Blobel, and N. H. Chua. NH2-terminal amino acid sequences of precursor and mature forms of the ribulose-1,5-bisphosphate carboxylase small subunit from Chlamydomonas reinhardtii. Journal of Cell Biology, 83(3):615–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/3/615.

Salmon:1975:NMH

[SE75] E. D. Salmon and G. W. Ellis. A new miniature hydrostatic pressure chamber for microscopy. Strain-free optical glass windows facilitate phase-contrast and polarized-light microscopy of living cells. Optional fixture permits simultaneous control of pressure and temperature. Journal of Cell Biology, 65(3):587–??, June 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/65/3/587.

Schmell:1977:ISR

[SEBL77] E. Schmell, B. J. Earles, C. Breaux, and W. J. Lennarz. Identification of a sperm receptor on the surface of the eggs of the sea urchin Arbacia punctulata. Journal of Cell Biology, 72(1):35–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
1/35.

Schliwa:1979:ERS

[SEHW79] M. Schliwa, U. Euteneuer, W. Herzog, and K. Weber. Evidence for rapid structural and functional changes of the melanophore microtubule-organizing center upon pigment movements. Journal of Cell Biology, 83(3):623–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/83/3/623.

Sato:1975:MOM

[SEI75] H. Sato, G. W. Ellis, and S. Inoué. Microtubular origin of mitotic spindle form birefringence. Demonstration of the applicability of Wiener’s equation. Journal of Cell Biology, 67(3):501–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/501.

Sharoni:1976:SOV

[SES76] Y. Sharoni, S. Eimerl, and M. Schramm. Secretion of old versus new exportable protein in rat parotid slics. Control by neurotransmitters. Journal of Cell Biology, 71(1):107–??, October 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/1/107.

Skutelsky:1976:VDC

[SF76] E. Skutelsky and M. G. Farquhar. Variations in distribution of con A receptor sites and anionic groups during red blood cell differentiation in the rat. Journal of Cell Biology, 71(1):218–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
1/218.

Schlaepfer:1978:NPR

[SF78] W. W. Schlaepfer and L. A. Freeman. Neurofilament proteins of rat peripheral nerve and spinal cord. Journal of Cell Biology, 78 (3):653–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/653.

Simmons:1975:MAG

[SFFB75] J. L. Simmons, P. H. Fishman, E. Freese, and R. O. Brady. Morphological alterations and ganglioside sialyltransferase activity induced by small fatty acids in HeLa cells. Journal of Cell Biology, 66(2):414–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/2/414.

Salpeter:1977:REM

[SFS77] M. M. Salpeter, H. C. Fertuck, and E. E. Salpeter. Resolution in electron microscope autoradiography. III. Iodine-125, the effect of heavy metal staining, and a reassessment of critical parameters. Journal of Cell Biology, 72(1):161–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/1/161.

Schwarz:1978:PAT

[SFSB78] R. I. Schwarz, D. A. Farson, W. J. Soo, and M. J. Bissell. Primary avian tendon cells in culture. An improved system for understanding malignant transformation. Journal of Cell Biology, 79(3):672–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/672.

Sale:1979:SMV

[SG79a] W. S. Sale and I. R. Gibbon. Study of the mechanism of vanadate inhibition of the dynein cross-bridge cycle in sea urchin sperm flagella. Journal of Cell Biology, 82(1):291–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/1/291.

Slot:1979:MSE

[SG79b] J. W. Slot and J. J. Geuze. A morphometrical study of the exocrine pancreatic cell in fasted and fed frogs. Journal of Cell Biology, 80(3):692–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/3/692.

Somlyo:1975:GOR

[SGCS75] A. P. Somlyo, R. E. Garfield, S. Chacko, and A. V. Somlyo. Golgi organelle response to the antibiotic X537A. Journal of Cell Biology, 66(2):425–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/66/2/425.

Salmon:1976:PID

[SGMB76] E. D. Salmon, D. Goode, T. K. Maugel, and D. B. Bonar. Pressure-induced depolymerization of spindle microtubules. III. Differential stability in HeLa cells. Journal of Cell Biology, 69(2):443–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/443.

Sandler:1975:ESC

[SGV75] J. A. Sandler, J. I. Gallin, and M. Vaughan. Effects of serotonin, carbamylcholine, and ascorbic acid on leukocyte cyclic GMP and chemotaxis. Journal of Cell Biology, 67(2):480–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/2/480.

Stossel:1976:IAM

[SH76] T. P. Stossel and J. H. Hartwig. Interactions of actin, myosin, and a new actin-binding protein of rabbit pulmonary macrophages. II. Role in cytoplasmic movement and phagocytosis. Journal of Cell Biology, 68(3):602–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/68/3/602.

Shore:1977:FPS

[SH77a] G. C. Shore and R. Harris. Fate of polypeptides synthesized on rough microsomal vesicles in a messenger-dependent rabbit reticulocyte system. Journal of Cell Biology, 74(1):315–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/1/315.

Steinkamp:1977:ALP

[SH77b] M. P. Steinkamp and L. L. Hoefert. Annulate lamellae in phloem cells of virus-infected Sonchus plants. Journal of Cell Biology, 74(1):111–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/111.

Sanes:1979:ABS

[SH79a] J. R. Sanes and Z. W. Hall. Antibodies that bind specifically to synaptic sites on muscle fiber basal lamina. Journal of Cell Biology, 83(2):357–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/83/2/357.

Slaughter:1979:MBR

[SH79b] S. R. Slaughter and D. E. Hultquist. Membrane-bound redox proteins of the murine Friend virus-induced erythroleukemia cell. Journal of Cell Biology, 83(1):231–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/1/231.

Saccomani:1979:CGM

[SHC^{+}79] G. Saccomani, H. F. Helander, S. Crago, H. H. Chang, D. W. Dailey, and G. Sachs. Characterization of gastric mucosal membranes. X. Immunological studies of gastric (H^{+} + K^{+})-ATPase. Journal of Cell Biology, 83(2):271–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/2/271.

Sheetz:1979:DDD

[She79] M. P. Sheetz. DNase-i-dependent dissociation of erythrocyte cytoskeletons. Journal of Cell Biology, 81(1):266–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/1/266.

Schacher:1976:SAF

[SHH76] S. Schacher, E. Holtzman, and D. C. Hood. Synaptic activity of frog retinal photoreceptors. A peroxidase uptake study. Journal of Cell Biology, 70(1):178–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/1/178.

Steven:1977:UPP

[SHM^{+}77] A. C. Steven, B. Heggeler, R. Müller, J. Kistler, and J. P. Rosenbusch. Ultrastructure of a periodic protein layer in the outer membrane of Escherichia coli. Journal of Cell Biology, 72(2):292–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
2/292.

Schachner:1977:ULG

[SHWH^{+}77] M. Schachner, E. T. Hedley-Whyte, D. W. Hsu, G. Schoonmaker, and A. Bignami. Ultrastructural localization of glial fibrillary acidic protein in mouse cerebellum by immunoperoxidase labeling. Journal of Cell Biology, 75(1):67–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/1/67.

Sheridan:1978:QAL

[SHWPJ78] J. D. Sheridan, M. Hammer-Wilson, D. Preus, and R. G. Johnson. Quantitative analysis of low-resistance junctions between cultured cells and correlation with gap-junctional areas. Journal of Cell Biology, 76(2):532–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/2/532.

Summers:1979:CPA

[SK79] K. Summers and M. W. Kirschner. Characteristics of the polar assembly and disassembly of microtubules observed in vitro by darkfield light microscopy. Journal of Cell Biology, 83(1):205–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/205.

Scheer:1976:EDN

[SKT^{+}76] U. Scheer, J. Kartenbeck, M. F. Trendelenburg, J. Stadler, and W. W. Franke. Experimental disintegration of the nuclear envelope. Evidence for pore-connecting fibrils. Journal of Cell Biology, 69(1):1–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/1.

Schlaepfer:1977:ISN

[SL77] W. W. Schlaepfer and R. G. Lynch. Immunofluorescence studies of neurofilaments in the rat and human peripheral and central nervous system. Journal of Cell Biology, 74(1):241–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/1/241.

Sarthy:1978:BSI

[SL78] P. V. Sarthy and D. M. K. Lam. Biochemical studies of isolated glial (muller) cells from the turtle retina. Journal of Cell Biology, 78(3):675–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/3/675.

Sherline:1977:BMP

[SLJ77] P. Sherline, Y. C. Lee, and L. S. Jacobs. Binding of microtubules to pituitary secretory granules and secretory granule membranes. Journal of Cell Biology, 72(2):380–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/2/380.

Suyama:1977:UCI

[SLR77] A. T. Suyama, J. A. Long, and J. Ramachandran. Ultrastructural changes induced by ACTH in normal adrenocortical cells in culture. Journal of Cell Biology, 72(3):757–??, March 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/72/3/757.

Sluder:1976:EMA

[Slu76] G. Sluder. Experimental manipulation of the amount of tubulin available for assembly into the spindle of dividing sea urchin eggs. Journal of Cell Biology, 70(1):75–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/70/1/75.

Sluder:1979:RSM

[Slu79] G. Sluder. Role of spindle microtubules in the control of cell cycle timing. Journal of Cell Biology, 80(3):674–??, March 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/3/674.

Shea:1975:SSG

[SM75a] S. M. Shea and A. B. Morrison. A stereological study of the glomerular filter in the rat. Morphometry of the slit diaphragm and basement membrane. Journal of Cell Biology, 67(2):436–??, November 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
2/436.

Shore:1975:SCS

[SM75b] G. Shore and G. A. Maclachlan. The site of cellulose synthesis. Hormone treatment alters the intracellular location of alkaliinsoluble beta-1,4-glucan (cellulose) synthetase activities. Journal of Cell Biology, 64(3):557–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/3/557.

Snyder:1975:IGM

[SM75c] J. A. Snyder and J. R. McIntosh. Initiation and growth of microtubules from mitotic centers in lysed mammalian cells. Journal of Cell Biology, 67(3):744–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/744.

Swift:1976:DFS

[SM76] J. G. Swift and T. M. Mukherjee. Demonstration of the fuzzy surface coat of rat intestinal microvilli by freeze-etching. Journal of Cell Biology, 69(2):491–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/491.

Sherline:1977:RTM

[SM77] P. Sherline and G. R. Mundy. Role of the tubulin-microtubule system in lymphocyte activation. Journal of Cell Biology, 74 (2):371–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/2/371.

Schlaepfer:1978:CSC

[SM78a] W. W. Schlaepfer and S. Micko. Chemical and structural changes of neurofilaments in transected rat sciatic nerve. Journal of Cell Biology, 78(2):369–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/78/2/369.

Seeds:1978:PMD

[SM78b] N. W. Seeds and R. B. Maccioni. Proteins from morphologically differentiated neuroblastoma cells promote tubulin polymerization. Journal of Cell Biology, 76(2):547–??, February 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/2/547.

Sleytr:1978:FFN

[SM78c] U. B. Sleytr and P. Messner. Freeze-fracturing in normal vacuum reveals ringlike yeast plasmalemma structures. Journal of Cell Biology, 79(1):276–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/1/276.

Schloss:1977:MSB

[SMG77] J. A. Schloss, A. Milsted, and R. D. Goldman. Myosin subfragment binding for the localization of actin-like microfilaments in cultured cells. A light and electron microscope study. Journal of Cell Biology, 74(3):794–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/3/794.

Smith:1977:IOS

[Smi77] G. L. Smith. Increased ouabain-sensitive 86Rubidium uptake after mitogenic stimulation of quiescent chicken embryo fibroblasts with purified multiplication-stimulating activity. Journal of Cell Biology, 73(3):761–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/3/761.

Schmucker:1978:SAH

[SMJ78] D. L. Schmucker, J. S. Mooney, and A. L. Jones. Stereological analysis of hepatic fine structure in the Fischer 344 rat. Influence of sublobular location and animal age. Journal of Cell Biology, 78(2):319–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/2/319.

Sanes:1978:RMF

[SMM78] J. R. Sanes, L. M. Marshall, and U. J. McMahan. Reinnervation of muscle fiber basal lamina after removal of myofibers. Differentiation of regenerating axons at original synaptic sites. Journal of Cell Biology, 78(1):176–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/1/176.

Silva:1975:MSS

[SMPGR75] P. P. Silva, A. Martínez-Palomo, and A. Gonzalez-Robles. Membrane structure and surface coat of Entamoeba histolytica. Topochemistry and dynamics of the cell surface: cap formation and microexudate. Journal of Cell Biology, 64(3):538–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/64/3/538.

Salpeter:1978:REM

[SMS78] M. M. Salpeter, F. A. McHenry, and E. E. Salpeter. Resolution of electron microscope autoradiography. IV. Application to analysis of autoradiographs. Journal of Cell Biology, 76(1):127–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
1/127.

Snell:1976:MCSa

[Sne76a] W. J. Snell. Mating in Chlamydomonas: a system for the study of specific cell adhesion. I. Ultrastructural and electrophoretic analyses of flagellar surface components involved in adhesion. Journal of Cell Biology, 68(1):48–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/1/48.

Snell:1976:MCSb

[Sne76b] W. J. Snell. Mating in Chlamydomonas: a system for the study of specific cell adhesion. II. A radioactive flagella-binding assay for quantitation of adhesion. Journal of Cell Biology, 68(1):70–??, January 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
1/70.

Szego:1977:EIR

[SNP77] C. M. Szego, M. B. Nazareno, and D. D. Porter. Estradiolinduced redistribution of lysosomal proteins in rat preputial gland. Evidence from immunologic probes. Journal of Cell Biology, 73(2):354–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/2/354.

Schofield:1975:RGH

[SO75] J. G. Schofield and L. Orci. Release of growth hormone from ox pituitary slices after pronase treatment. Journal of Cell Biology, 65(1):223–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/1/223.

Shimada:1977:PAF

[SO77] Y. Shimada and T. Obinata. Polarity of actin filaments at the initial stage of myofibril assembly in myogenic cells in vitro. Journal of Cell Biology, 72(3):777–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/3/777.

Sakiyama:1979:CIB

[SOK79] H. Sakiyama, H. Otsu, and S. Kanegasaki. Cellular interaction between fixed and living cells. Transfer of radioactive materials from living cells to fixed cells. Journal of Cell Biology, 81(3):498– ??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/498.

Somlyo:1979:BSS

[Som79] A. V. Somlyo. Bridging structures spanning the junctioning gap at the triad of skeletal muscle. Journal of Cell Biology, 80(3):743–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/743.

Schliwa:1978:MSI

[SOW78] M. Schliwa, M. Osborn, and K. Weber. Microtubule system of isolated fish melanophores as revealed by immunofluorescence microscopy. Journal of Cell Biology, 76(1):229–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/1/229.

Soderstrom:1976:IUC

[SP76] K. O. Söderström and M. Parvinen. Incorporation of (3H)uridine by the chromatoid body during rat spermatogenesis. Journal of Cell Biology, 70(1):239–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/70/1/239.

Sardet:1979:SET

[SPM79] C. Sardet, M. Pisam, and J. Maetz. The surface epithelium of teleostean fish gills. Cellular and junctional adaptations of the chloride cell in relation to salt adaptation. Journal of Cell Biology, 80(1):96–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/96.

Sheetz:1976:BMB

[SPS76] M. P. Sheetz, R. G. Painter, and S. J. Singer. Biological membranes as bilayer couples. III. Compensatory shape changes induced in membranes. Journal of Cell Biology, 70(1):193–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/70/1/193.

Scheele:1978:CFS

[SPT78] G. A. Scheele, G. E. Palade, and A. M. Tartakoff. Cell fractionation studies on the guinea pig pancreas. Redistribution of exocrine proteins during tissue homogenization. Journal of Cell Biology, 78(1):110–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/1/110.

Schaeffer:1978:MRC

[SR78a] S. F. Schaeffer and E. Raviola. Membrane recycling in the cone cell endings of the turtle retina. Journal of Cell Biology, 79(3):802–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
3/802.

Schaffner:1978:GPB

[SR78b] A. Schaffner and R. Rodewald. Glomerular permeability in the bullfrog Rana catesbeiana. Journal of Cell Biology, 79(2):314–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
2/314.

Salpeter:1978:AFE

[SRKM78] M. M. Salpeter, A. W. Rogers, H. Kasprzak, and F. A. McHenry. Acetylcholinesterase in the fast extraocular muscle of the mouse by light and electron microscope autoradiography. Journal of Cell Biology, 78(1):274–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/1/274.

Schengrund:1976:EGS

[SRR76] C. L. Schengrund, A. Rosenberg, and M. A. Repman. Ectoganglioside-sialidase activity of herpes simplex virus-transformed hamster embryo fibroblasts. Journal of Cell Biology, 70(3): 555–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/555.

Strader:1979:DTN

[SRR79] C. B. Strader, J. P. Revel, and M. A. Raftery. Demonstration of the transmembrane nature of the acetylcholine receptor by labeling with anti-receptor antibodies. Journal of Cell Biology, 83 (2):499–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/2/499.

Sale:1976:STC

[SS76a] W. S. Sale and P. Satir. Splayed Tetrahymena cilia. A system for analyzing sliding and axonemal spoke arrangements. Journal of Cell Biology, 71(2):589–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/2/589.

Sheetz:1976:EKE

[SS76b] M. P. Sheetz and S. J. Singer. Equilibrium and kinetic effects of drugs on the shapes of human erythrocytes. Journal of Cell Biology, 70(1):247–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/1/247.

Simionescu:1976:GLMa

[SS76c] N. Simionescu and M. Simionescu. Galloylglucoses of low molecular weight as mordant in electron microscopy. I. Procedure, and evidence for mordanting effect. Journal of Cell Biology, 70(3): 608–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/70/3/608.

Simionescu:1976:GLMb

[SS76d] N. Simionescu and M. Simionescu. Galloylglucoses of low molecular weight as mordant in electron microscopy. II. The moiety and functional groups possibly involved in the mordanting effect. Journal of Cell Biology, 70(3):622–??, September 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/70/3/622.

Sheetz:1977:MAI

[SS77a] M. P. Sheetz and S. J. Singer. On the mechanism of ATP-induced shape changes in human erythrocyte membranes. I. The role of the spectrin complex. Journal of Cell Biology, 73(3):638–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/3/638.

Simionescu:1977:OCJ

[SS77b] M. Simionescu and N. Simionescu. Organization of cell junctions in the peritoneal mesothelium. Journal of Cell Biology, 74(1):98– ??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/74/
1/98.

Savion:1978:MCR

[SS78a] N. Savion and Z. Selinger. Morphological changes in rat pancreatic slices associated with inhibition of enzyme secretion by high concentrations of secretagogues. Journal of Cell Biology, 76 (2):467–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/76/2/467.

Shaver:1978:OBR

[SS78b] J. L. Shaver and C. Stirling. Ouabain binding to renal tubules of the rabbit. Journal of Cell Biology, 76(2):278–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/278.

Sherline:1978:HMW

[SS78c] P. Sherline and K. Schiavone. High molecular weight MAPs are part of the mitotic spindle. Journal of Cell Biology, 77(1):??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/77/1/R9.

Stoner:1978:MII

[SS78d] C. D. Stoner and H. D. Sirak. Magnesium-induced inner membrane aggregation in heart mitochondria: prevention and reversal by carboxyatractyloside and bongkrekic acid. Journal of Cell Biology, 77(2):417–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/417.

Sugi:1978:UPS

[SS78e] H. Sugi and S. Suzuki. Ultrastructural and physiological studies on the longitudinal body wall muscle of Dolabella auricularia. I. Mechanical response and ultrastructure. Journal of Cell Biology, 79(2):454–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/454.

Suzuki:1978:UPS

[SS78f] S. Suzuki and H. Sugi. Ultrastructural and physiological studies on the longitudinal body wall muscle of Dolabella auricularia. II. Localization of intracellular calcium and its translocation during mechanical activity. Journal of Cell Biology, 79(2):467–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
2/467.

Saling:1979:MGI

[SS79a] P. M. Saling and B. T. Storey. Mouse gamete interactions during fertilization in vitro. Chlortetracycline as a fluorescent probe for the mouse sperm acrosome reaction. Journal of Cell Biology, 83 (3):544–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/3/544.

Scales:1979:MSS

[SS79b] D. J. Scales and R. A. Sabbadini. Microsomal T system: a stereological analysis of purified microsomes derived from normal and dystrophic skeletal muscle. Journal of Cell Biology, 83(1):33–??, October 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
1/33.

Selden:1979:CBI

[SS79c] S. C. Selden and S. M. Schwartz. Cytochalasin B inhibition of endothelial proliferation at wound edges in vitro. Journal of Cell Biology, 81(2):348–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/2/348.

Spudich:1979:ATT

[SS79d] A. Spudich and J. A. Spudich. Actin in triton-treated cortical preparations of unfertilized and fertilized sea urchin eggs. Journal of Cell Biology, 82(1):212–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/1/212.

Siu:1976:CCNa

[SSC76a] C. Siu, H. Swift, and K. Chiang. Characterization of cytoplasmic and nuclear genomes in the colorless alga Polytoma. I. Ultrastructural analysis of organelles. Journal of Cell Biology, 69(2):352–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/69/
2/352.

Siu:1976:CCNb

[SSC76b] C. Siu, H. Swift, and K. Chiang. Characterization of cytoplasmic and nuclear genomes in the colorless alga Polytoma. II. General characterization of organelle nucleic acids. Journal of Cell Biology, 69(2):371–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/371.

Setterfield:1978:SIN

[SSD^{+}78] G. Setterfield, R. Sheinin, I. Dardick, G. Kiss, and M. Dubsky. Structure of interphase nuclei in relation to the cell cycle.

Chromatin organization in mouse L cells temperature-sensitive for DNA replication. Journal of Cell Biology, 77(1):246–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/1/246.

Silverberg:1976:ICC

[SSF76] B. A. Silverberg, P. M. Stokes, and L. B. Ferstenberg. Intranuclear complexes in a copper-tolerant green alga. Journal of Cell Biology, 69(1):210–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/210.

Slot:1979:EFF

[SSG79] J. W. Slot, G. J. Strous, and J. J. Geuze. Effect of fasting and feeding on synthesis and intracellular transport of proteins in the frog exocrine pancreas. Journal of Cell Biology, 80(3):708–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/708.

Simionescu:1975:SDC

[SSP75a] M. Simionescu, N. Simionescu, and G. E. Palade. Segmental differentiations of cell junctions in the vascular endothelium. The microvasculature. Journal of Cell Biology, 67(3):863–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/67/3/863.

Simionescu:1975:PMC

[SSP75b] N. Simionescu, M. Siminoescu, and G. E. Palade. Permeability of muscle capillaries to small heme-peptides. Evidence for the existence of patent transendothelial channels. Journal of Cell Biology, 64(3):586–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/586.

Simionescu:1976:SDC

[SSP76] M. Simionescu, N. Simionescu, and G. E. Palade. Segmental differentiations of cell junctions in the vascular endothelium. Arteries and veins. Journal of Cell Biology, 68(3):705–??, March 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/68/3/705.

Simionescu:1978:OJE

[SSP78] N. Simionescu, M. Simionescu, and G. E. Palade. Open junctions in the endothelium of the postcapillary venules of the diaphragm. Journal of Cell Biology, 79(1):27–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/1/27.

Smith-Sonneborn:1976:LEC

[SSR76] J. Smith-Sonneborn and S. R. Rodermel. Loss of endocytic capacity in aging Paramecium. The importance of cytoplasmic organelles. Journal of Cell Biology, 71(2):575–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/2/575.

Somlyo:1977:EDS

[SSS77] A. V. Somlyo, H. Shuman, and A. P. Somlyo. Elemental distribution in striated muscle and the effects of hypertonicity: Electron probe analysis of cryo sections. Journal of Cell Biology, 74 (3):828–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/828.

Somlyo:1979:EPA

[SSS79] A. P. Somlyo, A. V. Somlyo, and H. Shuman. Electron probe analysis of vascular smooth muscle. Composition of mitochondria, nuclei, and cytoplasm. Journal of Cell Biology, 81(2):316–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/316.

Staubli:1977:PRH

[SSSW77] W. Stäubli, W. Schweizer, J. Suter, and E. R. Weibel. The proliferative response of hepatic peroxidomes of neonatal rats to treatment with SU-13 437 (nafenopin). Journal of Cell Biology, 74 (3):665–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/665.

Schwab:1979:SRT

[SST79] M. E. Schwab, K. Suda, and H. Thoenen. Selective retrograde transsynaptic transfer of a protein, tetanus toxin, subsequent to its retrograde axonal transport. Journal of Cell Biology, 82(3): 798–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/3/798.

Sasaki:1976:SPSa

[ST76a] S. Sasaki and Y. Tashiro. Studies on the posterior silk gland of the silkworm Bombyx mori. V. Electron microscope localization of fibroin in the posterior silk gland at the later stage of the fifth instar. Journal of Cell Biology, 70(3):648–??, September 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/70/3/648.

Sasaki:1976:SPSb

[ST76b] S. Sasaki and Y. Tashiro. Studies on the posterior silk gland of the silkworm Bombyx mori. VI. Distribution of microtubules in the posterior silk gland cells. Journal of Cell Biology, 71(2):565–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/71/
2/565.

Shore:1977:TFRa

[ST77a] G. C. Shore and J. R. Tata. Two fractions of rough endoplasmic reticulum from rat liver. I. Recovery of rapidly sedimenting endoplasmic reticulum in association with mitochondria. Journal of Cell Biology, 72(3):714–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/3/714.

Shore:1977:TFRb

[ST77b] G. C. Shore and J. R. Tata. Two fractions of rough endoplasmic reticulum from rat liver. II. Cytoplasmic messenger RNA’s which code for albumin and mitochondrial proteins are distributed differently between the two fractions. Journal of Cell Biology, 72(3):726–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/3/726.

Schatten:1978:NSC

[ST78a] G. Schatten and M. Thoman. Nuclear surface complex as observed with the high resolution scanning electron microscope. Visualization of the membrane surfaces of the neclear envelope and the nuclear cortex from Xenopus laevis oocytes. Journal of Cell Biology, 77(2):517–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/77/2/517.

Schwab:1978:SBU

[ST78b] M. E. Schwab and H. Thoenen. Selective binding, uptake, and retrograde transport of tetanus toxin by nerve terminals in the rat iris. An electron microscope study using colloidal gold as a tracer. Journal of Cell Biology, 77(1):1–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/1/1.

Staehelin:1976:RPM

[Sta76] L. A. Staehelin. Reversible particle movements associated with unstacking and restacking of chloroplast membranes in vitro. Journal of Cell Biology, 71(1):136–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/1/136.

Schneider:1979:FPMb

[STdDT79a] Y. Schneider, P. Tulkens, C. de Duve, and A. Trouet. Fate of plasma membrane during endocytosis. II. Evidence for recycling (shuttle) of plasma membrane constituents. Journal of Cell Biology, 82(2):466–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/82/2/466.

Schneider:1979:FPMa

[STdDT79b] Y. J. Schneider, P. Tulkens, C. de Duve, and A. Trouet. Fate of plasma membrane during endocytosis. I. Uptake and processing of anti-plasma membrane and control immunoglobulins by cultured fibroblasts. Journal of Cell Biology, 82(2):449–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/2/449.

Stephens:1975:BAM

[Ste75] R. E. Stephens. The basal apparatus. Mass isolation from the molluscan ciliated gill epithelium and a preliminary characterization of striated rootlets. Journal of Cell Biology, 64(2):408–??, February 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/64/
2/408.

Scheer:1975:EAD

[STF75] U. Scheer, F. Trendelenburg, and W. W. Franke. Effects of actinomycin D on the association of newly formed ribonucleoproteins with the cistrons of ribosomal RNA in Triturus oocytes. Journal of Cell Biology, 65(1):163–??, April 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/65/1/163.

Scheer:1976:RTG

[STF76] U. Scheer, M. F. Trendelenburg, and W. W. Franke. Regulation of transcription of genes of ribosomal rna during amphibian oogenesis. A biochemical and morphological study. Journal of Cell Biology, 69(2):465–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/465.

Storrie:1975:ADA

[Sto75] B. Storrie. Antagonism by dibutyryl adenosine cyclic 3’,5’- monophosphate and testololactone of concanavalin A capping. Journal of Cell Biology, 66(2):392–??, August 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/2/392.

Shotton:1978:ADS

[STWB78] D. Shotton, K. Thompson, L. Wofsy, and D. Branton. Appearance and distribution of surface proteins of the human erythrocyte membrane. An electron microscope and immunochemical labeling study. Journal of Cell Biology, 76(2):512–??, February 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/76/2/512.

Sakharov:1976:NDG

[SV76] V. N. Sakharov and N. Voronkova. Nucleolus degradation and growth induced by uv-microbeam irradiation of interphase cells grown in culture. Journal of Cell Biology, 71(3):963–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/3/963.

Sanderson:1977:MFT

[SV77] R. J. Sanderson and A. E. Vatter. A mode of formation of tubular myelin from lamellar bodies in the lung. Journal of Cell Biology, 74(3):1027–??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/74/3/1027.

Svoboda:1978:URR

[Svo78] D. J. Svoboda. Unusual responses of rat hepatic and renal peroxisomes to RMI 14, 514, a new hypolipidemic agent. Journal of Cell Biology, 78(3):810–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/3/810.

Steck:1979:GCC

[SVW79] P. A. Steck, P. G. Voss, and J. L. Wang. Growth control in cultured 3T3 fibroblasts. Assays of cell proliferation and demonstration of a growth inhibitory activity. Journal of Cell Biology, 83 (3):562–??, December 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/83/3/562.

Sweeney:1976:FFS

[Swe76] B. M. Sweeney. Freeze-fracture studies of the thecal membranes of Gonyaulax polyedra: circadian changes in the particles of one membrane face. Journal of Cell Biology, 68(3):451–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/68/3/451.

Saito:1978:MAE

[SWF78] A. Saito, C. T. Wang, and S. Fleischer. Membrane asymmetry and enhanced ultrastructural detail of sarcoplasmic reticulum revealed with use of tannic acid. Journal of Cell Biology, 79(3):601–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/79/
3/601.

Stenman:1977:CDM

[SWV77] S. Stenman, J. Wartiovaara, and A. Vaheri. Changes in the distribution of a major fibroblast protein, fibronectin, during mitosis and interphase. Journal of Cell Biology, 74(2):453–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/74/2/453.

Takeichi:1977:FCB

[Tak77] M. Takeichi. Functional correlation between cell adhesive properties and some cell surface proteins. Journal of Cell Biology, 75 (2):464–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/464.

Taljedal:1978:CFC

[Täl78] I. B. Täljedal. Chlorotetracycline as a fluorescent Ca^{2+} probe in pancreatic islet cells. Journal of Cell Biology, 76(3):652–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
3/652.

Tamm:1978:LMS

[Tam78] S. L. Tamm. Laser microbeam study of a rotary motor in termite flagellates. Evidence that the axostyle complex generates torque. Journal of Cell Biology, 78(1):76–??, July 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/78/1/76.

Tamm:1979:MMF

[Tam79] S. L. Tamm. Membrane movements and fluidity during rotational motility of a termite flagellate. A freeze-fracture study. Journal of Cell Biology, 80(1):141–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/1/141.

Tojo:1978:FBP

[TBZW78] S. Tojo, T. Betchaku, V. J. Ziccardi, and G. R. Wyatt. Fat body protein granules and storage proteins in the silkmoth, Hyalophora cecropia. Journal of Cell Biology, 78(3):823–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/78/3/823.

Tuan:1975:IIS

[TC75] R. S. Tuan and K. P. Chang. Isolation of intracellular symbiotes by immune lysis of flagellate protozoa and characterization of their DNA. Journal of Cell Biology, 65(2):309–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/65/2/309.

Tennekoon:1977:MON

[TCPM77] G. I. Tennekoon, S. R. Cohen, D. L. Price, and G. M. McKhann. Myelinogenesis in optic nerve. A morphological, autoradiographic, and biochemical analysis. Journal of Cell Biology, 72(3):604–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/72/
3/604.

Tilney:1979:MEA

[TCT79] L. G. Tilney, J. G. Clain, and M. S. Tilney. Membrane events in the acrosomal reaction of Limulus sperm. Membrane fusion, filament-membrane particle attachment, and the source and formation of new membrane surface. Journal of Cell Biology, 81 (1):229–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/1/229.

Tilney:1975:AEG

[TD75] L. G. Tilney and P. Detmers. Actin in erythrocyte ghosts and its association with spectrin. Evidence for a nonfilamentous form of these two molecules in situ. Journal of Cell Biology, 66(3): 508–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/508.

Theodosis:1976:SRU

[TDHO76] D. T. Theodosis, J. Dreifuss, M. C. Harris, and L. Orci. Secretionrelated uptake of horseradish peroxidase in neurohypophysial axons. Journal of Cell Biology, 70(2):294–??, August 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/70/2/294.

Theodosis:1978:FFS

[TDO78] D. T. Theodosis, J. J. Dreifuss, and L. Orci. A freeze-fracture study of membrane events during neurohypophysial secretion. Journal of Cell Biology, 78(2):542–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/78/2/542.

Tyrrell:1979:TCC

[TE79] D. L. Tyrrell and A. Ehrnst. Transmembrane communication in cells chronically infected with measles virus. Journal of Cell Biology, 81(2):396–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/2/396.

Tamm:1976:ADR

[THC76] I. Tamm, R. Hand, and L. A. Caliguiri. Action of dichlorobenzimidazole riboside on RNA synthesis in L-929 and HeLa cells. Journal of Cell Biology, 69(2):229–??, May 1976. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/69/2/229.

Teichberg:1975:CTS

[THCP75] S. Teichberg, E. Holtzman, S. M. Crain, and E. R. Peterson. Circulation and turnover of synaptic vesicle membrane in cultured fetal mammalian spinal cord neurons. Journal of Cell Biology, 67(1):215–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/67/1/215.

Tilney:1975:AFA

[Til75] L. G. Tilney. Actin filaments in the acrosomal reaction of Limulus sperm. Motion generated by alterations in the packing of the filaments. Journal of Cell Biology, 64(2):289–??, February 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/2/289.

Tilney:1976:PAIa

[Til76a] L. G. Tilney. The polymerization of actin: II. how nonfilamentous actin becomes nonrandomly distributed in sperm: evidence for the association of this actin with membranes. Journal of Cell Biology, 69(1):51–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/69/1/51.

Tilney:1976:PAIb

[Til76b] L. G. Tilney. The polymerization of actin. III. Aggregates of nonfilamentous actin and its associated proteins: a storage form of actin. Journal of Cell Biology, 69(1):73–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/1/73.

Tilney:1978:PAV

[Til78] L. G. Tilney. Polymerization of actin. V. A new organelle, the actomere, that initates the assembly of actin filaments in Thyone sperm. Journal of Cell Biology, 77(2):551–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/551.

Tait:1977:CMC

[TK77] A. Tait and J. K. Knowles. Characterization of mitochondrial and cytoplasmic ribosomes from Paramecium aurelia. Journal of Cell Biology, 73(1):139–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/1/139.

Tilney:1979:PAV

[TK79] L. G. Tilney and N. Kallenbach. Polymerization of actin. VI. The polarity of the actin filaments in the acrosomal process and how it might be determined. Journal of Cell Biology, 81(3):608–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
3/608.

Thoma:1979:IHH

[TKK79] F. Thoma, T. Koller, and A. Klug. Involvement of histone H1 in the organization of the nucleosome and of the salt-dependent superstructures of chromatin. Journal of Cell Biology, 83(2):403–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
2/403.

Tilney:1978:PAI

[TKST78] L. G. Tilney, D. P. Kiehart, C. Sardet, and M. Tilney. Polymerization of actin. IV. Role of Ca++ and H+ in the assembly of actin and in membrane fusion in the acrosomal reaction of echinoderm sperm. Journal of Cell Biology, 77(2):536–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/536.

Therien:1976:ISJ

[TM76a] H. M. Therien and W. E. Mushynski. Isolation of synaptic junctional complexes of high structural integrity from rat brain. Journal of Cell Biology, 71(3):807–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/3/807.

Tilney:1976:AFM

[TM76b] L. G. Tilney and M. S. Mooseker. Actin filament-membrane attachment: are membrane particles involved? Journal of Cell Biology, 71(2):402–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/71/2/402.

Tucker:1976:MCI

[TM76c] J. B. Tucker and M. Meats. Microtubules and control of insect egg shape. Journal of Cell Biology, 71(1):207–??, October 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/71/1/207.

Tkacz:1979:SCY

[TM79] J. S. Tkacz and V. L. MacKay. Sexual conjugation in yeast. Cell surface changes in response to the action of mating hormones. Journal of Cell Biology, 80(2):326–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/80/2/326.

Tanphaichitr:1976:RBC

[TMGC76] N. Tanphaichitr, K. C. Moore, D. K. Granner, and R. Chalkley. Relationship between chromosome condensation and metaphase lysine-rich histone phosphorylation. Journal of Cell Biology, 69(1):43–??, April 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/1/43.

Toylor:1976:QSP

[Toy76] D. L. Toylor. Quantitative studies on the polarization optical properties of striated muscle. I. Birefringence changes of rabbit psoas muscle in the transition from rigor to relaxed state. Journal of Cell Biology, 68(3):497–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/68/3/497.

Tandler:1976:FEM

[TP76] B. Tandler and J. H. Poulsen. Fusion of the envelope of mucous droplets with the luminal plasma membrane in acinar cells of the cat submandibular gland. Journal of Cell Biology, 68(3):775–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/68/
3/775.

Tippit:1977:MPD

[TPH77] D. H. Tippit and J. D. Pickett-Heaps. Mitosis in the pinnate diatom surirella ovalis. Journal of Cell Biology, 73(3):705–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/73/3/705.

Telzer:1979:CCD

[TR79a] B. R. Telzer and J. L. Rosenbaum. Cell cycle-dependent, in vitro assembly of microtubules onto pericentriolar material of HeLa cells. Journal of Cell Biology, 81(3):484–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/81/3/484.

Teng:1979:FCE

[TR79b] M. H. Teng and D. B. Rifkin. Fibronectin from chicken embryo fibroblasts contains covalently bound phosphate. Journal of Cell Biology, 80(3):784–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/80/3/784.

Taylor:1976:CBA

[TRH76] D. L. Taylor, J. A. Rhodes, and S. A. Hammond. The contractile basis of ameboid movement. II. Structure and contractility of motile extracts and plasmalemma-ectoplasm ghosts. Journal of Cell Biology, 70(1):123–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/70/1/123.

Tokuyasu:1976:IPI

[TS76] K. T. Tokuyasu and S. J. Singer. Improved procedures for immunoferritin labeling of ultrathin frozen sections. Journal of Cell Biology, 71(3):894–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/3/894.

Trenkner:1977:HMC

[TS77] E. Trenkner and R. L. Sidman. Histogenesis of mouse cerebellum in microwell cultures. Cell reaggregation and migration, fiber and synapse formation. Journal of Cell Biology, 75(3):915–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/915.

Tuan:1978:CBPa

[TSC78a] R. S. Tuan, W. A. Scott, and Z. A. Cohn. Calcium-binding protein of the chick chorioallantoic membrane. I. Immunohistochemical localization. Journal of Cell Biology, 77(3):743–??, June 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/3/743.

Tuan:1978:CBPb

[TSC78b] R. S. Tuan, W. A. Scott, and Z. A. Cohn. Calcium-binding protein of the chick chorioallantoic membrane. II. Vitamin K-dependent expression. Journal of Cell Biology, 77(3):752–??, June 1978. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/77/3/752.

Tamm:1975:RCO

[TSD75] S. L. Tamm, T. M. Sonneborn, and R. V. Dippell. The role of cortical orientation in the control of the direction of ciliary beat in Paramecium. Journal of Cell Biology, 64(1):98–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/64/1/98.

Tippit:1978:ADS

[TSPH78] D. H. Tippit, D. Schulz, and J. D. Pickett-Heaps. Analysis of the distribution of spindle microtubules in the diatom Fragilaria. Journal of Cell Biology, 79(3):737–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/79/3/737.

Tokuyasu:1979:DRM

[TSS79] K. T. Tokuyasu, R. Schekman, and S. J. Singer. Domains of receptor mobility and endocytosis in the membranes of neonatal human erythrocytes in the membranes of neonatal human erythrocytes and reticulocytes are deficient in spectrin. Journal of Cell Biology, 80(2):481–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/2/481.

Tartakoff:1979:PCI

[TV79] A. Tartakoff and P. Vassalli. Plasma cell immunoglobulin M molecules. Their biosynthesis, assembly, and intracellular transport. Journal of Cell Biology, 83(2):284–??, November 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/83/2/284.

Tartakoff:1978:CSI

[TVD78] A. Tartakoff, P. Vassalli, and M. Détraz. Comparative studies of intracellular transport of secretory proteins. Journal of Cell Biology, 79(3):694–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/79/3/694.

Tice:1975:CTJ

[TWC75] L. W. Tice, S. H. Wollman, and R. C. Carter. Changes in tight junctions of thyroid epithelium with changes in thyroid activity. Journal of Cell Biology, 66(3):657–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/66/3/657.

Tarlow:1977:LSS

[TWR^{+}77] D. M. Tarlow, P. A. Watkins, R. E. Reed, R. S. Miller, E. E. Zwergel, and M. D. Lane. Lipogenesis and the synthesis and secretion of very low density lipoprotein by avian liver cells in nonproliferating monolayer culture. Hormonal effects. Journal of Cell Biology, 73(2):332–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/2/332.

Tupper:1977:CCD

[TZ77] J. T. Tupper and F. Zorgniotti. Calcium content and distribution as a function of growth and transformation in the mouse 3T3 cell. Journal of Cell Biology, 75(1):12–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/1/12.

Ueda:1979:SDC

[UGB^{+}79] T. Ueda, P. Greengard, K. Berzins, R. S. Cohen, F. Blomberg, D. J. Grab, and P. Siekevitz. Subcellular distribution in cerebral cortex of two proteins phosphorylated by a cAMP-dependent protein kinase. Journal of Cell Biology, 83(2):308–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/83/2/308.

Ueda:1976:CBM

[UIOO76] M. J. Ueda, T. Ito, T. S. Okada, and S. I. Ohnishi. A correlation between membrane fluidity and the critical temperature for cell adhesion. Journal of Cell Biology, 71(2):670–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/2/670.

Uchida:1979:RLV

[UKY^{+}79] T. Uchida, J. Kim, M. Yamaizumi, Y. Miyake, and Y. Okada. Reconstitution of lipid vesicles associated with HVJ (Sendai virus) sikes. Purification and some properties of vesicles containing nontoxic fragment A of diphtheria toxin. Journal of Cell Biology, 80(1):10–??, January 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/80/1/10.

Underhill:1979:BHS

[UT79] C. B. Underhill and B. P. Toole. Binding of hyaluronate to the surface of cultured cells. Journal of Cell Biology, 82(2):475–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
2/475.

Vorhaben:1977:SLF

[VC77] J. E. Vorhaben and J. W. Campbell. Submitochondrial localization and function of enzymes of glutamine metabolism in avian liver. Journal of Cell Biology, 73(2):300–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/73/2/300.

Vogel:1976:UAR

[VD76] Z. Vogel and M. P. Daniels. Ultrastructure of acetylcholine receptor clusters on cultured muscle fibers. Journal of Cell Biology, 69(2):501–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/69/2/501.

vanDeurs:1979:TJC

[vDK79] B. van Deurs and J. K. Koehler. Tight junctions in the choroid plexus epithelium. A freeze-fracture study including complementary replicas. Journal of Cell Biology, 80(3):662–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/3/662.

Vermorken:1978:LDC

[VHvdVB78] A. J. Vermorken, J. M. Hilderink, W. J. van de Ven, and H. Bloemendal. Lens differentiation. Crystallin synthesis in isolated epithelia from calf lenses. Journal of Cell Biology, 76(1):175–??, January 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
1/175.

Vlodavsky:1979:VEC

[VJGG79] I. Vlodavsky, L. K. Johnson, G. Greenburg, and D. Gospodarowicz. Vascular endothelial cells maintained in the absence of fibroblast growth factor undergo structural and functional alterations that are incompatible with their in vivo differentiated properties. Journal of Cell Biology, 83(2):468–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/83/2/468.

Viamontes:1977:CSC

[VK77] G. I. Viamontes and D. L. Kirk. Cell shape changes and the mechanism of inversion in Volvox. Journal of Cell Biology, 75 (3):719–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/3/719.

Verma:1978:ICM

[VKZB78] D. P. Verma, V. Kazazian, V. Zogbi, and A. K. Bal. Isolation and characterization of the membrane envelope enclosing the bacteroids in soybean root nodules. Journal of Cell Biology, 78(3): 919–??, September 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/3/919.

Volsky:1978:RCV

[VL78] D. J. Volsky and A. Loyter. Role of Ca++ in virus-induced membrane fusion. Ca++ accumulation and ultrastructural changes induced by Sendai virus in chicken erythrocytes. Journal of Cell Biology, 78(2):465–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/2/465.

Vorbrodt:1979:RAP

[VMR79] A. Vorbrodt, P. Meo, and G. Rovera. Regulation of acid phosphatase activity in human promyelocytic leukemic cells induced to differentiate in culture. Journal of Cell Biology, 83(2):300–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
2/300.

vanOs:1977:DBS

[vOMW77] C. H. van Os, A. K. Mircheff, and E. M. Wright. Distribution of bicarbonate-stimulated ATPase in rat intestinal epithelium. Journal of Cell Biology, 73(1):257–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/1/257.

Vonderhaar:1979:LSA

[Von79] B. K. Vonderhaar. Lactose synthetase activity in mouse mammary glands is controlled by thyroid hormones. Journal of Cell Biology, 82(3):675–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/82/3/675.

Vial:1975:SMD

[VP75] J. Vial and K. R. Porter. Scanning microscopy of dissociated tissue cells. Journal of Cell Biology, 67(2):345–??, November 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/2/345.

Vujicic:1976:OAR

[VRN76] R. Vujicić, L. J. Radojević, and M. Nesković. Orderly arrangement of ribosomes in the embryogenic callus tissue of Corylus avellana L. Journal of Cell Biology, 69(3):686–??, June 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/3/686.

Villegas:1976:SCS

[VV76] G. M. Villegas and J. Villegas. Structural complexes in the squid giant axon membrane sensitive to ionic concentrations and cardiac glycosides. Journal of Cell Biology, 69(1):19–??, April 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/69/1/19.

VonDerMark:1977:IBS

[VV77] K. Von Der Mark and H. Von Der Mark. Immunological and biochemical studies of collagen type transition during in vitro chondrogenesis of chick limb mesodermal cells. Journal of Cell Biology, 73(3):736–??, June 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/73/3/736.

vanVeen:1976:ALI

[vVRN76] J. van Veen, R. M. Roberts, and K. D. Noonan. An analysis of lectin-initiated cell agglutination in a series of CHO subclones which respond morphologically to growth in dibutyryl cyclic AMP. Journal of Cell Biology, 70(1):204–??, July 1976. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/70/1/204.

vanVliet:1976:ICS

[vVSL^{+}76] H. H. van Vliet, F. Spies, W. A. Linnemans, A. Klepke, J. A. Op den Kamp, and L. L. van Deenen. Isolation and characterization of subcellular membranes of Entamoeba invadens. Journal of Cell Biology, 71(2):357–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/71/2/357.

Wey:1978:BIL

[WAC78] C. L. Wey, P. L. Ahl, and R. A. Cone. Bacteriorhodopsin induces a light-scattering change in Halobacterium halobium. Journal of Cell Biology, 79(3):657–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/79/3/657.

Warner:1978:CIA

[War78] F. D. Warner. Cation-induced attachment of ciliary dynein crossbridges. Journal of Cell Biology, 77(3):??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/R19.

Warner:1979:DNF

[War79] J. R. Warner. Distribution of newly formed ribosomal proteins in HeLa cell fractions. Journal of Cell Biology, 80(3):767–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/80/3/767.

Warren:1978:MCM

[WB78a] R. H. Warren and B. Brunside. Microtubules in cone myoid elongation in the teleost retina. Journal of Cell Biology, 78(1):247–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/78/
1/247.

Willison:1978:CWS

[WB78b] J. H. Willison and R. M. Brown. Cell wall structure and deposition in Glaucocystis. Journal of Cell Biology, 77(1):103–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/77/1/103.

Wallace:1975:CTC

[WBE75] D. C. Wallace, C. L. Bunn, and J. M. Eisenstadt. Cytoplasmic transfer of chloramphenicol resistance in human tissue culture cells. Journal of Cell Biology, 67(1):174–??, October 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/67/1/174.

Widdicombe:1979:LPT

[WBY79] J. H. Widdicombe, C. B. Basbaum, and J. Y. Yee. Localization of Na pumps in the tracheal epithelium of the dog. Journal of Cell Biology, 82(2):380–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/82/2/380.

Warner:1975:EPA

[WC75] R. R. Warner and J. R. Coleman. Electron probe analysis of calcium transport by small intestine. Journal of Cell Biology, 64(1):54–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/1/54.

Wilson:1977:TPC

[WC77] R. Wilson and K. S. Chiang. Temporal programming of chloroplast and cytoplasmic ribosomal RNA transcription in the synchronous cell cycle of Chlamydomonas reinhardtii. Journal of Cell Biology, 72(2):470–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/2/470.

Wang:1979:IMN

[WCC79] E. Wang, R. K. Cross, and P. W. Choppin. Involvement of microtubules and 10-nm filaments in the movement and positioning of nuclei in syncytia. Journal of Cell Biology, 83(2):320–??, November 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/83/
2/320.

Wattiaux:1978:IRL

[WCRdD78] R. Wattiaux, S. Wattiaux-De Coninck, M. F. Ronveaux-dupal, and F. Dubois. Isolation of rat liver lysosomes by isopycnic centrifugation in a metrizamide gradient. Journal of Cell Biology, 78(2):349–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/78/2/349.

Wieslander:1977:DBR

[WD77] L. Wieslander and B. Daneholt. Demonstration of Balbiani ring RNA sequences in polysomes. Journal of Cell Biology, 73(1):260– ??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
1/260.

Wylie:1979:SFC

[WDB79] D. E. Wylie, C. H. Damsky, and C. A. Buck. Studies on the function of cell surface glycoproteins. I. Use of antisera to surface membranes in the identification of membrane components relevant to cell-substrate adhesion. Journal of Cell Biology, 80(2):385–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
2/385.

Welsh:1979:TCE

[WDBM79] M. J. Welsh, J. R. Dedman, B. R. Brinkley, and A. R. Means. Tubulin and calmodulin. Effects of microtubule and microfilament inhibitors on localization in the mitotic apparatus. Journal of Cell Biology, 81(3):624–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/81/3/624.

Wanson:1975:ICP

[WDM^{+}75] J. C. Wanson, P. Drochmans, C. May, W. Penasse, and A. Popowski. Isolation of centrolobular and perilobular hepatocytes after phenobarbital treatment. Journal of Cell Biology, 66(1):23–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/1/23.

Wanson:1977:ARH

[WDMR77] J. C. Wanson, P. Drochmans, R. Mosselmans, and M. F. Ronveaux. Adult rat hepatocytes in primary monolayer culture. Ultrastructural characteristics of intercellular contacts and cell membrane differentiations. Journal of Cell Biology, 74(3):858– ??, September 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/3/858.

Weidner:1976:MSI

[Wei76a] E. Weidner. The microsporidian spore invasion tube. The ultrastructure, isolation, and characterization of the protein comprising the tube. Journal of Cell Biology, 71(1):23–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/1/23.

Weihing:1976:CBI

[Wei76b] R. R. Weihing. Cytochalasin B inhibits actin-related gelation of HeLa cell extracts. Journal of Cell Biology, 71(1):303–??, October 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/1/303.

Weihing:1977:EMH

[Wei77a] R. R. Weihing. Effects of myosin and heavy meromyosin on actinrelated gelation of HeLa cell extracts. Journal of Cell Biology, 75(1):95–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/1/95.

Weiss:1977:SSM

[Wei77b] R. L. Weiss. Site-specific membrane particle arrays in magnesiumdepleted Escherichia coli. Journal of Cell Biology, 73(2):505–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/73/
2/505.

Wallace:1975:SMC

[WF75] R. B. Wallace and K. B. Freeman. Selection of mammalian cells resistant to a chloramphenicol analog. Journal of Cell Biology, 65(2):492–??, May 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/65/2/492.

Wolstenholme:1976:PMC

[WF76] D. R. Wolstenholme and C. M. Fauron. A partial map of the circular mitochondrial genome of Drosophila melanogaster. Location of EcoRI-sensitive sites and the adenine-thymine-rich region. Journal of Cell Biology, 71(2):434–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/71/2/434.

Willinger:1979:FSPa

[WF79] M. Willinger and F. R. Frankel. Fate of surface proteins of rabbit polymorphonuclear leukocytes during phagocytosis. I. Identification of surface proteins. Journal of Cell Biology, 82(1):32–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/82/1/32.

Wang:1978:FCF

[WG78] E. Wang and R. D. Goldman. Functions of cytoplasmic fibers in intracellular movements in BHK-21 cells. Journal of Cell Biology, 79(3):708–??, December 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/3/708.

Wiseman:1977:NMA

[WGB77] A. Wiseman, N. W. Gillham, and J. E. Boynton. Nuclear mutations affecting mitochondrial structure and function in Chlamydomonas. Journal of Cell Biology, 73(1):56–??, April 1977. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/73/1/56.

Wunderlich:1978:EAF

[WGB78] F. Wunderlich, G. Giese, and C. Bucherer. Expansion and apparent fluidity decrease of nuclear membranes induced by low Ca/Mg. Modulation of nuclear membrane lipid fluidity by the membraneassociated nuclear matrix proteins? Journal of Cell Biology, 79 (2):479–??, November 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/2/479.

Whitlock:1976:VAH

[WGC76] J. P. Whitlock, H. V. Gelboin, and H. G. Coon. Variation in aryl hydrocarbon (benzo(a)pyrene) hydroxylase activity in heteroploid and predominantly diploid rat liver cells in culture. Journal of Cell Biology, 70(1):217–??, July 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/70/1/217.

Wang:1975:ICM

[WGE75] J. L. Wang, G. R. Gunther, and G. M. Edelman. Inhibition by colchicine of the mitogenic stimulation of lymphocytes prior to the S phase. Journal of Cell Biology, 66(1):128–??, July 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/66/1/128.

Willinger:1979:FSPb

[WGF79] M. Willinger, N. Gonatas, and F. R. Frankel. Fate of surface proteins of rabbit polymorphonuclear leukocytes during phagocytosis. II. Internalization of proteins. Journal of Cell Biology, 82(1):45–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/82/1/45.

Weiss:1977:MDS

[WGG77a] R. L. Weiss, D. A. Goodenough, and U. W. Goodenough. Membrane differentiations at sites specialized for cell fusion. Journal of Cell Biology, 72(1):144–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/72/1/144.

Weiss:1977:MPA

[WGG77b] R. L. Weiss, D. A. Goodenough, and U. W. Goodenough. Membrane particle arrays associated with the basal body and with contractile vacuole secretion in Chlamydomonas. Journal of Cell Biology, 72(1):133–??, January 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/72/1/133.

Wunderlich:1977:RCN

[WH77] F. Wunderlich and G. Herlan. Reversibly contractile nuclear matrix. Its isolation, structure, and composition. Journal of Cell Biology, 73(2):271–??, May 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/73/2/271.

Wiche:1979:MPP

[WHC79] G. Wiche, L. S. Honig, and R. D. Cole. Microtubule protein preparations from C6 glial cells and their spontaneous polymer formation. Journal of Cell Biology, 80(3):553–??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/80/3/553.

Willard:1977:ITI

[Wil77a] M. Willard. The identification of two intra-axonally transported polypeptides resembling myosin in some respects in the rabbit visual system. Journal of Cell Biology, 75(1):1–??, October 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/1/1.

Williams:1977:CLB

[Wil77b] M. C. Williams. Conversion of lamellar body membranes into tubular myelin in alveoli of fetal rat lungs. Journal of Cell Biology, 72(2):260–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/72/2/260.

Wallace:1976:PII

[WJ76] R. A. Wallace and D. W. Jared. Protein incorporation by isolated amphibian oocytes. V. Specificity for vitellogenin incorporation. Journal of Cell Biology, 69(2):345–??, May 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/69/2/345.

Wakshull:1978:PAU

[WJB78] E. Wakshull, M. I. Johnson, and H. Burton. Persistence of an amine uptake system in cultured rat sympathetic neurons which use acetylcholine as their transmitter. Journal of Cell Biology, 79(1):121–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/79/1/121.

Whitesell:1977:MSG

[WJTR77] R. R. Whitesell, R. A. Johnson, H. L. Tarpley, and D. M. Regen. Mitogen-stimulated glucose transport in thymocytes. Possible role of Ca++ and antagonism by adenosine 3’: 5’-monophosphate. Journal of Cell Biology, 72(2):456–??, February 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/2/456.

Wolin:1979:EMN

[WK79] S. L. Wolin and R. S. Kucherlapati. Expression of microtubule networks in normal cells, transformed cells, and their hybrids. Journal of Cell Biology, 82(1):76–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/1/76.

Williams:1975:QSPa

[WKBL75a] K. E. Williams, E. M. Kidston, F. Beck, and J. B. Lloyd. Quantitative studies of pinocytosis. I. Kinetics of uptake of (125I)polyvinylpyrrolidone by rat yolk sac cultured in vitro. Journal of Cell Biology, 64(1):113–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/64/1/113.

Williams:1975:QSPb

[WKBL75b] K. E. Williams, E. M. Kidston, F. Beck, and J. B. Lloyd. Quantitative studies of pinocytosis. II. Kinetics of protein uptake and digestion by rat yolk sac cultured in vitro. Journal of Cell Biology, 64(1):123–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/1/123.

Wallimann:1977:LCKb

[WKP^{+}77] T. Wallimann, H. J. Kuhn, G. Pelloni, D. C. Turner, and H. M. Eppenberger. Localization of creatine kinase isoenzymes in myofibrils. II. Chicken heart muscle. Journal of Cell Biology, 75 (2):318–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/318.

Williams:1976:MPA

[WL76] J. A. Williams and M. Lee. Microtubules and pancreatic amylase release by mouse pancreas in vitro. Journal of Cell Biology, 71 (3):795–??, December 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/3/795.

Wirth:1979:RIS

[WLR79] D. F. Wirth, H. F. Lodish, and P. W. Robbins. Requirements for the insertion of the Sindbis envelope glycoproteins into the endoplasmic reticulum membrane. Journal of Cell Biology, 81(1):154– ??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/81/
1/154.

Weatherbee:1978:VPM

[WLW78] J. A. Weatherbee, R. B. Luftig, and R. R. Weihing. In vitro polymerization of microtubules from HeLa cells. Journal of Cell Biology, 78(1):47–??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/78/1/47.

Wang:1976:TSM

[WM76] Y. J. Wang and H. R. Mahler. Topography of the synaptosomal membrane. Journal of Cell Biology, 71(2):639–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/71/2/639.

West:1977:ICR

[WM77] C. M. West and D. McMahon. Identification of concanavalin A receptors and galactose-binding proteins in purified plasma membranes of Dictyostelium discoideum. Journal of Cell Biology, 74(1):264–??, July 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/74/1/264.

Warner:1978:SCC

[WM78] F. D. Warner and D. R. Mitchell. Structural conformation of ciliary dynein arms and the generation of sliding forces in Tetrahymena cilia. Journal of Cell Biology, 76(2):261–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/76/2/261.

Willingham:1979:AMB

[WMP79] M. C. Willingham, F. R. Maxfield, and I. H. Pastan. alpha 2 Macroglobulin binding to the plasma membrane of cultured fibroblasts. Diffuse binding followed by clustering in coated regions. Journal of Cell Biology, 82(3):614–??, September 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/82/3/614.

Weisenseel:1975:LEC

[WNJ75] M. H. Weisenseel, R. Nuccitelli, and L. F. Jaffe. Large electrical currents traverse growing pollen tubes. Journal of Cell Biology, 66 (3):556–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/556.

Weiser:1978:SPG

[WNQK78] M. M. Weiser, M. M. Neumeier, A. Quaroni, and K. Kirsch. Synthesis of plasmalemmal glycoproteins in intestinal epithelial cells. Separation of Golgi membranes from villus and crypt cell surface membranes; glycosyltransferase activity of surface membrane. Journal of Cell Biology, 77(3):722–??, June 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/3/722.

Woolley:1977:ETP

[Woo77] D. M. Woolley. Evidence for “twisted plane” undulations in golden hamster sperm tails. Journal of Cell Biology, 75(3):851–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/851.

Wade:1979:MCM

[WOPB79] J. B. Wade, R. G. O’Neil, J. L. Pryor, and E. L. Boulpaep. Modulation of cell membrane area in renal collecting tubules by corticosteroid hormones. Journal of Cell Biology, 81(2):439–??, May 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/81/2/439.

Willingham:1975:CAC

[WP75] M. C. Willingham and I. Pastan. Cyclic AMP and cell morphology in cultured fibroblasts. Effects on cell shape, microfilament and microtubule distribution, and orientation to substratum. Journal of Cell Biology, 67(1):146–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/1/146.

Weibel:1978:ISB

[WP78] E. R. Weibel and D. Paumgartner. Integrated stereological and biochemical studies on hepatocytic membranes. II. Correction of section thickness effect on volume and surface density estimates. Journal of Cell Biology, 77(2):584–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/584.

Wolosewick:1979:MLC

[WP79] J. J. Wolosewick and K. R. Porter. Microtrabecular lattice of the cytoplasmic ground substance. Artifact or reality. Journal of Cell Biology, 82(1):114–??, July 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/82/1/114.

Weiss:1975:GMC

[WPMW75] L. Weiss, G. Poste, A. MacKearnin, and K. Willett. Growth of mammalian cells on substrates coated with cellular microexudates. I. Effect on cell growth at low population densities. Journal of Cell Biology, 64(1):135–??, January 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/64/1/135.

Witman:1978:CFM

[WPS78] G. B. Witman, J. Plummer, and G. Sander. Chlamydomonas flagellar mutants lacking radial spokes and central tubules. Structure, composition, and function of specific axonemal components. Journal of Cell Biology, 76(3):729–??, March 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/76/3/729.

Weisenberg:1975:VPM

[WR75a] R. C. Weisenberg and A. C. Rosenfeld. In vitro polymerization of microtubules into asters and spindles in homogenates of surf clam eggs. Journal of Cell Biology, 64(1):146–??, January 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/1/146.

Wight:1975:PPAa

[WR75b] T. N. Wight and R. Ross. Proteoglycans in primate arteries. I. Ultrastructural localization and distribution in the intima. Journal of Cell Biology, 67(3):660–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/67/3/660.

Wight:1975:PPAb

[WR75c] T. N. Wight and R. Ross. Proteoglycans in primate arteries. II. Synthesis and secretion of glycosaminoglycans by arterial smooth muscle cells in culture. Journal of Cell Biology, 67(3):675–??, December 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/67/
3/675.

Weissfeld:1977:ADKa

[WR77a] A. S. Weissfeld and H. Rouse. Arginine deprivation in KB cells: I. Effect on cell cycle progress. Journal of Cell Biology, 75(3):881–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/75/
3/881.

Weissfeld:1977:ADKb

[WR77b] A. S. Weissfeld and H. Rouse. Arginine deprivation in KB cells. II. Characterization of the DNA synthesized during starvation. Journal of Cell Biology, 75(3):889–??, December 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/75/3/889.

Weissfeld:1977:CID

[WR77c] A. S. Weissfeld and H. Rouse. Continued initiation of DNA synthesis in arginine-deprived Chinese hamster ovary cells. Journal of Cell Biology, 73(1):200–??, April 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/73/1/200.

Woodrum:1975:EBB

[WRP75] D. T. Woodrum, S. A. Rich, and T. D. Pollard. Evidence for biased bidirectional polymerization of actin filaments using heavy meromyosin prepared by an improved method. Journal of Cell Biology, 67(1):231–??, October 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/67/1/231.

Walter:1978:CCC

[WS78a] M. F. Walter and P. Satir. Calcium control of ciliary arrest in mussel gill cells. Journal of Cell Biology, 79(1):110–??, October 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/79/1/110.

Wasserman:1978:CBC

[WS78b] W. J. Wasserman and L. D. Smith. The cyclic behavior of a cytoplasmic factor controlling nuclear membrane breakdown. Journal of Cell Biology, 78(1):??, July 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/78/1/R15.

Wang:1979:DFL

[WT79] Y. L. Wang and D. L. Taylor. Distribution of fluorescently labeled actin in living sea urchin eggs during early development. Journal of Cell Biology, 81(3):672–??, June 1979. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/81/3/672.

Wallimann:1977:LCKa

[WTE77] T. Wallimann, D. C. Turner, and H. M. Eppenberger. Localization of creatine kinase isoenzymes in myofibrils. I. Chicken skeletal muscle. Journal of Cell Biology, 75(2):297–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/75/2/297.

Williams:1975:PMC

[WW75] M. C. Williams and S. L. Wissig. The permeability of muscle capillaries to horseradish peroxidase. Journal of Cell Biology, 66 (3):531–??, September 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/66/3/531.

Wissig:1978:PMC

[WW78] S. L. Wissig and M. C. Williams. Permeability of muscle capillaries to microperoxidase. Journal of Cell Biology, 76(2):341–??, February 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/76/
2/341.

Willard:1979:ATA

[WWLS79] M. Willard, M. Wiseman, J. Levine, and P. Skene. Axonal transport of actin in rabbit retinal ganglion cells. Journal of Cell Biology, 81(3):581–??, June 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/3/581.

Willingham:1978:MDP

[WY78] M. C. Willingham and S. S. Yamada. A mechanism for the destruction of pinosomes in cultured fibroblasts. Piranhalysis. Journal of Cell Biology, 78(2):480–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/78/2/480.

Yamada:1978:ICM

[Yam78] K. M. Yamada. Immunological characterization of a major transformation-sensitive fibroblast cell surface glycoprotein. Localization, redistribution, and role in cell shape. Journal of Cell Biology, 78(2):520–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/78/2/520.

Yu:1975:DRP

[YB75] R. C. Yu and R. P. Bunge. Damage and repair of the peripheral myelin sheath and node of Ranvier after treatment with trypsin. Journal of Cell Biology, 64(1):1–??, January 1975. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/64/1/1.

Young:1979:MSE

[YBB79] R. B. Young, W. G. Bergen, and P. B. Blauwiekel. Myosin synthesis in embryonic chicken fibroblasts. Journal of Cell Biology, 81(1):115–??, April 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/81/1/115.

Yeoh:1978:DPA

[YGH78] G. C. Yeoh, D. Greenstein, and H. Holtzer. DNA polymerase activity in muscle cultures. Journal of Cell Biology, 77(1):99–??, April 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/77/1/99.

Yamada:1979:FCP

[YK79] K. M. Yamada and D. W. Kennedy. Fibroblast cellular and plasma fibronectins are similar but not identical. Journal of Cell Biology, 80(2):492–??, February 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/80/2/492.

Yee:1975:ECJ

[YR75] A. G. Yee and J. P. Revel. Endothelial cell junctions. Journal of Cell Biology, 66(1):200–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/200.

Yee:1978:LRG

[YR78] A. G. Yee and J. P. Revel. Loss and reappearance of gap junctions in regenerating liver. Journal of Cell Biology, 78(2):554–??, August 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/78/
2/554.

Yamada:1977:QTS

[YYP77] K. M. Yamada, S. S. Yamada, and I. Pastan. Quantitation of a transformation-sensitive, adhesive cell surface glycoprotein. Decrease of several untransformed permanent cell lines. Journal of Cell Biology, 74(2):649–??, August 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/74/2/649.

Zaar:1979:VPE

[Zaa79] K. Zaar. Visualization of pores (export sites) correlated with cellulose production in the envelope of the gram-negative bacterium Acetobacter xylinum. Journal of Cell Biology, 80(3):773– ??, March 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/80/
3/773.

Zigmond:1977:APL

[Zig77] S. H. Zigmond. Ability of polymorphonuclear leukocytes to orient in gradients of chemotactic factors. Journal of Cell Biology, 75 (2):606–??, November 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/75/2/606.

Zigmond:1978:CPL

[Zig78] S. H. Zigmond. Chemotaxis by polymorphonuclear leukocytes. Journal of Cell Biology, 77(2):269–??, May 1978. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/77/2/269.

Zeilig:1976:AMC

[ZJSF76] C. E. Zeilig, R. A. Johnson, E. W. Sutherland, and D. L. Friedman. Adenosine 3’: 5’-monophosphate content and actions in the division cycle of synchronized HeLa cells. Journal of Cell Biology, 71(2):515–??, November 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/71/2/515.

Ziegler:1976:CCB

[ZM76] D. Ziegler and Y. Masui. Control of chromosome behavior in amphibian oocytes. II. The effect of inhibitors of RNA and protein synthesis on the induction of chromosome condensation in transplanted brain nuclei by oocyte cytoplasm. Journal of Cell Biology, 68(3):620–??, March 1976. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/68/3/620.

Zanetti:1979:EDC

[ZMW79] N. C. Zanetti, D. R. Mitchell, and F. D. Warner. Effects of divalent cations on dynein cross bridging and ciliary microtubule sliding. Journal of Cell Biology, 80(3):573–??, March 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/3/573.

Zigmond:1979:SAL

[ZS79a] S. H. Zigmond and S. J. Sullivan. Sensory adaptation of leukocytes to chemotactic peptides. Journal of Cell Biology, 82(2):517–??, August 1979. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/82/
2/517.

Zweig:1979:CIE

[ZS79b] S. Zweig and S. J. Singer. Concanavalin A-induced endocytosis in rabbit reticulocytes, and its decrease with reticulocyte maturation. Journal of Cell Biology, 80(2):487–??, February 1979. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/80/2/487.

Zimmerman:1975:MEP

[ZT75] B. L. Zimmerman and M. O. Tso. Morphologic evidence of photoreceptor differentiation of pinealocytes in the neonatal rat. Journal of Cell Biology, 66(1):60–??, July 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/66/1/60.

Zampighi:1975:CBT

[ZVR75] G. Zampighi, J. Vergara, and F. Ramon. On the connection between the transverse tubules and the plasma membrane in frog semitendinosus skeletal muscle. Journal of Cell Biology, 64(3):734–??, March 1975. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/64/3/734.

Zeligs:1977:UBP

[ZW77] J. D. Zeligs and S. H. Wollman. Ultrastructure of blebbing and phagocytosis of blebs by hyperplastic thyroid epithelial cells in vivo. Journal of Cell Biology, 72(3):584–??, March 1977. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/72/3/584.